A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1990–1994

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.07

Title word cross-reference

(1 + 1) [913, 1687, 1722, 2004, 142]. (1|1) [2026]. (2 + 1) [1654, 1293, 2215, 1414, 2039, 426, 1789, 1482, 2175, 1197, 1422, 1177, 397, 955]. (2, 0) [1462]. (3 + 1) [1268, 2182]. (k_{1} − k_{n}) [676]. (N, M) [2234]. (N ≥ 3) [949]. (SU_{2})^{3} [491]. (x · ∇)_{ord}^{2} [577]. ∗ [845, 595, 1385]. 0 < c < 1 [1646]. 1 [826, 1078, 539, 1534, 82]. 1 + 1 [2152, 688, 1613, 2050, 1508]. 1 + 2 [1702, 2271, 711]. 1/2 [82, 362, 1263, 1303, 1534]. 2 [882, 1053, 1396, 82, 630, 1886]. 2 + 1 [469, 473, 1421, 1508]. 2 + 2 [2056]. 2 × 2 [2063, 1587]. 2 × 3 × n [1480]. 3 [317, 2295, 2065, 1862, 1015]. 3 + 1 [1522, 870]. 3/2 [82]. 4p [1527]. 5 [1349]. 6 [2196, 1343, 2150]. [sl(2, C)] [1054].

∗ [2061, 484]. ^{3} [492]. ^{∘} [2169, 904]. [1489]. _{2} [492]. _{60} [1571]. _{7} [492]. _{q} [370]. A^{(1)}(1, 0) [1191]. A_{∞} [161]. A_{l} [1359]. A_{n−1} [886, 927]. B_{q}(n) [1589]. C [566, 814, 1066, 1630, 1299, 1494, 1813]. C(n + 1) [277]. C(SU_{q}(2)) [1426]. C^{(0)} [1841]. C^{∗} [546, 835, 2146, 1384, 696, 1133, 866]. C^{r} [987]. c_{1} [1367]. C_{n} [886]. C_{q}(n) [1589]. H [1211]. Ř [561]. CP [1347, 1949]. CP^{n} [228]. D [79, 1173, 1544, 113, 1270, 464, 2060]. D(2, 1; α) [971]. d = 1 + 1 [841]. D = 4 [2028]. D^{+}D^{−} [2058]. d^{2}ψ/dz^{2} − 4A^{2}z^{n−2}(z^{n} + c)ψ = 0 [1018]. D_{5} [1245]. δ [1071, 398, 2122]. df(x) = [f(x + b) − f(x)]_{ν}^{μ}da_{μ}^{ν} [1514]. D ≥ 2 [1500]. diff(S^{1}) [1907]. E(2) [1063]. E(3)_{q} [603, 1547]. E(3|2) [1782]. E_{(}n) [528]. E_{6} [306, 488, 339]. E_{7} [561, 562]. E_{8} [981]. E_{q}(2) [1976]. ϵ(ω)μ(ω) = 1 [217]. η [2202, 2322]. F (4) [971]. F (A_{1}) [2197]. f(∞) [1971]. f(x) [1971]. F_{4} [561, 562]. Fun_{q}(SU(2)) [1384]. G [1206, 1285]. G(3) [971]. G^{∞} [1638]. G_{2} [1823, 488]. G_{2}II [1209]. G_{q}(2) [1589]. GD(SL_{n}) [2071]. gl(2) ⊕ gl(2) [148]. gl(2/2) [148]. gl(∞) [99]. gl(m|n) [406, 371]. GL(N) [849]. gl(N, R) [2115]. gl(n|1) [230]. gl_{∞} [94]. GL_{p,q}(2) [1287]. GL_{p,q}(2, C) [1298]. GL_{q}(N) [2032]. GL_{q}(r, C) [2254]. H(1)_{q} [602]. H = P^{2} + 2X^{2} + 22λX^{3} + λX^{4} [177]. H_{n}(q) [927]. h_{n,2n−i}(LM|l)/(2n − i)! [335]. R̂ [562]. IU(n) [610]. j [643, 2196, 1343, 2150]. J_{ν}(x) [1198]. jl(2, C) [1237, 1523]. K [552, 145, 1090, 1370, 1830, 2240]. κ [2195]. L [2259, 1144]. L^{2} [623]. L^{2}(R) [65]. λ ≪ 1 [177]. Λ ≠ 0 [581]. l ≥ 2 [1359]. m [1939, 94, 1495, 2033]. M^{2} [1901]. M_{q}(2) [1205]. m × n [709]. N [501, 949, 1382, 12, 1305, 1467, 1942, 688, 780, 700, 589, 965, 134, 868, 208, 1251, 285, 1439, 675, 268, 1271, 1104, 1572, 866, 1805, 1885, 1567, 77, 1908, 528, 1705, 222, 2122, 1103, 250]. N + 1 [1066, 1630]. n − 1 [97]. n = 1 [954]. N = 2 [785, 1483, 2056, 1901, 1598, 147, 2026, 1006, 1579, 571, 1273, 818, 1693]. N = 3 [1623]. N ≥ 2 [52]. O(2) [216]. O(2, 1) [1276]. O(3) [195]. O^{∗} [738, 2061]. osp(1, 2) [1526, 2075]. OSp(1/2) [1825]. osp(1|2) [2021, 2104]. OSp(2m ∗ /2n) [527]. osp(2|2, R) [1498]. osp(3, 2) [642]. osp(3/2) [1134]. OSP(3|2) [1701, 2021]. OSp(3|2)/OSp(2|2) [1396]. osp(∞|∞) [331]. osp(m/2n) [21]. OSP(N, 4) [52]. OSp(n|1) [559]. P [1151, 1548, 1549, 1924, 828, 829, 1140, 39, 414, 2168, 573, 1108, 1109, 1645, 460, 948]. P(n) [890, 1250, 1353]. p4gm [1698]. p_{1}^{2}p_{2}^{2} ⋯ [102]. pg [1698]. Φ^{4} [917]. Q [197, 2198, 1593, 2074, 2187, 1253, 610, 2142, 2076, 886, 2143, 642, 1905, 2252, 1395, 1256, 1074, 974, 2106, 1205, 1186, 1784, 1975, 2260, 2305, 2027, 1232, 1437, 1552, 2193, 1319, 2041, 1673, 526, 783, 2032, 1648, 2307, 1535, 1869, 1354]. Q(n) [1353]. q^{p} = 1 [974]. Q_{p} [1487]. R [2170, 603, 1327, 536, 2024, 2259, 1094, 1029, 783, 1647, 1494, 889, 1858, 93, 1487]. R + R^{2} + Q^{2} [2248]. R^{3} [342]. R^{n} [1159]. K [868]. S [2077, 400, 351, 1180, 736]. S^{2} [1521]. S^{7} [969]. S^{D} [1605]. S_{N} [236, 197]. SA(4, R) [1219]. A/G [143]. S [1440]. sd [682]. σ [1649, 115, 142, 326]. SIM(d − 1, 1) [480]. sl(1|2) [2104]. SL(2, C) [297, 441, 1054, 1269, 1237, 1523]. SL(2, R) [302, 1342, 2100, 1684, 1106]. SL(2, R)/U(1) [1816, 1817]. sl(3) [437, 95]. sl(3, C) [769]. SL(3, R) [1099, 275]. sl(m/n) [979, 333]. SL(N) [2236]. SL(n + 1, C) [1772]. SL(N, R) [1311]. sl_{3} [160]. SL_{q}(2) [2295, 783]. sl_{q}(n + 1) [783]. so(1, 2)∗ [2114]. SO(2, 1) [1303]. SO(2, 1)/SO(2) [1187]. SO(3) [1602, 525, 928, 373, 637, 2069]. so(3) ⊕ sp(2) [1134]. SO(3, 1)_{q} [1547]. so(4)^{∗} [1760]. SO(4)_{q} [1547]. so(5) [2069]. SO(5)_{q} [2142]. SO(8) [605, 1489]. so(n) [1158]. SO(n, 1) [607]. SO_{0}(1, 4) [1783]. SO_{7} [491]. SO_{n+2} [978]. SO_{n+2} ⊃ SO_{n} × SO_{2} [978]. SO_{n} × SO_{2} [978]. so_{q}(3)

[1076]. Sp(2) [396, 515, 797]. sp(2n, R) [1064]. Sp(4) ⊃ U(2) [1286, 1853]. Sp(6, R) [1716]. SU(1, 1) [1279, 258, 1326, 1495]. su(1, 1)^{∗} [1760]. su(1, 1)_{q} [2258]. SU(2) [370, 144, 145, 2042, 786, 1343, 397, 1609]. su(2)^{∗} [1760]. SU(2)_{q} [1786, 162]. SU(2, 1) [234, 1249]. SU(2, 2) [2045]. SU(2, C) [187]. SU(3) [199, 1146, 2296, 1824, 558, 227, 118, 550, 1249]. SU(3) ⊃ SU(2) × U(1) [404]. SU(3/1) [8, 598]. SU(∞) [1449]. SU(N) [1730, 2150, 97]. SU(n) ⊃ SO(n) [522, 1286]. SU(N/M) [524]. SU(p, q) [1770]. SU(u, v) [726]. SU ∗ (∞) [2056]. SU_{2} [1989]. SU_{n} [58]. su_{q}(1, 1) [1546, 1074, 805]. SU_{q}(2) [2256, 1232, 1437, 1319, 1521, 1076, 805, 1831]. SU_{q}(4) [2151]. su_{q}(m|n) [1327]. SU_{q}(N) [1709, 1710]. T^{3} × R [92]. τ [852, 2051]. Θ [1020, 1019]. U [536]. U(1) [1536]. U(1, 1) [1394]. U(1/1) [1588]. U(2) [1992, 2311, 295]. U(3) [404, 1097]. u(5) ⊃ so(5) ⊃ so(3) [2070]. U(N) [1409, 1408, 596, 236, 850, 608, 237]. U(n, 1) [610, 607]. U(p, q|N) [728]. U(N) [1940]. U_{2} × U_{2} [1145, 200]. U_{4} [1145, 200]. U_{p+q} ⊇ U_{p} × U_{q} [200]. U_{p+q} ⊃ U_{p} × U_{q} [1145]. U_{q} [1077]. U_{q}(C_{2}) [1911]. U_{q}(gl(m/n)) [1829, 1144, 1496]. U_{q}(gl(n)) [1519]. U_{q}(osp(1|2)) [2067]. U_{q}(sl(3)) [2252]. U_{q}(so_{2,1}) [2106]. U_{q}(so_{3,1}) [2106]. U_{q}(su(1, 1)) [2307]. U_{q}(su_{2}) [403]. U_{q}(u_{r+s}) [1141]. U_{q}(u_{r,s}) [1141]. U_{q}[gl(2/2)] [2027]. U_{q}[gl(n)] [1036, 1037]. U_{q}[osp(1, 2)] [1526]. U_{q}[osp(3/2)] [1751]. U_{q}sl(3) [729]. U_{q}sl(n, m) [2024]. u_{t} = u_{xxx} + G(u, u_{x}, u_{xx}) [985]. u_{x} = (v_{xx})^{2} [733]. u_{xt} = F (u, u_{x}) [209]. u_{xy} = f(u, u_{x}) [877]. u ∧ k ∧ l = 0 [1551]. u ∧ k ∧ l ≠ 0 [1551]. V (r) = ar^{2} + br^{−4} + cr^{−6} [32]. v_{tt} = f(x, v_{x})v_{xx} + g(x, v_{x}) [851, 1666]. φ^{3} [675]. φ_{x} = λJφ + P φ [240]. W [2056, 2234, 1017, 1468, 1826, 2026]. W(D_{n}) [303]. W^{∗} [2270]. W_{4} [1693]. W_{n} [2071]. W_{q}(1/1) [1751]. Ŵ_{∞} [1816, 1817]. X [1963]. x = 0 [1971]. XY [1044, 1357]. y′′ = F (x)y^{n} [1361]. z [1407]. Z_{2} [408]. Z_{3} [2020, 972]. Z_{p} [1044]. |kq{K̃.}(k_{1} − k_{n}): [λ̃], S_{n}⟩⟩ [676]. |q| = 1 [1394]. |t|^{2} − |x|^{2} [1970].

-adic [1548, 1549, 1924, 573, 1108, 1109, 1645, 460, 948]. -adics [2168]. -algebra [2074, 1186, 1784, 1975, 2234, 1384, 1494, 1133]. -algebras [845, 595, 484, 738, 2061, 1826, 2146, 866]. -analog [1232, 1437, 526, 2032]. -analogs [610]. -analytic [2198]. -automorphisms [2061]. -Bessel [1256]. -beta [2143]. -body [1805, 1885, 1467, 949, 12, 1305, 1942, 134, 868, 866]. -boson [974, 1205]. -branes [414]. -Campbell [2260]. -constrained [2240]. -covariant [396, 515, 797]. -D [1078, 317, 826, 882, 1053, 1349, 1015, 630]. -Deformation [2076, 1869]. -deformed [1253, 886, 2106, 2041, 783, 1648, 1535]. -dependent [623]. -difference [1905, 2252]. -dimensional [1654, 1293, 1268, 2215, 913, 1687, 1414, 1722, 2182, 528, 1705, 2039, 2004, 222, 426, 2122, 142, 2175, 1103, 955, 250, 1270, 464]. -dimensions [1177]. -dynamical [696]. -elliptic [1206]. -equivalence [736]. -Euclidean [2193]. -expansion [398, 2122]. -fold [780]. -forms [828, 829, 1140]. -function [1939, 852, 1019, 1020]. -gamma [2143]. -gauge [145]. -Gegenbauer [1074]. -generic [2142]. -graded [2020, 972, 408]. -group [726]. -integrable [566, 814, 1066, 1813, 1630]. -integral [1975]. -integration [2305]. -invariance [1523, 1237]. -invariant [1523, 1054, 1269, 1237, 1311]. -invariants [2202, 2322]. -led [1963]. -like [1276]. -matrices [1094, 783]. -matrix [2170, 603, 1327, 868, 2259, 400, 1029, 1180]. -multiplication [2143]. -operator [2259]. -orbifold [2065]. -order [285]. -oscillator [642, 1784]. -perturbation [1071]. -Poincaré [1552, 1830, 2195]. -point [589, 965, 1104, 1572]. -principal [94]. -representations [1385]. -rung [675]. -self-dual [1367]. -sheeted [1495, 2033]. -soliton [688, 780]. -space [501, 552]. -sphere [1396]. -statistical [1299]. -symmetric [1063]. -symmetries [2187]. -symmetry [2071]. -symplectic [1370]. -symplecton [1319]. -systems [546, 835, 1090]. -term [93]. -theorem [1211]. -torus [268].

1408-1426 [2113].

26 [3, 1317]. 27 [57]. 28 [219, 1317, 2264]. 29 [56].

30 [1467, 123, 520, 166, 598, 226, 1672, 521, 337, 153]. 31 [1145, 723, 677, 1658, 1251, 597, 532]. 32 [1938, 1140, 1250, 1098]. 33 [1466, 1522, 1715, 1372, 1468, 2320, 1523, 1870, 1282]. 34 [1714, 2318, 2112, 2309, 1789, 1755, 1871]. 35 [2317, 2316, 2313, 2322, 2319, 2318, 2310, 2311, 2312, 2265, 2314, 2113].

41 [2321].

58204a [166].

83046a [219]. 86h [3]. 87f [57]. 88a [1317]. 89c [2264].

90a [598, 56]. 90d [1672]. 90e [219]. 90h [1467]. 90i [520]. 90j [337]. 90m [166]. 91d [1145, 597]. 91e [723]. 91h [677, 532]. 91m [1658]. 92a [1251]. 92e [1938]. 92m [1372, 1468, 1282]. 93e [1523]. 93h [1715]. 93j [1714, 1870, 1755]. 93m [2318]. 94e [2309]. 94g [2112]. 94h [1789]. 95d [2317, 2316]. 95e [2311, 2312]. 95f [2313]. 95g [2318, 2310]. 95h [2319]. 95i [2314].

Abel [973]. Abelian [1091, 1769, 141, 514, 1912, 2125, 487, 1799, 982, 549, 1999, 938, 63]. Abelization [188]. Ablowitz [1654, 1848]. absence [1741, 664]. absolute [1690]. absorption [1407]. abstract [2262, 1936, 1970, 2113]. Abundant [2004]. ac [315]. accelerated [5, 2264]. accelerations [131, 1818]. Accuracy [378, 1490, 613]. Accuracy-weighted [378]. acoustic [1884, 2088, 937]. acoustics [1138]. acting [2259]. action [106, 635, 2203, 1317, 130, 879, 896, 1734, 171, 538]. action-at-a-distance [1734]. actions [2189, 394, 3, 2221, 1]. Addendum [2112, 2309, 226, 1789, 56, 1672]. Addition [1256, 1781, 1975, 445, 2029]. Additive [154, 2083]. adèles [2168]. Adiabatic [1873, 639, 1436, 1921, 247, 1835, 827, 2121]. adic [1548, 1549, 1924, 573, 1108, 1109, 1645, 460, 948]. adics [2168]. Adjoint [227, 2150, 1644, 530]. adjointness [739, 1321]. Adler [1541]. admitting [2095, 107, 1478, 184, 470, 1365]. advection [882]. Affine [1122, 1926, 2244, 472, 115, 2111, 2187, 1184, 1202, 1466, 2297, 272, 1245, 107, 2145, 2302, 257, 752, 1326, 278, 544, 1218, 1191, 617, 1391, 269, 600, 1344, 1832]. Aharonov [1940, 1759, 744]. Akhiezer [1920]. AKNS [655, 313, 1322]. albedo [482]. ALE [1123]. algebra [1546, 2074, 1483, 1904, 2234, 155, 1857, 296, 672, 2075, 319, 2102, 2019, 647, 523, 536, 1905, 1515, 1906, 844, 2254, 1973, 1017, 1468, 160, 2257, 2258, 1205, 1473, 1559, 403, 369, 441, 1666, 1783, 1244, 1054, 258, 1186, 1784, 1975, 733, 1667, 1706, 1786, 437, 95, 1094, 90, 597, 1491, 636, 1830, 1976, 1384, 1102, 1101, 1605, 99, 161, 1751, 1709, 1710, 2068, 1303, 928, 1046, 1831, 2071, 1428, 962, 526, 783, 1494, 1648, 1739, 676, 600, 301, 1535, 2002, 298, 809, 1133, 1911, 2151, 1816, 1817, 1693, 1869, 806, 194, 914]. Algebraic [2189, 871, 1297, 2078, 1054, 1269, 2083, 2305, 1913, 1102, 1962, 2134, 1729, 1935, 2239, 286, 2279, 1279, 2220, 1548, 413, 864, 147, 754, 830, 1142, 1815, 1685, 338, 324]. algebraically [2057]. Algebras [340, 268, 1470, 2267, 738, 2061, 564, 1382, 2062, 2187, 1184, 1822, 2297, 1100, 270, 272, 968, 782, 1359, 1485, 1486, 565, 2036, 1663, 227, 1929, 66, 845, 19, 2221, 1973, 595, 163, 201, 2066, 62, 405, 974, 1254, 2106, 1826, 1105, 2302, 246, 560, 1076, 1075, 1722, 694, 1383, 1655, 120, 766, 2146, 484, 1861, 777, 686, 926, 980, 1186, 972, 927, 1141, 1961, 976, 1463, 2028, 1196, 734, 1299, 1553, 1249, 1670, 1107, 1357, 507, 1554, 1635, 93, 899, 678, 679, 99, 161, 2197, 442, 2306, 1674, 2206, 198, 1980, 408, 1344, 1206, 202, 1868, 2263]. algebras [730, 2111, 1832, 1593, 1928, 866]. algorithm [311, 2086, 654, 557, 2068, 1962, 2134, 2149]. Algorithmic [648]. Aligned [354]. allied [1995]. allowing [1209]. Almost [1904, 2241]. almost-Robertson [2241]. along [1381]. alpha [335, 1399]. already [548]. alternate [381]. Alternative [1646, 822, 688, 1781, 19]. ambiguity [1877, 101]. Amir [2316]. among [1180, 1734, 1343]. amplitude [579]. Amplitudes [743, 838, 998, 1562]. analog [2088, 2274, 1232, 1437, 526, 2032]. analogs [610]. Analysis [2255, 1808, 2003, 1108, 1109, 388, 2091, 2233, 448, 934, 1792, 311, 35, 1124, 2135, 1009, 1014, 36, 1650, 2145, 389, 1587, 23, 1355, 1219, 1505, 222, 636, 648, 64, 108, 2175, 898, 574, 307]. Analytic [841, 994, 760, 1686, 1871, 127, 2100, 412, 1604, 368, 1332, 2198, 1968, 1040, 32, 1928]. Analytical [1513, 1288, 989, 1368, 1182, 1058, 1909]. Anandan [1940]. angles [247]. Angular [190, 1586, 445]. anharmonic [2268, 1253, 1401, 1149, 2005, 906, 1443, 1233, 1386, 1765, 951]. anharmonic-oscillator [2268]. anharmonicities [1149, 32]. Anisotropic [352, 218, 75, 1765]. anisotropy [1009]. annihilation [1095, 1341, 1235, 1106]. Anomalies [343, 1265, 880, 1153, 1681, 2245].

Anomalous [489, 392, 909]. Anomaly [1966, 725, 1034, 1912, 2138, 1237, 1523, 1006, 1056, 2042, 1968]. Anosov [2221]. Ansatz [397, 718, 1913, 2236, 774]. Ansätze [865]. anti [329, 649, 2114, 829, 1140, 1644, 52, 427]. anti- [2114, 427]. anti-de [1140, 829]. anti-de-Sitter [52]. anti-self-adjoint [1644]. anti-self-dual [329, 649]. antiferromagnets [753]. antifield [1560]. antilinear [365]. antipode [1428]. antisymmetrized [1917]. antisymmetry [282]. any [2271, 990, 2191, 1160, 2068, 1936, 2113]. anyspace [2110]. aperiodic [1351]. apparent [503]. Application [1566, 189, 159, 16, 521, 388, 739, 1198, 685, 1920, 1665, 1355, 1060, 528, 1850, 1147, 932, 1429, 711, 1302]. Applications [580, 350, 1744, 1629, 1737, 2103, 1257, 2255, 1258, 1285, 1726, 1068, 2197, 975, 910, 526, 494, 2045]. applied [645, 1591, 2290]. approach [517, 1540, 1279, 1620, 1621, 1556, 856, 1161, 2170, 1091, 2208, 804, 1596, 1327, 1346, 1719, 761, 2252, 2253, 868, 891, 1840, 375, 1808, 1732, 2003, 1991, 2214, 414, 1815, 251, 750, 626, 1914, 2275, 1296, 619, 1887, 307, 364, 2227, 2314, 744, 1739, 1891, 762, 1917, 96]. approaches [864]. appropriately [2146]. approximant [1198, 651]. Approximate [579, 1332, 1138, 895, 616, 1801, 1713]. approximated [2146]. approximately [282]. approximation [1639, 376, 869, 179, 2162, 693, 1998, 2290, 802, 1950]. approximations [1828, 1591, 611, 1363]. Araki [815]. arbitrarily [952]. arbitrariness [1866]. Arbitrary [1444, 700, 2282, 2319, 1165, 2137, 1926, 1544, 2092, 2294, 2030, 280, 1441, 1511, 1087, 1354]. Area [563, 768]. Area-preserving [563, 768]. arising [1354]. Arithmetic [1053, 708]. arithmetical [519]. Arnol′d [1090, 1659]. array [1814]. article [1812]. ary [1908]. Ashtekar [1171, 1240]. Askey [737, 773]. aspect [1690]. aspects [1034, 125, 911, 933, 1129, 1056, 2042, 1204, 1189]. Associated [570, 519, 297, 2100, 2234, 771, 462, 1199, 1018, 1833, 168, 622, 205, 655, 660, 973, 2053, 1016, 2206, 1197, 496, 617, 1065, 1693]. associativity [1432]. Asymmetric [951, 2168]. Asymptotic [1641, 2094, 1735, 1984, 1278, 1529, 463, 837, 759, 388, 1971, 1152, 215, 1934, 937, 533, 409, 574, 1591, 1112, 249]. Asymptotically [1217, 1742, 1192]. Asymptotics [1085, 2261, 2051, 692, 2153, 361, 2023, 740, 461, 1882]. Atiyah [1915, 1916]. atmosphere [1743]. atom [1758]. atomic [1399]. atoms [1137, 2225, 1113]. attractors [1844]. atypical [406]. austerity [876]. Autocatalytic [1777]. automata [1690]. automorphism [1718, 2034, 1936, 2113]. Automorphisms [372, 2061, 1670, 1593]. autonomous [254]. Auxiliary [206]. average [300]. averages [72]. Averaging [167, 812]. Avoided [853, 1835]. axes [2253]. axial [394]. Axially [1527, 1030]. axis [931, 2294]. axisymmetric [1173, 540, 2139, 2099, 1927].

B [57, 219]. background [1433, 2014, 549, 1465]. backgrounds [1620, 1621]. Bäcklund [2017, 440, 2172, 2008, 2174, 380, 817, 419, 507, 415, 1360, 169, 877, 313]. bag [149, 49, 759, 810]. Baker [2260, 1920, 1424, 497]. Balance [310]. ball [1936, 2113]. balls [1581]. Banach [1548]. Band [1698]. bands [1472, 1331, 483]. bang [665]. bar [1944]. Bardeen [1473, 1686, 1871]. Bargmann [146, 1474]. barotropic [152]. barrier [1121]. Barut [940]. baryon [748, 1300, 1301, 1302]. based [2179, 2241, 2105, 1615, 1107, 1338, 600, 1512, 1185, 1770]. bases [1286, 1853, 2201, 446]. Basic [46]. basis [1145, 2093, 2142, 163, 2223, 1134, 489, 148, 200, 491, 492, 978, 1709, 1710, 608, 2070, 2069]. Batalin [2036]. Baxter [782, 2018, 1359, 622, 1530, 756, 1455, 1463, 1107, 1636, 1637, 34, 1494]. Baxterization [2258, 1454]. be [582, 1328, 1090]. Beals [2107]. Becchi [2036, 1929]. behavior [2129, 758, 908, 463, 2147, 1963, 947]. behaviors [2094]. Bekenstein [707]. Bel [92]. Belavin [2259]. Bell [1756, 1164]. Bellert [1204]. Beltrami [1790, 369, 1462, 555]. Benney [87]. Bergmann [2287]. Bernstein [1263]. Berry [1034, 1033, 1203, 1940, 284, 63]. Bertrand [677, 204]. Bessel [1047, 930, 2103, 1256, 1587, 531, 1147, 295, 550]. beta [2143]. Bethe [592, 362, 2236]. between [1972, 1940, 1857, 588, 2253, 1105, 1652, 1436, 1153, 997, 236, 241, 2162, 1683, 532, 860]. Beyond [194, 2252, 1473, 781, 811, 931]. bi [1157, 1228, 129, 68, 33, 1504, 2010, 732, 1740, 1333, 1275]. bi-Hamiltonian [1157, 1228, 129, 68, 33, 2010, 732, 1740, 1333, 1275]. bi-isotropic [1504]. Bianchi [1898, 467, 254, 1124, 2135, 1551, 752, 1022, 789]. biased [1258]. Bicovariant [2295, 1287]. bidimensional [654]. bidirectional [509]. Biedenharn [2196]. Bifurcation [2269, 1844]. big [665]. big-bang [665]. bilateral [1982, 2313]. bilinear [2132, 845, 375, 2174, 1252, 808]. bilinears [1608]. billiard [1567, 1995, 2120, 1878]. billiards [1568, 249]. binomial [1749]. Biorthogonal [483, 1286, 983, 1974, 233]. biquadratic [1129]. Birkhoff [1578, 220]. Birman [1094]. Bispinor [292]. bivectors [1503]. BKP [916]. black [2057, 2181, 114, 2293, 541, 1349, 1511, 954, 707]. Bleuler [776, 274]. Bloch [518]. Bloch-wave [518]. Block [132, 881]. Block-diagonalization [132]. Blow [187]. Blow-up [187]. Bluman [1736]. bodies [253, 1039, 994, 1282, 961]. body [2219, 949, 594, 1805, 1885, 12, 1305, 1467, 1942, 293, 134, 1414, 1214, 868, 1477, 251, 461, 1235, 827, 445, 921, 1180, 1565, 1533, 494, 1097, 866]. Bogoliubov [701, 1837, 1762]. Bogomol′nyıequations [1392]. Bohm [1759, 744]. Bokhari [2316]. Boltzmann [2179, 2241, 55, 1211, 984, 913, 430, 1127]. Bonnet [1488, 183]. Bonnet-type [183]. Borel [1110, 713, 1865, 1711]. Borel-like [1865]. Born [1033, 1352, 179]. Bose [231, 2006, 524, 246, 1794, 1768, 2261, 1913, 1320, 336]. Boson [1526, 1906, 748, 1300, 1301, 1302, 488, 388, 2102, 2034, 2021, 1095, 62, 405, 551, 974, 2258, 1205, 2098, 1837, 1762, 1255, 2163, 526, 783, 2150, 429, 2044]. Boson-fermion [1526, 748, 1300, 1301, 1302, 551, 2258, 2044]. bosonic [838, 2226, 2071, 1133]. bosonization [1010, 1433, 754]. bosons [122]. bottom [949]. bottomed [2127]. Bound [2266, 1403, 1500, 1597, 122, 178, 1969, 267, 1950, 1234, 620]. boundaries [130]. Boundary [2126, 1809, 1083, 1182, 1151, 1629, 1699, 453, 1941, 580, 876, 2262, 691, 556, 1217, 922, 137, 496, 778, 342]. Bounded [772, 2049, 280]. Bounding [1207]. Bounds [867, 1137, 2225, 178, 1401, 1956, 997, 895, 1828, 101, 906]. Boussinesq [205, 2291, 1740]. Boussinesq-like [2291]. box [1643]. brachistochrone [873]. bracket [516]. brackets [717, 716, 1915, 1154, 1809]. Braid [2062, 622, 1354, 2302, 1455, 95, 2206, 1220]. Braided [1552, 1661, 2025, 2305, 887, 1749, 1491, 2263]. braiding [2083]. braidings [2030]. Branching [406, 303, 1483, 1278, 1105]. branes [414]. Bravais [722]. breaking [1925, 757, 591, 306, 1404, 360, 1096, 1189]. breathing [49]. Breit [922]. Bretón [1755, 1755]. bridge [841]. brightness [1058]. Brillouin [2275]. Broer [2238]. broken [505, 1246]. Brownian [61, 1271]. BRS [188, 83]. BRST [901, 224, 20, 723, 776, 1238, 2278, 1683, 1022, 2043, 2239]. BRST-like [1022]. building [397]. built [525]. bulk [218]. Bundle [982, 2186, 1860, 2245, 206, 90, 597, 511, 798, 13, 153, 1377, 2178, 433]. bundles [724, 24, 1940, 1153, 1521, 1329, 1675, 1676, 1475]. Burgers [453, 205, 1894].

C [226, 1571]. C. [2053]. calculate [2068]. calculating [557, 335]. calculation [1974, 1341, 1518, 1798, 335, 1399, 286]. calculations [1207]. calculus [2295, 1110, 1904, 929, 1585, 2299, 421, 1318, 2109, 1269, 1667, 1749, 1101, 184, 1287]. Calogero [2237, 2000, 2054]. Campbell [2260, 497]. can [1090]. cancellation [1966]. Canonical [476, 1547, 649, 1701, 717, 1259, 1734, 1801, 1064, 1072, 1470, 1146, 2249, 1292, 396, 20, 2296, 448, 2201, 1902, 150, 196, 1960, 2133, 1531, 1890, 2134, 2289, 2072]. Cantor [1331, 1944]. Cantor-set-like [1331]. caricature [1843]. Carnevale [1894]. carry [1651]. Cartan [116, 394, 1461, 1269, 670, 1171, 1042, 493, 1754, 1512, 183]. Cartesian [1330, 555, 878]. case [2129, 2087, 1574, 844, 1031, 98, 407, 2071, 1816, 1817, 1004, 430]. Casimir [1823, 1378, 217, 10, 123, 2302, 1219, 1829, 879, 729, 1143, 464, 429, 1501]. Casorati [1774]. cat [1090]. categories [2201]. cationic [2128]. Cauchy [337, 17, 584, 266, 1012, 1148]. Caudrey [2107]. causal [204, 677, 1002, 136, 884]. causality [261]. causally [137]. caused [568]. Caustics [2050]. cavities [952]. cavity [321, 1802]. Cayley [1972, 53, 1747, 156, 157, 158, 560, 1075, 676]. Cech [1153]. Cech-cohomology [1153]. Celestial [1427]. cell [1243]. cellular [1690]. censorship [1172]. Central [644, 1635, 2153, 2297, 885, 1159, 1232, 1437, 2163, 822, 623, 1565]. centrifugal [207]. centro [310]. certain [697, 2115, 450, 70, 1832, 1225]. chain [1009, 1845, 1357, 1129, 2236]. chains [2128, 279, 81]. Chandra [1980]. Chandrasekhar [36, 1065]. change [2209, 2215]. chaos [1009, 1921, 2227, 2314, 803]. chaotic [2290]. Character [333, 109, 118, 715, 1343]. Characteristic [399, 767, 1761, 995]. Characterization [564, 2097, 618, 1756, 1000, 381]. Characterizations [1385, 1646]. characters [601]. Charge [545, 372, 885, 629]. charged [334, 1735, 2229, 1173, 540, 1235, 658, 1726, 1738, 1180, 1734, 1511, 638, 1038]. charges [1013]. Chern [24, 1900, 2203, 1488, 752, 2125, 473, 1879, 330, 549, 397, 1888, 1083]. Chevalley [2142]. Chiral [880, 1034, 1314, 149, 1376, 800, 1237, 1523, 1398, 817, 1311, 2126, 1177, 516]. chiral-bag-like [149]. Choubey [57]. chromodynamics [1730]. circle [588, 23]. Circular [1995, 999]. Circular-sector [1995]. Clarkson [1736]. class [1513, 2130, 297, 2100, 348, 1202, 1466, 1804, 2285, 916, 1124, 867, 1649, 1381, 858, 2105, 935, 1195, 820, 406, 1361, 2283, 1908, 795, 875, 48, 1149, 206, 439, 262, 76, 1015, 145, 52, 1405, 1581, 70, 2306, 1039, 877, 1218, 338, 65, 1040, 471, 1225]. classes [1288, 2049, 823]. Classical [334, 314, 205, 1481, 859, 2236, 1479, 63, 695, 2152, 718, 618, 504, 1805, 1885, 886, 2128, 1766, 163, 1953, 2310, 1158, 1688, 915, 2026, 652, 2231, 1068, 1128, 2004, 225, 1249, 1914, 1259, 1997, 1684, 1129, 397, 778, 762, 359, 1816, 2085]. Classification [455, 2207, 126, 230, 371, 1418, 338, 277, 1209, 2189, 1537, 2258, 15, 443, 520, 103, 1021, 851, 2274, 1707, 985, 640, 678]. classifications [2166]. Clebsch [1546, 199, 1342, 536, 403, 2067, 1709, 1710, 162]. Clifford [710, 1822, 124, 1100, 319, 523, 120, 90, 597, 1554, 678, 679, 1751, 198, 2002]. Clifton [1754]. Closed [2284, 466, 2049, 1352, 75, 1113, 344, 811]. Closed-spaced [811]. clouds [110, 625, 583]. cluster [1836]. coadjoint [2152, 1447, 1023]. coagulation [764]. Cocycles [1837]. coefficient [1335, 1375, 781, 811, 931, 1896, 1899, 118, 715]. coefficients [1546, 199, 522, 1146, 536, 1771, 1036, 1037, 403, 404, 596, 557, 379, 805, 2067, 1709, 1710, 237, 162, 1399, 1407, 786, 1868, 1190, 2150, 2151]. Coherent [604, 528, 1561, 1495, 2033, 1917, 1678, 940, 1342, 2100, 1573, 1274, 1715, 1825, 869, 1721, 404, 1326, 423, 1252, 2175, 1998, 807, 2069, 1609, 784, 681]. coherent-state [1678, 681]. Cohomology [408, 188, 1790, 1722, 1153]. Coifman [2107]. Coinciding [548]. cold [509]. Cole [1736]. collapse [1543, 290, 1085, 751, 1031, 538]. collapses [2233]. Collective [1716, 1300, 1301, 1302, 1798, 1213]. Colliding [2320, 543, 584, 585, 2321, 393, 1658, 138, 468, 793]. collineation [256, 1088]. Collineations [1657, 1174, 2055, 2317, 2316, 2180, 831, 832, 2244, 472, 37]. Collision [328, 1694, 510, 1013, 1040, 860]. collision-dominated [510]. collisions [55]. color [970]. come [1692]. Comment [1145, 2316, 2018, 1372, 3, 1317, 2320, 1870, 166, 2321, 1755, 1032]. Comments [1714, 1584, 1877, 2315, 532, 423]. commutants [1385]. commutation [1470, 1292, 1321, 2043, 2072]. commutations [948]. Commutative [916, 1904, 1269, 485]. Commutator [1796, 497]. comoving [1003]. Compact [1158, 2183, 1236, 1035, 340, 174, 1522, 1268, 836, 2065, 389, 1732, 1750, 1707, 1270, 436, 1069]. compactified [2224]. compactly [278]. Compactness [1305]. comparison [1490, 2099, 186]. Compatibility [862, 116]. Compilation [2119]. Complete [412, 539, 641, 139, 1049, 121, 813, 1822, 2016, 15, 520, 2231, 450, 1616, 269, 1832, 349]. Completely [845, 430, 1502, 205, 660, 1892, 1945, 1625]. Completeness [2107, 435, 1410, 1065, 1671, 1069]. completing [1384]. Complex [2300, 655, 1750, 1928, 296, 1839, 3, 1317, 170, 1902, 185, 1602, 1827, 1958, 939, 1561, 1437, 1775, 1121, 1788, 1271, 1617, 1753, 1407, 1936, 1970, 2113, 2033, 1]. complex-time [1121]. complex-valued [1271]. complexes [2032, 2263]. component [44]. Composite [1243, 2140, 1423, 894, 1235, 2161]. composition [1138, 979, 1981]. Compressional [2005]. Computation [1467, 1142, 12]. compute [780]. conception [2123]. Concerning [503]. Concomitants [1045]. condensation [429]. condition [839, 2087, 580, 844, 2159, 281, 207]. conditional [1073, 1668]. conditionally [304]. Conditions [1328, 1460, 2047, 1182, 467, 1412, 1699, 453, 1941, 1305, 862, 582, 661, 1724, 2223, 2262, 446, 691, 556, 1217, 922, 1363, 342]. conductance [2300, 2212]. Conduction [720, 858]. conductivity [300, 1839, 1956]. conductors [894]. cone [2139, 911, 434]. cones [1427]. configuration [719, 1430]. configurations [1995]. confined [133, 1687]. confinement [568]. confocal [655]. Conformal [710, 130, 181, 702, 794, 704, 220, 1132, 434, 1364, 1460, 1314, 107, 1122, 523, 754, 185, 1285, 908, 731, 980, 1006, 1544, 1646, 941, 164, 614, 1788, 1056, 662, 703, 1660, 977, 1968, 1087, 1816, 1817]. Conformally [42, 255, 121, 957, 1080, 1819, 912, 427, 1082]. congruence [650]. congruences [1964, 1320]. conjecture [860]. conjectures [2120]. conjugate [845]. conjugate-bilinear [845]. conjugations [372]. connected [280, 2073, 294, 1627, 342, 802]. connection [1033, 1203, 1857, 1413, 1328, 1926, 257, 1281, 544, 2162, 1512]. connections [1940, 1381, 1216, 2210, 1779, 1850, 590, 1240, 1367, 982, 1475]. consequence [475]. consequences [1438, 137, 359]. Conservation [1789, 1482, 419, 1979, 774, 1955]. Conserved [214, 1260, 1953, 2310, 512, 712]. Consideration [1439]. considering [429]. consistent [2260]. Constant [1425, 1393, 321, 1003, 2157, 465, 1530, 1958, 1281, 572, 98, 407, 628, 93, 1364, 2060]. Constants [902, 1838, 588, 1018, 1019, 1225]. Constrained [1854, 1445, 2091, 85, 257, 71, 2240]. Constraint [2289, 2086, 779, 1841, 1733, 2276]. Constraints [1336, 901, 396, 1804, 2173, 1537, 1372, 2105, 820, 821, 1323, 2287, 996, 119, 379, 1405, 1887, 27, 338]. construct [1157]. constructed [344]. constructing [446]. Construction [1301, 1271, 1221, 688, 2207, 308, 1862, 225, 1300, 1302, 679, 212, 1133]. constructions [404, 137, 600, 849]. Contact [734, 1167]. contacts [816]. content [1097]. contextual [864]. Continual [1249]. continuation [1628]. Continued [1855, 1083, 287, 32]. Continuity [2077, 546, 835, 696, 59]. Continuous [749, 1492, 1154, 422, 1857, 983, 935, 1276, 307, 233, 737]. continuous-discrete [233]. Continuum [1374, 949, 836, 1009, 378, 2222, 1129]. contracted [37]. Contraction [452, 809, 2142, 458, 1553]. Contractions [1076, 370, 769, 730, 1859]. contribution [1284]. control [189]. Controlling [1987]. convenient [1737]. Convergence [836, 1931, 1763, 1753, 839, 69, 1843, 645, 1775, 70, 964]. convergences [659]. convergent [1639, 1487]. convex [787, 1640, 1793, 2120, 442]. convolution [595]. Cooper [1473, 1686, 1871, 1061]. coordinate [1412, 592, 1187]. coordinates [1663, 1330, 259, 379, 175, 878, 1025, 1867, 395]. coproducts [536]. Corepresentations [365, 196]. Correction [219]. Corrections [423, 386]. correlated [2128, 1258]. correlation [801, 1989, 486, 1570, 1890, 962, 359]. correlations [1555, 2164]. correspondence [606, 1277, 72]. corresponding [2173]. coset [919, 977]. cosets [605]. Cosmic [1172, 1910]. Cosmological [1126, 1358, 925, 1948, 1949, 321, 1062, 74, 2135, 218, 109, 2184, 628, 789, 1270, 1364, 2060]. cosmologies [1898, 74, 352, 1124, 539, 641, 803, 1479, 1627]. cosmology [1183, 254, 765, 872, 705, 910, 353]. cotangent [2186, 1860]. Coulomb [989, 1079, 2076, 1167, 824, 1267, 1882, 216, 250, 1950]. Coulomb-like [1950]. Coulombic [1717]. countable [1666]. countable-dimensional [1666]. counterexample [764]. Coupled [247, 172, 1411, 1919, 1242, 420, 957, 2118, 173, 912, 928, 1787, 1059, 2291, 1087, 232]. coupled-wave [1787]. coupling [667, 1146, 699, 659, 1048, 1902, 1264, 1042]. couplings [394]. Courant [1229]. covariance [1460, 1202, 1466]. Covariant [224, 723, 1898, 565, 1339, 594, 2094, 396, 515, 797, 293, 779, 1826, 1077, 2304, 258, 1681, 2245, 262, 411, 941]. covariantization [1907, 2026]. Covariantly [465]. covering [1495]. coverings [1451, 1166]. CP [1948]. creation [1095, 1341, 1106, 925]. creation-annihilation [1095]. Criteria [134]. criterion [1921, 631, 1767]. Critical [2006, 710, 781, 811, 1031, 1810, 178]. crossed [1296]. crossings [853, 1835]. crystal [634, 518, 932]. crystallographic [1698]. crystals [505, 1374]. cubes [1956]. cubic [686, 972]. Cummings [907]. Curiosities [708]. curl [1867, 1065]. current [1488, 441, 337]. currents [1231, 1162, 894]. Curvature [1257, 831, 832, 1819, 896, 353, 1969, 1460, 2093, 131, 1003, 2017, 312, 1337, 582, 1403, 1123, 1926, 1080, 1818, 1958, 1002, 1281, 2039, 1916, 1800, 1046, 538, 1069, 1322]. Curvature-squared [353, 538]. curve [916, 192, 213, 896]. curved [106, 1178, 1208, 1489, 828, 1988, 995, 151]. curves [754, 1835, 280]. curvilinear [1025, 395]. Curzon [1742]. Cusson [865]. cut [464]. cut-off [464]. Cutkosky [592]. cutoffs [1617]. cuts [911, 1427]. cyclic [519, 974, 1722, 1385]. cylinder [390, 289, 2118, 1773]. cylindrical [2079, 721, 379, 709, 1677, 395, 1065]. Cylindrically [1903, 1545, 665, 1543].

D [317, 826, 1078, 2295, 882, 1053, 1349, 1015, 630]. d’Adhémar [1032]. D’Alembertians [1444]. damped [1919, 1680, 71, 1230, 16, 521, 657]. d’arbitraire [493]. Darboux [780, 455, 1508]. data [2137, 266]. Daubechies [1484]. Davey [1847, 1092, 1252, 1789, 1482]. day [925]. deal [1936, 2113]. decay [2037, 2038, 178, 2037]. decaying [1403]. decoherence [2274, 2214]. decompactification [880]. Decomposing [1926]. Decomposition [160, 1371, 2157]. decompositions [1483, 1650, 1753]. Deconfinement [810]. defect [279]. defects [2157, 572, 518, 640]. defined [1249]. Defining [124, 1103]. Definition [1356]. Deformation [2076, 2254, 1830, 1792, 2019, 1394, 1794, 1673, 1869]. Deformations [2131, 201, 185, 830, 1932, 1706, 1648, 1937, 1359, 2115, 1395, 1463, 2028, 408, 849]. deformed [967, 1253, 886, 2102, 2106, 1786, 1751, 2197, 1977, 2041, 783, 1648, 1535]. deforming [536]. degeneracies [1423, 658, 950, 1243]. degeneracy [504, 1402, 1159, 1616]. degenerate [378, 1532]. Degenerations [1674]. degré [493]. degree [1862]. degrees [1005, 102, 1807, 450, 1373, 664, 948]. delay [2035]. Demiański [321]. densities [1260, 243]. Density [1680, 1984, 1596, 1381, 1488, 2079, 1876, 218, 629, 2007, 1469, 765, 1213, 1042, 1113]. density-functional [1876]. dependence [952, 2092, 921]. dependent [826, 1875, 1182, 396, 515, 2086, 1346, 1476, 554, 1771, 38, 576, 218, 1815, 1149, 113, 1695, 1128, 1842, 1296, 577, 1763, 511, 1887, 623, 1563, 1445]. deposition [841]. Derivation [592, 1180, 962, 516, 787, 963, 1773, 2084, 162]. derivations [204, 677]. derivative [1481, 409]. derivatives [1632, 1720, 387, 9, 1672, 1065]. derived [1349, 2228]. deriving [165, 648]. describe [925, 209]. describing [231, 1543, 1505, 474]. description [1964, 1849, 170, 1330, 2210, 1852, 1422]. descriptions [112]. design [999]. determinant [2202, 2322, 1774, 951]. determinants [1151, 2190, 2308, 149, 2148, 692]. determinate [949]. Determination [1971, 1556, 1534]. determine [1103]. determining [582, 2105]. deterministic [475]. detonation [1355]. deviations [345]. diagonal [2135, 1565]. diagonalization [902, 2302, 132]. diagram [1221, 935, 675]. diagrams [965, 890, 1250, 1727, 2204, 294]. diamond [720]. diamond-shaped [720]. diatomic [1368]. Dickey [2071]. dielectric [440, 1743]. Diffeomorphism [1790, 968, 231, 724, 649, 1675]. diffeomorphisms [1412, 1001, 268, 563, 768, 987]. difference [2074, 1905, 2252, 1668, 1147, 2054]. differences [659]. different [965, 1649, 654, 717, 1187]. differentiability [350]. differentiable [1631]. Differential [1503, 2007, 550, 1524, 632, 856, 466, 1904, 653, 2017, 929, 1066, 1737, 1813, 1630, 1618, 916, 1549, 1847, 647, 2251, 413, 1701, 1515, 2299, 66, 67, 1018, 405, 551, 2105, 1254, 740, 167, 1361, 103, 1021, 990, 1269, 1862, 629, 206, 1668, 734, 1749, 2194, 973, 648, 291, 1977, 2306, 34, 1979, 1389, 1866, 493, 1287, 1362, 848, 96, 1225, 1109]. differential-difference [1668]. differentials [1579]. differentiation [566]. difficulties [964]. diffraction [1566, 1614]. diffuse [1465]. Diffusion [1192, 1016, 1182, 1898, 1195, 1060, 939, 1785, 456, 1456, 1324]. diffusions [1295, 1539, 2162, 1763]. Diffusive [1089, 556, 1217]. diffusivity [882]. Dilation [1231]. dilaton [2181, 1511]. dimension [1345, 2266, 1856, 2188, 1448, 318, 202]. dimensional [1099, 1120, 1594, 1236, 2129, 305, 1157, 1654, 1289, 1293, 2187, 2013, 480, 956, 1897, 2189, 1522, 1268, 1737, 2215, 871, 370, 603, 322, 1497, 417, 418, 203, 913, 2201, 292, 1687, 2144, 1489, 45, 163, 1414, 149, 201, 2066, 722, 609, 1118, 2258, 1912, 606, 1472, 749, 1459, 1461, 1722, 660, 2048, 1666, 1808, 2003, 575, 2182, 1530, 1356, 1652, 528, 908, 512, 139, 794, 1451, 1705, 1352, 1909, 2223, 727, 2211, 1323, 1634, 2039, 2027, 991, 2004, 1669, 917, 222, 1255, 1311, 1927, 973, 950, 2184, 1243, 8, 598, 1520, 2167, 426, 2122, 221]. dimensional [1010, 937, 1799, 899, 148, 1248, 1223, 142, 1129, 2052, 2175, 941, 1046, 164, 1442, 1831, 614, 910, 2070, 307, 2035, 1882, 1270, 464, 1542, 1888, 1391, 1535, 46, 1050, 2044, 216, 730, 1103, 955, 1233, 1386, 1765, 250, 640, 1911, 1496, 324, 232, 265]. dimensionality [1725, 2309]. dimensionally [1083]. dimensions [501, 1639, 2152, 880, 454, 1527, 1066, 1630, 1045, 1901, 1567, 688, 2282, 2319, 2207, 1966, 1061, 263, 1840, 1702, 2271, 1350, 1277, 2119, 1815, 469, 473, 1421, 41, 857, 1508, 113, 220, 1708, 1707, 883, 1007, 490, 870, 1789, 1482, 666, 774, 1511, 478, 1197, 1422, 1177, 397, 711, 1087, 1686, 1871]. dipole [867, 721, 1416, 360, 1440]. Dirac [760, 1471, 2279, 1178, 2091, 174, 1047, 1849, 319, 2014, 1985, 293, 779, 1718, 320, 1687, 1167, 1048, 918, 2089, 1214, 85, 365, 1131, 855, 795, 627, 959, 363, 945, 260, 47, 48, 1049, 2287, 1604, 750, 222, 636, 1132, 1554, 2134, 2084, 1729, 2124, 1565, 1533, 364, 878, 960, 1025, 1764, 1366, 223, 117, 395, 2002, 471, 250, 516, 207, 267, 621]. Direct [1373, 1732, 208, 1251, 1976, 1838]. directed [1729]. directing [35]. directions [1810]. directly [2132]. Dirichlet [1941, 580]. disconnected [1999]. discontinuities [314, 1376]. discontinuity [1854]. discontinuous [2167]. Discrete [2299, 949, 55, 29, 1855, 913, 1498, 1459, 1106, 416, 233, 773]. disk [762]. Dispersion [204, 677, 2280, 1181, 1577, 1479]. dispersive [134, 1400, 883]. Dissipation [1943]. dissipative [1279, 656, 923, 1030, 1732, 1261, 307, 310]. distance [1734]. Distorted [1389]. distributed [1243]. distribution [595, 1108, 384, 191, 745]. distributions [2128, 746, 190, 101]. divergence [2243, 1660]. divergence-free [1660]. division [227, 1670]. Dixon [1820, 1156]. do [1692]. does [1651]. Dolan [844]. domain [1262, 391, 923, 1732, 937, 2073, 1506, 342]. dominated [510]. dominating [815]. Double [2176, 1401, 761, 958, 1647, 1494, 324]. double-well [1401]. doubling [1008]. doubly [342]. Draayer [522]. drift [1735]. driven [1880]. driving [1474]. DSCR [1054]. dual [1413, 1449, 1901, 329, 649, 1919, 844, 2048, 451, 469, 1171, 1634, 1240, 1367, 490, 581, 245, 186, 1542, 676, 427]. Duality [1298, 615, 1889, 1360, 874, 1597, 1143, 799]. duals [1698]. Duffin [2163]. Duistermaat [105]. Dust [1003, 357, 358, 110, 625, 583]. Dym [1895]. Dynamical [1874, 2233, 547, 1446, 1815, 1535, 106, 396, 348, 1573, 697, 50, 1920, 1807, 656, 2049, 686, 258, 139, 1373, 973, 1456, 696, 490, 1405, 1223, 1962, 2134, 1777, 9, 1672, 112, 1189]. dynamically [308]. Dynamics [1539, 1230, 537, 1872, 901, 1731, 843, 2156, 966, 1414, 1026, 1027, 170, 1473, 821, 943, 1768, 1114, 875, 71, 1068, 2162, 2235, 842, 1917]. Dynkin [1250, 890]. Dynkin-like [1250, 890]. dyon [542, 267]. Dyonic [2181]. dyons [1136]. Dyson [2269, 2036, 486].

É. [1171]. early [925]. Eddington [517]. Editorial [1283]. effect [2213, 1759, 400, 1728, 744, 464, 193]. Effective [300, 682, 394, 681, 82, 780, 130, 1696, 1348, 802]. Effects [109, 1142, 741]. Eguchi [859]. eigenequation [2268]. Eigenfunction [1065, 894, 159]. Eigenfunctions [1055, 1867, 787, 1556, 1985, 2107, 1981, 1065]. Eigenstates [1608, 2000]. eigenvalue [1459, 1625, 1896, 1899, 815, 410, 1767]. Eigenvalues [1985, 1829, 787, 701, 1556, 700, 1490, 350, 997, 613, 1981, 207]. eigenvectors [505, 350]. eight [1489, 926]. eight-dimensional [1489]. eight-vertex [926]. Einstein [706, 1209, 2179, 2241, 874, 466, 1578, 1413, 1543, 1778, 78, 957, 958, 1967, 393, 1658, 627, 144, 857, 585, 266, 1708, 426, 2164, 392, 581, 705, 493, 427, 223, 954, 1348, 471, 1087]. Einsteinian [1350]. Elastic [2167, 1566, 989, 554, 384, 385, 2127]. elastodynamics [1247]. electric [352, 867, 315, 1330, 1296]. Electrically [1764]. electrodynamics [1378, 1529, 2082, 1723, 1734]. Electromagnetic [79, 1627, 638, 1639, 1612, 1411, 1163, 322, 314, 554, 2088, 923, 271, 2161, 1879, 98, 407, 2126, 1763, 574, 1042, 568, 267]. Electron [1248, 758, 1858, 319, 1876, 1796, 1356, 113, 1368, 1128, 1917]. electrons [1914, 1113]. electroweak [366, 1992, 2311, 433]. element [1909]. elementary [1234, 1517]. elements [2296, 1036, 403, 1784, 596, 225, 1137, 2225, 608, 1097, 267]. elevation [1114]. Élie [1269]. Elliott [522, 2196]. ellipses [1193]. Elliptic [1920, 2297, 930, 1206]. Elliptical [1567, 1568]. embedded [701, 40, 1125]. Embedding [321, 953, 1412, 1510, 2104, 1783]. embeddings [710]. Emden [172, 1713]. emergence [572]. endowed [1416]. Energy [1402, 700, 586, 1233, 1386, 1765, 1955, 826, 1612, 1522, 1268, 1941, 952, 867, 361, 10, 123, 2079, 829, 1140, 2116, 830, 999, 1796, 218, 1331, 92, 998, 2007, 544, 921, 675, 1532, 385, 142, 177, 2035, 1882, 1291, 906, 1443, 1348, 1501, 483, 1942]. energy-dependent [826]. Energy-momentum [586, 1955, 830, 544, 1348]. Enhanced [882]. enhancing [348]. ensemble [2023]. ensembles [1570, 1555]. entropic [1090]. entropies [847]. entropy [127, 2153, 1618, 685, 2301, 2146, 346, 895, 616, 2007, 1828, 696, 2227, 2314, 2290, 1429, 2199, 2149, 762]. enumerating [294]. Envelope [1601]. enveloping [1359, 1973, 160, 437, 95]. Epstein [45, 2255, 854, 1863]. equal [1105, 1385]. equal-rank [1105]. equation [1071, 1745, 317, 826, 1345, 1594, 1524, 2266, 787, 1640, 1875, 1178, 1182, 1556, 172, 1432, 422, 55, 1623, 1573, 1161, 1731, 1624, 2171, 1780, 448, 1247, 453, 1630, 386, 1596, 1278, 1335, 412, 503, 2269, 916, 2014, 782, 2018, 2077, 1781, 1758, 647, 1346, 1359, 779, 1011, 1018, 508, 1626, 1138, 918, 2089, 1214, 622, 740, 1453, 2116, 369, 1419, 885, 1655, 1030, 1530, 283, 1229, 1131, 1130, 239, 756, 2174, 795, 627, 346, 1312, 959, 1352, 363, 945, 733, 557, 652, 1645, 401, 939, 48, 1049, 571, 1682, 2112, 1267, 368, 1092, 88, 1455, 1332, 1368]. equation [1893, 2009, 2161, 1785, 592, 2110, 1934, 222, 636, 1252, 1015, 419, 145, 1230, 973, 2051, 1147, 1181, 280, 1520, 2281, 2122, 937, 1916, 2147, 932, 1774, 1636, 1637, 430, 1689, 577, 1696, 2084, 533, 1166, 1576, 898, 1685, 1393, 2206, 1866, 2124, 1565, 1533, 774, 1713, 878, 960, 1025, 2073, 1494, 1407, 1577, 1935, 550, 715, 848, 223, 395, 2002, 1127, 250, 1592, 1148, 1507, 1740, 1950, 1334, 684, 1686, 1871]. Equations [254, 242, 1209, 1653, 1288, 1157, 1641, 1873, 2130, 632, 1620, 1621, 813, 856, 874, 1035, 124, 1260, 1957, 466, 2101, 653, 2017, 2063, 440, 116, 2220, 1194, 566, 1066, 984, 1813, 1941, 1413, 1163, 240, 329, 968, 2173, 1847, 913, 2251, 293, 892, 420, 312, 1687, 413, 1543, 852, 2252, 861, 361, 327, 1952, 1778, 1833, 1313, 78, 792, 683, 1771, 958, 1254, 749, 1375, 1461, 1195, 2080, 167, 2081, 375, 1200, 2303, 655, 689, 2048, 382, 2097, 325, 1051, 1666, 1383, 752, 146, 1361, 438, 103, 1021, 990, 851, 894, 1448, 1222, 627, 1579]. equations [259, 539, 1310, 1109, 2230, 380, 444, 1421, 895, 616, 629, 2007, 1634, 173, 1425, 1959, 1726, 206, 1668, 585, 14, 1400, 2009, 309, 1196, 734, 266, 2194, 140, 1226, 1179, 1311, 456, 1772, 883, 937, 2276, 648, 1773, 291, 226, 1012, 2126, 2041, 2247, 64, 108, 34, 2052, 209, 985, 2175, 1787, 164, 1442, 1979, 304, 549, 364, 2177, 877, 1324, 663, 493, 496, 337, 186, 1739, 1542, 808, 1059, 1656, 509, 427, 760, 1362, 117, 46, 762, 1348, 1506, 471, 1325, 96, 2291, 193, 1692, 1225]. equilibria [645]. Equilibrium [843, 431, 839, 1843, 452, 510]. Equivalence [1652, 736, 1683, 1641, 1972, 1210, 1666, 1092, 816, 2010, 1493, 2178]. equivalent [1253, 308, 232, 1099]. Equivariance [2278]. Equivariant [244, 14, 1772, 226]. ergodic [1880]. Ergodicity [915]. Errata [2318]. Erratum [2317, 723, 1466, 57, 6, 1522, 1467, 1715, 2313, 2322, 2319, 1938, 677, 1317, 2308, 123, 2310, 1468, 1140, 1250, 1098, 520, 1523, 2311, 2312, 1251, 597, 18, 598, 2264, 2265, 2314, 521, 337, 2113, 153, 1282, 1871, 219]. Error [895, 1828, 1142, 651]. Essential [1175, 739, 2262, 931]. Establishing [742]. Estévez [1755, 1755]. Estévez-Bretón [1755, 1755]. estimate [1865]. estimates [2121]. estimation [2149]. Estimations [1472]. Euclidean [881, 1432, 1595, 603, 2270, 2108, 1001, 528, 1070, 1517, 1518, 1634, 750, 2193, 581, 1677, 1188]. Euler [1313, 752]. evaluating [1399]. Evaluation [1165, 1147, 1303, 1513, 989, 2250, 1162, 531, 1591, 1097]. Evaluations [191]. Even [1797, 442, 480, 1045, 292, 1277, 716, 285]. even-dimensional [480, 292]. Event [2272]. everywhere [1487]. Evolution [92, 2247, 1442, 1874, 1260, 2155, 240, 53, 1771, 2139, 655, 1051, 109, 1436, 639, 1959, 88, 1785, 1196, 1406, 2052, 304, 574, 2177, 1838]. evolve [925]. evolving [74, 243]. Exact [501, 1209, 2014, 1212, 420, 634, 1138, 2300, 2023, 1043, 1423, 1356, 627, 1835, 1910, 1349, 1785, 1311, 1181, 279, 1696, 1787, 1369, 1764, 356, 223, 250, 630, 1182, 1413, 1847, 589, 2105, 38, 1195, 1743, 2161, 1217, 145, 456, 1137, 2225, 674, 1581, 1223, 2176, 2177, 907, 117, 395, 1282, 961, 471, 2248, 1058]. exactly [1557, 1780, 894]. Exactness [424]. example [2028, 137, 429, 914, 1301]. Examples [2279, 887, 1708, 155, 1510, 1978]. exceptional [1823, 782, 1095, 264, 488, 1868, 971]. Exchange [1989]. excitation [189, 1798]. excited [192, 213, 279]. exciton [1909]. exciton-phonon [1909]. Existence [2169, 569, 1017, 1468, 1435, 1532, 1160, 637, 55, 1062, 1502, 203, 1892]. Exotic [2216, 2182, 1647]. expanding [1417]. expansion [398, 69, 1796, 2191, 2260, 18, 11, 166, 1865, 1520, 2122, 1638, 249, 1399, 118, 715, 815]. expansions [1071, 650, 1184, 1550, 894, 379, 1259, 409, 1981, 1065]. Explicit [199, 1146, 1286, 225, 54, 688, 1510, 2210, 679]. exponential [1749, 2121, 1440, 1753]. Exponentially [243]. exponents [2006, 1761, 1776]. expression [1781, 709, 1480]. Expressions [10, 123, 1047]. Extended [796, 20, 1569, 473, 251, 511, 2206, 224, 723, 1202, 1466, 334, 757, 1052, 1967, 795, 52]. extending [1500]. Extension [293, 2136, 484, 1525, 2279, 1783, 447, 1115, 295, 550]. Extensions [529, 777, 2297, 1780, 644, 2066, 49, 1635, 1338, 1582, 1166, 530]. exterior [2254, 1269, 1902]. external [881, 320, 289, 1988, 840, 2082, 1796, 552, 2161, 878, 960, 1025, 1764, 337, 267]. extrapolation [127, 779, 964]. Extremal [416]. extremality [599]. extremals [794]. eye [402].

F [1145]. face [756]. factor [1916]. factorial [1700]. Factorization [1030, 501, 317, 1345, 2268, 1665]. Factorizations [351, 1704]. factors [1145, 517, 979, 236, 200, 978]. facts [387]. Faddeev [1726]. Families [1194, 737, 773]. family [1642, 874, 1194, 958, 1200, 393, 1658, 1934, 1197, 2150]. far [452]. FBIS [482]. features [1631, 1053]. Fermi [760, 43, 1596, 58, 524, 840, 368, 44, 2122, 1320, 1441, 760]. Fermion [552, 1564, 784, 630, 1034, 1556, 1526, 1906, 2021, 1095, 551, 2258, 2098, 1837, 748, 1300, 1301, 1302, 2276, 2042, 2044, 267, 1803, 1010]. fermion-boson [1095]. Fermionic [149, 1595, 819, 2223, 343]. Fermions [667, 881, 1595, 1043, 1722, 122, 590, 914]. ferromagnet [2282, 2319]. ferromagnetic [1845, 842]. Ferromagneticity [1236]. ferromagnets [1889]. few [251]. few-body [251]. Feynman [779, 31, 589, 965, 2273, 1604, 1259, 502, 70, 2148, 2204, 994, 745]. fiber [1153, 798, 433]. fibered [993]. Fibonacci [1908]. field [967, 542, 305, 758, 739, 632, 881, 51, 398, 1411, 2154, 1091, 836, 1413, 352, 322, 319, 1245, 1242, 2251, 2201, 1009, 320, 1613, 2207, 966, 633, 1953, 2310, 754, 315, 828, 957, 668, 391, 840, 1560, 2081, 185, 1477, 1285, 1051, 1783, 2119, 1244, 271, 908, 731, 1054, 1269, 1237, 980, 1523, 2083, 876, 1645, 40, 1125, 1910, 658, 1819, 585, 879, 2004, 691, 228, 1879, 76, 544, 917, 1357, 86, 1388, 1798, 2281, 221, 1010, 360, 1440, 1799, 2163, 98, 407, 1160, 1056, 2123, 2289, 343, 977, 1954, 328, 276, 248, 676, 1365, 1339, 112]. field [537, 433, 344, 1050, 1087, 1081, 1082, 2165, 207, 267, 802, 1886]. field-antifield [1560]. fields [1470, 1236, 1639, 1628, 43, 1208, 135, 480, 1162, 296, 672, 270, 1047, 1163, 1849, 35, 1489, 477, 1988, 615, 772, 2082, 2271, 2117, 915, 1435, 702, 1030, 669, 1651, 115, 150, 1724, 690, 260, 1032, 817, 144, 612, 750, 775, 2161, 145, 555, 2224, 1296, 912, 791, 1441, 2306, 581, 1303, 513, 245, 574, 1418, 878, 960, 1025, 1155, 1369, 1764, 336, 1042, 1156, 1802, 298, 833, 187, 436, 2085]. Fierz [388, 1886]. filaments [231]. film [883]. films [1951, 193]. final [936]. Finite [801, 2258, 1459, 2027, 1670, 8, 598, 148, 142, 198, 1911, 1496, 1157, 1737, 601, 361, 609, 2301, 668, 660, 281, 678, 679, 2052, 1564]. Finite-dimensional [2258, 1459, 2027, 8, 598, 148, 1157, 1737, 609, 660, 2052]. finite-periodic [281]. Finite-size [801]. Finiteness [949, 1438]. Finslerian [1754]. First [1476, 792, 1641, 1378, 1804, 240, 1965, 820, 308, 1060, 1705, 1149, 262, 1842, 2306, 353, 338, 987]. first-class [1804, 820, 262]. first-order [1641, 240, 308, 1705, 353]. fish [402]. fish-eye [402]. Fisher [1534]. Five [1927, 2184, 2013, 306, 1046, 2070, 1888]. five- [2070]. Five-dimensional [1927, 2184, 2013, 1046, 1888]. fixed [181, 2244, 999, 998, 72, 1810]. fixed-energy [998]. fixed-isospin [72]. fixed-seniority [72]. flag [105]. flags [1277]. Flat [1475, 724, 1272, 1080, 1742, 2210, 567, 464, 255]. flattening [652]. Flessas [316]. flexible [191]. flip [2189, 1836]. flow [57, 1655, 1086, 2]. flows [1573, 155, 2229, 1833, 1776, 1740]. fluctuating [1388]. Fluctuation [1721, 843]. Fluctuations [2318, 1450, 2059, 587, 537]. Fluid [872, 74, 352, 1371, 254, 1124, 2135, 289, 1543, 1257, 1173, 540, 1694, 1352, 1420, 42, 1773, 1997, 1581, 705, 1039, 256, 2127, 630]. fluids [354, 152, 2229, 38, 1551, 704, 1023, 1042, 255]. flux [1292, 1898]. flux-limited [1898]. Fock [1941, 441, 331, 1474, 1913, 526, 1803]. Focused [1689]. Fokker [346, 1312, 1520]. fold [780]. foliations [392]. force [1161, 1731, 217, 1159, 612]. force-free [612]. forces [1474]. Form [65, 448, 503, 1758, 102, 576, 196, 1352, 2276, 2235, 1788, 1112, 1365, 1970]. Formal [1306, 1260]. formalism [1525, 127, 20, 580, 1953, 2310, 1902, 1560, 265, 414, 2046, 1252, 1734, 511, 9, 1672, 353, 2290, 1936, 2113, 516, 993]. formalisms [1099, 1683]. formally [1182]. formation [1261]. forms [738, 1289, 24, 466, 1618, 828, 829, 1140, 126, 1461, 740, 1775, 1616, 1083, 344, 790]. formula [1514, 31, 2143, 558, 424, 421, 2174, 1603, 1032, 1424, 1456, 922, 497, 105, 2204, 744, 825]. formulas [367, 1256, 1703, 425, 72, 1406, 333]. formulation [1452, 1372, 312, 819, 604, 1026, 2222, 776, 2223, 996, 206, 357, 1852, 1999, 479, 1135, 1339, 1512]. formulations [2177]. Four [2187, 2201, 614, 1639, 871, 1901, 649, 2119, 1448, 2223, 727, 1276, 1634, 857, 2009, 1669, 941, 2176, 936, 2200]. Four-dimensional [2187, 2201, 614, 871, 2223, 727, 1634, 1669, 941]. four-manifold [2200]. four-manifolds [649]. four-momenta [1276]. four-point [2119]. four-vectors [936]. Fourier [2140, 1587, 271, 2305, 379, 143, 373]. Fourier-wavelet [2140]. fractal [861, 494, 1891]. fractal-shaped [861]. fractals [988, 2283]. fraction [1855]. Fractional [1044, 2251, 1312, 1788, 1271, 1475]. fractional-linear [1788]. fractions [1982, 2313, 287, 32]. frame [5, 2264, 1754]. frames [788, 1932, 2230, 1478, 278, 1024]. framework [224, 723, 328]. Fredholm [895, 616, 2148]. Free [2271, 1749, 742, 650, 135, 1964, 454, 2006, 174, 354, 1380, 170, 2192, 1032, 2223, 612, 113, 1128, 1320, 1660, 336]. freedom [102, 1807, 450, 1373, 664, 948]. frequencies [1288, 1775]. Frequency [113, 1520, 1163, 1387, 252, 1415, 533, 2005]. Frequency-dependent [113]. Friedmann [321]. front [1606]. fronted [110, 625]. frustrated [1452]. fuel [1738]. fuel/shell [1738]. Fuerth [939]. Fueter [1001]. full [1567]. Fully [2280]. function [1116, 1939, 801, 1699, 1047, 1163, 1757, 1700, 2064, 852, 2253, 10, 123, 45, 1632, 2255, 1988, 1019, 1020, 743, 1376, 1808, 1732, 2003, 1356, 2280, 2261, 461, 1368, 1914, 2051, 1147, 44, 1296, 409, 416, 249, 335, 1399, 295, 550, 286, 1008, 216, 374, 2291]. function-series [2291]. Functional [1110, 2190, 2308, 2109, 513, 1344, 1380, 421, 1876, 1562, 2236, 1802, 692]. functionals [1861, 2274, 2214]. Functions [722, 2232, 2249, 1208, 2100, 1079, 1202, 1466, 1162, 2154, 2141, 129, 1224, 2143, 1568, 930, 2103, 204, 677, 1920, 168, 1256, 829, 1140, 1169, 1989, 858, 2270, 2108, 1043, 486, 2259, 1030, 1356, 751, 539, 641, 1474, 573, 1108, 854, 1863, 463, 399, 176, 400, 101, 411, 531, 1915, 1570, 932, 362, 693, 169, 1563, 409, 197, 2198, 1846, 2000, 359, 191, 1611, 267, 630]. Fundamental [94, 1605, 1965, 1516, 339]. fusion [2083, 1094]. Future [182, 2246].

G [226, 1755]. Gabor [1932]. gain [823]. Galerkin [646]. Galilean [1293]. Galilei [273]. galore [449]. gamma [2143]. gap [1686, 1871]. gas [231, 2047, 1211, 2006, 1843, 1768, 113, 1128, 299, 510, 44]. gases [708, 2098]. Gaudin [2075, 1958, 1281]. Gauge [671, 1804, 2034, 2157, 85, 1092, 228, 798, 13, 153, 2185, 796, 1292, 1471, 881, 919, 1452, 224, 20, 515, 723, 797, 934, 593, 1449, 1433, 503, 2036, 3, 1317, 2299, 67, 1953, 2310, 1600, 1840, 1453, 1860, 909, 2159, 690, 1398, 1992, 2311, 1994, 775, 145, 86, 1093, 434, 674, 1605, 1160, 513, 245, 1999, 27, 338, 586, 276, 248, 118, 715, 2178, 433, 1536, 1189, 1, 232, 1955]. gauge-invariant [934]. gauge/parametrization [27]. gauged [265, 971, 1396]. Gauss [1488, 1733, 183]. Gaussian [1108, 1915, 1810, 194]. Gegenbauer [1074, 1946]. Gel′fand [767, 560, 1075, 1709, 1710, 608, 2071, 676]. Gell [558]. Gell-Mann [558]. geminal [1917]. General [1432, 1687, 965, 38, 1214, 1703, 494, 1507, 453, 354, 1297, 1242, 372, 293, 355, 289, 2229, 506, 2137, 1390, 831, 832, 465, 702, 2244, 654, 1743, 733, 1903, 557, 704, 1086, 2099, 1349, 1439, 1852, 507, 937, 1581, 2235, 288, 1039, 1418, 1866, 27, 942, 663, 2032, 1088, 1282, 961, 342, 1710, 1366]. Generalization [1371, 1168, 1551, 631, 22, 1756, 1211, 1435, 670, 1786, 1357, 104, 1891]. generalizations [204, 677, 609, 1123, 1742]. Generalized [1412, 1274, 1715, 1769, 1597, 1337, 1986, 767, 885, 1285, 694, 150, 1308, 690, 1863, 1836, 176, 450, 709, 1480, 5, 2264, 2197, 2001, 1148, 1872, 1540, 1874, 2062, 1875, 396, 20, 1162, 1957, 2189, 1578, 448, 1247, 1527, 2298, 1199, 930, 1747, 2096, 2255, 1720, 1347, 1721, 1473, 772, 15, 520, 2008, 1277, 2214, 258, 1032, 380, 1421, 939, 1373, 439, 241, 1368, 2175, 532, 1752, 2084, 1674, 2000, 1739, 799, 1656, 429, 1392, 216, 1127, 1693, 287, 1334, 530, 1108]. generally [941]. generate [270]. generated [1548, 1751]. Generating [129, 930, 983, 1993, 2312, 957, 997, 1562, 186]. Generation [1420, 1951, 858, 1564, 433]. generator [771, 2302, 608]. generators [515, 1320, 1103]. generic [2142, 2027]. genus [1382]. Geodesic [542, 1851, 106, 323]. geodesical [1069]. geodesics [1477, 794, 1671, 2165]. Geometric [821, 945, 1273, 1945, 2267, 386, 1503, 633, 414, 1921, 591, 2287, 1959, 1667, 775, 1887, 343, 479, 1135, 900, 1475, 1188]. geometrical [1690, 819, 1026, 1681, 90, 597, 1999]. geometrically [2228]. geometries [198, 1065]. Geometrization [1688, 1088, 590]. geometrodynamics [1412]. Geometry [188, 716, 73, 1890, 245, 2209, 2179, 2241, 2213, 2203, 1242, 292, 66, 67, 147, 1926, 1601, 1277, 2304, 1237, 1523, 25, 727, 136, 485, 1107, 2030, 1750, 555, 1007, 1977, 1788, 416, 1389, 1096, 789, 1410]. Gibbs [943]. Ginzburg [193]. Girardello [940]. Glansdorff [631]. glasses [802]. Glauber [875]. Gleason [2274]. Global [1005, 55, 2229, 911, 834, 1431, 992, 2042, 307, 1013]. globally [359]. go [1640, 2260]. Gödel [1174, 884]. Gödel-type [1174, 884]. golden [588]. Good [2315]. Gordan [1546, 199, 1342, 536, 403, 2067, 1709, 1710, 162]. Gordon [2112, 1484, 1624, 420, 1011, 2116, 1200, 1655, 1130, 1682, 1226, 364, 1739, 1656]. Gorringe [316]. Goursat [283, 438, 870]. governing [1381]. Gowdy [92]. grade [2104, 230, 277]. grade-star [277]. Graded [769, 1107, 806, 1859, 866, 2018, 2020, 486, 1587, 1588, 686, 972, 1249, 52, 2306, 408]. gradient [1962, 2134]. gradient-holonomic [1962, 2134]. gradings [717]. Grady [844]. Graph [439, 1285]. Graphs [2299, 2189, 174, 885, 2119, 709, 620]. Grassman [194]. Grassmanian [1649]. Grassmann [1663, 1667, 1669, 914]. Grassmannian [1649, 1224]. gravitating [1900, 38, 1307]. gravitation [1210, 17, 1171, 586, 1583]. Gravitational [2138, 1310, 2060, 1041, 834, 1005, 1431, 1543, 1806, 1084, 2318, 79, 391, 2058, 1085, 2217, 1694, 1477, 138, 468, 1450, 1759, 909, 2245, 40, 1125, 1744, 543, 584, 793, 1388, 1139, 1659, 387, 64, 538, 1155, 1369, 110, 625, 583, 1040, 433, 833, 1697]. gravitationally [1969]. Graviton [587, 111]. gravity [2093, 2215, 1449, 2056, 2181, 2293, 1414, 1347, 957, 1350, 1415, 139, 794, 1451, 2245, 192, 213, 1851, 572, 426, 628, 666, 703, 2177, 1391, 1083, 790, 1087, 324, 2248]. gravity-gyroscopic [192, 213]. greater [1448]. Green [865, 1163, 204, 677, 1169, 2270, 2108, 1030, 1376, 1808, 1732, 2003, 1352, 1474, 1368, 1913, 1915, 932, 1296, 693, 286]. Greenberg [1486]. Gribov [1877]. Ground [1941, 177, 867, 2015, 997, 1532, 184, 906, 1443]. Ground-state [177]. Group [1574, 1316, 903, 2008, 1505, 2073, 242, 1099, 231, 724, 2186, 2140, 2152, 2062, 881, 1661, 1823, 2095, 273, 593, 602, 603, 649, 1274, 1715, 647, 558, 341, 1014, 1298, 458, 58, 524, 402, 1989, 622, 1453, 1702, 726, 1588, 1446, 851, 1355, 303, 2304, 1222, 1153, 1474, 2160, 2223, 1276, 673, 487, 2231, 1850, 1481, 1992, 2311, 1706, 626, 1455, 159, 933, 1786, 95, 1552, 1976, 2195, 1384, 1181, 360, 489, 1319, 1709, 1710, 1684, 64, 108, 1442, 898, 2206, 1999, 1954, 526, 1287, 2032, 1675, 786, 1343, 617, 1677, 848, 1936, 1970, 2113, 1190, 1891]. group [1495, 1103, 96, 1354]. group- [1153]. Group-invariant [1316].

Group-theoretic [2073, 1274, 1715]. Groups [1772, 61, 2185, 901, 1698, 671, 956, 1897, 340, 601, 370, 886, 1548, 1924, 125, 523, 3, 1317, 1664, 644, 78, 389, 683, 606, 767, 1318, 156, 157, 158, 1589, 2025, 1447, 528, 1862, 561, 562, 1517, 1518, 487, 1425, 225, 887, 1299, 1708, 1707, 488, 1635, 98, 407, 164, 975, 1712, 338, 1143, 1656, 1220, 28, 1, 1859]. Gupta [1714, 776, 274]. Gürses [1778]. gyroscopic [192, 213].

H [2316, 532]. hadron [759]. Hahn [1857, 983, 1492]. hair [1307, 1544]. Half [1896, 1899, 1380, 1864, 279, 435, 1055, 2265, 2031]. half-Hartley [2265, 2031]. half-Hilbert [2265, 2031]. half-infinite [279]. half-plane [1380, 1864, 435, 1055]. Half-solitons [1896, 1899]. Hall [2213, 2212]. Hamburger [2301, 2199]. Hamilton [1972]. Hamiltonian [1372, 2129, 1157, 1654, 682, 1228, 1502, 129, 2173, 417, 418, 902, 862, 455, 779, 2218, 1199, 1476, 1167, 1201, 102, 1165, 681, 170, 68, 820, 660, 382, 1142, 33, 1815, 996, 400, 450, 1373, 2133, 88, 2046, 309, 1852, 1842, 2010, 1945, 1683, 732, 1333, 1809, 1291, 1275, 993, 410, 1740, 1679]. Hamiltonian-free [170]. Hamiltonians [2219, 739, 1939, 476, 432, 991]. Hamiltonization [1161, 992]. handlebody [2206]. Hannay [247]. Hanson [859]. hard [1352, 762]. hard-disk [762]. hard-rod [1352]. hard-sphere [762]. Harish [1980]. Harish-Chandra [1980]. Harmonic [35, 2145, 30, 747, 1951, 865, 189, 1253, 680, 1792, 1919, 1434, 389, 1118, 1559, 2303, 1376, 575, 350, 1680, 1579, 71, 241, 1927, 279, 244, 175, 532, 2070, 657, 1611, 1755]. harmonics [1586, 1677]. Harry [1895]. Hartley [2265, 2031]. Hartmann [286]. Hartree [1941]. Hausdorff [2260, 1424, 497, 1811]. having [137]. Head [1040, 2127]. Head-on [1040]. Hearing [342, 2090]. Heat [740, 61, 2094, 1152, 2065, 858, 1086, 2110, 18, 11, 166, 1865, 2147, 435, 1055]. Hecke [927, 1357, 1709, 1710]. Heckman [105]. hedgehog [718]. Heiles [242, 1957, 1575]. Heisenberg [2074, 1661, 1857, 602, 1009, 2282, 2319, 966, 2257, 1249, 842, 1129, 676, 617]. Helicity [1136]. Helicity-supersymmetry [1136]. Hellmann [1259]. Helmholtz [1781, 1138, 2088, 1934, 932]. Hénon [242, 1957, 1575]. hereditary [732]. Hermite [2156, 2144, 1946]. Hermitian [1289, 1035, 1746, 1587, 1830, 1197]. heterogeneous [440]. Hida [745]. Hidden [1745, 402, 950, 619, 232, 864, 2192, 216, 1536, 2083]. hidden-variables [864]. hierarchic [1814]. Hierarchical [2168]. hierarchies [771, 913, 852, 1313, 432, 1323, 2009, 1196, 1422, 313, 1692, 1693, 712]. Hierarchy [1653, 1654, 2234, 1336, 1212, 1337, 1990, 205, 660, 331, 512, 1848, 444, 87, 214, 490, 2235, 1894, 2052, 2053, 1895, 2237, 1509, 276, 1275, 1816, 1817, 2240]. Higgs [2034, 361, 757, 957, 755, 1993, 2312, 266, 549, 938, 1135]. High [329, 552, 675, 533, 1882, 1415, 1619, 1980, 1291, 1869]. High-energy [675, 1882]. High-frequency [533, 1415]. high-noise [1619]. High-order [552, 1869]. Higher [1265, 322, 821, 816, 812, 2154, 457, 1632, 432, 1277, 308, 669, 908, 1114, 41, 2287, 236, 220, 76, 490, 387, 9, 1672, 1059, 2286]. Higher-dimensional [322, 908]. Higher-order [1265, 821, 816, 812, 2154, 457, 308, 1114, 2287, 236, 76, 1059]. Highest [99, 161, 126, 21]. highly [1616]. Hilbert [1467, 1098, 2265, 2113, 1594, 2100, 1662, 1983, 593, 12, 1305, 1942, 761, 924, 1229, 1644, 484, 251, 626, 933, 456, 1230, 1397, 2031, 1266, 154, 1936, 324]. Hill [2077, 951]. Hill-determinant [951]. Hirota [808]. histories [1991, 2214]. hole [1566, 2057, 541, 1349, 707]. holes [2181, 114, 2293, 1511, 954]. holomorphic [51, 1936, 2113]. holomorphicity [929]. holonomic [1962, 2134]. Holonomy [1707, 465, 1733, 567, 1708]. Holstein [1715, 1274]. Homogeneous [689, 1241, 2130, 1411, 1211, 352, 1242, 1825, 2145, 38, 15, 520, 283, 2183, 1779, 244, 2126, 1334, 1770]. homothetic [78]. homotopy [834, 225]. Hooft [713]. hopelessly [170]. Hopf [317, 2201, 1973, 1748, 1786, 1428, 1494, 1067]. horizons [1724, 584]. Hosotani [306]. Hove [1050]. Hubbard [1469, 232]. Hurwitz [1747, 45, 1632, 2255, 409]. Huygens [2131, 1444, 995]. hydrodynamic [382]. hydrodynamical [656]. hydrodynamics [1126, 1316]. hydrogen [1758, 315]. hydrogenic [1137, 2225]. hydrogenlike [2062]. hyper [1123, 1750, 276]. hyper-Kähler [1123, 276]. hyper-Kählerian [1750]. Hyperbolic [1827, 1280, 1522, 1268, 2141, 2065, 293, 3, 1317, 858, 2283, 2147, 1]. Hyperbolicity [892, 1854]. hyperboloid [1141, 1187]. hyperelliptic [1568]. hypergeometric [2078, 374, 96]. hyperheavenly [2247, 108]. Hyperinflation [81]. hyperplane [1904]. hypersphere [824]. hyperspherical [1835, 827, 241, 532]. hypersurfaces [661, 1366]. hypervirial [1259].

I-PI [620]. Ianiro [1507]. ibid [2318]. Ichikawa [169]. icosahedral [159]. icosahedron [1571]. ideal [231, 2229, 2098]. ideals [466]. identical [384]. identically [344]. identities [1972, 467, 2285, 1516, 1019, 1020, 1551, 767, 752, 60, 909, 497, 34]. identity [1872, 1650, 643, 2196, 1022, 711]. II [1640, 682, 2241, 1621, 699, 195, 116, 934, 814, 354, 1549, 418, 1020, 2016, 157, 832, 2003, 2038, 811, 971, 1109, 1845, 1518, 40, 626, 222, 1301, 1102, 1010, 2225, 1916, 1150, 2246, 407, 679, 1637, 953, 148, 1710, 2134, 1896, 1899, 198, 1675, 186, 625, 1970, 706, 1817, 32]. III [1099, 1813, 354, 1924, 1516, 1741, 158, 138, 1021, 931, 1396, 636, 1302, 335, 1676]. Illustration [2179, 2241]. images [2217]. imbedding [2298, 1538, 923]. immersed [1676]. immersions [735]. impedance [314, 342]. imperfect [2006, 254, 1124, 2135, 1768]. implications [104]. important [823]. imposing [1699]. imprimitivity [273]. Improved [1858, 1724]. improvement [1198]. Improving [613]. impulsive [2318, 1450, 2059, 110, 625]. Ince [1745]. incidence [1504]. Including [508]. inclusion [387]. incomplete [1978]. increasing [461]. indecomposable [62, 551]. independent [374]. index [2103, 2245, 1748, 1915, 1916]. indices [2156, 1333, 295, 550]. individual [211]. Induced [480, 1874, 1973, 100, 1861, 1134, 1879, 2195, 2184, 807]. induction [1709, 1710]. inequalities [1756, 1858, 763, 1164]. inequality [1930]. inertia [1820]. Infinite [2232, 163, 609, 231, 710, 417, 418, 314, 201, 2066, 749, 1931, 1356, 777, 279, 937, 899, 99, 161, 1896, 1899, 112, 1050, 2127]. Infinite-dimensional [163, 609, 417, 418, 201, 2066, 749, 1356, 899, 1050]. infinitely [1532, 948, 1548]. Infinitesimal [507]. Infinitesimally [967]. infinity [891, 1651, 182, 2246, 1500, 178]. Inflation [75]. inflationary [1430]. influence [743]. information [452, 823, 1534, 2149]. Infrared [1470, 221, 1010, 360, 1014]. inherent [676]. inheritance [2180, 1257, 662, 1364]. inhibited [1485]. Inhomogeneous [62, 551, 854, 2282, 2319, 2255, 405, 2028, 533, 248]. Initial [138, 468, 357, 556, 55, 2137, 576, 1039, 496, 936, 401]. initial-boundary [496]. initial-data [2137]. inner [906]. Inönü [1553]. Inönü-Wigner [1553]. instability [701, 897, 2011, 954]. Instanton [1314, 834, 1489, 64]. Instanton-like [1314]. instantons [1005, 1431, 1527, 2056, 2058]. Integrability [1540, 1260, 467, 1373, 281, 1691, 549, 2054, 121, 813, 412, 1618, 2282, 2319, 381, 1201, 139, 1481, 450, 1129]. Integrable [440, 462, 1771, 686, 1895, 410, 1770, 570, 2140, 1157, 1620, 1621, 1957, 2170, 1502, 2075, 566, 814, 1066, 1813, 1630, 897, 819, 314, 852, 1920, 1814, 1459, 375, 205, 660, 2048, 2008, 2050, 1421, 1892, 1959, 1960, 1893, 165, 1945, 490, 1154, 1160, 1962, 2134, 1625, 2053, 1922, 2288, 304, 1541, 1509, 1197, 1422, 1177, 711, 128]. Integral [1474, 101, 297, 2100, 1079, 500, 2208, 697, 1380, 133, 1497, 1919, 1719, 2273, 535, 894, 414, 1975, 895, 616, 1604, 1149, 423, 531, 180, 2126, 70, 1591, 786, 1343, 1409, 1408, 1730, 1802, 373]. Integrals [1714, 1458, 1159, 1678, 989, 1612, 1595, 905, 203, 1598, 589, 1476, 869, 792, 1721, 1060, 1842, 502, 2148, 494, 1838, 460, 825, 194]. integrand [745]. integration [1178, 467, 726, 2305, 1373, 105, 692]. integrations [190]. integrodifferential [984, 2009, 419]. intelligent [2001]. interacting [758, 1925, 1043, 122, 44, 488, 833]. Interaction [2303, 110, 625, 1167, 966, 1168, 366, 208, 1251, 1264, 1879, 1619, 921, 387, 2176, 994, 568, 583, 1041, 620]. interactions [813, 505, 801, 1335, 50, 141, 293, 461, 922, 499, 82, 530]. intercell [882]. interchange [763]. interesting [137]. interior [540, 1282, 961]. intermediate [1337, 1980]. Internal [289, 2096, 792, 1420, 222]. interpolation [106, 1456, 1737, 2194]. interpretation [1574, 2273, 656, 541, 90, 597, 2122]. interpretations [2212]. interrupted [1785]. intertwining [1905]. interval [1777]. Intrinsic [1000, 1629, 1656]. introduce [2064]. introduction [2105, 1070, 411]. Invariance [1361, 1804, 503, 1718, 2034, 1052, 1860, 1237, 1523, 794, 1132, 774, 1970, 860, 232]. Invariant [546, 1779, 1850, 544, 61, 121, 2189, 653, 195, 934, 1316, 558, 1538, 559, 892, 915, 923, 1908, 1447, 1054, 1269, 1237, 1523, 1159, 1819, 1226, 1311, 648, 2277, 982, 818, 164, 1442, 666, 1846, 1013, 1082]. Invariants [2166, 1685, 242, 2093, 2208, 1194, 871, 2202, 2322, 102, 1559, 767, 2302, 239, 253, 1815, 399, 487, 1819, 1959, 991, 236, 1829, 1933, 143, 2068, 1223, 676, 889, 1144, 183]. Inverse [317, 1078, 1387, 855, 999, 624, 401, 826, 1120, 1524, 2266, 1884, 1622, 1812, 2180, 311, 252, 1808, 1732, 2003, 1340, 176, 14, 419, 870, 226, 1306, 623, 1592]. inverses [1608]. inversion [1378, 1703, 1610]. Investigation [1227]. Involutive [493, 771, 655, 2052, 2053, 1866]. involving [82]. irrational [885]. Irreducibility [1244]. Irreducible [1220, 234, 1184, 2258, 1702, 1076, 196, 850, 607, 8, 598, 99, 161, 333, 1911, 1496]. irrep [339]. irreps [522, 230, 371, 197]. irreversible [547, 1618, 639]. Ishimori [401, 1092]. Ising [1452, 1889, 81, 1571, 1351, 802]. isolated [1436]. Isometric [1510, 735]. isometries [1634, 1241, 154]. Isometry [956, 1897]. isoscalar [1145, 236, 200, 978]. isosceles [1806]. Isospectral [1655, 1196, 313]. isospin [72]. isotopic [279]. Isotropic [2128, 1277, 1411, 1504, 1255, 1129]. isovectors [800]. issue [876]. Itô [1514]. iteration [778]. iterations [1392]. Ito [421]. IV [1746, 468, 1399].

J [1145, 2317, 723, 2316, 1466, 57, 219, 1522, 1714, 1467, 1715, 2313, 1372, 2322, 2319, 1938, 677, 3, 1317, 123, 2318, 2310, 1468, 1140, 1250, 1098, 520, 1658, 2320, 1523, 1870, 2311, 2312, 2112, 2309, 1251, 597, 166, 598, 2264, 226, 1789, 2265, 56, 532, 1672, 2314, 521, 337, 2113, 153, 1282, 2321, 1755, 1871]. Jacobi [1872, 1157, 1371, 930, 211, 439]. Jacobi-elliptic [930]. Jaynes [907]. Jeffreys [2275]. Jet [1111, 2185, 206]. jet-bundle [206]. Jones [2204, 806, 287]. Jordan [156, 157, 158, 849]. Jost [624, 463, 400]. Jump [2047]. jumps [749]. Junction [661]. JWKB [424].

Kac [1184, 31, 272, 1486, 565, 1843, 441, 979, 1553, 507, 899, 269, 600, 324]. Kadomtsev [2171, 448, 1335, 1336, 1323, 2235, 1509, 1577, 1816, 1817]. Kähler [615, 1123, 90, 597, 1367, 276, 1935]. Kählerian [1750, 1797]. Kaluza [542, 2013, 2292, 114, 2096, 1850, 1349, 1695, 1696, 1270, 1348]. Kantowski [1819, 1081, 1082]. Kasner [84]. Kato [134]. Kaup [1654, 1848, 1625, 169, 2238]. Kaya [2318]. KdV [1260, 412, 689, 1419, 2009, 1148]. KdV-like [1260]. KdV-type [412]. Kemmer [2163]. Kendall [1065]. Kepler [50, 1446, 2231, 207]. Kepler-type [50]. kernel [2094, 1152, 2065, 1267, 1726, 18, 11, 166, 1865, 456, 482]. kernels [61, 595, 1474, 624, 435, 1055, 1303]. Kerr [2015, 2016, 39, 959, 541, 40, 1416, 2294]. Killing [56, 1209, 219, 1579, 584, 581, 7, 1046, 910, 662, 703, 663]. Kilmister [664]. kinds [1243]. kinematic [487]. kinematical [606, 98, 407, 288, 1859]. kinematics [1539]. kinetic [867, 1988, 1569, 1217]. Kink [506, 955, 2136, 2303]. kink-type [2303]. Kinking [1748]. kinks [1084]. Kirchhoff [2087, 1032]. Klein [2112, 542, 2013, 1484, 2292, 114, 420, 53, 1747, 1011, 2096, 156, 157, 158, 560, 1075, 2116, 1200, 1277, 1130, 1850, 1682, 1349, 1695, 1226, 1696, 364, 1270, 1348]. KMS [2270]. knot [143]. knots [2205].

Knowledge [2164]. Kolmogorov [839, 2087]. Konno [169]. Korteweg [1653, 172, 1623, 2234, 1375, 1990, 2174, 571, 1508, 2051, 2010, 2053, 2175, 1693]. Kostant [1541]. Koszul [2263]. KP [2173, 331, 444, 214, 2237, 1275, 2240]. Kramers [2275]. Kreın [1244, 744]. Krichever [1382, 93, 1674]. Kronecker [60, 197]. Kruskal [1736]. Kubo [2108]. Kulshreshta [1372]. Kummer [2261]. Kupershmidt [1893].

L. [423]. ladder [1394, 675, 275]. Lagrange [752, 183]. Lagrangian [1099, 797, 2036, 1840, 821, 459, 17, 1760, 309, 76, 1683, 511, 1887, 9, 1672, 433, 993, 128, 1445, 2286]. Lagrangians [2285, 2086, 457, 2166, 325, 308, 990, 1447, 2046, 736, 822, 282]. Laguerre [2023, 1570, 1112]. Lamé [1780]. Lämmerzahl [2309]. Landau [195, 593, 1603, 1576, 774, 193, 684]. Lane [1713]. Laplace [1558, 1226, 1015, 280]. Laplacian [1169, 65]. Laplacians [435, 1055]. Large [393, 1658, 345, 659, 1259, 675, 1113]. large-momentum-transfer [675]. laser [1016]. Lattice [2132, 926, 1577, 1639, 1513, 1540, 2250, 2170, 836, 1834, 2222, 1043, 1356, 512, 1323, 1799, 1774, 281, 1571, 294, 118, 715, 1409, 1408, 712, 914]. Lattices [1019, 1020, 722, 668, 1908, 1243, 1480, 513, 802, 1923]. law [1806, 963, 1733, 1188]. laws [876, 419, 1789, 1482, 1979, 774, 310]. Lax [449, 1221, 565, 1199, 1196, 648, 1962, 2134, 1575, 2176]. layer [2127]. layered [2127]. layers [882]. Leach [316]. lead [172]. leaves [650]. Lebowitz [1507]. led [1963]. left [61, 484]. Legendre [1224, 1930, 168, 76]. lengths [1117]. Lennard [287]. Lennard-Jones [287]. lens [2217, 1744, 1697]. lensing [1744, 1659]. less [1708]. level [1378, 1984, 1402, 53, 853, 781, 811, 1603, 2159, 639]. levels [886, 2116, 1436, 1233, 1386, 1765]. Levinson [47, 2035]. Levy [311]. Li [1870]. Liapunov [1776]. Lie [242, 61, 2186, 564, 2295, 1279, 698, 699, 529, 340, 296, 672, 2297, 440, 2220, 782, 1924, 558, 1359, 413, 1663, 1664, 163, 201, 62, 405, 551, 918, 1931, 2302, 443, 1285, 1783, 1142, 979, 1447, 686, 2260, 487, 450, 1373, 1706, 1463, 225, 1976, 1554, 1635, 899, 2040, 148, 99, 161, 2068, 442, 2306, 1360, 1303, 898, 1685, 1673, 1218, 1191, 808, 408, 202, 1868, 333, 2263, 809, 730, 1832]. Lie-consistent [2260]. Lie-supertriple [2040]. Lieb [806]. Lifshitz [1576, 774, 684]. light [742, 1276, 911, 1427, 434, 1606]. light-cone [434]. light-front [1606]. like [1260, 1314, 149, 890, 1250, 1331, 1276, 41, 1865, 266, 1022, 2290, 2291, 1950, 714]. limit [667, 388, 880, 836, 1346, 721, 963, 1723, 2211, 1845, 875, 1070, 1232, 1437, 1264, 485, 1671, 1213, 2085, 920]. limited [1898]. Limiting [1473, 943, 1768]. limits [1942, 1914, 2147, 1089]. Lindblad [746, 1680]. Line [1335, 1884, 1805, 855, 781, 811, 461, 994, 1521]. Linear [1573, 2251, 1058, 1531, 706, 1099, 1288, 632, 1279, 2172, 1194, 413, 314, 3, 1317, 894, 258, 1590, 247, 206, 552, 1491, 140, 556, 1217, 1280, 1788, 245, 1673, 2032, 359, 1]. linearizable [858]. Linearization [1290, 456, 1745, 1505, 427]. linearized [248, 1148]. linearly [396, 515, 657]. lines [2202, 2322]. link [864, 1559, 1933, 437, 2204, 2206, 1354]. linking [2205, 2204, 936]. links [2208]. Liouville [1502, 205, 2004, 1625, 1056, 676]. Liouville-space [676]. Lippmann [1267]. liquid [1374]. little [1276]. Liu [532]. Lobachevskian [260]. Local [1537, 2034, 893, 1959, 119, 98, 407, 332, 947, 1525, 1292, 270, 1718, 3, 1317, 795, 627, 1054, 2245, 222, 950, 221, 1010, 1493, 1013, 1156, 1, 232]. locality [1560]. Localization [1633, 1528, 215]. Localized [122, 1252, 1276, 1997, 2175]. locally [1750, 442, 344]. located [2294]. Loewner [1421]. logarithmic [1192]. logic [475, 1991, 2214, 619]. logics [1164, 1947]. Long [2219, 2129, 758, 1879, 1963, 801, 50, 1337, 1889, 898, 1592]. Long-range [2219, 1879, 801, 50, 1889]. Long-time [758]. long-wave [1337, 898]. look [1607]. Loop [1739, 671, 694, 1383, 1655, 1562]. loops [1772]. López [226]. Lorentz [2264, 1161, 1731, 523, 1905, 2089, 1670, 5, 171, 936, 1936, 1970, 2113, 498]. Lorentzian [1897, 871, 788, 795, 1978]. Lorenz [1714, 1458]. loss [440, 1738]. Lotka [1194]. Lovelock [1083]. Low [252, 753, 938, 1289, 1387, 455, 1105, 1213]. low-density [1213]. low-dimensional [1289]. Low-frequency [252, 1387]. low-order [455]. low-rank [1105]. Low-temperature [753]. Low-velocity [938]. Löwdin [335, 1399]. Lower [1401, 620, 1653, 997, 906]. lowering [1519]. lowest [527]. Lubanski [879]. Lucas [1752]. Luneburg [150]. Lüscher [1435]. Luttinger [1457]. Lyra [789].

M [2316, 2318, 226]. M.I.T. [759]. Maass [435, 1055]. Macías [2309]. Mack [1435]. macromolecules [191]. Madelung [964]. Madelung-sum [964]. magic [710]. magnet [1958, 1281]. Magnetic [90, 597, 305, 1292, 334, 1558, 840, 1060, 658, 1416, 1357, 1296, 1441, 2236, 207]. magnetized [541]. magnetogas [57, 2]. magnetohydrostatic [645]. Magnus [69]. Majid [2263]. Majorana [297, 1595, 2182]. Mandelstam [1972]. Maneff [1806]. manifold [1649, 1989, 444, 222, 636, 1810, 171, 1270, 2200]. manifolds [1236, 106, 1573, 649, 394, 1902, 615, 15, 520, 689, 1216, 1850, 1159, 567, 1750, 1708, 1707, 1029, 1367, 105, 1998, 2011, 900, 987, 993, 1770]. Mann [558]. many [594, 1114, 309, 1302, 1180, 1734, 494, 948]. many-baryon [1302]. many-body [594, 1180, 494]. many-particle [1114, 1734]. many-time [309]. map [80, 588, 2303, 670, 1749, 1830]. mapping [2192, 748, 1300, 1301, 1302, 434]. mappings [1793]. maps [495, 536, 845, 694, 1383, 1908, 1893, 1927, 1090, 244]. Marchenko [1247]. Marder [41]. Marder-like [41]. Marinari [2158]. Markov [534, 2162, 1050]. Martin [2108]. Maslov [1333]. Mass [2081, 1849, 2034, 2137, 2138, 253, 751, 1993, 2312, 1417, 1564, 1838, 1697]. mass-momentum [2137]. masses [965]. massive [1161, 320, 589, 965, 1215, 668, 1722, 1651, 122, 1049, 82, 637]. massless [480, 2114, 669, 1032, 221, 360, 1810, 1954, 82, 328, 1013, 1087, 1081]. master [1833, 885]. Matching [39, 709, 1480]. material [1538, 858]. materials [1505, 104]. Math [1145, 2317, 723, 2316, 1466, 57, 219, 1522, 1714, 1467, 1715, 2313, 1372, 2322, 2319, 1938, 677, 3, 1317, 123, 2318, 2310, 1468, 1140, 1250, 1098, 520, 1658, 2320, 1523, 1870, 2311, 2312, 2112, 2309, 1251, 597, 166, 598, 2264, 226, 1789, 2265, 56, 532, 1672, 2314, 521, 337, 2113, 153, 1282, 2321, 1755, 1871]. mathematical [1162, 575, 411, 387]. mathematically [954]. Mathieu [535]. matrices [519, 536, 1381, 2300, 1587, 847, 1094, 887, 2029, 1570, 99, 161, 783, 889, 1315, 1767]. Matrix [327, 1036, 1383, 1784, 978, 1345, 1145, 1573, 2296, 2170, 466, 2063, 1737, 603, 1327, 780, 2077, 852, 1298, 66, 2023, 1205, 868, 403, 2024, 2259, 1704, 1680, 1909, 596, 561, 562, 400, 21, 485, 1029, 1137, 2225, 1555, 351, 1180, 200, 491, 492, 608, 1647, 1494, 1097, 374, 1325, 267]. matrix-valued [466]. Matter [2180, 742, 2320, 1398, 2184, 833, 1041, 2321, 1004]. Maurer [1461]. Maximal [1424]. maximally [1169, 1511, 111]. Maximum [127, 2301, 2199, 2149, 762, 685, 346, 895, 616, 1828, 2290, 1429]. Maximum-entropy [762]. Maxwell [1211, 2101, 284, 420, 1687, 861, 78, 402, 1843, 1967, 393, 1658, 438, 2230, 585, 937, 496, 1506, 706]. McVittie [1417]. mean [867, 588, 1798, 59, 112]. mean-field [112]. Meaning [1892]. means [742, 1114, 922, 2164]. measure [1619, 513]. measurement [422]. measurements [29, 1791, 475, 823, 1801, 1947]. measures [915, 65]. mechanical [1719, 1766, 1118, 1977, 1895, 1969]. mechanics [785, 51, 698, 699, 944, 29, 1793, 386, 1242, 417, 418, 1598, 779, 1766, 210, 1498, 1986, 347, 215, 1341, 1688, 1111, 1990, 2272, 1119, 1304, 1602, 2037, 2038, 1073, 2160, 1499, 573, 1070, 261, 629, 2007, 1667, 1179, 2277, 1296, 1532, 1605, 387, 1563, 1811, 1648, 548, 947, 1610, 695, 349]. mechanism [188, 1135]. media [1078, 1120, 1839, 861, 1276, 1504, 140, 2167, 2127]. mediated [505]. medium [533, 248]. Meinardus [2232]. Meissner [193]. Mellin [18, 11, 166, 1591]. meltdown [725]. membranes [1501]. Memory [554]. Memory-dependent [554]. Meromorphic [785, 1986, 298]. meson [476]. mesoscopic [2300]. metaplectic [642, 1074]. Method [1198, 611, 501, 787, 1178, 2250, 1033, 1812, 476, 2064, 2298, 2132, 780, 1538, 1665, 2105, 1214, 776, 1355, 1060, 1447, 176, 159, 1147, 165, 1520, 648, 1180, 1121, 1846, 1521, 464, 1754, 191, 2291, 1363]. Methods [567, 1103, 1736, 413, 327, 1111, 1490, 1142, 997, 1591]. Metric [2294, 874, 1433, 322, 523, 2096, 582, 1328, 540, 752, 1742, 39, 1416, 544, 549, 1365, 1391, 1936, 1970, 2113]. metric-affine [752, 1391]. Metrics [451, 219, 1003, 956, 1897, 558, 289, 661, 1123, 2015, 2016, 1390, 1841, 795, 1819, 857, 960, 955, 1081, 1082]. MHD [683]. microlensing [1139]. microscopic [682, 1716, 681]. microstretch [1374]. Mie [26]. Mills [361, 266, 337, 1578, 1900, 1527, 2056, 329, 1126, 1489, 477, 615, 2048, 2118, 1448, 1681, 2211, 1634, 755, 144, 14, 222, 837, 2224, 490, 226, 1012, 1393, 186, 1542, 397, 954, 2200, 187]. Milnor [2204]. minima [700]. Minimal [1398, 1042, 1329]. minimally [912]. Minimization [195]. minimizes [1439]. minimum [285, 1439, 1410]. Minisuperspaces [912]. Minkowski [1629, 2172, 477, 2089, 1032, 2224, 415]. Minkowskian [1649]. minuscule [60]. missing [864]. Miura [1653, 1893]. Mixed [691, 1182]. mixing [348, 1212]. mixing-enhancing [348]. mixture [1738]. Mixtures [1119, 1089]. MKDV [660]. Möbius [2064, 1464, 1788]. mode [1795, 1063]. model [718, 388, 682, 1716, 801, 843, 1901, 1769, 897, 913, 1649, 1084, 841, 1215, 681, 2258, 431, 1044, 1723, 1238, 2037, 2038, 756, 800, 1355, 115, 926, 1398, 2211, 1481, 1836, 481, 1092, 1469, 1961, 592, 1505, 419, 1301, 1798, 488, 1800, 81, 1532, 1997, 362, 1338, 91, 941, 545, 1016, 741, 1571, 759, 810, 2000, 1270, 907, 938, 1177, 356, 537, 1457, 1050, 2044, 516, 1536, 1816, 1817, 1507, 232, 1185, 1679]. modeled [189]. Modeling [1738]. Models [1186, 1784, 1975, 1067, 1566, 882, 919, 1411, 1526, 1062, 2075, 74, 2135, 867, 1212, 1613, 1663, 67, 757, 1057, 1834, 1814, 858, 2105, 1473, 1307, 218, 265, 971, 1396, 753, 875, 2039, 299, 1249, 2184, 884, 221, 1010, 430, 1280, 288, 142, 1646, 1606, 925, 789, 1197, 1040, 1351, 326, 714]. Modes [1705, 1189]. modification [315]. Modified [825, 1745, 172, 689, 2174, 1508, 2053, 938, 295, 377, 1740]. Modular [2267, 1052, 1861, 1846]. modulated [1513, 2177]. modules [124, 1486, 979, 1134, 1266, 1980, 333, 1832]. moduli [1839, 2210]. Mohapatra [1486]. molecular [1068]. molecules [189, 190, 1368]. Moment [239, 867, 2301, 694, 1383, 1171, 1416]. momenta [102, 1276, 638]. moments [1858, 1387, 685, 252, 179, 1811, 1429, 2199]. Momentum [1292, 1586, 2137, 830, 445, 1552, 544, 675, 577, 586, 1348, 1955]. momentum-dependent [577]. monodromy [848]. Monogenic [1628]. Monopole [674, 334, 50, 1558, 658, 266, 207]. monopole-like [266]. monopoles [90, 597]. monotone [1392]. Moody [1184, 272, 1486, 565, 441, 1553, 507, 899, 269, 600, 324]. morphism [1751]. Morphology [1008]. Morse [998, 1139]. Moser [2237, 2000, 2054]. Moshinsky [241, 532]. Motion [2013, 542, 1035, 1194, 1714, 902, 78, 1987, 792, 1458, 859, 1388, 1773, 91, 1271, 994, 470, 1838]. motions [61]. Movable [1057]. moves [2189]. moving [1643, 1599, 705, 1754]. Moyal [1293, 1983, 389]. MR [1145, 2317, 723, 2316, 57, 219, 1714, 1467, 1715, 2313, 1372, 2319, 1938, 677, 3, 1317, 2318, 2310, 1468, 520, 1658, 1523, 1870, 2311, 2312, 2112, 2309, 1251, 597, 166, 598, 2264, 1789, 56, 532, 1672, 2314, 337, 1282, 1755]. Mullins [652]. multi [1489, 431, 2270, 1154, 382]. Multi-Hamiltonian [382]. multi-instanton [1489]. multi-Poisson [1154]. multi-time [2270]. multi-urn [431]. Multibody [596]. multicenter [1550]. Multicentered [1511]. multichannel [1873]. Multicomponent [1509]. multiconfiguration [604]. Multidimensional [383, 311, 1316, 854, 1863, 755, 872]. multidimensions [1737, 1885, 1201]. Multifold [1343]. multigroup [569]. multimomentum [2289]. Multiparameter [1359, 1455, 1933, 849].

Multiparametric [1590]. Multiphase [763]. Multiple [2218, 384, 1182, 55, 1376, 1149, 385, 1744]. multiplets [1234]. multiplication [2143, 1851]. multipliers [183]. multiply [2073, 1627]. multipolar [1125]. multipole [650, 1330]. multipoles [298]. multipositon [2051]. multipositon-soliton [2051]. multiquark [748]. Multiresolution [646]. multisolenoid [744]. multispeed [913]. Multivariable [233, 930, 983, 737, 773]. multivariate [346]. Multivectorial [670, 363]. Murakami [1094].

N [57, 1714]. N.P [36]. Nahm [968]. naked [751]. Nambu [1026, 1027]. narrow [1338]. Natanzon [1497, 2078]. Natural [1329, 1675, 1676]. nature [846]. near [1502, 588, 891, 463, 247, 2147]. nearest [2128, 1423, 781, 811, 1243]. nearest-neighbor [1423, 1243]. Necessary [446, 839]. need [2217]. Negative [687, 1472]. neighbor [2128, 1423, 1243]. neighbors [874]. Nelson [1763]. nested [497]. net [2124]. network [875]. Neumann [1939, 463, 1945, 2053]. Neumann-type [1945]. neural [875]. Neutral [857, 1764, 2005]. Neutrino [1155, 1369, 290]. neutrinos [2014]. neutron [569]. Neveu [1937]. Newell [1654, 1848, 1625, 169]. Newman [959, 541, 1416, 2294]. Newton [1957, 1247]. Newtonian [1062, 804, 1114]. Nielsen [1652]. Nijenhuis [1022]. nilpotent [61, 2114, 818, 1937]. Nizhnik [1961]. No [459, 1145, 1640, 2317, 723, 2316, 57, 219, 1714, 1467, 1715, 2313, 1372, 2319, 1938, 677, 3, 1317, 2318, 2310, 1468, 520, 1658, 1307, 1523, 2260, 1870, 2311, 2312, 2112, 2309, 1251, 597, 166, 598, 2264, 226, 1789, 56, 532, 1672, 2314, 337, 1282, 1755, 459]. no-go [1640, 2260]. no-hair [1307]. Nodal [2120]. Nodal-surface [2120]. Noether [2285, 326]. Noetherian [1615]. noise [1341, 1619]. Noisy [988, 80]. Non [514, 487, 1799, 741, 1964, 1091, 141, 844, 918, 1840, 1350, 1615, 2125, 222, 636, 1999, 2005, 63]. Non-Abelian [514, 487, 1799, 1091, 141, 2125, 1999, 63]. non-Einsteinian [1350]. non-Lagrangian [1840]. non-Lie [918]. non-neutral [2005]. non-Noetherian [1615]. non-null [1964]. Non-Riemannian [741, 222, 636]. non-self-dual [844]. nonabelian [1189]. Nonassociative [766, 264]. nonautonomous [1505, 1223]. noncentral [697, 1297]. Nonclassical [1736, 1400, 618]. noncoinciding [548]. noncollinear [543, 793, 1040]. Noncommutative [66, 67, 1237, 1523, 2030, 1977, 1209, 2213, 2203, 1337, 1201, 2221, 389, 2304, 2192]. noncommuting [2164]. noncompact [174, 126, 527, 487, 1092]. nonconfocal [771]. noncovariant [779]. nondegenerate [1436]. nondemolition [422]. Nondiagonal [958, 960]. Nondispersive [1131, 1130, 364]. nonequilibrium [1618]. Nonexistence [203]. nongeneric [783]. nonholonomic [1111]. nonhomogeneous [1078, 1120, 1574]. Nonimpulsive [793]. noninteger [1856, 2188, 2119]. Nonintegrability [1009]. noninteracting [552]. noninteraction [208, 1251]. nonintertwining [850, 607]. Noninvertible [701]. nonisospectral [1335, 1196, 313]. Nonlinear [386, 240, 569, 897, 1952, 2080, 2092, 164, 1886, 1071, 1639, 2129, 2130, 121, 1951, 367, 856, 1855, 440, 566, 814, 1066, 984, 1813, 1630, 1574, 1771, 858, 2105, 918, 1195, 1200, 1383, 1142, 1307, 2174, 115, 17, 751, 380, 173, 1481, 450, 1373, 206, 1226, 456, 1261, 973, 883, 165, 1181, 2281, 52, 648, 490, 430, 362, 1962, 2134, 732, 2052, 1166, 304, 2177, 1324, 568, 1059, 1443, 1325, 1536, 1507, 326, 1593]. nonlinearities [686]. nonlinearization [1459]. Nonlocal [1419, 684, 2093, 554, 2046]. nonlocality [503]. nonminimal [740]. nonminimally [1411]. nonperfect [74]. nonperiodic [812]. Nonperturbative [376, 1414, 1996, 1713, 398]. nonpolynomial [501]. nonprincipal [1994]. nonprobabilistic [847]. nonreciprocal [1376]. nonregular [1470, 76, 993]. Nonrelativistic [920, 388, 739, 944, 1805, 1885, 180, 407, 332]. Nonrenormalization [1912]. nonsemisimple [2028]. Nonsingular [747, 2293, 1434, 1755]. nonsmooth [861]. Nonspurious [175]. Nonstandard [2117, 1011, 1529, 2272, 502, 1284]. nonsymmetric [924, 1098]. nonsymmorphic [1698]. Nontransitive [273]. Nontrivial [1177, 813, 225, 137, 1494]. Nontypical [148]. nonunimodular [28]. Nonuniqueness [1884]. Nonuniversal [1570, 1555]. nonvanishing [1620, 1621]. Nonzero [1123, 1168]. norm [546, 835, 659]. norm-continuity [546, 835]. Normal [738, 1365, 576, 1775]. normalizers [1517, 1518]. Normally [58, 524]. Norms [362]. Note [2057, 1658, 1114, 2230, 850, 409, 647, 908, 1711, 1583, 936, 1133]. notes [523, 555]. Nöther [2310, 1953]. Novel [1118, 1350, 1421, 72, 456, 364]. Novikov [1901, 1961, 1382, 93, 1674]. Novikov-type [93]. nuclear [682, 1716, 681, 1917]. nuclei [682]. Null [2139, 1477, 1651, 1964, 661, 2320, 182, 2246, 1012, 110, 625, 583, 2321]. number [863, 506, 2217, 2007, 1748, 1480, 178, 1156, 620]. numbers [1900, 2253, 1967, 1645, 2246, 709, 2204]. Numerical [2281, 291, 327].

obey [632]. object [40, 1125]. objects [861, 872]. oblate [259]. oblique [1504]. observables [1793, 1762, 2164, 59, 349]. observations [841]. observers [1295]. obstacle [2126]. obtained [437, 95, 1384, 1671]. Obtaining [1558, 179]. octet [756, 437]. octic [1443, 1386]. octonion [679]. Octonionic [969, 605, 981, 1636]. octonions [636]. odd [2217, 716, 1797]. ODEs [2238]. off [838, 464]. off-shell [838]. Okubo [558]. old [1812]. Olver [687]. One [1223, 1099, 1345, 1120, 1594, 2266, 2129, 1382, 454, 1858, 1497, 320, 1503, 1061, 1118, 606, 1808, 2003, 575, 1652, 1352, 991, 1368, 1311, 1520, 279, 1562, 1248, 1129, 318, 307, 2035, 663, 482, 1097, 2044, 216, 1697, 232, 1686, 1871, 1808]. one- [1311, 1097, 1697]. One-dimensional [1223, 1099, 1120, 1594, 2129, 1497, 1118, 606, 1808, 2003, 575, 1652, 1352, 991, 1520, 1129, 307, 2035, 216, 232]. one-electron [1858, 1368]. one-Killing-vector [663]. one-speed [482]. Onsager [844]. onto [1751]. Open [1727, 838, 2036, 746, 1049, 1501]. opening [1115]. operations [1428]. Operator [2072, 317, 1471, 863, 135, 1699, 174, 916, 1985, 1199, 1515, 659, 134, 58, 524, 1138, 1516, 1472, 365, 2259, 1796, 1907, 2026, 115, 980, 596, 211, 1682, 2112, 646, 1733, 879, 1196, 850, 607, 1338, 732, 1882, 962, 815, 1838, 1867, 1366, 1206, 1144, 1322, 684, 498]. operators [949, 1546, 1146, 1292, 881, 1622, 2296, 1823, 1662, 2155, 1502, 1735, 1737, 1221, 1717, 1549, 455, 1905, 565, 1095, 893, 458, 238, 1516, 1834, 1341, 1826, 740, 1318, 1077, 451, 1644, 1861, 259, 47, 1892, 2262, 1219, 1519, 1106, 1107, 525, 434, 1562, 2226, 1321, 1022, 2306, 975, 729, 1608, 977, 1143, 1753, 1981, 1097, 687, 621]. Oppenheimer [1033]. optical [1951, 1247, 999]. Optics [2277, 1279, 1833, 1340, 1188]. optimal [189, 1316]. Optimized [1550]. orbifold [2065]. Orbit [1105, 2152, 699, 623]. Orbit-orbit [1105]. Orbital [2012, 1399]. orbits [2284, 1184, 124, 235, 2114, 1447, 1023, 1521]. Order [2116, 1524, 1265, 1641, 2093, 1260, 2154, 653, 240, 203, 1965, 455, 457, 77, 1254, 821, 325, 308, 669, 1956, 990, 1312, 1705, 1114, 2287, 236, 552, 285, 1439, 76, 734, 816, 1259, 2306, 812, 1271, 353, 1059, 1869, 1225, 1072, 2286]. ordered [58, 524, 1440]. ordering [1165]. orders [781, 811, 931]. ordinary [653, 1847, 1254, 1587, 734, 1443, 1225]. origin [453]. orthogonal [2153, 156, 1075, 331, 608, 1712, 737, 773, 395, 1637]. orthogonality [2256]. Orthogonalization [522, 1286, 1853]. Orthomodularity [1379]. orthonormal [199, 1146, 97, 446, 1024]. orthoposets [599]. orthosymplectic [265]. Oscillating [1821]. oscillation [2005]. oscillations [1615, 192, 213]. Oscillator [1353, 30, 747, 1288, 2268, 1792, 886, 1919, 1718, 642, 1048, 1434, 1395, 1118, 1559, 575, 350, 1680, 258, 1186, 1784, 658, 71, 1255, 1132, 175, 16, 521, 657, 2005, 1535, 1233, 1765, 1611, 1755, 621]. oscillators [2129, 865, 1294, 1253, 1077, 247, 1149, 279, 1880, 2070, 906, 1443, 1386, 951]. other [753, 1593]. outflow [290]. output [624]. overdetermined [799]. Overhauser [843]. overmoded [1616].

P [57]. packet [364]. Padé [2290, 997]. Padé-like [2290]. Painlevé [1745, 448, 1741, 1195, 1227, 375, 103, 1021, 1150, 648, 1894, 898, 2177, 907, 2238]. Painlevé-Ince [1745]. pair [1473]. pairing [1061]. pairs [449, 862, 1464, 648, 1575, 245]. Palatini [538]. Pan [1145]. Para [553, 43, 1498, 1913, 246]. para-Bose [1913, 246]. para-Fermi [43]. Para-squeezed [553]. parabolic [647, 2147]. paraboson [1608]. parabosonic [1984]. parafermionic [1485]. parallaxes [288]. parallel [1381, 1733, 1365]. parameter [321, 2095, 2298, 782, 2019, 1574, 1337, 1808, 1357, 1287]. parameterization [1747]. parameters [1538, 53, 2003, 837, 1315, 1767]. parametric [522]. Parametrization [97, 1790, 1462, 27, 374]. parametrized [974, 1981, 1339]. paraoperators [1751]. pararelativistic [1294]. Parastatistics [229, 2040, 1794]. Parasupercoherent [970, 229]. parasupersymmetric [698, 699]. Parasupersymmetry [1499]. Parisi [2158]. part [1522, 1268]. Partial [1982, 2313, 595, 629, 738, 2061, 632, 856, 2017, 566, 1066, 1737, 1813, 1630, 647, 413, 845, 103, 1021, 484, 206, 2194, 179, 648, 1979, 1866]. Partially [1226, 1160]. Particle [282, 1161, 334, 1735, 1380, 133, 779, 2034, 210, 1643, 1052, 170, 1989, 1114, 1993, 2312, 1738, 2161, 180, 896, 1800, 1417, 175, 1734, 925, 1881, 1922, 2288, 1089]. particles [682, 2013, 969, 320, 2114, 1263, 792, 1038, 2118, 211, 25, 727, 2125, 1235, 2039, 1726, 208, 1251, 1243, 1388, 1180, 1734, 362, 2084, 1764, 82, 328, 1013, 638, 474]. particular [2014]. Partition [1208, 2232, 1043, 1846, 249]. parton [1278]. parts [1937]. passive [882]. patchwork [2243]. Path [500, 697, 905, 133, 1497, 1919, 1598, 1678, 2208, 869, 1721, 726, 2273, 346, 414, 1604, 423, 180, 502, 70, 494, 460]. path-integral [2208]. pattern [1318]. patterns [676]. Pauli [388, 1485, 879, 2276]. PDEs [814]. Peierls [1915]. penetrationlike [1121]. perfect [354, 74, 352, 152, 289, 38, 1173, 1551, 540, 1694, 704, 1773, 709, 1480, 1581, 1023, 1042]. period [1008]. period-doubling [1008]. Periodic [1246, 170, 176, 1914, 555, 281, 513, 318]. periodical [1472]. permutation [519]. Pertinent [474]. Perturbation [835, 2179, 2241, 1560, 1071, 321, 2094, 504, 2057, 647, 355, 36, 79, 1665, 575, 997, 645, 358, 765, 1259, 16, 521, 1443, 1233, 1765]. perturbations [1260, 772, 539, 641, 65, 2060, 1386, 706]. perturbative [30, 856, 1596, 1766]. Perturbed [2268, 1062, 2077, 167, 915, 350, 330]. Petiau [2163]. Petrov [1209, 467, 354, 1741]. Petviashvili [2171, 448, 1335, 1336, 1323, 2235, 1509, 1577, 1816, 1817]. Pfaffians [1595]. Phase [1719, 2156, 458, 1650, 2125, 942, 112, 1351, 428, 1513, 863, 1034, 1432, 1731, 934, 284, 2145, 935, 720, 727, 1340, 775, 976, 1977, 810, 1339, 63, 1296]. Phase-integral [1719]. phase-invariant [934]. phase-modulated [1513]. Phase-space [2156, 942, 1296]. Phase-spaces [112]. Phaseless [1340]. phases [1104]. phenomena [1599, 1121]. phenomenological [1127]. phonon [1909]. photodissociation [190]. photon [190]. photons [284, 943]. Phys [1145, 2317, 723, 2316, 1466, 57, 219, 1522, 1714, 1467, 1715, 2313, 1372, 2322, 1938, 677, 3, 1317, 123, 2318, 2310, 1468, 1140, 1250, 1098, 520, 1658, 2320, 1523, 1870, 2311, 2312, 2112, 2309, 1251, 597, 166, 598, 2264, 226, 1789, 2265, 56, 532, 1672, 521, 337, 2113, 153, 1282, 2321, 1755, 1871]. Phys. [2319, 2314]. Physical [379, 632, 863, 189, 2103, 2255, 664, 2226, 942]. physics [2209, 2213, 2205, 2216, 2208, 2215, 1198, 2203, 1631, 2201, 2202, 2218, 2207, 2217, 1827, 2214, 876, 2210, 2211, 1379, 1438, 411, 651, 2204, 2206, 494, 2212, 2200]. PI [620]. picture [951]. piezoelectric [1787]. Planar [2099, 1806, 1104, 1572, 987]. Planck [2157, 963, 346, 1312, 572, 1358, 1520]. Plane [833, 787, 1640, 1698, 1380, 462, 254, 420, 391, 1080, 138, 468, 720, 1864, 1561, 1870, 445, 543, 1309, 572, 280, 435, 1055, 1918, 1407, 110, 625, 1677, 583, 1040, 2033, 1041, 706]. plane-autonomous [254]. plane-symmetric [1870, 1309]. plane-wave [420]. planes [2253, 1744, 1365]. plasma [897, 2280, 509, 2005]. plasmas [2261]. plate [1566]. Plateau [104]. Podolsky [2164]. Poincaré [677, 2152, 593, 523, 204, 1702, 1783, 673, 1864, 1552, 1830, 2195, 86, 360, 489, 674, 435, 1055, 164]. Poincaré-Bertrand [204]. Point [304, 1198, 1766, 589, 965, 829, 1140, 1844, 2118, 2244, 923, 2119, 651, 1810, 1143, 1891, 1050, 191, 1103, 1697, 1104, 1572, 530]. pointlike [861]. points [891, 181, 781, 811, 253, 463]. Poisson [2186, 1678, 2295, 1514, 770, 1503, 1929, 2066, 1807, 717, 1943, 1976, 1029, 1154, 1809, 1648, 276, 900, 1891]. polarizability [300]. polarization [1265, 1954, 1040, 1188]. polarizations [543, 793]. polarized [706]. poles [1021, 2235]. Polyakov [478]. polycrystals [300]. polygon [1878]. polygons [787, 1640]. Polynomial [2257, 1148, 672, 2208, 203, 2252, 1929, 60, 437, 1946, 2204, 1362, 410, 806]. polynomials [2153, 519, 1857, 1194, 2156, 2144, 1930, 983, 1290, 1074, 246, 1492, 316, 1752, 2206, 1112, 233, 737, 773, 1354]. polyspheroidal [176]. polytope [2253]. Poncelet [1567]. Pontryagin [530]. Pöschl [998, 377]. position [457, 215]. Positive [1498, 1163, 845, 1181, 1577]. positive-frequency [1163]. Positon [1624]. positron [319]. possibility [1997, 568]. Possible [1347, 1438, 1708]. post [1873]. post-adiabatic [1873]. posterior [422]. potential [1471, 1716, 500, 2173, 1297, 867, 700, 1207, 1167, 2157, 402, 957, 1330, 1796, 1835, 1604, 552, 1961, 1619, 1914, 179, 1180, 1734, 577, 1393, 1882, 286, 216, 96, 377, 207, 1950]. Potentials [2228, 501, 826, 1157, 2153, 1641, 1292, 880, 1939, 1253, 195, 1401, 1497, 203, 2078, 2229, 1119, 1304, 463, 755, 1996, 1192, 1605, 1728, 70, 1763, 822, 623, 318, 1500, 1565, 178, 586, 287, 32, 1896, 1899]. Potts [519]. Power [1487, 1917]. powers [1444, 687, 339]. Precise [2121]. predator [1691]. predator-prey [1691]. predictive [347]. Preliminary [851]. prescription [1994, 433]. presence [231, 130, 1155, 1369, 1764, 207]. present [925]. present-day [925]. Presentation [93, 1673]. preserving [2131, 2223, 563, 768]. pressure [753]. prey [1691]. Prigogine [631]. Primakoff [1715, 1274]. Primary [977, 1435]. principal [724, 2186, 24, 94, 800, 1153, 982, 433]. principle [635, 2131, 1444, 1485, 378, 1027, 167, 346, 876, 612, 995]. principles [1839, 1258]. prior [1534]. priori [2149]. Probabilistic [1406, 1379, 750]. probabilities [618, 1756, 1631, 1073, 1534]. probability [135, 1528, 215, 1436, 629]. problem [22, 317, 1594, 1853, 1566, 1525, 305, 594, 1342, 1622, 55, 1378, 2180, 2076, 462, 685, 1806, 1210, 1347, 2301, 1459, 924, 1098, 138, 468, 923, 283, 438, 1876, 1446, 1060, 17, 192, 213, 1340, 401, 827, 626, 1878, 14, 917, 2163, 870, 226, 953, 873, 992, 1625, 1039, 623, 1896, 1899, 104, 964, 496, 337, 1429, 2199, 778, 2149, 1282, 961, 1572, 1148, 207, 410, 1322, 866]. problems [1162, 1629, 1735, 1805, 1885, 569, 1839, 1766, 2103, 891, 576, 1808, 1732, 2003, 654, 781, 811, 931, 251, 1159, 556, 870, 1306, 1979, 482, 324]. procedure [698, 699, 1505, 2281, 2134, 335, 1399]. procedures [1642, 291]. process [1678, 904, 1619, 1050]. processes [1279, 1849, 902, 345, 1263, 534, 2168, 1456, 1137, 2225, 1880, 675, 1016, 1891]. produce [2217]. produced [1162, 1625]. Product [680, 710, 1483, 272, 1931, 1439, 850, 607, 531, 1266, 1190]. products [2232, 1011, 100, 60, 1186, 1426, 285, 1636, 1637, 197, 1648]. profile [1138]. projected [1853]. Projection [1699, 1926, 544, 906, 784]. Projective [731, 1702, 1054, 1269, 1237, 1523, 2083, 1797]. projector [1874].

Prolongation [1576]. Proof [1872, 645, 208, 1251, 652, 1915, 1069]. propagating [2014]. propagation [1484, 1376, 1504, 1404, 1997, 1617, 995]. propagator [1272, 1165, 1052, 668, 1326, 1733, 1093, 248, 1609, 111]. Propagators [141, 1729]. Proper [703, 1660]. Properties [719, 1471, 519, 881, 2171, 653, 616, 316, 546, 835, 1202, 1466, 2296, 1792, 1793, 1272, 431, 1633, 1644, 1420, 444, 646, 2133, 44, 883, 577, 1978, 416, 1820, 774, 586, 1679]. property [168, 1741, 2080, 1227, 652, 995]. proposal [2315]. proposed [71]. proton [1757]. proven [548]. pseudo [1514, 679]. pseudo-octonion [679]. pseudo-Poisson [1514]. pseudoclassical [1667]. pseudoconvexity [1172]. pseudodifferential [1907, 1682, 2112]. pseudodipole [721]. pseudopotential [1168]. pseudospheres [1985]. pseudospherical [209, 985]. pseudosymmetric [2242]. Pulse [1261]. punctures [1382, 672].

Qadir [2316]. QCD [1278, 1409, 1408]. QCD-parton [1278]. QED [1014]. Quadratic [770, 19, 2066, 129, 1165, 1138, 325, 686, 1379, 1264, 1762, 1616, 703, 13, 153, 1377, 2178]. quadratic-Hamiltonian [129]. Quadrics [1958, 1281, 1788]. quadrupling [838]. quadrupole [190]. Qualitative [1124, 2135]. quantities [29, 1953, 2310, 512, 214, 712]. Quantization [2152, 2114, 33, 1405, 1684, 941, 2084, 390, 1265, 2267, 1292, 2279, 43, 1293, 135, 224, 396, 20, 515, 723, 797, 1983, 697, 319, 424, 320, 2036, 1011, 633, 1600, 389, 1038, 257, 656, 776, 459, 1238, 1073, 1590, 487, 824, 262, 1976, 1273, 1945, 2277, 2315, 2134, 545, 2043, 343, 657, 132, 28, 900, 377, 1572, 207, 1475, 1072]. Quantized [2025, 1090, 758, 1208, 1378, 1789, 1482, 1832]. Quantum [901, 504, 1242, 2020, 417, 418, 536, 1643, 1599, 475, 1833, 828, 1394, 1559, 2302, 2082, 1794, 2304, 1991, 2214, 847, 1760, 1474, 561, 562, 827, 976, 2028, 1491, 1944, 946, 387, 1947, 1204, 2227, 2314, 1521, 2263, 1610, 1765, 1189, 1803, 1872, 1546, 428, 2249, 1525, 785, 739, 2187, 1279, 51, 2093, 135, 698, 699, 1294, 944, 2213, 29, 348, 398, 376, 1904, 2155, 1791, 929, 270, 1824, 1793, 386, 370, 602, 603, 12, 1305, 1467, 1942, 886, 1274, 1327, 1715, 235, 782, 1598, 125, 2201, 642, 1905, 1719, 1766, 1498, 1298, 1986, 1528]. quantum [421, 1973, 1403, 1720, 2115, 215, 1987, 1529, 1988, 1118, 974, 263, 1989, 749, 1205, 767, 1318, 1990, 1076, 403, 534, 2272, 2117, 1589, 2024, 1119, 1304, 1051, 1530, 1602, 1723, 2037, 2038, 1073, 746, 731, 1054, 1269, 1237, 980, 1523, 2083, 2159, 1724, 2305, 2160, 1499, 573, 2211, 1141, 1070, 261, 629, 939, 1561, 487, 1425, 1481, 461, 1836, 176, 2027, 823, 1232, 1437, 1995, 2120, 1878, 1933, 879, 691, 1786, 437, 95, 1761, 1299, 2029, 2193, 1179, 1384, 859, 2067, 2162, 221, 1010, 1296, 2164, 1319, 1532, 1605, 1801, 1709, 1710, 1977, 1406, 1890, 1606, 1164, 619]. quantum [1831, 975, 162, 741, 2123, 574, 1712, 2124, 2198, 27, 249, 526, 783, 1287, 1647, 1494, 1730, 2212, 803, 1479, 548, 1344, 112, 1050, 947, 1911, 2151, 1351, 1475, 1817, 1611, 806, 889, 1354, 1144, 1496, 1593, 1928, 349]. quantum-billiard [1995, 1878]. quantum-mechanical [1719, 1118]. quantumelectrodynamics [594]. quark [1301]. quarkonium [867]. quarter [496]. quartet [188]. quartic [1401, 102, 1443, 1765]. Quasi [1557, 1780, 1447, 870, 1248]. Quasi-exactly [1780]. Quasi-exactly-solvable [1557]. quasi-Goursat [870]. Quasi-invariant [1447]. quasi-one-dimensional [1248]. quasicaustics [150]. quasiclassical [207]. quasicrystals [640]. Quasiexactly [2063]. quasifractional [1198, 651]. quasifree [1038, 1724]. Quasiperiodic [1331, 1325, 1655, 1356, 1652, 81]. Quasiprobability [746]. quasitriangular [1494]. quaternion [1266]. Quaternionic [775, 615, 1644, 2037, 2038, 1367, 1636, 614, 154]. quaternions [946, 636]. quintic [557]. quommutators [1720].

R [57, 2318, 226]. Racah [1585, 805, 1101, 2196, 237, 786, 233, 1868, 1190]. radar [101]. Radial [86, 1093]. radially [842, 1324]. Radiating [2292, 2047, 1477, 1086]. Radiation [1176, 758, 665, 751, 1137, 2225, 1627]. Radiative [151, 46]. radius [867]. Raising [1519]. Raman [761]. Ramond [1937]. Random [1856, 2110, 1838, 22, 300, 2188, 1952, 2023, 772, 1956, 1570, 1555, 26, 1248, 1729, 1891]. randomly [915, 1785]. range [2219, 1566, 801, 50, 1168, 247, 1879, 1889]. rank [1105, 306, 1980]. rapid [652]. Rarita [1607]. rate [1505]. rate-type [1505]. Rational [2235, 916, 1138, 1896, 1899, 620, 1021]. ray [1614]. Raychaudhuri [508]. rays [742]. reactions [1777, 1963]. reactor [569]. Real [1100, 1163, 678, 679, 1121, 503, 126, 931, 986, 1281, 1831, 2072, 1936, 1970, 2113, 730, 1942]. Real-energy [1942]. realistic [475]. Realization [1095, 1106, 1218, 2074, 1515, 94, 58, 524, 1353, 62, 551, 974, 1205, 1704, 258, 1255, 1319, 728, 783, 1344, 2150]. realizations [886, 1274, 1715, 642, 1701, 1906, 2021, 405, 2258, 21, 52, 98, 407, 982, 1928]. Reasons [632]. reciprocal [1504, 1539]. Reciprocity [385]. Reconstruction [2108, 1538, 1808, 2003]. recovery [1078]. rectangular [1616]. rectilinear [2090]. recurrence [1632]. Recursion [1892, 118, 1409, 1408, 1730, 1502, 687, 684]. Recursive [715]. Redefinition [457]. redshift [288]. Reduced [1037, 1330, 2151, 481, 324]. reducibility [448, 1832]. Reducible [2021, 54, 2036, 1600, 223]. reducing [1214, 1244]. Reduction [1841, 28, 1822, 457, 918, 265, 731, 481, 1068, 1029, 72, 2134, 212, 663, 1888, 2286]. Reductions [1634, 2186, 1847, 1424, 1323, 1400, 1509, 1542, 1847]. Redundancy [2159]. reexamined [619]. reference [1478, 5, 2264]. Refined [249]. Reflection [1258, 781, 811, 931, 2258, 891, 1425, 1896, 1899, 2206, 1407]. regime [398]. region [133]. regions [280, 2090]. Regular [534, 324, 1983]. Regularization [952, 1678, 1717, 45, 1912]. regularizations [1968]. Regularized [1592, 838, 10, 123, 898]. Regulated [1272]. relate [654]. related [1641, 2062, 1957, 2075, 240, 1618, 893, 686, 2083, 1141, 1481, 400, 2092, 1893, 1895, 2288, 409, 954, 410, 1322]. relation [1470, 1940, 2253, 1153, 2160, 997, 118, 1730]. Relations [588, 241, 532, 1292, 1547, 204, 677, 2144, 983, 2256, 2257, 1989, 1321, 435, 385, 2043, 2072, 1409, 1408, 799, 1103, 948]. Relationship [236]. Relative [863, 1595, 171]. Relativistic [1758, 1876, 1028, 358, 299, 922, 803, 2047, 1875, 1620, 1621, 1556, 116, 1792, 54, 141, 210, 2229, 1613, 347, 327, 1295, 366, 1446, 2280, 1114, 25, 727, 261, 2039, 208, 1251, 765, 879, 1842, 1179, 510, 2276, 1946, 98, 896, 1800, 1774, 288, 2041, 1039, 356, 282, 1611]. relativistically [2230]. relativity [354, 355, 289, 506, 2137, 1390, 831, 832, 465, 702, 2244, 1743, 1903, 704, 1086, 2099, 1349, 1852, 507, 1581, 1418, 171, 663, 1936, 2113, 1282, 961, 695]. relaxation [299]. relevant [1810]. Reliability [355]. remainder [1207]. Remarks [797, 238, 173, 696, 2185, 31, 1818, 2262, 274]. removal [921]. Renormalization [919, 881, 2250, 1045, 341, 1014, 1453, 1222, 1891]. Renormalized [140, 1193, 906, 1443]. reparametrization [713, 2277]. reparametrization-invariant [2277]. Representation [2252, 2070, 1145, 1640, 1292, 302, 863, 297, 1983, 680, 1737, 642, 1337, 69, 477, 1074, 1721, 2082, 756, 331, 980, 1114, 1171, 827, 626, 437, 502, 932, 200, 491, 492, 1406, 526]. Representations [231, 272, 1824, 1486, 2106, 1205, 441, 927, 1141, 991, 95, 1831, 2043, 1712, 948, 1470, 2185, 2140, 1873, 2062, 1698, 234, 1184, 529, 1079, 480, 2017, 1100, 593, 2142, 579, 2104, 94, 605, 966, 890, 1250, 62, 551, 974, 2258, 1394, 126, 622, 406, 1702, 156, 157, 158, 560, 1076, 1075, 403, 527, 1589, 1304, 60, 115, 1244, 1186, 1784, 1975, 1760, 998, 2027, 21, 1706, 1426, 1455, 1196, 1106, 850, 607, 1669, 1830, 2195, 837, 8, 598, 489, 678, 679, 148, 99, 161, 978, 728, 1684, 1064, 2176, 807, 2069, 2072, 275, 942, 1385, 1191, 1647, 1220, 617]. representations [1190, 809, 1911, 277, 806, 1354, 1496]. representing [1581]. Reshetikhin [2204]. Resolution [1881, 1329]. resolvent [659]. Resolving [1676]. Resonance [2127, 463, 247, 1235]. resonances [2035, 1882, 1617]. resonant [897, 721]. resonator [2088, 1338]. resonators [1599]. respect [134, 1312]. respects [713]. Response [2321, 858, 552, 630]. responsible [741]. rest [1865]. restoration [757]. restricted [1740]. restriction [681]. results [455, 589, 965, 2300, 2023, 1551, 18, 11, 166]. reversible [1456]. revisited [305, 635, 1585, 2015, 2016, 1913, 498]. Rham [1169, 2032]. Riccati [1573]. Ricci [2317, 2316, 2055, 1657, 2097, 567, 37, 256, 255]. Ricci-flat [567, 255]. Ricci-type [2097]. Riemann [1098, 1594, 1382, 672, 116, 871, 394, 761, 1986, 147, 924, 185, 369, 2097, 853, 1356, 626, 933, 89, 1012, 1042, 298, 768, 324, 183]. Riemannian [742, 1289, 956, 1257, 582, 472, 1958, 1281, 222, 636, 567, 1707, 741, 1389, 695]. Right [558]. Right-invariant [558]. rigid [253, 1282, 961]. Rigidity [1367]. Rigorous [963, 1798]. ring [1554]. Rings [1616, 1201]. Robertson [2179, 2241, 539, 641, 48, 1049, 182, 37, 803, 1479, 1627, 117]. Robinson [650, 1741, 92]. Robustness [1791]. rod [1352]. role [2096]. root [2142, 972, 981, 1682, 2112, 2307, 269, 1188]. Roots [1712, 557]. Rosen [998, 2164]. rotating [2318, 1477, 2059, 1420, 2230, 1773, 5, 2264, 1282, 961]. rotation [321, 1371, 1515, 1585, 1906, 1171, 2193, 1102, 1101, 2196, 1360]. Rotational [1295, 795, 627, 190, 1023]. Rotations [211, 2028, 1593]. rototranslational [1182]. Rouet [2036, 1929]. Rough [341]. Ruelle [815]. rule [2196, 1343]. rules [1483, 1105, 406, 303, 2315, 1752, 786]. rung [675]. running [837].

S [57, 423, 1755]. Sachs [1819, 1081, 1082]. saddle [191]. saddle-point [191]. Salpeter [592]. sample [291]. sandpile [1769, 1690]. sawtooth [964]. Scalar [1262, 836, 272, 1242, 1048, 1123, 957, 668, 1085, 100, 2117, 1544, 1870, 1910, 1819, 1309, 850, 607, 734, 221, 1916, 912, 1689, 791, 1266, 1418, 2073, 676, 1087, 1081, 1082, 2085]. Scale [666, 1358]. Scale-invariant [666]. Scaling [1723, 480, 801, 588, 1652, 875]. Scattering [826, 1120, 1524, 121, 26, 318, 530, 2169, 2219, 542, 317, 1345, 1078, 2266, 1884, 1812, 334, 904, 1735, 1849, 12, 1305, 1467, 1942, 1387, 311, 897, 902, 1207, 580, 579, 578, 1938, 252, 761, 868, 923, 855, 999, 2037, 2038, 624, 1340, 49, 1235, 1836, 176, 827, 1726, 1944, 419, 179, 2167, 1562, 518, 870, 2126, 384, 385, 1306, 2035, 938, 1177, 1883, 888, 1187, 1610, 2228, 2045, 1592, 2038]. scheme [964, 294]. schemes [30]. Schifini [226]. Schild [2015, 2016]. Schlesinger [1150]. Schrieffer [1473, 1686, 1871]. Schrödinger [949, 317, 826, 1345, 1594, 2266, 787, 1640, 1641, 1873, 2130, 422, 440, 2220, 386, 1717, 503, 462, 1574, 1346, 345, 134, 2022, 69, 1472, 2080, 576, 1383, 380, 1645, 939, 1481, 400, 1181, 1916, 2276, 362, 184, 577, 1442, 1166, 1685, 1882, 1059, 1410, 1325, 1950, 1234]. Schrödinger-Fuerth [939]. Schrödinger-type [1645]. Schulman [423]. Schur [1700, 197, 2000]. Schwarz [1937]. Schwarzian [848]. Schwarzschild [2057]. Schwinger [2269, 1597, 2036, 633, 2108, 156, 157, 158, 1722, 1238, 908, 1398, 1267, 1303, 1607, 537, 849, 516]. Schwinger-type [1267]. screened [500, 1882]. SDIFF [1901]. Search [1544, 1842, 375, 1051]. Searchlight [1465]. Second [1951, 325, 1524, 653, 1965, 1038, 734, 1405, 1607, 132, 1225, 1072]. Second-order [325, 1524, 653, 734, 1225, 1072]. Sectional [131, 1080, 1818]. sector [1995]. sectors [2190, 2308, 757, 442]. seed [958]. seen [1295]. Segur [1654, 1848]. Selberg [1156]. selective [189]. Self [874, 1900, 1240, 490, 1321, 739, 2205, 653, 1413, 804, 1449, 1901, 329, 649, 844, 1925, 38, 2048, 1307, 1644, 1796, 2283, 1171, 1634, 1367, 387, 581, 245, 470, 186, 1542, 799, 427, 611, 1363, 530]. self-adjoint [530]. Self-adjointness [1321, 739]. Self-dual [1240, 490, 1413, 1449, 1901, 2048, 1171, 1634, 581, 245, 186, 1542, 427]. Self-duality [874, 799]. self-energy [1796]. Self-gravitating [1900, 38, 1307]. self-interacting [1925]. self-interaction [387]. self-linking [2205]. self-similar [653, 2283, 470, 611, 1363]. self-similarity [804]. semi [1914]. Semiclassical [824, 425, 1563, 907, 1950, 1641, 1346, 1528, 1760, 2211, 1070, 1998]. semiconductor [1071]. semidirect [564]. semigroup [846]. semigroups [348, 184, 1320]. semiintegrable [87]. semiline [453]. semisimple [1463]. semitransparent [1115]. seniority [72]. Separability [1575, 1373]. Separated [781]. separates [627, 592]. Separation [795, 1281, 2236, 960, 1025, 1155, 117, 395, 2080, 259, 1958, 878]. sequence [1443]. sequences [964]. Series [959, 1071, 1971, 680, 1498, 1487, 45, 1020, 265, 971, 1106, 409, 1591, 1980, 815, 1981, 1443, 2291]. set [1331, 1049]. sets [2137, 539, 641, 483]. Setting [2192, 2287, 49]. seven [1708, 1241]. seventh [1260]. seventh-order [1260]. several [1021, 1243]. sextic [16, 521, 1443, 1233, 1386]. Shabat [1622, 1896, 1899]. shallow [1400]. shape [342, 2090]. shaped [952, 861, 720]. Shear [354, 650, 1964, 109]. Shear-free [354, 650, 1964]. sheared [1063]. sheet [853]. sheeted [1495, 2033]. shell [217, 838, 1167, 1738, 1301, 1113]. shells [2320, 1041, 2321, 1821]. Shift [621, 80, 1846]. Shifted [197, 1206]. showing [2128]. sigma [718, 1307, 1536, 714, 1057]. signature [1584]. similar [653, 1839, 2283, 470, 611, 1363]. Similarity [196, 663, 2095, 804, 683, 1195]. similitude [164]. Simmons [2125]. Simons [549, 24, 1900, 2203, 752, 473, 1879, 330, 397, 1888, 1083]. Simple [986, 335, 1399, 2074, 340, 1359, 1515, 1632, 1105, 1355, 299, 1243, 1301, 1881, 1776, 137, 162, 1191, 1069, 128, 1679, 914, 1258]. simplest [1494]. simplicial [1236]. simplicity [1002]. simply [802]. Simultaneous [1538, 2003]. Sinaı [22]. sine [1230, 1624, 1655, 1739, 1656]. sine-Hilbert [1230]. Singer [1915, 1916]. Singh [57, 57]. single [841, 170, 1876, 1114, 280, 175]. single-connected [280]. single-electron [1876]. single-particle [1114]. single-step [841]. singletons [514]. Singular [645, 32, 1470, 1641, 1471, 2268, 2285, 916, 781, 811, 931, 1267, 1996, 499, 1728, 9, 1672]. Singularities [1215, 1697, 1004, 1717, 1057, 1227, 931, 751, 1239, 1002, 360, 1175, 1881, 1676]. Singularity [1009, 2175, 444, 1744, 1358, 1659, 1691]. Singularly [330, 167]. sintered [104]. sinusoidally [279]. site [22, 841]. Sitter [1140, 587, 2114, 2136, 829, 146, 52]. six [1966, 643, 7, 56, 2070, 202]. six- [643]. six-dimensional [2070]. size [801]. sKdV [1228, 312]. skewed [2283]. Skyrme [719, 1057, 1215, 1332, 714]. Skyrme-like [714]. slab [923, 1030, 1376, 1732, 482]. Slater [1399]. Slater-type [1399]. slices [1003]. slit [1338]. slowly [1403]. Small [1615, 321, 1566, 891]. smooth [1776]. smoothness [2216, 134, 753]. Smorodinsky [903]. Snell [1188]. solid [1371, 104, 2127]. solids [554, 1787, 483]. solitary [2012, 860]. Soliton [1620, 1621, 1059, 2178, 719, 1157, 834, 813, 771, 1335, 688, 780, 634, 958, 1695, 2051, 1772, 91, 169, 1177, 1577, 1377, 1692, 712]. solitonic [77]. Solitons [735, 383, 1900, 2132, 1230, 1799, 1896, 1899]. soluble [2037, 2038]. Solution [346, 280, 2163, 964, 1182, 1432, 1278, 2057, 133, 2078, 1687, 1543, 327, 1118, 1347, 38, 958, 2016, 290, 1615, 1743, 1352, 1909, 652, 1086, 41, 84, 368, 357, 1785, 937, 674, 1774, 2235, 533, 1571, 1866, 1713, 1511, 2073, 186, 1935, 778, 760, 223, 117, 395, 250, 1334, 1686, 1871]. Solutions [718, 172, 782, 432, 1626, 756, 1577, 397, 1392, 2240, 501, 1209, 1640, 1288, 1736, 1620, 1621, 1624, 1262, 2171, 2101, 1246, 1413, 54, 1163, 1314, 329, 688, 412, 2269, 916, 2014, 2018, 1847, 1781, 1316, 1212, 420, 1687, 413, 1649, 2252, 634, 361, 1489, 1215, 1138, 1558, 683, 2105, 958, 1173, 1741, 622, 1195, 1200, 2303, 205, 655, 660, 1655, 1229, 1131, 1130, 654, 1448, 627, 959, 469, 624, 895, 616, 48, 1049, 1910, 1425, 585, 1455, 1332, 1368, 1349, 1695, 2161, 1934, 1226, 1015, 1311, 1927, 145, 456, 883, 1181, 279, 2147, 932, 1012, 1581, 2235, 142]. solutions [1696, 1306, 169, 1787, 2176, 1896, 1899, 1685, 1393, 364, 1369, 1764, 2177, 907, 994, 470, 1059, 1362, 1377, 2178, 1282, 961, 471, 1325, 1507, 2248]. Solvable [984, 1805, 1885, 202, 1557, 1780, 2063, 601, 1613, 1834, 894, 426, 430]. Solving [1530, 2281]. Some [2185, 2101, 155, 31, 1413, 125, 1510, 589, 263, 1589, 1818, 1742, 1415, 1644, 1448, 1420, 823, 2262, 18, 11, 166, 555, 1012, 274, 387, 1820, 1422, 1191, 737, 773, 1377, 949, 1145, 1286, 1162, 2220, 1737, 902, 1631, 201, 643, 915, 777, 469, 487, 445, 1134, 1726, 870, 200, 978, 237, 712, 1516]. Soul [1638]. source [476, 1538, 540, 923, 1819, 1081, 1082]. Sources [1417, 2294]. Space [2209, 2154, 2095, 107, 25, 727, 1870, 1309, 1582, 501, 428, 587, 742, 1178, 650, 1289, 2187, 1208, 2013, 1432, 1731, 1262, 2292, 1662, 4, 6, 219, 1000, 467, 1629, 1412, 1983, 2172, 1522, 1268, 354, 871, 804, 593, 1183, 235, 1122, 1584, 292, 1687, 2156, 2114, 2242, 323, 1503, 2157, 1257, 1489, 1925, 1650, 79, 477, 828, 1988, 1170, 2089, 1967, 2139, 1001, 441, 181, 2259, 752, 146, 2182, 1841, 1644, 92, 1244, 777, 627, 539, 959, 1239, 1724, 1474, 1032, 1310, 1580, 2125, 1049, 1634, 1478, 251, 1725, 2309, 2092, 552, 136, 976, 1913, 816]. space [2193, 1669, 222, 1174, 859, 2224, 182, 37, 1728, 392, 415, 873, 1890, 7, 56, 1175, 1881, 1046, 1241, 164, 1442, 910, 137, 741, 1389, 171, 154, 1660, 942, 1042, 995, 526, 464, 496, 1676, 676, 1339, 2060, 1391, 256, 1936, 2113, 223, 117, 955, 1821, 1087, 436, 1069, 695, 1803, 183, 1859, 1296]. space-splitting [1728]. spaced [811]. spacelike [1276, 470]. spaces [519, 2100, 1035, 480, 116, 2141, 12, 1305, 1467, 1942, 2020, 788, 1257, 2145, 829, 1140, 1169, 472, 1958, 2305, 2210, 48, 1779, 1281, 544, 244, 1977, 581, 2198, 1197, 1542, 111, 112, 530]. spacetimes [2055, 2317, 2316]. spacings [2116]. Spatial [2061, 2153, 1522, 1268]. Spatially [1411, 1211, 1242, 1687, 75]. Special [2249, 1024, 2113, 2057, 2087, 107, 1186, 1590, 439, 411, 1673, 703, 695, 1709, 1936]. spectra [2062, 2081, 962]. Spectral [1792, 2141, 1793, 1614, 1679, 1622, 383, 462, 782, 1574, 1337, 1472, 1601, 1644, 998, 401, 2092, 44, 2149]. Spectrum [1611, 949, 1436, 2262, 1798, 759]. speculation [423]. speed [1078, 482]. speeds [1212]. sphaleron [1557]. sphere [2292, 340, 672, 644, 1396, 1141, 1635, 1831, 298, 762]. spheres [834, 1985, 2190, 2308, 1427, 42, 384]. Spherical [2191, 801, 217, 1781, 1586, 2079, 2318, 290, 1415, 1450, 531, 1147, 26, 1867, 2005]. Spherically [2188, 114, 1086, 144, 765, 791, 470, 2055, 2317, 2316, 4, 6, 219, 1183, 1510, 2282, 2319, 1167, 661, 357, 358, 7, 56, 1175]. spheroidal [1626, 259, 1368]. spheroids [1390]. spiked [30, 747, 1434, 1755]. Spin [1453, 572, 1677, 784, 388, 519, 699, 1161, 505, 1413, 2251, 2128, 1009, 320, 1263, 1664, 508, 1044, 2082, 2271, 669, 2191, 259, 539, 2223, 1845, 362, 842, 1129, 1303, 623, 197, 1534, 1042, 82, 1609, 548, 802, 232, 1886, 1675]. spin- [1263, 362, 1303, 1534, 1886]. spin-boson [388]. spin-one [320]. spin-orbit [699, 623]. Spin-plane [572]. spin-statistics [548]. spin-wave [259].

Spin-weighted [1677]. spinflip [1619]. spinless [210]. spinning [1390, 2084]. Spinor [84, 590, 1669, 1179, 1262, 934, 1757, 523, 1589, 1277, 1910, 180, 1935]. spinorial [1152]. Spinors [1366, 2182, 986, 1729]. splitting [1728, 1506]. spontaneous [1096]. spontaneously [1246]. Square [1304, 2140, 710, 376, 1682, 2112, 1243, 1500]. square-cell [1243]. square-integrable [2140]. Square-well [1304, 376]. squared [353, 538]. squeezed [1874, 1346, 53, 1795, 1213, 553]. Squeezing [336, 211, 2001]. Stability [311, 1880, 1291, 1363, 55, 2233, 1557, 1792, 547, 569, 1844, 1031, 307, 631, 496, 427, 2012]. Stable [495, 2011, 2284, 987]. Stably [884]. stacking [1956]. standard [698]. Star [390, 1011, 2104, 230, 371, 1428, 1648, 356, 277]. Star-products [1011]. Stark [2219, 315, 400]. State [44, 1678, 1874, 863, 1573, 1984, 1941, 1597, 235, 254, 1346, 604, 869, 2015, 404, 997, 1439, 423, 279, 184, 1890, 1696, 177, 807, 2069, 1534, 906, 762, 1917, 1443, 681]. states [2266, 522, 546, 940, 1342, 2100, 229, 905, 1274, 1715, 1690, 867, 378, 53, 970, 604, 1825, 1403, 315, 1721, 1795, 943, 122, 528, 1724, 1063, 1326, 1561, 748, 1302, 696, 2226, 175, 1532, 608, 1998, 1500, 553, 1954, 178, 599, 1609, 1410, 2001, 1495, 2033, 784, 267, 1950, 1234]. Static [7, 56, 2055, 2317, 2316, 4, 6, 219, 476, 2098, 1049, 1478, 144, 42, 1004]. static-source [476]. Stationarity [359]. Stationary [1173, 540, 1125, 1910, 1619, 1927, 873, 1050]. Statistical [2123, 1051, 629, 2007, 1299, 1811, 942, 494]. statistics [1044, 548, 1891, 1475, 474]. Steady [894, 57, 1152, 1690, 279, 2]. steady-state [279]. step [841, 318]. step-periodic [318]. Stephani [1510, 1308]. Stephenson [664]. Sternberg [212]. Stewartson [1847, 1092, 1252, 1789, 1482]. Stieltjes [2301]. stiff [1543, 1694]. stimulated [761]. Stochastic [106, 43, 51, 210, 2273, 656, 1073, 1117, 1456, 184, 1891, 2129, 421, 1341, 749, 167, 2168, 140, 291, 91, 1089, 825]. Stokes [1091, 1018, 1225, 1315, 1767]. Stora [2036, 1929]. straight [994]. straight-line [994]. strange [890, 1250, 1353, 717]. strategy [12, 1467]. stratified [1030, 1376, 1732, 1420]. Stratonovich [606, 825]. strength [1239, 1441, 493]. Strengths [751]. stress [829, 1140, 1968, 1193]. stress-energy [829, 1140]. String [1183, 967, 188, 1910, 1836, 90, 597, 1727, 2226, 479, 478]. string-flip [1836]. strings [1984, 2056, 838, 1053, 1133]. stringy [2209]. strip [1883]. Strong [1002, 2153, 2233, 1347, 732, 1385]. strong/weak [2233]. strongly [1471, 1478]. Structural [1868, 1190, 2150]. Structure [672, 1315, 428, 1653, 1654, 742, 1514, 348, 1757, 1547, 1690, 1924, 417, 418, 1009, 77, 1929, 1488, 792, 2270, 2108, 382, 1807, 465, 2244, 853, 1860, 1681, 733, 591, 2223, 2125, 2210, 444, 88, 425, 1913, 1976, 1102, 2224, 221, 1010, 1532, 1248, 1962, 2134, 732, 1576, 1691, 1999, 338, 2239, 676, 1275, 1740, 712]. structures [2186, 2295, 1370, 1289, 1228, 698, 699, 770, 649, 1965, 2218, 1199, 893, 68, 821, 185, 2272, 1827, 1054, 1269, 2083, 1462, 2168, 1797, 986, 1252, 291, 1012, 2175, 614, 1754, 324]. Study [1720, 469, 2044, 682, 2220, 904]. Sub [1973]. Sub-Hopf-algebra-induced [1973]. subalgebra [1666]. subalgebras [1105, 443]. subclass [1616]. subduction [596]. subgraph [246]. subgroup [306]. subgroups [605, 443, 1001]. subject [1405]. submanifolds [1760, 1676, 1445]. subspaces [1244, 1311, 1300, 1301, 1302]. subsystems [949]. successively [335]. Sufficient [582, 1724, 839, 446]. Sugawara [600]. Sugawara-type [600]. Sum [1752, 786, 564, 2196, 964, 1343]. summability [713, 1711]. summation [1156, 1443]. Summational [253]. summations [1982, 2313, 190]. Summing [998]. sums [1513, 2250]. Super [1116, 1586, 728, 901, 1623, 124, 1449, 1901, 1924, 1515, 1585, 1906, 147, 1826, 331, 1396, 1462, 88, 89, 1102, 1101, 2196, 2010, 435, 1055, 1980, 848, 298, 1275, 1693, 1923]. super- [1396, 1826]. Super-angular [1586]. super-Beltrami [1462]. super-Lie [1924]. super-rotation [1515, 1585, 1906, 1102, 1101, 2196]. super-spherical [1586]. Super-Toda [1923]. super-Virasoro [1980]. super-Wess [1901]. Superadiabatic [1436]. superalgebra [1701, 2104, 890, 1250, 1590, 945, 2027, 2067, 8, 598, 148, 1751, 1191]. superalgebras [529, 1327, 1548, 2021, 1095, 2022, 1353, 551, 2024, 365, 2026, 979, 717, 525, 442, 1673, 1218, 269, 600, 333, 1937]. superanalysis [1548, 1549, 1924, 1284]. supercoherent [905]. superconducting [193]. superconductivity [554]. superconductors [1997]. Superconformal [120, 2026, 1007, 298, 1483, 559, 846, 885, 766, 1006, 1462, 1579]. Superconformally [1826]. Supercritical [1844]. superdeterminants [1915]. superdiffeomorphism [1901]. superdiffeomorphisms [1461]. superdifferential [2026]. superequations [367]. superfiber [24]. Superfield [21, 1314]. superfluity [36]. superfunctions [1638]. supergenerator [676]. supergravity [1901, 1006, 473, 1512]. supergroup [901, 124, 527, 1144, 1496, 1185]. Supergroups [1704, 125, 970, 265, 971, 1396, 1933, 806, 889, 1354]. Supermanifolds [1549]. supermatrices [399]. Supermultiplets [1902]. superoperators [1102]. superparticle [262, 89]. Superposition [367, 2174]. superpotentials [1986]. superprojective [818]. supersigma [1185]. superspace [1782, 1825, 264, 2026, 1108, 1109, 716, 2110, 1397, 2278, 1329]. superspaces [1746, 1797]. superstring [1431, 1430, 2158, 2223, 1273]. Supersymmetric [1990, 1166, 2200, 785, 865, 1598, 424, 1498, 1048, 1986, 1602, 87, 571, 481, 879, 1532, 1541, 947, 920, 1072]. Supersymmetries [1782, 944]. supersymmetrization [698, 699, 1642]. Supersymmetry [1915, 1916, 501, 1653, 51, 229, 1136, 972, 52, 332, 1189]. supersymplectic [1825]. supertrace [1152]. supertransformation [1531]. supertransparency [2051]. supertriple [2040]. superversion [1038]. support [1896, 1899]. supported [278]. Surface [1612, 838, 578, 1938, 2120, 89, 26, 1810, 1115, 985]. surfaces [1382, 2172, 780, 341, 1986, 147, 185, 369, 1908, 266, 415, 209, 151, 768]. susceptibility [113, 1128]. Sutherland [2000]. Symbols [2155, 643, 2196, 1343]. Symes [1541]. Symmetric [788, 2055, 2317, 2188, 1035, 2316, 4, 6, 219, 1000, 1746, 1527, 1183, 114, 1510, 2282, 2319, 665, 1543, 1167, 1328, 661, 1169, 391, 1030, 404, 1474, 1903, 1063, 1086, 1870, 144, 1309, 357, 358, 765, 42, 791, 1545, 842, 7, 56, 1175, 1324, 1197, 470, 583, 111]. symmetrical [320]. Symmetries [4, 6, 219, 323, 347, 1375, 2058, 1795, 1222, 1848, 981, 658, 1668, 2009, 1166, 276, 712, 242, 1745, 2284, 2187, 1260, 440, 697, 50, 1476, 1778, 609, 263, 1419, 2008, 181, 1080, 1615, 2192, 512, 450, 1373, 119, 2004, 214, 950, 732, 1881, 898, 1979, 304, 1660, 1324, 808, 684]. Symmetrized [339]. Symmetry [1265, 2296, 757, 683, 1254, 259, 481, 1961, 1404, 883, 662, 888, 1187, 2045, 1099, 1540, 1736, 1005, 224, 723, 505, 2180, 969, 329, 1847, 647, 2036, 3, 1317, 2253, 1257, 846, 1925, 389, 402, 1840, 1350, 486, 369, 366, 795, 627, 258, 2160, 733, 591, 1420, 1323, 704, 2231, 1960, 2133, 1400, 734, 2294, 2278, 2071, 1096, 1364, 2239, 1656, 395, 216, 1610, 2228, 282, 1536, 1, 2286]. symmetry-breaking [591]. Symplectic [2091, 235, 864, 1860, 1068, 1998, 2286, 1370, 495, 893, 15, 520, 158, 331, 1797, 2210, 439, 1750, 1637, 1712]. symplecton [1319]. Syms [1541]. synchronizable [1478]. Synge [416]. synthesis [364]. system [2152, 2179, 2241, 135, 1452, 1957, 1557, 1578, 1714, 54, 240, 569, 2173, 1574, 417, 418, 1847, 36, 1920, 903, 2107, 1044, 1453, 1458, 656, 855, 981, 2231, 173, 1604, 461, 87, 639, 208, 552, 1251, 592, 1261, 973, 490, 499, 1789, 1482, 1890, 1684, 842, 732, 1625, 2053, 1922, 2288, 1575, 1691, 9, 1672, 2238, 1104, 2012, 474, 349]. systematic [804, 1518]. systems [1099, 570, 106, 546, 835, 1157, 1641, 1293, 863, 396, 1573, 189, 2170, 2017, 1502, 1804, 839, 771, 129, 697, 273, 2285, 886, 1567, 462, 1316, 254, 141, 2218, 819, 53, 1476, 475, 102, 149, 1158, 1600, 170, 2300, 85, 1459, 935, 820, 167, 821, 924, 1098, 205, 655, 660, 2048, 1807, 2008, 257, 1142, 253, 1331, 1652, 1861, 746, 2049, 1603, 686, 1705, 33, 1815, 452, 2050, 1421, 824, 450, 1373, 1892, 1943, 1508, 481, 2092, 1960, 1893, 228, 1842, 950, 165, 921, 696, 1945, 1405, 1090, 2277, 1154, 1160, 1636, 1637, 2040, 1223, 1406, 1962, 2134]. systems [2052, 1895, 511, 1887, 169, 1777, 812, 307, 2289, 2035, 1333, 1541, 1617, 1534, 2290, 493, 1969, 1291, 269, 1206, 711, 112, 332, 954, 359, 1187, 638, 948, 410, 712, 128, 1072, 1445, 2286, 1770, 2054, 310]. Szego [447]. Szekeres [1551].

’t [713]. tableau [737, 773]. tableau-discrete [773]. tableaux [197]. Tabor [1894]. tadpoles [620]. Talmi [241, 532]. tangency [987]. tangent [511]. target [1465]. Taub [1967, 2098, 220, 2246]. tautochrone [1028]. Taylor [1071, 1591, 815, 1981]. technique [1629, 1699, 681, 957, 1030, 1376, 1389, 962]. techniques [754, 186]. teleparallel [1852]. Teller [998, 377]. temperature [753, 1128, 1564]. temperature-dependent [1128]. Temperley [806]. temporal [2214, 1994]. ten [1707]. Tensor [1546, 1483, 1516, 1186, 1426, 975, 1190, 1146, 2296, 1657, 1629, 871, 320, 582, 238, 1318, 1926, 830, 736, 525, 1102, 1266, 1820, 1660, 586, 1968, 1348]. tensors [1000, 829, 1140, 2097, 465, 404, 544, 928, 910, 676, 1193]. term [93, 397, 516, 326]. terms [1413, 1596, 1095, 633, 1722, 752, 908, 991, 1106, 937, 1213, 338, 1648, 1083, 714, 1067, 1886]. ternary [1423, 1637]. Test [436, 2154, 792, 1195, 1388, 1894, 907]. tetrahedron [563]. th [77, 2234, 285, 1439]. th-order [77]. their [2186, 2062, 1005, 2284, 1184, 383, 1900, 1163, 2166, 3, 1317, 2066, 551, 1200, 1001, 2244, 404, 528, 1850, 823, 991, 456, 2147, 1635, 72, 2197, 2235, 1608, 1422, 1647, 730, 1, 1234]. them [893]. theorem [2232, 1640, 739, 650, 1211, 1378, 1062, 1091, 1578, 1247, 129, 1567, 1371, 455, 204, 677, 2243, 1912, 1435, 1307, 2274, 2260, 1724, 2245, 47, 1232, 1437, 208, 1251, 1469, 1748, 1749, 447, 1915, 1916, 1606, 2035, 1533, 1541, 548, 326, 22]. theorems [2169, 1781, 1975, 445, 1726, 220, 1259, 385]. theoretic [1274, 1715, 108, 2073]. theoretical [1452, 2008, 64]. theories [967, 796, 121, 671, 224, 20, 515, 723, 797, 2154, 1246, 1449, 1314, 2036, 2096, 864, 2207, 1925, 633, 1953, 2310, 754, 1210, 185, 909, 908, 876, 473, 755, 1850, 1379, 76, 544, 837, 86, 2010, 1799, 343, 1999, 977, 942, 494, 1888, 1339, 993, 1189, 920, 1886, 1955]. Theory [2103, 835, 2061, 739, 188, 1431, 2179, 2241, 51, 2091, 2094, 2093, 135, 398, 189, 376, 1898, 1091, 764, 1527, 284, 504, 836, 1045, 12, 1305, 1467, 1942, 1597, 1245, 1372, 2201, 355, 578, 1938, 252, 314, 2299, 2252, 1201, 67, 1614, 1374, 903, 1665, 828, 1988, 85, 840, 1560, 576, 1453, 1569, 1844, 2118, 393, 1658, 1601, 1285, 1051, 1876, 2119, 999, 746, 1680, 1991, 2214, 17, 731, 1054, 1269, 1237, 980, 1523, 2083, 1681, 2305, 2223, 645, 1517, 939, 2287, 996, 49, 1851, 1092, 1744, 439, 358, 1349, 879, 2004, 691, 411, 664, 1217, 917]. theory [2121, 567, 2184, 2275, 221, 1010, 1029, 2276, 674, 330, 675, 628, 1280, 1139, 1659, 1441, 1696, 1115, 1338, 1887, 807, 1056, 1204, 2123, 1607, 1820, 703, 27, 1954, 479, 798, 16, 521, 328, 1013, 586, 494, 118, 715, 397, 1583, 1366, 408, 1050, 2200, 1765, 2165, 2127, 1803, 1300]. thermal [943]. thermo [966]. Thermodynamic [510, 547, 963, 1845]. thermodynamical [517]. Thermodynamics [2098, 1918, 547, 1618, 1176]. Theta [1020, 2259, 1356, 1019]. thick [391]. thin [1903, 883, 1821]. thin-film [883]. Third [1956, 1254, 2093]. Third-order [1956, 1254]. Thomas [1596, 840, 368, 2122, 1441, 760]. Three [370, 461, 445, 2167, 2219, 454, 956, 1897, 340, 672, 2172, 771, 2095, 603, 203, 53, 1061, 1472, 1840, 1807, 827, 450, 1373, 1726, 1255, 883, 1007, 937, 430, 1922, 2288, 2176, 171, 774, 1882, 1391, 1535, 46, 730, 1233, 1765, 790, 1686, 1871]. Three-body [461, 445, 2219, 827]. Three-dimensional [370, 2167, 956, 1897, 603, 203, 1472, 1255, 937, 1882, 1391, 1535, 46, 730, 1233, 1765]. three-forms [790]. three-level [53]. three-manifold [171]. three-parameter [2095]. three-particle [1922, 2288]. three-space [2172]. three-sphere [340]. three-wave [2176]. tidal [131, 1818]. Tight [278]. tiling [787, 720]. tilted [1926]. Time [1182, 576, 1695, 1887, 1506, 2129, 1288, 758, 2209, 742, 1875, 1178, 2187, 2013, 1262, 2292, 1000, 653, 1412, 1522, 1268, 804, 1183, 2086, 1122, 1584, 1346, 292, 1687, 2114, 323, 1476, 2157, 1925, 828, 1988, 1771, 1170, 38, 2270, 1967, 1227, 1001, 181, 923, 1376, 1732, 752, 1841, 777, 627, 1312, 959, 1239, 1815, 1032, 1310, 25, 727, 1725, 2309, 1149, 2092, 309, 816, 1842, 1456, 1296, 37, 1406, 873, 387, 1763, 1582, 1121, 1881, 511, 164, 1442, 1563, 1963, 137, 741, 2035, 464, 1838, 1391, 359, 955, 1087, 695, 1445, 1859]. Time-dependent [1182, 576, 1695, 1887, 1875, 2086, 1346, 1476, 1771, 38, 1815, 1149, 1842, 1296, 1763, 511, 1563, 1445]. Time-domain [1506, 1732]. time-harmonic [1376]. time-reversible [1456]. times [1208, 4, 6, 219, 467, 354, 2095, 107, 2242, 1257, 79, 2139, 2182, 92, 539, 1724, 1580, 1870, 1478, 1309, 1174, 182, 392, 7, 56, 1175, 1241, 910, 1660, 1042, 995, 2060, 256, 117, 436, 1069]. Tjon [774]. Toda [1540, 1245, 2218, 432, 512, 1323, 1774, 281, 2288, 1923]. Toda-related [2288]. Toda-type [2218]. Todorov [779]. top [2087]. Topological [1034, 1488, 1925, 2146, 1056, 2042, 2212, 1391, 640, 1091, 2201, 2207, 548, 1104, 2165, 1189, 714, 1727]. topological-field-theory [1091]. topologies [2299]. Topology [2215, 2209, 284, 1170, 1068, 2209, 2213, 2205, 2216, 2208, 2215, 2203, 2201, 2202, 2218, 2207, 2217, 2214, 2210, 2211, 2204, 2206, 2212, 2200]. tops [694]. tori [1978, 1674]. toroidal [1925]. torsion [1210, 139, 794, 1451, 1310, 1850, 2039, 1240, 896, 1800]. torus [296, 1502, 1272, 268, 1270, 2111]. Total [2137]. Totally [1580]. Toy [1084]. trace [2065, 1908, 1861, 425, 711]. traces [927]. trains [2177]. trajectories [2049, 1121]. trajectory [778]. transcendental [1223]. Transcendentally [891]. transfer [675, 151, 815, 46]. transform [2140, 1855, 1110, 2064, 1558, 673, 401, 2092, 18, 11, 166, 419, 1147, 1464, 2031, 1591, 1592]. Transformation [577, 476, 653, 293, 1747, 3, 1317, 1974, 2174, 271, 211, 76, 1762, 419, 5, 2264, 1360, 586, 509, 1936, 1970, 2113, 1803, 1]. Transformations [1134, 1099, 2249, 1653, 701, 2017, 2172, 130, 85, 1931, 1227, 2008, 1361, 1652, 196, 2223, 380, 1379, 1837, 439, 241, 933, 1670, 507, 1150, 818, 415, 169, 532, 1788, 798, 877, 313, 936, 13, 153, 1443, 1593]. transforming [1315, 1767]. transforms [297, 2100, 383, 1703, 1633, 2109, 535, 1975, 817, 1508, 143, 2265]. Transient [1504, 1538]. transition [882, 1436, 810, 1351]. transitively [1862]. Translation [1517, 1518, 653, 1371]. Translationally [1013]. translations [972]. Transmission [1407]. transport [1071, 1735, 569, 1381, 2262, 140, 556, 1248, 1280, 1089]. Transversal [257]. transverse [57, 1009, 2]. Trapped [266]. Tratnik [983]. Trautman [1741]. Treatment [308, 500, 1297, 1497, 1058, 575, 1996]. trial [1699]. triangular [168]. tridimensional [1810]. trigonometric [163]. trilinear [2238]. Triple [1636, 1637]. triplet [274]. Trivector [271]. trivialization [27]. truncated [1571]. Truncation [651]. Tsetlin [560, 1075]. tube [2145]. tube-type [2145]. tunneling [1719, 1407]. turbulence [91]. turning [891]. twisted [2072, 600]. twisting [2193]. Twistor [1731, 477, 136, 2281, 728, 1422]. twistorial [1161, 1849]. twistors [1289, 1277]. twists [1973]. Two [54, 1198, 2019, 829, 1140, 660, 2048, 2049, 1235, 1726, 651, 973, 1728, 1565, 994, 1542, 1236, 1209, 522, 305, 1641, 682, 1556, 880, 2189, 340, 12, 1305, 1467, 1942, 203, 293, 779, 53, 2156, 2103, 2144, 2207, 45, 1061, 102, 1414, 149, 722, 1214, 1461, 1490, 1350, 1795, 2003, 720, 1530, 2119, 1603, 1436, 512, 139, 794, 1451, 1815, 1909, 997, 1063, 2211, 1323, 639, 190, 2294, 1357, 1311, 531, 950, 44, 1243, 221, 1010, 2010, 921, 2276, 1799, 1997, 1320, 1810, 164, 1442, 1831, 910, 666, 1533, 1287, 295, 550, 1097, 374, 1697, 1386, 640, 324]. two [1686, 1871]. two- [203, 53]. Two-body [1235, 1565, 293, 1414, 1214, 921, 1533, 1097]. two-component [44]. Two-dimensional [2048, 973, 1542, 1236, 305, 2189, 2144, 45, 149, 722, 1461, 1530, 512, 139, 794, 1451, 1909, 2211, 1323, 1311, 950, 1243, 221, 1010, 1799, 164, 1442, 1831, 910, 1386, 640, 324]. two-fermion [1556, 2276]. two-fluid [1997]. two-index [2103]. two-level [1603]. two-matrix [374]. two-mode [1795, 1063]. two-parameter [1357, 1287]. two-phase [720]. two-photon [190]. Two-point [1198, 829, 1140, 651, 2119, 1697]. two-sphere [340]. type [1209, 467, 1746, 984, 354, 412, 1690, 2135, 50, 2218, 45, 2145, 79, 1173, 1741, 2058, 2303, 382, 2097, 2174, 539, 927, 463, 1645, 1267, 1505, 222, 1174, 884, 1440, 1945, 1727, 93, 1532, 209, 985, 789, 877, 2238, 1399, 2060, 600, 287, 183]. type- [2060]. type-I [467, 789]. types [886, 830]. typical [406, 2027]. Tyutin [2036, 1929]. Tzitzeica [1812].

U [1372]. Ultralocal [1412, 2085]. unbounded [484, 1861, 59, 249]. Uncertainty [2160, 417, 418, 285, 1439, 73, 1410]. underdetermined [733]. underlying [1825]. underwater [1138]. undulator [1058]. unidimensional [1493]. unification [114]. Unified [265, 971, 1396, 433, 1620, 1621, 1514, 1964, 1663]. Uniform [869, 1469, 891]. Unimodular [628, 1347, 1583]. Uniqueness [870, 1266, 301, 1203, 1017, 1468, 1724, 14, 226, 707]. unit [680, 1936, 2113]. Unitarity [1832, 926, 1646]. Unitarizable [669, 161]. unitarization [100]. Unitary [29, 505, 527, 682, 302, 234, 2207, 681, 126, 1702, 157, 560, 885, 1186, 211, 1106, 850, 607, 99, 728, 1753, 1593, 265]. unity [2142, 1712, 2307]. Universal [2297, 1630, 2024, 2245, 889, 1144, 1359, 588, 1313, 1017, 1468, 2023, 743, 908, 1451, 1326, 1647, 1494, 1609, 1008]. universe [321, 1510, 665, 2098, 84, 358, 1417, 75, 925, 705]. universes [2079, 1308, 2183, 1545]. unoriented [838]. Unstable [2165, 1810, 2011, 987]. Untenability [71]. untruncated [2158]. upon [423]. Upper [906, 997, 435, 1055]. urn [431]. use [1558]. using [501, 1260, 1538, 204, 677, 2097, 923, 1142, 2109, 346, 1094, 222, 636, 1997, 608, 906, 1443, 516, 530].

V [1467, 12, 254, 1124, 2135, 1150, 1306]. vacua [1189]. Vacuum [1430, 1522, 1268, 1209, 467, 466, 952, 2079, 2016, 2139, 221, 1010, 663, 2060]. Vaidya [322, 2014]. valence [682]. value [55, 1629, 576, 138, 468, 401, 1994, 357, 1039, 496, 778]. valued [466, 1901, 1489, 573, 1271]. values [1151, 59, 1809]. Variable [878, 1335, 1503, 1375, 478]. variable-coefficient [1335]. variables [1412, 457, 2156, 864, 1720, 795, 259, 1958, 211, 1281, 439, 1240, 510, 619, 2289, 2236, 960, 1025, 1155, 1809, 117, 395, 374, 1886]. Variation [778, 538]. Variation-iteration [778]. Variational [30, 1839, 1773, 1113, 1540, 378, 1207, 1027, 612, 1979]. Variations [1733]. varieties [150]. various [203]. varying [1288]. Vector [404, 807, 2069, 599, 1628, 296, 672, 680, 394, 702, 122, 150, 736, 1669, 2163, 2306, 581, 1846, 663, 298, 2240]. vector-axial-vector [394]. Vector-coherent-state [807, 2069]. vector-tensor [736]. vectors [1209, 97, 219, 107, 1122, 7, 56, 1046, 1213, 662, 703, 256, 936]. velocities [152, 430, 171]. velocity [1735, 1181, 938, 310]. Veronese [2192]. version [396, 1269, 1148]. versus [493, 548]. Vertex [1662, 1834, 1562, 2226, 2258, 1077, 756, 926, 434, 1206]. Veselov [1961]. VI [1305]. via [1387, 458, 308, 122, 1323, 206, 1976, 1915, 2134, 1392]. vibrational [189]. vicious [1580]. view [1766, 1050]. viewpoint [1631, 2010, 992]. VII [1942]. Vilkovisky [2036]. violation [1485, 261, 1948, 1949]. Virasoro [2019, 1486, 1722, 926, 626, 1980, 600, 301, 1869]. Virial [1533]. virtual [949]. viscoelasticity [1839]. viscosity [218, 109]. viscous [1545, 255]. Vlasov [239]. Volterra [1194, 595]. vortex [2169, 231, 880, 904, 549, 1104, 1572]. vortex-vortex [2169, 904]. vortices [1918, 938, 637]. Vries [1653, 172, 1623, 2234, 1375, 1990, 2174, 571, 1508, 2051, 2010, 2053, 2175, 1693]. Vries-type [2174].

Wadati [169]. Wake [454]. Wake-free [454]. walk [2110, 1729]. Walker [2179, 2241, 539, 641, 48, 1049, 182, 37, 803, 1479, 1627, 117]. Walker-type [539]. walks [1856, 2188, 1258]. wall [1262]. walls [1643, 1599, 391, 1903, 2099, 833]. Ward [968, 909]. water [1400, 419]. Wave [1758, 461, 499, 1617, 1883, 267, 1345, 1078, 1875, 2233, 505, 2101, 1699, 116, 1163, 1757, 1781, 892, 420, 1337, 327, 1952, 2318, 1626, 1030, 1376, 283, 1229, 1450, 2059, 2191, 1356, 259, 173, 1504, 445, 585, 1368, 1400, 179, 937, 2276, 518, 1997, 1689, 362, 2041, 169, 1787, 533, 898, 2176, 364, 2177, 2073, 1656, 509, 1506, 1592, 1611, 1334]. wave-mediated [505]. waveguides [1403, 999, 1616]. Wavelet [23, 2140, 1974, 1703, 1633, 2109, 673, 646, 446]. Wavelets [1864, 617, 1484, 278]. Wavelike [2248]. waves [1640, 742, 1854, 454, 891, 1085, 1694, 393, 1658, 1080, 138, 468, 39, 1415, 192, 213, 1420, 1276, 1870, 543, 584, 793, 1309, 419, 2167, 26, 2126, 1689, 384, 385, 2176, 995, 110, 625, 1040, 833, 1041, 2012, 706, 2127, 860]. weak [667, 2233, 315, 1264, 1385]. Weight [1221, 126, 527, 2083, 21, 99, 161]. weighted [378, 1677]. Weingarten [2172]. Weiss [1894]. Weizsäcker [760]. well [376, 1401, 700, 1304, 1500]. Wentzel [2275]. Wenzl [1094]. Wess [785, 1901, 481, 1532, 516, 326]. wetted [104]. Weyl [1471, 1184, 1272, 1138, 606, 1216, 303, 1299, 1751, 545, 2042, 1660, 617, 117, 1821]. Wheeler [779, 994]. which [949, 1162, 627, 192, 213, 1439, 209]. White [1341]. White-noise [1341]. Whittaker [1758, 1147]. whose [2049]. Wick [592, 2193, 1440]. Wick-ordered [1440]. Wiener [317, 1678]. Wightman [328]. Wigner [1466, 1116, 1939, 1202, 1585, 1988, 1036, 1037, 404, 146, 196, 463, 1943, 1219, 101, 805, 1553, 1914, 1101, 922, 1321, 1563]. Wigner-function [1914]. Wilson [233]. Winternitz [903]. Wishart [2300]. with/without [891]. within [682, 1716, 681, 1912, 746, 1680, 810]. without [546, 835, 2293, 2034, 891, 1331, 1993, 2312, 90, 597, 620]. Witt [1303]. Witten [1901, 1602]. WKB [1253, 613, 693]. Woltjer [612]. working [27]. world [1493, 896, 288, 416]. wormholes [77]. Wright [774].

X [532, 1614]. X-ray [1614].

Yagle [311]. Yang [1578, 144, 266, 337, 954, 1900, 1527, 2056, 329, 1126, 782, 2018, 1359, 361, 1489, 477, 2258, 615, 622, 2048, 2118, 1530, 756, 1448, 1681, 2211, 1634, 755, 1455, 1454, 14, 1463, 664, 1107, 222, 145, 837, 2224, 490, 226, 1012, 1636, 1637, 34, 1393, 1494, 186, 1542, 397, 2200, 187]. yielding [1881]. Yvon [1352].

Z [3, 1317, 1870]. Zakharov [1622, 499, 1896, 1899, 2012]. Zassenhaus [825]. Zener [1603]. Zero [1460, 1566, 1382, 2017, 312, 1337, 445, 917, 1532, 2035, 1069, 1189, 1322]. zero-curvature [2017, 1337]. zero-dimensional [917]. zero-energy [1532, 2035]. zero-range [1566]. zeros [1752, 1868, 1190, 2150, 1611]. zeta [2141, 2253, 10, 123, 45, 1632, 2255, 854, 1863, 409]. zeta-function [10, 123, 45]. Zhong [1317, 3]. zone [937, 2011]. ZS [1322]. ZS-AKNS [1322]. Zumino [1901, 785, 481, 1532, 516, 326]. Zumino-type [1532].

References

Zhong:1985:HCL

[1] Zai Zhe Zhong. On the hyperbolic complex linear symmetry groups and their local gauge transformation actions. Journal of Mathematical Physics, 26(3):404–406, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [3] and comment [1317].

Singh:1986:STM

[2] S. N. Singh, K. R. Choubey, and B. P. Singh. On steady transverse magnetogas flow. Journal of Mathematical Physics, 27(5):1466–1470, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [57].

Demys:1987:CHC

[3] K. Demys. Comment: “On the hyperbolic complex linear symmetry groups and their local gauge transformation actions”, [J. Math. Phys.

26 (1985), no. 3, 404–406; MR 86h:83055] by Z. Z. Zhong. Journal of Mathematical Physics, 28(2):339, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1] and erratum [1317].

Bokhari:1987:SSS

[4] Ashfaque H. Bokhari and Asghar Qadir. Symmetries of static, spherically symmetric space–times. Journal of Mathematical Physics, 28(5): 1019–1022, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [6] and correction-Bokhari:1990:KVS.

Nelson:1987:GLT

[5] Robert A. Nelson. Generalized Lorentz transformation for an accelerated, rotating frame of reference. Journal of Mathematical Physics, 28(10): 2379–2383, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2264].

Bokhari:1988:ESS

[6] Ashfaque H. Bokhari and Asghar Qadir. Erratum: “Symmetries of static, spherically symmetric space–times”. Journal of Mathematical Physics, 29(2):525, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [4, 219].

Qadir:1988:SSS

[7] Asghar Qadir and M. Ziad. Static spherically symmetric space–times with six Killing vectors. Journal of Mathematical Physics, 29(11):2473– 2474, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [56].

Mitra:1989:FDI

[8] Bani Mitra. Finite-dimensional irreducible representations of the SU(3/1) superalgebra. Journal of Mathematical Physics, 30(3):719–731, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [598].

Saito:1989:DFS

[9] Yoshihiko Saito, Reiji Sugano, Tadayuki Ohta, and Toshiei Kimura. A dynamical formalism of singular Lagrangian system with higher derivatives. Journal of Mathematical Physics, 30(5):1122–1132, May 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [1672].

Elizalde:1989:EZF

[10] E. Elizalde and A. Romeo. Expressions for the zeta-function regularized Casimir energy. Journal of Mathematical Physics, 30(5):1133–1139, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [123].

Malbouisson:1989:SRM

[11] A. P. C. Malbouisson, F. R. A. Simão, and A. Ferraz de Camargo Filho. Some results from a Mellin transform expansion for the heat kernel. Journal of Mathematical Physics, 30(6):1226–1229, June 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [18] and comment [166].

Chandler:1989:BQS

[12] Colston Chandler and A. G. Gibson. N-body quantum scattering theory in two Hilbert spaces. V. Computation strategy. Journal of Mathematical Physics, 30(7):1533–1544, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1467].

Vaklev:1989:GTQ

[13] Y. S. Vaklev. Gauge transformations for the quadratic bundle. Journal of Mathematical Physics, 30(8):1744–1755, August 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [153].

Lopez:1989:EIP

[14] M. C. López, R. J. Noriega, and C. G. Schifini. The equivariant inverse problem and the uniqueness of the Yang–Mills equations. Journal of Mathematical Physics, 30(10):2382–2387, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [226].

Grigore:1989:CCG

[15] D. R. Grigore and O. T. Popp. The complete classification of generalized homogeneous symplectic manifolds. Journal of Mathematical Physics, 30 (11):2476–2483, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [520].

Srivastava:1989:APT

[16] Sunita Srivastava and Vishwamittar. Application of perturbation theory to the damped sextic oscillator. Journal of Mathematical Physics, 30 (12):2815–2818, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [521].

Jakubiec:1989:CPT

[17] Andrzej Jakubiec and Jerzy Kijowski. On the Cauchy problem for the theory of gravitation with nonlinear Lagrangian. Journal of Mathematical Physics, 30(12):2923–2924, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [337].

Malbouisson:1989:ESR

[18] A. P. C. Malbouisson, F. R. A. Simão, and A. F. de Camargo Filho. Erratum: “Some results from a Mellin transform expansion for the heat kernel”. Journal of Mathematical Physics, 30(12):3008, December 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [11, 166].

Elduque:1990:QAA

[19] Alberto Elduque. Quadratic alternative algebras. Journal of Mathematical Physics, 31(1):1–5, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i1/p1_s1.

Batalin:1990:EBQ

[20] I. A. Batalin, P. M. Lavrov, and I. V. Tyutin. Extended BRST quantization of gauge theories in the generalized canonical formalism. Journal of Mathematical Physics, 31(1):6–13, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i1/p6_s1.

LeBlanc:1990:SMR

[21] R. Le Blanc and D. J. Rowe. Superfield and matrix realizations of highest weight representations for osp(m/2n). Journal of Mathematical Physics, 31(1):14–36, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i1/p14_s1.

Ahmed:1990:GST

[22] E. Ahmed and A. S. Hegazi. On generalization of the Sinaĭ theorem in the random site problem. Journal of Mathematical Physics, 31(1): 37–38, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i1/p37_s1.

Holschneider:1990:WAC

[23] Matthias Holschneider. Wavelet analysis on the circle. Journal of Mathematical Physics, 31(1):39–44, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i1/p39_s1.

Bartocci:1990:CSF

[24] Claudio Bartocci, Ugo Bruzzo, and Giovanni Landi. Chern–Simons forms on principal superfiber bundles. Journal of Mathematical Physics, 31(1): 45–54, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i1/p45_s1.

Kim:1990:STG

[25] Y. S. Kim and E. P. Wigner. Space–time geometry of relativistic particles. Journal of Mathematical Physics, 31(1):55–60, January 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p55_s1.

Ogura:1990:SWR

[26] Hisanao Ogura and Nobuyuki Takahashi. Scattering of waves from a random spherical surface — Mie scattering. Journal of Mathematical Physics, 31(1):61–75, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i1/p61_s1.

Shiekh:1990:TCQ

[27] A. Y. Shiekh. The trivialization of constraints in quantum theory (working in a general gauge/parametrization). Journal of Mathematical Physics, 31(1):76–82, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i1/p76_s1.

Tuynman:1990:RQN

[28] G. M. Tuynman. Reduction, quantization, and nonunimodular groups. Journal of Mathematical Physics, 31(1):83–90, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p83_s1.

Beltrametti:1990:UMD

[29] Enrico G. Beltrametti, Gianni Cassinelli, and Pekka J. Lahti. Unitary measurements of discrete quantities in quantum mechanics. Journal of Mathematical Physics, 31(1):91–98, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i1/p91_s1.

Aguilera-Navarro:1990:VPS

[30] V. C. Aguilera-Navarro, G. A. Estévez, and R. Guardiola. Variational and perturbative schemes for a spiked harmonic oscillator. Journal of Mathematical Physics, 31(1):99–104, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i1/p99_s1.

Brzezniak:1990:SRF

[31] Zdzisław Brzeźniak. Some remarks on the Feynman–Kac formula. Journal of Mathematical Physics, 31(1):105–107, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p105_s1.

Znojil:1990:SAA

[32] M. Znojil. Singular anharmonicities and the analytic continued fractions. II. The potentials V (r) = ar^{2} + br^{−4} + cr^{−6}. Journal of Mathematical Physics, 31(1):108–112, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i1/p108_s1.

Kaup:1990:QBH

[33] D. J. Kaup and Peter J. Olver. Quantization of bi-Hamiltonian systems. Journal of Mathematical Physics, 31(1):113–117, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p113_s1.

Pu:1990:YBE

[34] Fu-Cho Pu and D. H. Sattinger. The Yang–Baxter equations and differential identities. Journal of Mathematical Physics, 31(1):118–126, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i1/p118_s1.

Coleman:1990:HAD

[35] Robert Alan Coleman and Herbert Korte. Harmonic analysis of directing fields. Journal of Mathematical Physics, 31(1):127–130, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i1/p127_s1.

Edgar:1990:CPA

[36] S. B. Edgar. On Chandrasekhar’s perturbation analysis. I. The superfluity of the N.P. system. Journal of Mathematical Physics, 31(1):131– 136, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i1/p131_s1.

Nunez:1990:RCR

[37] L. Núñez, U. Percoco, and V. M. Villalba. Ricci and contracted Ricci collineations of the Robertson–Walker space–time. Journal of Mathematical Physics, 31(1):137–139, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i1/p137_s1.

Gaete:1990:GES

[38] Patricio Gaete and Roberto Hojman. General exact solution for homogeneous time-dependent self-gravitating perfect fluids. Journal of Mathematical Physics, 31(1):140–143, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i1/p140_s1.

Hoenselaers:1990:MWK

[39] C. Hoenselaers and F. J. Ernst. Matching p p waves to the Kerr metric. Journal of Mathematical Physics, 31(1):144–146, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p144_s1.

Krori:1990:KOE

[40] K. D. Krori and R. Bhattacharjee. A Kerr object embedded in a gravitational field. II. Journal of Mathematical Physics, 31(1):147–148, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i1/p147_s1.

Krori:1990:MLS

[41] K. D. Krori, P. Borgohain, and Kanika Das. A Marder-like solution in higher dimensions. Journal of Mathematical Physics, 31(1):149–150, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i1/p149_s1.

Maartens:1990:CSS

[42] R. Maartens and M. S. Maharaj. Conformally symmetric static fluid spheres. Journal of Mathematical Physics, 31(1):151–155, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/
p151_s1.

Balakrishnan:1990:SQP

[43] Janaki Balakrishnan, S. N. Biswas, Ashok Goyal, and S. K. Soni. Stochastic quantization of para-Fermi fields. Journal of Mathematical Physics, 31(1):156–163, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i1/p156_s1.

McGuire:1990:SFS

[44] J. B. McGuire. State function and spectral properties of a two-component interacting Fermi gas. Journal of Mathematical Physics, 31(1):164– 169, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i1/p164_s1.

Elizalde:1990:ZFR

[45] E. Elizalde. On the zeta-function regularization of a two-dimensional series of Epstein–Hurwitz type. Journal of Mathematical Physics, 31(1):170–174, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i1/p170_s1.

Wang:1990:BET

[46] Alan P. Wang. Basic equations of three-dimensional radiative transfer. Journal of Mathematical Physics, 31(1):175–181, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p175_s1.

Klaus:1990:LTD

[47] Martin Klaus. On the Levinson theorem for Dirac operators. Journal of Mathematical Physics, 31(1):182–190, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i1/p182_s1.

Kovalyov:1990:DER

[48] M. Kovalyov and M. Légaré. The Dirac equation in Robertson–Walker spaces: A class of solutions. Journal of Mathematical Physics, 31(1):191– 198, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i1/p191_s1.

Kuperin:1990:ETS

[49] Yu. A. Kuperin, K. A. Makarov, and B. S. Pavlov. An extensions theory setting for scattering by breathing bag. Journal of Mathematical Physics, 31(1):199–201, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i1/p199_s1.

Cordani:1990:KTD

[50] B. Cordani, L. Gy. Fehér, and P. A. Horváthy. Kepler-type dynamical symmetries of long-range monopole interactions. Journal of Mathematical Physics, 31(1):202–211, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i1/p202_s1.

Bandyopadhyay:1990:SFT

[51] Pratul Bandyopadhyay, Kohinur Hajra, and Pradip Ghosh. Stochastic field theory, holomorphic quantum mechanics, and supersymmetry. Journal of Mathematical Physics, 31(1):212–220, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p212_s1.

Mozrzymas:1990:NCN

[52] M. Mozrzymas. New class of nonlinear realizations of extended graded anti-de-Sitter supersymmetry OSP(N, 4), N ≥ 2. Journal of Mathematical Physics, 31(1):221–235, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i1/p221_s1.

Dattoli:1990:CKP

[53] G. Dattoli and A. Torre. Cayley–Klein parameters and evolution of twoand three-level systems and squeezed states. Journal of Mathematical Physics, 31(1):236–240, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i1/p236_s1.

Carbonaro:1990:TES

[54] P. Carbonaro. Two explicit solutions in a reducible relativistic system. Journal of Mathematical Physics, 31(1):241–244, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/p241_s1.

Bellomo:1990:DBE

[55] N. Bellomo and S. Kawashima. The discrete Boltzmann equation with multiple collisions: Global existence and stability for the initial value problem. Journal of Mathematical Physics, 31(1):245–253, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i1/
p245_s1.

Qadir:1990:ASS

[56] Asghar Qadir and M. Ziad. Addendum: “Static spherically symmetric space–times with six Killing vectors”, [J. Math. Phys. 29 (1988), no. 11, 2473–2474; MR 90a:83015]. Journal of Mathematical Physics, 31(1): 254, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i1/p254_s1. See [7].

Biedenharn:1990:EST

[57] L. C. Biedenharn and H. van Dam. Erratum: “On steady transverse magnetogas flow”, [J. Math. Phys. 27 (1986), no. 5, 1466–1470; MR 87f:76128] by S. N. Singh, K. R. Choubey and B. P. Singh. Journal of Mathematical Physics, 31(1):v, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2].

Fan:1990:NOF

[58] Hong Yi Fan. Normally ordered Fermi operator realization of the SU_{n} group. Journal of Mathematical Physics, 31(2):257–260, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/
p257_s1.

Rusinek:1990:CMV

[59] Jan Rusinek. On continuity of mean values of unbounded observables. Journal of Mathematical Physics, 31(2):261–263, February 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/p261_
s1.

Iosifescu:1990:KPM

[60] Mircea Iosifescu and Horia Scutaru. Kronecker products, minuscule representations, and polynomial identities. Journal of Mathematical Physics, 31(2):264–277, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i2/p264_s1.

Albeverio:1990:LIB

[61] Sergio Albeverio, Teresa Arede, and Zbigniew Haba. On left invariant Brownian motions and heat kernels of nilpotent Lie groups. Journal of Mathematical Physics, 31(2):278–286, February 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i2/p278_s1.

Fu:1990:IBR

[62] Hong Chen Fu and Chang Pu Sun. Inhomogeneous boson realization of indecomposable representations of Lie algebras. Journal of Mathematical Physics, 31(2):287–293, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i2/p287_s1.

Wu:1990:CNA

[63] Yihren Wu. Classical non-Abelian Berry’s phase. Journal of Mathematical Physics, 31(2):294–299, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i2/p294_s1.

Przanowski:1990:GTAa

[64] M. Przanowski. The group theoretical analysis of gravitational instanton equations. Journal of Mathematical Physics, 31(2):300–307, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p300_s1.

Tip:1990:FPL

[65] A. Tip. Form perturbations of the Laplacian on L^{2}(R) by a class of measures. Journal of Mathematical Physics, 31(2):308–315, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p308_s1.

Dubois-Violette:1990:NDGa

[66] Michel Dubois-Violette, Richard Kerner, and John Madore. Noncommutative differential geometry of matrix algebras. Journal of Mathematical Physics, 31(2):316–322, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i2/p316_s1.

Dubois-Violette:1990:NDGb

[67] Michel Dubois-Violette, Richard Kerner, and John Madore. Noncommutative differential geometry and new models of gauge theory. Journal of Mathematical Physics, 31(2):323–330, February 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i2/p323_s1.

Frauendiener:1990:BHS

[68] Jörg Frauendiener and Ezra T. Newman. On bi-Hamiltonian structures. Journal of Mathematical Physics, 31(2):331–337, February 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/p331_
s1.

Fernandez:1990:CME

[69] Francisco M. Fernández, Julián Echave, and Eduardo A. Castro. On the convergence of the Magnus expansion in the Schrödinger representation. Journal of Mathematical Physics, 31(2):338–341, February 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/p338_
s1.

Papanicolaou:1990:CFP

[70] Vassilis G. Papanicolaou. On the convergence of the Feynman path integral for a certain class of potentials. Journal of Mathematical Physics, 31(2):342–347, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i2/p342_s1.

Lemos:1990:UPC

[71] Nivaldo A. Lemos. Untenability of a proposed constrained dynamics for the damped harmonic oscillator. Journal of Mathematical Physics, 31(2):348–350, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i2/p348_s1.

Nomura:1990:NCF

[72] Masao Nomura. Novel correspondence of the fixed-seniority and the fixed-isospin averages in their reduction formulas. Journal of Mathematical Physics, 31(2):351–353, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i2/p351_s1.

Narcowich:1990:GU

[73] Francis J. Narcowich. Geometry and uncertainty. Journal of Mathematical Physics, 31(2):354–364, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i2/p354_s1.

Carot:1990:NFC

[74] J. Carot. On nonperfect fluid cosmologies evolving towards perfect fluid cosmological models. Journal of Mathematical Physics, 31(2):365–370, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i2/p365_s1.

PoncedeLeon:1990:ISC

[75] J. Ponce de León. Inflation in a spatially closed anisotropic universe. Journal of Mathematical Physics, 31(2):371–377, February 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/p371_
s1.

Magnano:1990:LTC

[76] Guido Magnano, Marco Ferraris, and Mauro Francaviglia. On the Legendre transformation for a class of nonregular higher-order Lagrangian field theories. Journal of Mathematical Physics, 31(2):378–387, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p378_s1.

Degen:1990:OSS

[77] F. Degen. On the nth-order structure of solitonic wormholes. Journal of Mathematical Physics, 31(2):388–400, February 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i2/p388_s1.

Faridi:1990:EME

[78] Abbas M. Faridi. Einstein–Maxwell equations and the groups of homothetic motion. Journal of Mathematical Physics, 31(2):401–409, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p401_s1.

Fayos:1990:EGP

[79] F. Fayos, J. J. Ferrando, and X. Jaén. Electromagnetic and gravitational perturbation of type D space–times. Journal of Mathematical Physics, 31(2):410–415, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i2/p410_s1.

Carlson:1990:NSM

[80] Lee Carlson and W. C. Schieve. The noisy shift map. Journal of Mathematical Physics, 31(2):416–420, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i2/p416_s1.

Odagaki:1990:HIM

[81] T. Odagaki. Hyperinflation in the Ising model on quasiperiodic chains. Journal of Mathematical Physics, 31(2):421–422, February 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/p421_
s1.

Spehler:1990:EII

[82] D. Spehler. Effective interactions involving massive and massless particles of spin 1/2, 1, 3/2, and 2. Journal of Mathematical Physics, 31(2):423–427, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i2/p423_s1.

Lusanna:1990:B

[83] Luca Lusanna. On the BRS’s. Journal of Mathematical Physics, 31(2):428–445, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i2/p428_s1.

Krori:1990:SSK

[84] K. D. Krori, P. Borgohain, P. Kar, and Dipali Das (Kar). Spinor solution in a Kasner universe. Journal of Mathematical Physics, 31(2):446– 447, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i2/p446_s1.

Galvao:1990:GTD

[85] Carlos A. P. Galvão and João Batista T. Boechat. Gauge transformations in Dirac theory of constrained systems. Journal of Mathematical Physics, 31(2):448–451, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i2/p448_s1.

Modanese:1990:RGP

[86] G. Modanese and M. Toller. Radial gauge in Poincaré gauge field theories. Journal of Mathematical Physics, 31(2):452–458, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i2/
p452_s1.

Kupershmidt:1990:SBH

[87] B. A. Kupershmidt. A supersymmetric Benney hierarchy: A semiintegrable system. Journal of Mathematical Physics, 31(2):459–469, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p459_s1.

Li:1990:HSS

[88] Yi Shen Li and Li Ning Zhang. Hamiltonian structure of the super evolution equation. Journal of Mathematical Physics, 31(2):470–475, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p470_s1.

Matsumoto:1990:SSR

[89] Shuji Matsumoto, Shozo Uehara, and Yukinori Yasui. A superparticle on the super Riemann surface. Journal of Mathematical Physics, 31(2):476– 501, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i2/p476_s1.

Maia:1990:MMS

[90] Adolfo Maia, Jr., Erasmo Recami, Waldyr A. Rodrigues, Jr., and Marcio A. F. Rosa. Magnetic monopoles without string in the Kähler–Clifford algebra bundle: A geometrical interpretation. Journal of Mathematical Physics, 31(2):502–505, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i2/p502_s1. See erratum [597].

Qian:1990:TMS

[91] S. Qian, H. H. Chen, and Y. C. Lee. A turbulence model with stochastic soliton motion. Journal of Mathematical Physics, 31(2):506–516, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i2/p506_s1.

Isenberg:1990:EBR

[92] James Isenberg, Martin Jackson, and Vincent Moncrief. Evolution of the Bel–Robinson energy in Gowdy T^{3} × R space–times. Journal of Mathematical Physics, 31(2):517–519, February 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i2/p517_s1.

Nuyts:1990:PCT

[93] Jean Nuyts and Isabelle Platten. Presentation of the constant r-term Krichever–Novikov-type algebras. Journal of Mathematical Physics, 31(3):527–532, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i3/p527_s1.

Dodd:1990:FRP

[94] R. K. Dodd. Fundamental representations of the m-principal realization of gl_{∞}. Journal of Mathematical Physics, 31(3):533–549, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i3/p533_s1.

Ma:1990:RBG

[95] Zhong Qi Ma. Representations of the braid group obtained from quantum sl(3) enveloping algebra. Journal of Mathematical Physics, 31(3):550– 556, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p550_s1.

Wu:1990:PGA

[96] J. Wu and Y. Alhassid. The potential group approach and hypergeometric differential equations. Journal of Mathematical Physics, 31(3): 557–562, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p557_s1.

Bincer:1990:POV

[97] Adam M. Bincer. Parametrization of SU(n) with n − 1 orthonormal vectors. Journal of Mathematical Physics, 31(3):563–567, March 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p563_
s1.

Negro:1990:LRKa

[98] Javier Negro and Mariano A. del Olmo. Local realizations of kinematical groups with a constant electromagnetic field. I. The relativistic case. Journal of Mathematical Physics, 31(3):568–578, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p568_s1.

Palev:1990:HWIa

[99] Tchavdar D. Palev. Highest weight irreducible unitary representations of Lie algebras of infinite matrices. I. The algebra gl(∞). Journal of Mathematical Physics, 31(3):579–586, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i3/p579_s1.

Gruber:1990:ISP

[100] Bruno Gruber, Romuald Lenczewski, and Miguel Lorente. On induced scalar products and unitarization. Journal of Mathematical Physics, 31(3):587–593, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i3/p587_s1.

Lieb:1990:IBR

[101] Elliott H. Lieb. Integral bounds for radar ambiguity functions and Wigner distributions. Journal of Mathematical Physics, 31(3):594– 599, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p594_s1.

Evans:1990:HST

[102] N. W. Evans. On Hamiltonian systems in two degrees of freedom with invariants quartic in the momenta of form p_{1}^{2}p_{2}^{2} · · ·. Journal of Mathematical Physics, 31(3):600–604, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i3/p600_s1.

Hlavaty:1990:PCP

[103] Ladislav Hlavatý. On the Painlevé classification of partial differential equations. Journal of Mathematical Physics, 31(3):605–609, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/
p605_s1.

Salamon:1990:WSG

[104] Peter Salamon, J. Bernholc, R. Stephen Berry, Martin E. Carrera-Patiño, and Bjarne Andresen. The wetted solid—a generalization of the Plateau problem and its implications for sintered materials. Journal of Mathematical Physics, 31(3):610–615, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i3/p610_s1.

Picken:1990:DHI

[105] R. F. Picken. The Duistermaat–Heckman integration formula on flag manifolds. Journal of Mathematical Physics, 31(3):616–638, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/
p616_s1.

Aldrovandi:1990:SAD

[106] Ettore Aldrovandi, Daniela Dohrn, and Francesco Guerra. Stochastic action of dynamical systems on curved manifolds. The geodesic interpolation. Journal of Mathematical Physics, 31(3):639–648, March 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p639_
s1.

Coley:1990:STA

[107] A. A. Coley and B. O. J. Tupper. Space–times admitting special affine conformal vectors. Journal of Mathematical Physics, 31(3):649– 652, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p649_s1.

Przanowski:1990:GTAb

[108] M. Przanowski. The group theoretic analysis of hyperheavenly equations. Journal of Mathematical Physics, 31(3):653–658, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p653_s1.

Huang:1990:ESV

[109] Wung Hong Huang. Effects of the shear viscosity on the character of cosmological evolution. Journal of Mathematical Physics, 31(3):659– 663, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p659_s1.

Taub:1990:IND

[110] A. H. Taub. Interaction of null dust clouds fronted by impulsive plane gravitational waves. Journal of Mathematical Physics, 31(3):664– 668, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p664_s1.

Turyn:1990:GPM

[111] Michael Turyn. The graviton propagator in maximally symmetric spaces. Journal of Mathematical Physics, 31(3):669–679, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p669_s1.

Unnerstall:1990:PSD

[112] T. Unnerstall. Phase-spaces and dynamical descriptions of infinite meanfield quantum systems. Journal of Mathematical Physics, 31(3):680– 688, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p680_s1.

Lee:1990:FDS

[113] M. Howard Lee and Jon T. Nelson. Frequency-dependent susceptibility of a free electron gas in D dimensions. Journal of Mathematical Physics, 31(3):689–694, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i3/p689_s1.

Cho:1990:SSB

[114] Y. M. Cho and D. H. Park. Spherically symmetric black holes in Kaluza–Klein unification. Journal of Mathematical Physics, 31(3):695– 698, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p695_s1.

Isham:1990:AFO

[115] Christopher J. Isham and John R. Klauder. Affine fields and operator representations for the nonlinear σ model. Journal of Mathematical Physics, 31(3):699–711, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i3/p699_s1.

Buchdahl:1990:CRW

[116] H. A. Buchdahl. On the compatibility of relativistic wave equations in Riemann–Cartan spaces. II. Journal of Mathematical Physics, 31(3): 712–714, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p712_s1.

Villalba:1990:SVEa

[117] V. M. Villalba and U. Percoco. Separation of variables and exact solution to Dirac and Weyl equations in Robertson–Walker space–times. Journal of Mathematical Physics, 31(3):715–720, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p715_s1.

Tanimura:1990:RRC

[118] N. Tanimura and O. Tanimura. Recursion relation of character expansion coefficient in SU(3) lattice gauge theory. Journal of Mathematical Physics, 31(3):721–724, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i3/p721_s1.

Lee:1990:LSC

[119] Joohan Lee and Robert M. Wald. Local symmetries and constraints. Journal of Mathematical Physics, 31(3):725–743, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p725_s1.

Hasiewicz:1990:SAC

[120] Z. Hasiewicz, K. Thielemans, and W. Troost. Superconformal algebras and Clifford algebras. Journal of Mathematical Physics, 31(3):744– 756, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p744_s1.

Baez:1990:SCI

[121] John C. Baez. Scattering and complete integrability in conformally invariant nonlinear theories. Journal of Mathematical Physics, 31(3):757– 762, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p757_s1.

Ionescu:1990:LBS

[122] D. C. Ionescu, J. Reinhardt, B. Müller, W. Greiner, and G. Soff. Localized bound states of fermions interacting via massive vector bosons. Journal of Mathematical Physics, 31(3):763–770, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i3/p763_s1.

Elizalde:1990:EEZ

[123] E. Elizalde and A. Romeo. Erratum: “Expressions for the zeta-function regularized Casimir energy”, [J. Math. Phys. 30, 1133 (1989)]. Journal of Mathematical Physics, 31(3):771, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i3/p771_s1. See [10].

Bergvelt:1990:DES

[124] M. J. Bergvelt. Defining equations for supergroup orbits in super Clifford modules. Journal of Mathematical Physics, 31(4):773–775, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i4/p773_s1.

Corrigan:1990:SAQ

[125] E. Corrigan, D. B. Fairlie, P. Fletcher, and R. Sasaki. Some aspects of quantum groups and supergroups. Journal of Mathematical Physics, 31(4):776–780, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i4/p776_s1.

Garcia-Escudero:1990:CUH

[126] Juan Garcia-Escudero and Miguel Lorente. Classification of unitary highest weight representations for noncompact real forms. Journal of Mathematical Physics, 31(4):781–790, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i4/p781_s1.

Antolin:1990:MEF

[127] J. Antolín. Maximum entropy formalism and analytic extrapolation. Journal of Mathematical Physics, 31(4):791–797, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i4/p791_s1.

deAlmeidadaSilva:1990:SLI

[128] M. A. de Almeida da Silva and Ashok Das. A simple Lagrangian for integrable systems. Journal of Mathematical Physics, 31(4):798–800, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i4/p798_s1.

Carinena:1990:GFB

[129] José F. Cariñena and Manuel F. Rañada. Generating functions, bi-Hamiltonian systems, and the quadratic-Hamiltonian theorem. Journal of Mathematical Physics, 31(4):801–807, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i4/p801_s1.

Dowker:1990:CTE

[130] J. S. Dowker and J. P. Schofield. Conformal transformations and the effective action in the presence of boundaries. Journal of Mathematical Physics, 31(4):808–818, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i4/p808_s1.

Beem:1990:SCT

[131] John K. Beem and Phillip E. Parker. Sectional curvature and tidal accelerations. Journal of Mathematical Physics, 31(4):819–827, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i4/p819_s1.

Tarantelli:1990:BDS

[132] A. Tarantelli and L. S. Cederbaum. Block-diagonalization in second quantization. Journal of Mathematical Physics, 31(4):828–837, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i4/p828_s1.

Chetouani:1990:PIS

[133] L. Chetouani, A. Chouchaoui, and T. F. Hammann. Path integral solution for a particle confined in a region. Journal of Mathematical Physics, 31(4):838–841, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i4/p838_s1.

Derezinski:1990:CKS

[134] Jan Derezínski. Criteria for the Kato smoothness with respect to a dispersive N-body Schrödinger operator. Journal of Mathematical Physics, 31(4):842–850, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i4/p842_s1.

Basu:1990:POQ

[135] Sujit Basu and V. V. Kuryshkin. The probability operator in quantum theory and quantization of a system of free fields. Journal of Mathematical Physics, 31(4):851–862, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i4/p851_s1.

Low:1990:CGT

[136] Robert J. Low. The causal geometry of twistor space. Journal of Mathematical Physics, 31(4):863–867, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i4/p863_s1.

Rube:1990:ENC

[137] Petra Rübe. An example of a nontrivial causally simple space–time having interesting consequences for boundary constructions. Journal of Mathematical Physics, 31(4):868–870, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i4/p868_s1.

Hauser:1990:IVP

[138] Isidore Hauser and Frederick J. Ernst. Initial value problem for colliding gravitational plane waves. III. Journal of Mathematical Physics, 31(4): 871–881, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p871_s1.

Katanaev:1990:CIT

[139] M. O. Katanaev. Complete integrability of two-dimensional gravity with dynamical torsion. Journal of Mathematical Physics, 31(4):882– 891, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p882_s1.

Malvagi:1990:REL

[140] F. Malvagi and G. C. Pomraning. Renormalized equations for linear transport in stochastic media. Journal of Mathematical Physics, 31(4): 892–900, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p892_s1.

Corns:1990:PRS

[141] R. A. Corns and T. A. Osborn. Propagators for relativistic systems with non-Abelian interactions. Journal of Mathematical Physics, 31(4): 901–915, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p901_s1.

Piette:1990:FES

[142] Bernard Piette and Wojciech J. Zakrzewski. Finite energy solutions for (1 + 1)-dimensional σ models. Journal of Mathematical Physics, 31(4): 916–923, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p916_s1.

Majid:1990:FTK

[143] Shahn Majid. Fourier transforms on A/G and knot invariants. Journal of Mathematical Physics, 31(4):924–927, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i4/p924_s1.

Kunzle:1990:SSS

[144] H. P. Künzle and A. K. M. Masood ul Alam. Spherically symmetric static SU(2) Einstein–Yang–Mills fields. Journal of Mathematical Physics, 31(4):928–935, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i4/p928_s1.

Matsuno:1990:CES

[145] Y. Matsuno. A class of exact solutions of Yang’s K-gauge equation for SU(2) gauge fields. Journal of Mathematical Physics, 31(4):936– 938, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p936_s1.

Heidenreich:1990:BWE

[146] W. F. Heidenreich and M. Lorente. Bargmann–Wigner equations in de Sitter space. Journal of Mathematical Physics, 31(4):939–947, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i4/p939_s1.

Falqui:1990:SRS

[147] Gregorio Falqui and Cesare Reina. N = 2 super Riemann surfaces and algebraic geometry. Journal of Mathematical Physics, 31(4):948– 952, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p948_s1.

Palev:1990:FDR

[148] Tchavdar D. Palev and Nedjalka I. Stoilova. Finite-dimensional representations of the Lie superalgebra gl(2/2) in a gl(2) ⊕ gl(2) basis. II. Nontypical representations. Journal of Mathematical Physics, 31(4):953– 988, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p953_s1.

Falomir:1990:FDC

[149] H. Falomir, M. A. Muschietti, and E. M. Santangelo. Fermionic determinants for chiral-bag-like two-dimensional systems. Journal of Mathematical Physics, 31(4):989–996, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i4/p989_s1.

Janeczko:1990:GLC

[150] S. Janeczko. Generalized Luneburg canonical varieties and vector fields on quasicaustics. Journal of Mathematical Physics, 31(4):997–1009, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i4/p997_s1.

Tessendorf:1990:RTC

[151] J. Tessendorf. Radiative transfer on curved surfaces. Journal of Mathematical Physics, 31(4):1010–1019, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i4/p1010_s1.

Coll:1990:VBP

[152] Bartolomé Coll and Joan Josep Ferrando. On the velocities of the barotropic perfect fluids. Journal of Mathematical Physics, 31(4):1020– 1025, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p1020_s1.

Vaklev:1990:EGT

[153] Y. Vaklev. Erratum: “Gauge transformations for the quadratic bundle”, [J. Math. Phys. 30, 1744 (1989)]. Journal of Mathematical Physics, 31 (4):1026, April 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i4/p1026_s1. See [13].

Sharma:1990:AIQ

[154] C. S. Sharma and D. F. Almeida. Additive isometries on a quaternionic Hilbert space. Journal of Mathematical Physics, 31(5):1035–1041, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1035_s1.

Bose:1990:SEA

[155] S. K. Bose. Some examples of the algebra of flows. Journal of Mathematical Physics, 31(5):1042–1046, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i5/p1042_s1.

Gromov:1990:JSRa

[156] N. A. Gromov and V. I. Man′ko. The Jordan–Schwinger representations of Cayley–Klein groups. I. The orthogonal groups. Journal of Mathematical Physics, 31(5):1047–1053, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i5/p1047_s1.

Gromov:1990:JSRb

[157] N. A. Gromov and V. I. Man′ko. The Jordan–Schwinger representations of Cayley–Klein groups. II. The unitary groups. Journal of Mathematical Physics, 31(5):1054–1059, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i5/p1054_s1.

Gromov:1990:JSRc

[158] N. A. Gromov and V. I. Man′ko. The Jordan–Schwinger representations of Cayley–Klein groups. III. The symplectic groups. Journal of Mathematical Physics, 31(5):1060–1064, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i5/p1060_s1.

Liu:1990:AEM

[159] Fa Liu, Jia Lun Ping, and Jin Quan Chen. Application of the eigenfunction method to the icosahedral group. Journal of Mathematical Physics, 31(5):1065–1075, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i5/p1065_s1.

Flath:1990:DEA

[160] Daniel E. Flath. Decomposition of the enveloping algebra of sl_{3}. Journal of Mathematical Physics, 31(5):1076–1077, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1076_s1.

Palev:1990:HWIb

[161] Tchavdar D. Palev. Highest weight irreducible unitarizable representations of Lie algebras of infinite matrices. The algebra A_{∞}. Journal of Mathematical Physics, 31(5):1078–1084, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i5/p1078_s1.

Ruegg:1990:SDQ

[162] H. Ruegg. A simple derivation of the quantum Clebsch–Gordan coefficients for SU(2)_{q}. Journal of Mathematical Physics, 31(5):1085–1087, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1085_s1.

Fairlie:1990:IDA

[163] D. B. Fairlie, P. Fletcher, and C. K. Zachos. Infinite-dimensional algebras and a trigonometric basis for the classical Lie algebras. Journal of Mathematical Physics, 31(5):1088–1094, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1088_s1.

Rideau:1990:NEI

[164] G. Rideau and P. Winternitz. Nonlinear equations invariant under the Poincaré, similitude, and conformal groups in two-dimensional space– time. Journal of Mathematical Physics, 31(5):1095–1105, May 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1095_
s1.

Melnikov:1990:NMD

[165] V. K. Mel′nikov. New method for deriving nonlinear integrable systems. Journal of Mathematical Physics, 31(5):1106–1113, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1106_s1.

Malbouisson:1990:CSR

[166] A. P. C. Malbouisson, F. R. A. Simão, and A. Ferraz de Camargo Filho. Comment on: “Some results from a Mellin transform expansion for the heat kernel”, [J. Math. Phys. 30 (1989), no. 6, 1226–1229; MR 90m:58204a]. Journal of Mathematical Physics, 31(5):1114–1115, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1114_s1. See [11, 18].

Golec:1990:APS

[167] J. Golec and G. Ladde. Averaging principle and systems of singularly perturbed stochastic differential equations. Journal of Mathematical Physics, 31(5):1116–1123, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i5/p1116_s1.

Fineschi:1990:TPA

[168] S. Fineschi and E. Landi Degl’Innocenti. A triangular property of the associated Legendre functions. Journal of Mathematical Physics, 31(5): 1124–1125, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1124_s1.

Rangwala:1990:BTS

[169] Abbas A. Rangwala and Jyoti A. Rao. Bäcklund transformations, soliton solutions and wave functions of Kaup–Newell and Wadati–Konno– Ichikawa systems. Journal of Mathematical Physics, 31(5):1126–1132, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1126_s1.

Forest:1990:HFD

[170] Étienne Forest. A Hamiltonian-free description of single particle dynamics for hopelessly complex periodic systems. Journal of Mathematical Physics, 31(5):1133–1144, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i5/p1133_s1.

Sen:1990:LAS

[171] D. K. Sen. Lorentz action on the space of relative velocities and relativity on a three-manifold. Journal of Mathematical Physics, 31(5): 1145–1151, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1145_s1.

Basu:1990:SEL

[172] Ashis Basu and Dipankar Ray. Solutions of an equation that lead to solutions of a modified Emden equation and a coupled Korteweg–de Vries equation. Journal of Mathematical Physics, 31(5):1152–1154, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1152_s1.

Krishnan:1990:RSC

[173] E. V. Krishnan. Remarks on a system of coupled nonlinear wave equations. Journal of Mathematical Physics, 31(5):1155–1156, May 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1155_
s1.

Bulla:1990:FDO

[174] W. Bulla and T. Trenkler. The free Dirac operator on compact and noncompact graphs. Journal of Mathematical Physics, 31(5):1157–1163, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1157_s1.

Novoselsky:1990:NHO

[175] Akiva Novoselsky and Jacob Katriel. Nonspurious harmonic oscillator states in single particle coordinates. Journal of Mathematical Physics, 31(5):1164–1166, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i5/p1164_s1.

Kuznetsov:1990:GPP

[176] Vadim B. Kuznetsov. Generalized polyspheroidal periodic functions and the quantum inverse scattering method. Journal of Mathematical Physics, 31(5):1167–1174, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i5/p1167_s1.

Rath:1990:GSE

[177] Biswanath Rath. Ground-state energy of H = P^{2}+2X^{2}+2\sqrt{2λ}X^{3}+λX^{4} for λ ≪ 1. Journal of Mathematical Physics, 31(5):1175–1176, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1175_s1.

Stubbe:1990:BNB

[178] Joachim Stubbe. Bounds on the number of bound states for potentials with critical decay at infinity. Journal of Mathematical Physics, 31(5): 1177–1180, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1177_s1.

Mavromatis:1990:OSP

[179] H. A. Mavromatis and A. M. Al-Jalal. On obtaining the scattering potential and its moments in the partial wave Born approximation. Journal of Mathematical Physics, 31(5):1181–1188, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1181_s1.

Miura:1990:PIN

[180] Takashi Miura. A path integral for a nonrelativistic spinor particle. Journal of Mathematical Physics, 31(5):1189–1197, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1189_s1.

Hall:1990:CSF

[181] G. S. Hall. Conformal symmetries and fixed points in space–time. Journal of Mathematical Physics, 31(5):1198–1207, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1198_s1.

Moreschi:1990:FNI

[182] Osvaldo M. Moreschi. Future null infinity of Robertson–Walker space– times. Journal of Mathematical Physics, 31(5):1208–1216, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i5/
p1208_s1.

deAzeredoCampos:1990:LMG

[183] R. de Azeredo Campos. Lagrange multipliers and Gauss–Bonnet-type invariants in Riemann–Cartan space. Journal of Mathematical Physics, 31(5):1217–1220, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i5/p1217_s1.

Pantic:1990:SCS

[184] Dražen Pantić. Stochastic calculus and Schrödinger semigroups admitting a ground state. Journal of Mathematical Physics, 31(5):1221–1225, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1221_s1.

Guadagnini:1990:DCS

[185] E. Guadagnini, M. Martellini, and M. Mintchev. Deformations of complex structures and conformal field theories on Riemann surfaces. Journal of Mathematical Physics, 31(5):1226–1233, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1226_s1.

Tafel:1990:CSG

[186] Jacek Tafel. A comparison of solution generating techniques for the selfdual Yang–Mills equations. II. Journal of Mathematical Physics, 31(5): 1234–1236, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1234_s1.

Yang:1990:BYM

[187] Yisong Yang. Blow-up of the SU(2, C) Yang–Mills fields. Journal of Mathematical Physics, 31(5):1237–1239, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i5/p1237_s1.

Aratyn:1990:BCS

[188] Henrik Aratyn. BRS cohomology in string theory: Geometry of abelization and the quartet mechanism. Journal of Mathematical Physics, 31(5): 1240–1252, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1240_s1.

Beumee:1990:AOC

[189] J. G. B. Beumee and H. Rabitz. Application of optimal control theory for selective vibrational excitation in molecules modeled as harmonic physical systems. Journal of Mathematical Physics, 31(5):1253–1260, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i5/p1253_s1.

Lai:1990:ADR

[190] Shan Tao Lai and Ying Nan Chiu. Angular distributions, rotational summations, and integrations in the quadrupole and two-photon photodissociation of molecules. Journal of Mathematical Physics, 31(5): 1261–1273, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1261_s1.

Wei:1990:EDF

[191] Gaoyuan Wei and B. E. Eichinger. Evaluations of distribution functions for flexible macromolecules by the saddle-point method. Journal of Mathematical Physics, 31(5):1274–1279, May 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i5/p1274_s1.

Kharik:1990:PGGa

[192] V. M. Kharik and U. D. Pletner. The problem of gravity-gyroscopic waves, which are excited by the oscillations of a curve. Journal of Mathematical Physics, 31(5):1280–1283, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i5/p1280_s1.

Yang:1990:GLE

[193] Yisong Yang. The Ginzburg–Landau equations for superconducting films and the Meissner effect. Journal of Mathematical Physics, 31(5):1284– 1289, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i5/p1284_s1.

deSouza:1990:BGI

[194] S. M. de Souza and M. T. Thomaz. Beyond Gaussian integrals in Grassman algebra. Journal of Mathematical Physics, 31(6):1297–1299, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1297_s1.

Brihaye:1990:MLP

[195] Y. Brihaye, F. Delaunois, and J. Nuyts. Minimization of Landau potentials invariant under O(3). II. Journal of Mathematical Physics, 31(6): 1300–1303, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1300_s1.

Kasperkovitz:1990:STI

[196] P. Kasperkovitz and M. K. Peev. Similarity transformations of irreducible corepresentations in Wigner canonical form. Journal of Mathematical Physics, 31(6):1304–1309, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i6/p1304_s1.

Salam:1990:STS

[197] M. A. Salam and B. G. Wybourne. Shifted tableaux, Schur’s Q functions, and Kronecker products of S_{n} spin irreps. Journal of Mathematical Physics, 31(6):1310–1314, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i6/p1310_s1.

Shaw:1990:FGC

[198] R. Shaw and T. M. Jarvis. Finite geometries and Clifford algebras. II. Journal of Mathematical Physics, 31(6):1315–1324, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/p1315_s1.

Alisauskas:1990:EOC

[199] Sigitas Ališauskas. Explicit orthonormal Clebsch–Gordan coefficients of SU(3). Journal of Mathematical Physics, 31(6):1325–1332, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/
p1325_s1.

Pan:1990:MRB

[200] Feng Pan. The matrix representation of U_{4} in the U_{2} × U_{2} basis and some isoscalar factors for U_{p+q} ⊇ U_{p} × U_{q}. Journal of Mathematical Physics, 31(6):1333–1339, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i6/p1333_s1. See comment [1145].

Fialowski:1990:DSI

[201] Alice Fialowski. Deformations of some infinite-dimensional Lie algebras. Journal of Mathematical Physics, 31(6):1340–1343, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/p1340_s1.

Turkowski:1990:SLA

[202] P. Turkowski. Solvable Lie algebras of dimension six. Journal of Mathematical Physics, 31(6):1344–1350, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1344_s1.

Cleary:1990:NEV

[203] Paul W. Cleary. Nonexistence and existence of various order integrals for two- and three-dimensional polynomial potentials. Journal of Mathematical Physics, 31(6):1351–1355, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1351_s1.

Davies:1990:DRC

[204] K. T. R. Davies, R. W. Davies, and G. D. White. Dispersion relations for causal Green’s functions: derivations using the Poincaré-Bertrand theorem and its generalizations. Journal of Mathematical Physics, 31(6): 1356–1373, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1356_s1. See erratum [677].

Gu:1990:CLC

[205] Zhu Quan Gu. Classical Liouville completely integrable systems associated with the solutions of Boussinesq–Burgers’ hierarchy. Journal of Mathematical Physics, 31(6):1374–1380, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i6/p1374_s1.

Lee:1990:ALE

[206] You-Quan Lee. Auxiliary linear equations for a class of nonlinear partial differential equations via jet-bundle formulation. Journal of Mathematical Physics, 31(6):1381–1387, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1381_s1.

Yoshioka:1990:QCP

[207] Akira Yoshioka and Kiyotaka Ii. The quantization condition in the presence of a magnetic field and quasiclassical eigenvalues of the Kepler problem with a centrifugal potential and Dirac’s monopole field. Journal of Mathematical Physics, 31(6):1388–1394, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i6/p1388_s1.

Li:1990:PNT

[208] Fu Bin Li. Proof of the noninteraction theorem for a system of N relativistic particles in direct interaction. Journal of Mathematical Physics, 31(6):1395–1399, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i6/p1395_s1. See erratum [1251].

Rabelo:1990:ETW

[209] Mauro L. Rabelo and Keti Tenenblat. On equations of type u_{xt} = F (u, u_{x}) which describe pseudospherical surfaces. Journal of Mathematical Physics, 31(6):1400–1407, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1400_s1.

DeAngelis:1990:SMR

[210] Gian Fabrizio De Angelis. Stochastic mechanics of a relativistic spinless particle. Journal of Mathematical Physics, 31(6):1408–1412, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1408_s1.

Khanna:1990:RSU

[211] F. C. Khanna, A. Mann, and Y. Takahashi. Rotations, squeezing, and the unitary transformation operator from individual particles to Jacobi variables. Journal of Mathematical Physics, 31(6):1413–1415, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/
p1413_s1.

Razumov:1990:SCR

[212] A. V. Razumov. Sternberg construction and reduction. Journal of Mathematical Physics, 31(6):1416–1421, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1416_s1.

Kharik:1990:PGGb

[213] V. M. Kharik and U. D. Pletner. The problem of gravity-gyroscopic waves, which are excited by the oscillations of a curve. Journal of Mathematical Physics, 31(6):1422–1425, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1422_s1.

Matsukidaira:1990:CQS

[214] J. Matsukidaira, J. Satsuma, and W. Strampp. Conserved quantities and symmetries of KP hierarchy. Journal of Mathematical Physics, 31(6): 1426–1434, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1426_s1.

Farina:1990:ALP

[215] J. E. G. Farina and S. N. Mosley. The asymptotic localization of position probability in quantum mechanics. Journal of Mathematical Physics, 31(6):1435–1440, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i6/p1435_s1.

Weber:1990:ODC

[216] T. A. Weber and C. L. Hammer. The one-dimensional Coulomb potential as a generalized function and the hidden O(2) symmetry. Journal of Mathematical Physics, 31(6):1441–1444, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i6/p1441_s1.

Brevik:1990:CFS

[217] Iver Brevik and Roger Sollie. Casimir force on a spherical shell when ϵ(ω)μ(ω) = 1. Journal of Mathematical Physics, 31(6):1445–1455, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1445_s1.

Huang:1990:ACM

[218] Wung Hong Huang. Anisotropic cosmological models with energy density dependent bulk viscosity. Journal of Mathematical Physics, 31(6):1456– 1462, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1456_s1.

Bokhari:1990:KVS

[219] Ashfaque H. Bokhari and Asghar Qadir. Killing vectors of static spherically symmetric metrics. Correction to: “Symmetries of static, spherically symmetric space–times”, [J. Math. Phys. 28 (1987), no. 5, 1019– 1022; erratum; MR 90e:83046a b]. Journal of Mathematical Physics, 31 (6):1463, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1463_s1. See [4, 6].

Liang:1990:CBT

[220] Can Bin Liang. Conformal Birkhoff and Taub theorems in higher dimensions. Journal of Mathematical Physics, 31(6):1464–1466, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/
p1464_s1.

Morchio:1990:IVS

[221] G. Morchio, D. Pierotti, and F. Strocchi. Infrared and vacuum structure in two-dimensional local quantum field theory models. The massless scalar field. Journal of Mathematical Physics, 31(6):1467–1477, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1467_s1.

Marques-Bonham:1990:DEN

[222] S. Marques-Bonham. The Dirac equation in a non-Riemannian manifold. II. An analysis using an internal local n-dimensional space of the Yang–Mills type. Journal of Mathematical Physics, 31(6):1478–1482, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1478_s1.

Villalba:1990:ESD

[223] Víctor M. Villalba. Exact solution of the Dirac equation in a reducible Einstein space. Journal of Mathematical Physics, 31(6):1483–1486, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1483_s1.

Batalin:1990:CQG

[224] I. A. Batalin, P. M. Lavrov, and I. V. Tyutin. Covariant quantization of gauge theories in the framework of extended BRST symmetry. Journal of Mathematical Physics, 31(6):1487–1493, June 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i6/p1487_s1. See er-

ratum [723].

Lundell:1990:ECN

[225] Albert T. Lundell and Yasunari Tosa. Explicit construction of nontrivial elements for homotopy groups of classical Lie groups. Journal of Mathematical Physics, 31(6):1494–1502, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1494_s1.

Noriega:1990:AEI

[226] R. J. Noriega and C. G. Schifini. Addendum to: “The equivariant inverse problem and the uniqueness of the Yang–Mills equations”, [J. Math. Phys. 30 (1989), no. 10, 2382–2387] by M. C. López, R. J. Noriega and C. G. Schifini. Journal of Mathematical Physics, 31(6):1503, June 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i6/p1503_

s1. See [14].

Dixon:1990:ADA

[227] Geoffrey Dixon. Adjoint division algebras and SU(3). Journal of Mathematical Physics, 31(6):1504–1505, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i6/p1504_s1.

Ma:1990:GFS

[228] Zhong Qi Ma and D. H. Tchrakian. Gauge field systems on CP^{n}. Journal of Mathematical Physics, 31(6):1506–1512, June 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i6/p1506_s1.

Beckers:1990:PSP

[229] J. Beckers and N. Debergh. Parastatistics, supersymmetry and parasupercoherent states. Journal of Mathematical Physics, 31(6):1513–1523, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1513_s1.

Gould:1990:CASa

[230] M. D. Gould and R. B. Zhang. Classification of all star and grade star irreps of gl(n|1). Journal of Mathematical Physics, 31(6):1524–1534, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i6/p1524_s1.

Albertin:1990:RDG

[231] Uwe K. Albertin and Harry L. Morrison. Representations of the diffeomorphism group describing an infinite Bose gas in the presence of ideal vortex filaments. Journal of Mathematical Physics, 31(6):1535– 1543, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1535_s1.

Zhou:1990:HLG

[232] Huan Qiang Zhou, Lin Jie Jiang, and Ping Feng Wu. Hidden local gauge invariance in the one-dimensional Hubbard model and its equivalent coupled spin model. Journal of Mathematical Physics, 31(6):1544– 1550, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i6/p1544_s1.

Tratnik:1990:MBC

[233] M. V. Tratnik. Multivariable biorthogonal continuous-discrete Wilson and Racah polynomials. Journal of Mathematical Physics, 31(7):1559– 1575, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i7/p1559_s1.

Bars:1990:UIR

[234] Itzhak Bars and Zhong jian Teng. The unitary irreducible representations of SU(2, 1). Journal of Mathematical Physics, 31(7):1576–1586, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1576_s1.

Chruscinski:1990:SOQ

[235] Dariusz Chruściński. Symplectic orbits in quantum state space. Journal of Mathematical Physics, 31(7):1587–1588, July 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1587_s1.

Li:1990:RBI

[236] Xiang Zhu Li and Josef Paldus. Relationship between S_{N} and U(n) isoscalar factors and higher-order U(n) invariants. Journal of Mathematical Physics, 31(7):1589–1599, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i7/p1589_s1.

Quesne:1990:SRC

[237] C. Quesne. On some Racah coefficients of U(n). Journal of Mathematical Physics, 31(7):1600–1604, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i7/p1600_s1.

Flath:1990:RTO

[238] Dan Flath. Remarks on tensor operators. Journal of Mathematical Physics, 31(7):1605–1609, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i7/p1605_s1.

Holm:1990:MIV

[239] D. D. Holm, W. P. Lysenko, and J. C. Scovel. Moment invariants for the Vlasov equation. Journal of Mathematical Physics, 31(7):1610–1615, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1610_s1.

Cavalcante:1990:NEE

[240] J. A. Cavalcante and J. A. Silva. Nonlinear evolution equations related to the first-order system: φ_{x} = λJφ + P φ. Journal of Mathematical Physics, 31(7):1616–1620, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i7/p1616_s1.

Liu:1990:RBH

[241] Xian Hui Liu. Relations between hyperspherical harmonic transformations and generalized Talmi–Moshinsky transformations. Journal of Mathematical Physics, 31(7):1621–1626, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v31/i7/p1621_s1. See comment

[532].

Abraham-Shrauner:1990:LGS

[242] Barbara Abraham-Shrauner. Lie group symmetries and invariants of the Hénon–Heiles equations. Journal of Mathematical Physics, 31(7):1627– 1631, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i7/p1627_s1.

Verosky:1990:EED

[243] John M. Verosky. Exponentially evolving densities. Journal of Mathematical Physics, 31(7):1632–1634, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i7/p1632_s1.

Negreiros:1990:EHM

[244] Caio J. C. Negreiros. Equivariant harmonic maps into homogeneous spaces. Journal of Mathematical Physics, 31(7):1635–1642, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/
p1635_s1.

Roiesnel:1990:GLP

[245] C. Roiesnel and A. Chakrabarti. Geometry of linear pairs of self-dual gauge fields. Journal of Mathematical Physics, 31(7):1643–1651, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1643_s1.

Graudenz:1990:PBA

[246] Dirk Graudenz. Para-Bose algebras and subgraph polynomials. Journal of Mathematical Physics, 31(7):1652–1656, July 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1652_s1.

Koiller:1990:HAO

[247] Jair Koiller. Hannay’s angles outside the adiabatic range: Coupled linear oscillators near resonance. Journal of Mathematical Physics, 31(7):1657– 1660, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i7/p1657_s1.

Tang:1990:LGF

[248] P. Tang and L. Wilets. The linearized gauge field propagator in an inhomogeneous medium. Journal of Mathematical Physics, 31(7):1661– 1669, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i7/p1661_s1.

Steiner:1990:RAE

[249] F. Steiner and P. Trillenberg. Refined asymptotic expansion for the partition function of unbounded quantum billiards. Journal of Mathematical Physics, 31(7):1670–1676, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i7/p1670_s1.

Wong:1990:ESD

[250] M. K. F. Wong. Exact solution of the n-dimensional Dirac–Coulomb equation. Journal of Mathematical Physics, 31(7):1677–1680, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/
p1677_s1.

Kuperin:1990:EHS

[251] Yu. A. Kuperin, K. A. Makarov, S. P. Merkuriev, A. K. Motovilov, and B. S. Pavlov. Extended Hilbert space approach to few-body problems. Journal of Mathematical Physics, 31(7):1681–1690, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1681_s1.

Dassios:1990:LFM

[252] George Dassios. Low-frequency moments in inverse scattering theory. Journal of Mathematical Physics, 31(7):1691–1692, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1691_s1.

Huber:1990:SIS

[253] Wolfgang Huber. Summational invariants in systems of mass points and rigid bodies. Journal of Mathematical Physics, 31(7):1693–1697, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1693_s1.

Coley:1990:ESP

[254] Alan A. Coley. Equations of state and plane-autonomous systems in Bianchi V imperfect fluid cosmology. Journal of Mathematical Physics, 31(7):1698–1703, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i7/p1698_s1.

Tupper:1990:CRF

[255] B. O. J. Tupper. Conformally Ricci-flat viscous fluids. Journal of Mathematical Physics, 31(7):1704–1706, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i7/p1704_s1.

Tsamparlis:1990:RCV

[256] M. Tsamparlis and D. P. Mason. Ricci collineation vectors in fluid space–times. Journal of Mathematical Physics, 31(7):1707–1722, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1707_s1.

Hajicek:1990:TAC

[257] Petr Hájíček and Karel V. Kuchař. Transversal affine connection and quantization of constrained systems. Journal of Mathematical Physics, 31(7):1723–1732, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i7/p1723_s1.

Kagramanov:1990:CLO

[258] È. D. Kagramanov, R. M. Mir-Kasimov, and Sh. M. Nagiyev. The covariant linear oscillator and generalized realization of the dynamical SU(1, 1) symmetry algebra. Journal of Mathematical Physics, 31(7):1733–1738, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1733_s1.

Kalnins:1990:SOS

[259] E. G. Kalnins and G. C. Williams. Symmetry operators and separation of variables for spin-wave equations in oblate spheroidal coordinates. Journal of Mathematical Physics, 31(7):1739–1744, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1739_s1.

Kerimov:1990:LDF

[260] G. A. Kerimov. Lobachevskian Dirac fields. Journal of Mathematical Physics, 31(7):1745–1754, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i7/p1745_s1.

Kosinski:1990:VCR

[261] Piotr Kosiński and Paweł Maślanka. On the violation of causality in relativistic quantum mechanics. Journal of Mathematical Physics, 31(7): 1755–1760, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i7/p1755_s1.

Lindstrom:1990:CQF

[262] U. Lindström, M. Roček, W. Siegel, P. van Nieuwenhuizen, and A. E. van de Ven. On the covariant quantization of a first-class superparticle. Journal of Mathematical Physics, 31(7):1761–1769, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1761_s1.

Fuchs:1990:SSQ

[263] Jürgen Fuchs and Peter van Driel. Some symmetries of quantum dimensions. Journal of Mathematical Physics, 31(7):1770–1775, July 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i7/p1770_
s1.

Gunaydin:1990:ENS

[264] M. Günaydin. On an exceptional nonassociative superspace. Journal of Mathematical Physics, 31(7):1776–1782, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i7/p1776_s1.

Jarvis:1990:UMG

[265] P. D. Jarvis. Unified models from gauged supergroups. I. Dimensional reduction formalism. Unitary and orthosymplectic series. Journal of Mathematical Physics, 31(7):1783–1790, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i7/p1783_s1.

Malec:1990:TSM

[266] Edward Malec and Piotr Koc. Trapped surfaces in monopole-like Cauchy data of Einstein–Yang–Mills–Higgs equations. Journal of Mathematical Physics, 31(7):1791–1795, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i7/p1791_s1.

Zhang:1990:WFB

[267] Jian Zu Zhang and Yong Chang Qi. Wave functions of bound states of a fermion and a Dirac dyon and matrix elements in an external electromagnetic field. Journal of Mathematical Physics, 31(7):1796–1799, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i7/p1796_s1.

Ramos:1990:ADT

[268] E. Ramos, C. H. Sah, and R. E. Shrock. Algebras of diffeomorphisms of the N-torus. Journal of Mathematical Physics, 31(8):1805–1816, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i8/p1805_s1.

Tsohantjis:1990:CRS

[269] I. Tsohantjis and J. F. Cornwell. The complete root systems of the affine Kac–Moody superalgebras. Journal of Mathematical Physics, 31(8):1817–1838, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i8/p1817_s1.

Buchholz:1990:QFG

[270] Detlev Buchholz. On quantum fields that generate local algebras. Journal of Mathematical Physics, 31(8):1839–1846, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1839_s1.

Jancewicz:1990:TFT

[271] Bernard Jancewicz. Trivector Fourier transformation and electromagnetic field. Journal of Mathematical Physics, 31(8):1847–1852, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i8/p1847_s1.

Capps:1990:RAK

[272] Richard H. Capps. Representations of affine Kac–Moody algebras and the affine scalar product. Journal of Mathematical Physics, 31(8):1853– 1858, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1853_s1.

Cassinelli:1990:NIS

[273] Gianni Cassinelli and Pekka J. Lahti. Nontransitive imprimitivity systems for the Galilei group. Journal of Mathematical Physics, 31(8):1859– 1861, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1859_s1.

Pierotti:1990:SRG

[274] Dario Pierotti. Some remarks on the Gupta–Bleuler triplet. Journal of Mathematical Physics, 31(8):1862–1871, August 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1862_s1.

Sijacki:1990:ALR

[275] Dj. Šijački. All SL(3, R) ladder representations. Journal of Mathematical Physics, 31(8):1872–1876, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i8/p1872_s1.

Takasaki:1990:SHK

[276] K. Takasaki. Symmetries of hyper-Kähler (or Poisson gauge field) hierarchy. Journal of Mathematical Physics, 31(8):1877–1888, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/
p1877_s1.

Zhang:1990:CSG

[277] R. B. Zhang and M. D. Gould. Classification of star and grade-star representations of C(n+1). Journal of Mathematical Physics, 31(8):1889– 1897, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1889_s1.

Lawton:1990:TFC

[278] Wayne M. Lawton. Tight frames of compactly supported affine wavelets. Journal of Mathematical Physics, 31(8):1898–1901, August 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1898_
s1.

Mokole:1990:ESS

[279] E. L. Mokole, A. L. Mullikin, and M. B. Sledd. Exact and steady-state solutions to sinusoidally excited, half-infinite chains of harmonic oscillators with one isotopic defect. Journal of Mathematical Physics, 31(8):1902– 1913, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1902_s1.

Minotti:1990:SLE

[280] F. Minotti and C. Moreno. Solution of Laplace’s equation in plane singleconnected regions bounded by arbitrary single curves. Journal of Mathematical Physics, 31(8):1914–1918, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i8/p1914_s1.

Okubo:1990:ICF

[281] Susumu Okubo. Integrability condition and finite-periodic Toda lattice. Journal of Mathematical Physics, 31(8):1919–1928, August 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1919_
s1.

Woodcock:1990:PSA

[282] Harry W. Woodcock and Warren N. Herman. Particle symmetry and antisymmetry in approximately relativistic Lagrangians. Journal of Mathematical Physics, 31(8):1929–1938, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i8/p1929_s1.

Hillion:1990:GPH

[283] Pierre Hillion. The Goursat problem for the homogeneous wave equation. Journal of Mathematical Physics, 31(8):1939–1941, August 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1939_
s1.

Cai:1990:BPP

[284] Y. Q. Cai, G. Papini, and W. R. Wood. Berry’s phase for photons and topology in Maxwell’s theory. Journal of Mathematical Physics, 31(8):1942–1946, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i8/p1942_s1.

Lynch:1990:TEO

[285] Robert Lynch and Harry A. Mavromatis. Nth (even)-order minimum uncertainty products. Journal of Mathematical Physics, 31(8):1947– 1950, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1947_s1.

Vaidya:1990:ACG

[286] Arvind Narayan Vaidya and Henrique Boschi Filho. Algebraic calculation of the Green’s function for the Hartmann potential. Journal of Mathematical Physics, 31(8):1951–1954, August 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1951_s1.

Znojil:1990:GCF

[287] Miloslav Znojil. The generalized continued fractions and potentials of the Lennard-Jones type. Journal of Mathematical Physics, 31(8):1955– 1961, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1955_s1.

Perlick:1990:RPG

[288] V. Perlick. On redshift and parallaxes in general relativistic kinematical world models. Journal of Mathematical Physics, 31(8):1962–1971, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i8/p1962_s1.

Davidson:1990:IEM

[289] W. Davidson. Internal and external metrics for a perfect fluid cylinder in general relativity. Journal of Mathematical Physics, 31(8):1972– 1973, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p1972_s1.

Glass:1990:SCS

[290] E. N. Glass. A spherical collapse solution with neutrino outflow. Journal of Mathematical Physics, 31(8):1974–1977, August 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1974_s1.

Nakazawa:1990:NPS

[291] Hiroshi Nakazawa. Numerical procedures for sample structures on stochastic differential equations. Journal of Mathematical Physics, 31 (8):1978–1990, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i8/p1978_s1.

Crawford:1990:BGE

[292] J. P. Crawford. Bispinor geometry for even-dimensional space–time. Journal of Mathematical Physics, 31(8):1991–1997, August 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1991_
s1.

Crater:1990:ETB

[293] H. W. Crater and P. Van Alstine. Extension of two-body Dirac equations to general covariant interactions through a hyperbolic transformation. Journal of Mathematical Physics, 31(8):1998–2014, August 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p1998_
s1.

Song:1990:NSE

[294] Yongxin Song and Yuan Ben Dai. A new scheme for enumerating connected diagrams on the lattice. Journal of Mathematical Physics, 31(8):2015–2024, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i8/p2015_s1.

Tanimura:1990:EMB

[295] N. Tanimura and O. Tanimura. U(2) extension of modified Bessel function with two indices. Journal of Mathematical Physics, 31(8):2025– 2032, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i8/p2025_s1.

Bremner:1990:LAV

[296] Murray R. Bremner. On a Lie algebra of vector fields on a complex torus. Journal of Mathematical Physics, 31(8):2033–2034, August 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i8/p2033_
s1.

Basu:1990:CIT

[297] Debabrata Basu and T. K. Kar. A class of integral transforms associated with the Majorana representation of SL(2, C). Journal of Mathematical Physics, 31(8):2035–2039, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v31/i8/p2035_s1.

Wang:1990:SAM

[298] Shi Kun Wang and Kai Wen Xu. Superconformal algebra of meromorphic vector fields with multipoles on a super Riemann sphere. Journal of Mathematical Physics, 31(8):2040–2041, August 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i8/p2040_s1.

Majorana:1990:RRM

[299] A. Majorana. Relativistic relaxation models for a simple gas. Journal of Mathematical Physics, 31(8):2042–2046, August 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i8/p2042_s1.

Avellaneda:1990:ECA

[300] Marco Avellaneda and Oscar Bruno. Effective conductivity and average polarizability of random polycrystals. Journal of Mathematical Physics, 31(8):2047–2056, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i8/p2047_s1.

Uretsky:1990:UVA

[301] Jack L. Uretsky. Uniqueness of the Virasoro algebra. Journal of Mathematical Physics, 31(9):2059–2060, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i9/p2059_s1.

Bacry:1990:UR

[302] Henri Bacry. A unitary representation of SL(2, R). Journal of Mathematical Physics, 31(9):2061–2077, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i9/p2061_s1.

IommiAmunategui:1990:BRW

[303] G. Iommi Amunátegui. Branching rules for the Weyl group W(D_{n}). Journal of Mathematical Physics, 31(9):2078–2084, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2078_s1.

Rubin:1990:PSC

[304] J. Rubin and P. Winternitz. Point symmetries of conditionally integrable nonlinear evolution equations. Journal of Mathematical Physics, 31(9):2085–2090, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2085_s1.

Anghel:1990:TDM

[305] N. Anghel. The two-dimensional magnetic field problem revisited. Journal of Mathematical Physics, 31(9):2091–2093, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2091_s1.

McInnes:1990:HBS

[306] Brett McInnes. Hosotani breaking of E_{6} to a subgroup of rank five. Journal of Mathematical Physics, 31(9):2094–2104, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2094_s1.

Santamato:1990:NAG

[307] E. Santamato, G. Abbate, and P. Maddalena. A new approach to global stability analysis of one-dimensional continuous dissipative systems. Journal of Mathematical Physics, 31(9):2105–2115, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2105_s1.

Hebda:1990:THO

[308] Piotr W. Hebda. Treatment of higher-order Lagrangians via the construction of dynamically equivalent first-order Lagrangians. Journal of Mathematical Physics, 31(9):2116–2125, September 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2116_s1.

Lusanna:1990:LHM

[309] Luca Lusanna. Lagrangian and Hamiltonian many-time equations. Journal of Mathematical Physics, 31(9):2126–2135, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2126_s1.

vanGroesen:1990:BLC

[310] E. van Groesen and F. Mainardi. Balance laws and centro velocity in dissipative systems. Journal of Mathematical Physics, 31(9):2136–2140, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i9/p2136_s1.

Cheney:1990:SAY

[311] Margaret Cheney. Stability analysis of the Yagle–Levy multidimensional inverse scattering algorithm. Journal of Mathematical Physics, 31(9):2141–2144, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2141_s1.

Das:1990:ZCF

[312] Ashok Das and Shibaji Roy. The zero curvature formulation of the sKdV equations. Journal of Mathematical Physics, 31(9):2145–2149, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2145_s1.

Tian:1990:BTI

[313] C. Tian and Y. J. Zhang. Bäcklund transformations for the isospectral and nonisospectral AKNS hierarchies. Journal of Mathematical Physics, 31(9):2150–2154, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2150_s1.

DeFacio:1990:CLE

[314] Brian DeFacio. Classical, linear, electromagnetic impedance theory with infinite integrable discontinuities. Journal of Mathematical Physics, 31 (9):2155–2164, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2155_s1.

Florescu:1990:MSS

[315] Viorica Florescu and Suzana Patrascu. On the modification of the Stark states of hydrogen by a weak ac electric field. Journal of Mathematical Physics, 31(9):2165–2169, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i9/p2165_s1.

Pursey:1990:PLF

[316] D. L. Pursey. Properties of Leach–Flessas–Gorringe polynomials. Journal of Mathematical Physics, 31(9):2170–2171, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2170_s1.

Aktosun:1990:ISP

[317] Tuncay Aktosun and Cornelis van der Mee. Inverse scattering problem for the 3-D Schrödinger equation and Wiener–Hopf factorization of the scattering operator. Journal of Mathematical Physics, 31(9):2172–2180, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i9/p2172_s1.

Roberts:1990:SSP

[318] Thomas M. Roberts. Scattering for step-periodic potentials in one dimension. Journal of Mathematical Physics, 31(9):2181–2191, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2181_s1.

Cho:1990:CAQ

[319] H. T. Cho, Adel Diek, and R. Kantowski. A Clifford algebra quantization of Dirac’s electron positron field. Journal of Mathematical Physics, 31 (9):2192–2200, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2192_s1.

Darkhosh:1990:DQM

[320] Teymour Darkhosh. Dirac quantization of massive spin-one particles in an external symmetrical tensor field. Journal of Mathematical Physics, 31(9):2201–2203, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2201_s1.

Aguirregabiria:1990:EDC

[321] J. M. Aguirregabiria, A. Chamorro, J. R. Etxebarria, and M. Rivas. Embedding of a Demiański cavity with small rotation parameter in a perturbation of a Friedmann universe with cosmological constant. Journal of Mathematical Physics, 31(9):2204–2207, September 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2204_s1.

Chatterjee:1990:HDV

[322] S. Chatterjee, B. Bhui, and A. Banerjee. Higher-dimensional Vaidya metric with an electromagnetic field. Journal of Mathematical Physics, 31(9):2208–2210, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2208_s1.

Del-Castillo-Negrete:1990:SST

[323] Diego Del-Castillo-Negrete and Sergio Hojman. Symmetries of space– time and geodesic symmetries. Journal of Mathematical Physics, 31 (9):2211–2216, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2211_s1.

Zhong:1990:RDR

[324] Zai Zhe Zhong. Regular double Riemann–Hilbert problems and double Kac–Moody algebraic structures for two-dimensional reduced gravity. Journal of Mathematical Physics, 31(9):2217–2220, September 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2217_
s1.

Hammond:1990:SOE

[325] Richard T. Hammond. Second-order equations and quadratic Lagrangians. Journal of Mathematical Physics, 31(9):2221–2224, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2221_s1.

Zainuddin:1990:NTN

[326] Hishamuddin Zainuddin. Noether’s theorem in nonlinear σ models with a Wess–Zumino term. Journal of Mathematical Physics, 31(9):2225–2236, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i9/p2225_s1.

Durand:1990:MMN

[327] Loyal Durand and Alan Gara. Matrix methods for the numerical solution of relativistic wave equations. Journal of Mathematical Physics, 31(9):2237–2243, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2237_s1.

Strube:1990:CTM

[328] Dieter Strube. Collision theory for massless particles in the framework of a Wightman field theory. Journal of Mathematical Physics, 31(9):2244– 2252, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i9/p2244_s1.

Chakravarty:1990:HSS

[329] Sarbarish Chakravarty, S. L. Kent, and E. T. Newman. High symmetry solutions of the anti-self-dual Yang–Mills equations. Journal of Mathematical Physics, 31(9):2253–2257, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i9/p2253_s1.

Nash:1990:SPC

[330] Charles Nash. Singularly perturbed Chern–Simons theory. Journal of Mathematical Physics, 31(9):2258–2262, September 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2258_s1.

Ikeda:1990:FRO

[331] Kaoru Ikeda. The Fock representation of osp(∞|∞) and the orthogonal symplectic super KP hierarchy. Journal of Mathematical Physics, 31 (9):2263–2270, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i9/p2263_s1.

Urrutia:1990:LSN

[332] L. F. Urrutia and J. Zanelli. Local supersymmetry in nonrelativistic systems. Journal of Mathematical Physics, 31(9):2271–2277, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2271_s1.

VanderJeugt:1990:CFI

[333] J. Van der Jeugt, J. W. B. Hughes, R. C. King, and J. Thierry-Mieg. Character formulas for irreducible modules of the Lie superalgebras sl(m/n). Journal of Mathematical Physics, 31(9):2278–2304, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2278_s1.

Bowlin:1990:CSC

[334] James B. Bowlin and Alfred S. Goldhaber. Classical scattering of a charged particle on an extended magnetic monopole. Journal of Mathematical Physics, 31(9):2305–2313, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i9/p2305_s1.

Suzuki:1990:SCL

[335] Noboru Suzuki. Simple calculation of Löwdin’s alpha function. III. Procedure for calculating h_{n,2n−i}(LM|l)/(2n − i)! successively. Journal of Mathematical Physics, 31(9):2314–2326, September 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2314_s1.

Slowikowski:1990:SFB

[336] Wojtek Słowikowski and Klaus Mølmer. Squeezing of free Bose fields. Journal of Mathematical Physics, 31(9):2327–2333, September 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i9/p2327_
s1.

Swierczynski:1990:ECP

[337] Z. Świerczyński. Erratum: “On the Cauchy problem for Yang–Mills equations with external current”, [J. Math. Phys. 30 (1989), no. 8, 1699–1706; MR 90j:58023]. Journal of Mathematical Physics, 31(9):2334, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2334_s1. See [17].

Sugano:1990:CGG

[338] Reiji Sugano and Toshiei Kimura. Classification of gauge groups in terms of algebraic structure of first class constraints. Journal of Mathematical Physics, 31(10):2337–2344, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i10/p2337_s1.

Wybourne:1990:SPF

[339] B. G. Wybourne. Symmetrized powers of the fundamental irrep of E_{6}. Journal of Mathematical Physics, 31(10):2345, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i10/p2345_s1.

Bose:1990:ATS

[340] S. K. Bose and S. A. Bruce. Algebras for the two-sphere and the threesphere groups of compact simple Lie groups. Journal of Mathematical Physics, 31(10):2346–2351, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i10/p2346_s1.

Dashen:1990:RSR

[341] Roger Dashen and Gregory J. Orris. Rough surfaces and the renormalization group. Journal of Mathematical Physics, 31(10):2352–2360, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i10/p2352_s1.

Zayed:1990:HSG

[342] E. M. E. Zayed. Hearing the shape of a general doubly connected domain in R^{3} with impedance boundary conditions. Journal of Mathematical Physics, 31(10):2361–2365, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i10/p2361_s1.

Schaller:1990:AGQ

[343] P. Schaller and G. Schwarz. Anomalies from geometric quantization of fermionic field theories. Journal of Mathematical Physics, 31(10):2366– 2377, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2366_s1.

Wald:1990:ICF

[344] Robert M. Wald. On identically closed forms locally constructed from a field. Journal of Mathematical Physics, 31(10):2378–2384, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i10/p2378_s1.

Dawson:1990:SPL

[345] D. Dawson, L. Gorostiza, and A. Wakolbinger. Schrödinger processes and large deviations. Journal of Mathematical Physics, 31(10):2385– 2388, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2385_s1.

Jumarie:1990:SMF

[346] Guy Jumarie. Solution of the multivariate Fokker–Planck equation by using a maximum path entropy principle. Journal of Mathematical Physics, 31(10):2389–2392, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i10/p2389_s1.

Droz-Vincent:1990:SPR

[347] Philippe Droz-Vincent and Paweł Nurowski. Symmetries in predictive relativistic mechanics. Journal of Mathematical Physics, 31(10):2393– 2398, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2393_s1.

Benatti:1990:MES

[348] Fabio Benatti. On the mixing-enhancing structure of a class of quantum dynamical semigroups. Journal of Mathematical Physics, 31(10):2399– 2405, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2399_s1.

deOliveira:1990:CSO

[349] César R. de Oliveira. On the complete system of observables in quantum mechanics. Journal of Mathematical Physics, 31(10):2406–2409, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i10/p2406_s1.

Ifantis:1990:ADE

[350] E. K. Ifantis and P. N. Panagopoulos. Applications of the differentiability of eigenvectors and eigenvalues to a perturbed harmonic oscillator. Journal of Mathematical Physics, 31(10):2410–2413, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i10/p2410_s1.

Newton:1990:FM

[351] Roger G. Newton. Factorizations of the S matrix. Journal of Mathematical Physics, 31(10):2414–2424, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i10/p2414_s1.

Cataldo:1990:AHC

[352] Mauricio Cataldo and Nikolai V. Mitskiévic. Anisotropic homogeneous cosmologies with perfect fluid and electric field. Journal of Mathematical Physics, 31(10):2425–2428, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i10/p2425_s1.

Shahid-Saless:1990:CSC

[353] Bahman Shahid-Saless. Curvature-squared cosmology in the first-order formalism. Journal of Mathematical Physics, 31(10):2429–2433, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i10/p2429_s1.

Carminati:1990:SFP

[354] J. Carminati. Shear-free perfect fluids in general relativity. II. Aligned, Petrov type III space–times. Journal of Mathematical Physics, 31 (10):2434–2440, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i10/p2434_s1.

Damour:1990:RPT

[355] Thibault Damour and Bernd Schmidt. Reliability of perturbation theory in general relativity. Journal of Mathematical Physics, 31(10):2441– 2453, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2441_s1.

Tikekar:1990:EMR

[356] Ramesh Tikekar. Exact model for a relativistic star. Journal of Mathematical Physics, 31(10):2454–2458, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i10/p2454_s1.

Liu:1990:IVF

[357] Hong Ya Liu. Initial value formulation for the spherically symmetric dust solution. Journal of Mathematical Physics, 31(10):2459–2461, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i10/p2459_s1.

Liu:1990:RTS

[358] Hong Ya Liu. Relativistic theory of spherically symmetric perturbation in dust universe. Journal of Mathematical Physics, 31(10):2462– 2465, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2462_s1.

Wassam:1990:STC

[359] W. A. Wassam, Jr., J. A. Balderas-López, and Go. Torres-Vega. Stationarity of time correlation functions for globally linear classical systems and its consequences. Journal of Mathematical Physics, 31(10):2466– 2479, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2466_s1.

Moschella:1990:ISB

[360] Ugo Moschella. Infrared singularities and breaking of the Poincaré group: The massless dipole field. Journal of Mathematical Physics, 31(10):2480– 2489, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2480_s1.

Dostoglou:1990:AFE

[361] S. Dostoglou. On the asymptotics of finite energy solutions of the Yang– Mills–Higgs equations. Journal of Mathematical Physics, 31(10):2490– 2496, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2490_s1.

Pang:1990:NBW

[362] Gen Di Pang, Fu cho Pu, and Bao Heng Zhao. Norms of Bethe wave functions for the nonlinear Schrödinger model of spin-1/2 particles. Journal of Mathematical Physics, 31(10):2497–2500, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i10/p2497_s1.

Keller:1990:MDE

[363] Jaime Keller and Suemi Rodríguez-Romo. A multivectorial Dirac equation. Journal of Mathematical Physics, 31(10):2501–2510, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i10/
p2501_s1.

Shaarawi:1990:NAS

[364] Amr M. Shaarawi, Ioannis M. Besieris, and Richard W. Ziolkowski. A novel approach to the synthesis of nondispersive wave packet solutions to the Klein–Gordon and Dirac equations. Journal of Mathematical Physics, 31(10):2511–2519, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i10/p2511_s1.

Hannabuss:1990:CSA

[365] K. C. Hannabuss and W. R. Weiss. Corepresentations of superalgebras and an antilinear Dirac operator. Journal of Mathematical Physics, 31(10):2520–2524, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i10/p2520_s1.

Hannabuss:1990:EIR

[366] K. C. Hannabuss. The electroweak interaction as a relativistic symmetry. Journal of Mathematical Physics, 31(10):2525–2527, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i10/
p2525_s1.

Beckers:1990:SFN

[367] J. Beckers, L. Gagnon, V. Hussin, and P. Winternitz. Superposition formulas for nonlinear superequations. Journal of Mathematical Physics, 31(10):2528–2534, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i10/p2528_s1.

Laurenzi:1990:AST

[368] Bernard J. Laurenzi. An analytic solution to the Thomas–Fermi equation. Journal of Mathematical Physics, 31(10):2535–2537, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i10/
p2535_s1.

Guo:1990:BAS

[369] Han Ying Guo, Jian Min Shen, Shi Kun Wang, and Kai Wen Xu. Beltrami algebra and symmetry of the Beltrami equation on Riemann surfaces. Journal of Mathematical Physics, 31(11):2543–2547, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i11/p2543_s1.

Celeghini:1990:TDQ

[370] Enrico Celeghini, Riccardo Giachetti, Emanuele Sorace, and Marco Tarlini. Three-dimensional quantum groups from contractions of SU(2)_{q}. Journal of Mathematical Physics, 31(11):2548–2551, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/
p2548_s1.

Gould:1990:CASb

[371] M. D. Gould and R. B. Zhang. Classification of all star irreps of gl(m|n). Journal of Mathematical Physics, 31(11):2552–2559, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/
p2552_s1.

Cieciura:1990:AGC

[372] Grzegorz Cieciura and Igor Szczyrba. Automorphisms and general charge conjugations. Journal of Mathematical Physics, 31(11):2560–2566, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2560_s1.

Wei:1990:FI

[373] Gaoyuan Wei and B. E. Eichinger. On a Fourier integral over SO(3). Journal of Mathematical Physics, 31(11):2567–2569, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/
p2567_s1.

Wei:1990:PIV

[374] Gaoyuan Wei. A parametrization for independent variables of a twomatrix hypergeometric function. Journal of Mathematical Physics, 31 (11):2570–2571, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i11/p2570_s1.

Grammaticos:1990:SIB

[375] B. Grammaticos, A. Ramani, and J. Hietarinta. A search for integrable bilinear equations: The Painlevé approach. Journal of Mathematical Physics, 31(11):2572–2578, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i11/p2572_s1.

Boettcher:1990:NSW

[376] Stefan Boettcher and Carl M. Bender. Nonperturbative square-well approximation to a quantum theory. Journal of Mathematical Physics, 31 (11):2579–2585, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i11/p2579_s1.

Wu:1990:QMP

[377] Yihren Wu. On the quantization of the modified Pöschl–Teller potential. Journal of Mathematical Physics, 31(11):2586–2587, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i11/p2586_s1.

Danos:1990:AWV

[378] Michael Danos. Accuracy-weighted variational principle for degenerate continuum states. Journal of Mathematical Physics, 31(11):2588–2591, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2588_s1.

Lewis:1990:PCC

[379] H. Ralph Lewis and Paul M. Bellan. Physical constraints on the coefficients of Fourier expansions in cylindrical coordinates. Journal of Mathematical Physics, 31(11):2592–2596, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i11/p2592_s1.

Kingston:1990:BTG

[380] J. G. Kingston and C. Sophocleous. Bäcklund transformations for generalized nonlinear Schrödinger equations. Journal of Mathematical Physics, 31(11):2597–2602, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i11/p2597_s1.

Das:1990:ACI

[381] Ashok Das and Wen-Jui Huang. An alternate characterization of integrability. Journal of Mathematical Physics, 31(11):2603–2605, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i11/p2603_s1.

Gumral:1990:MHS

[382] H. Gümral and Y. Nutku. Multi-Hamiltonian structure of equations of hydrodynamic type. Journal of Mathematical Physics, 31(11):2606– 2611, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2606_s1.

Boiti:1990:MST

[383] M. Boiti, J. P. Leon, and F. Pempinelli. Multidimensional solitons and their spectral transforms. Journal of Mathematical Physics, 31(11):2612– 2618, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2612_s1.

Phanord:1990:MSE

[384] D. D. Phanord and N. E. Berger. Multiple scattering of elastic waves by a distribution of identical spheres. Journal of Mathematical Physics, 31(11):2619–2625, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i11/p2619_s1.

Phanord:1990:RRE

[385] D. D. Phanord and N. E. Berger. Reciprocity relations and energy theorems for multiple scattering of elastic waves. Journal of Mathematical Physics, 31(11):2626–2632, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i11/p2626_s1.

Castro:1990:NCS

[386] Carlos Castro. Nonlinear corrections to the Schrödinger equation from geometric quantum mechanics. Journal of Mathematical Physics, 31 (11):2633–2636, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i11/p2633_s1.

Popova:1990:QMH

[387] Alla D. Popova. Quantum mechanics with higher time derivatives: Some mathematical facts and the inclusion of gravitational self-interaction. Journal of Mathematical Physics, 31(11):2637–2652, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/
p2637_s1.

Arai:1990:AAA

[388] Asao Arai. An asymptotic analysis and its application to the nonrelativistic limit of the Pauli–Fierz and a spin-boson model. Journal of Mathematical Physics, 31(11):2653–2663, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i11/p2653_s1.

Figueroa:1990:MQC

[389] Héctor Figueroa, José M. Gracia-Bondía, and Joseph C. Várilly. Moyal quantization with compact symmetry groups and noncommutative harmonic analysis. Journal of Mathematical Physics, 31(11):2664–2671, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2664_s1.

Alcalde:1990:SQC

[390] Carlos Alcalde. Star quantization on the cylinder. Journal of Mathematical Physics, 31(11):2672–2682, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i11/p2672_s1.

Goetz:1990:GFP

[391] Guenter Goetz. The gravitational field of plane symmetric thick domain walls. Journal of Mathematical Physics, 31(11):2683–2687, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i11/p2683_s1.

Ove:1990:AFE

[392] Roger Ove. Anomalous foliations of Einstein space–times. Journal of Mathematical Physics, 31(11):2688–2693, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/p2688_s1.

Halilsoy:1990:LFC

[393] Mustafa Halilsoy. Large family of colliding waves in the Einstein–Maxwell theory. Journal of Mathematical Physics, 31(11):2694–2698, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i11/p2694_s1.

Cognola:1990:EAV

[394] Guido Cognola, Paola Giacconi, and Roberto Soldati. Effective actions with vector-axial-vector couplings on Riemann–Cartan manifolds. Journal of Mathematical Physics, 31(11):2699–2701, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/
p2699_s1.

Villalba:1990:SVEb

[395] Víctor M. Villalba. Separation of variables and exact solution to the Dirac equation in curvilinear orthogonal coordinates with cylindrical symmetry. Journal of Mathematical Physics, 31(11):2702–2707, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/
p2702_s1.

Batalin:1990:CVG

[396] I. A. Batalin, P. M. Lavrov, and I. V. Tyutin. An Sp(2)-covariant version of generalized canonical quantization of dynamical systems with linearly dependent constraints. Journal of Mathematical Physics, 31(11):2708– 2717, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2708_s1.

Teh:1990:SCY

[397] Rosy Teh. Solutions of the classical SU(2) Yang–Mills theory in (2 + 1) dimensions with the Chern–Simons term: Ansatz building. Journal of Mathematical Physics, 31(11):2718–2721, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/p2718_s1.

Bender:1990:EQF

[398] Carl M. Bender, Kimball A. Milton, Stephen S. Pinsky, and L. M. Simmons, Jr. δ-expansion for a quantum field theory in the nonperturbative regime. Journal of Mathematical Physics, 31(11):2722–2725, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i11/p2722_s1.

Kobayashi:1990:CFI

[399] Yuji Kobayashi and Shigeaki Nagamachi. Characteristic functions and invariants of supermatrices. Journal of Mathematical Physics, 31(11):2726– 2730, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i11/p2726_s1.

Kvitsinsky:1990:MJF

[400] Andrei A. Kvitsinsky and Vadim V. Kostrykin. S-matrix and Jost functions of Schrödinger Hamiltonian related to the Stark effect. Journal of Mathematical Physics, 31(11):2731–2736, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i11/p2731_s1.

Konopelchenko:1990:IST

[401] B. G. Konopel′chenko and B. T. Matkarimov. Inverse spectral transform for the Ishimori equation. I. Initial value problem. Journal of Mathematical Physics, 31(11):2737–2746, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i11/p2737_s1.

Frank:1990:HSP

[402] Alejandro Frank, François Leyvraz, and Kurt Bernardo Wolf. Hidden symmetry and potential group of the Maxwell fish-eye. Journal of Mathematical Physics, 31(12):2757–2768, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i12/p2757_s1.

Groza:1990:CGC

[403] V. A. Groza, I. I. Kachurik, and A. U. Klimyk. On Clebsch–Gordan coefficients and matrix elements of representations of the quantum algebra U_{q}(su_{2}). Journal of Mathematical Physics, 31(12):2769–2780, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2769_s1.

Hecht:1990:VCS

[404] K. T. Hecht and L. C. Biedenharn. Vector coherent state constructions of U(3) symmetric tensors and their SU(3) ⊃ SU(2) × U(1) Wigner coefficients. Journal of Mathematical Physics, 31(12):2781–2796, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2781_s1.

Fu:1990:NIB

[405] Hong Chen Fu and Chang Pu Sun. New inhomogeneous boson realizations and inhomogeneous differential realizations of Lie algebras. Journal of Mathematical Physics, 31(12):2797–2802, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i12/p2797_s1.

Gould:1990:BRC

[406] M. D. Gould, P. D. Jarvis, and A. J. Bracken. Branching rules for a class of typical and atypical representations of gl(m|n). Journal of Mathematical Physics, 31(12):2803–2810, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i12/p2803_s1.

Negro:1990:LRKb

[407] Javier Negro and Mariano A. del Olmo. Local realizations of kinematical groups with a constant electromagnetic field. II. The nonrelativistic case. Journal of Mathematical Physics, 31(12):2811–2821, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2811_s1.

Tripathy:1990:CTD

[408] K. C. Tripathy and M. K. Patra. Cohomology theory and deformations of Z_{2}-graded Lie algebras. Journal of Mathematical Physics, 31(12):2822– 2831, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p2822_s1.

Rudaz:1990:NAS

[409] S. Rudaz. Note on asymptotic series expansions for the derivative of the Hurwitz zeta function and related functions. Journal of Mathematical Physics, 31(12):2832–2834, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i12/p2832_s1.

Zeng:1990:IHS

[410] Yun Bo Zeng and Yi Shen Li. Integrable Hamiltonian systems related to the polynomial eigenvalue problem. Journal of Mathematical Physics, 31(12):2835–2839, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i12/p2835_s1.

Lucquiaud:1990:ICT

[411] J. C. Lucquiaud. Introduction to a covariant theory of special functions of mathematical physics. Journal of Mathematical Physics, 31(12):2840– 2850, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p2840_s1.

Chen:1990:CIA

[412] Zhi xiong Chen, Ben yu (Pen-yu) Guo, and Long wan Xiang. Complete integrability and analytic solutions of a KdV-type equation. Journal of Mathematical Physics, 31(12):2851–2855, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i12/p2851_s1.

Dattoli:1990:LAM

[413] G. Dattoli, M. Richetta, G. Schettini, and A. Torre. Lie algebraic methods and solutions of linear partial differential equations. Journal of Mathematical Physics, 31(12):2856–2863, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i12/p2856_s1.

Kafkalidis:1990:GAP

[414] Menelaos S. Kafkalidis. A geometric approach to the path integral formalism of p-branes. Journal of Mathematical Physics, 31(12):2864–2871, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p2864_s1.

Palmer:1990:BTS

[415] Bennett Palmer. Bäcklund transformations for surfaces in Minkowski space. Journal of Mathematical Physics, 31(12):2872–2875, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2872_s1.

Rylov:1990:EPS

[416] Yu. A. Rylov. Extremal properties of Synge’s world function and discrete geometry. Journal of Mathematical Physics, 31(12):2876–2890, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2876_s1.

Cirelli:1990:QMIa

[417] Renzo Cirelli, Alessandro Manià, and Livio Pizzocchero. Quantum mechanics as an infinite-dimensional Hamiltonian system with uncertainty structure. I. Journal of Mathematical Physics, 31(12):2891–2897, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2891_s1.

Cirelli:1990:QMIb

[418] Renzo Cirelli, Alessandro Manià, and Livio Pizzocchero. Quantum mechanics as an infinite-dimensional Hamiltonian system with uncertainty structure. II. Journal of Mathematical Physics, 31(12):2898–2903, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p2898_s1.

Matsuno:1990:BTC

[419] Y. Matsuno. Bäcklund transformation, conservation laws, and inverse scattering transform of a model integrodifferential equation for water waves. Journal of Mathematical Physics, 31(12):2904–2916, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2904_s1.

Das:1990:EPW

[420] A. Das, T. Biech, and D. Kay. Exact plane-wave solutions of the coupled Maxwell–Klein–Gordon equations. Journal of Mathematical Physics, 31 (12):2917–2920, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i12/p2917_s1.

Ekhaguere:1990:FIF

[421] G. O. S. Ekhaguere. The functional Ito formula in quantum stochastic calculus. Journal of Mathematical Physics, 31(12):2921–2929, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2921_s1.

Belavkin:1990:PSE

[422] V. P. Belavkin. A posterior Schrödinger equation for continuous nondemolition measurement. Journal of Mathematical Physics, 31(12):2930– 2934, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p2930_s1.

Marchioro:1990:CCS

[423] T. L. Marchioro II. Corrections to the coherent state path integral: Comments upon a speculation of L. S. Schulman. Journal of Mathematical Physics, 31(12):2935–2945, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i12/p2935_s1.

Crescimanno:1990:ESJ

[424] M. Crescimanno. Exactness of the supersymmetric JWKB quantization formula. Journal of Mathematical Physics, 31(12):2946–2951, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2946_s1.

Littlejohn:1990:SST

[425] Robert G. Littlejohn. Semiclassical structure of trace formulas. Journal of Mathematical Physics, 31(12):2952–2977, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i12/p2952_s1.

Moncrief:1990:HSD

[426] Vincent Moncrief. How solvable is (2 + 1)-dimensional Einstein gravity? Journal of Mathematical Physics, 31(12):2978–2982, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2978_s1.

Torre:1990:LSC

[427] C. G. Torre. On the linearization stability of the conformally (anti-) selfdual Einstein equations. Journal of Mathematical Physics, 31(12):2983– 2986, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p2983_s1.

Aldrovandi:1990:SQP

[428] R. Aldrovandi and D. Galetti. On the structure of quantum phase space. Journal of Mathematical Physics, 31(12):2987–2995, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2987_s1.

Voigt:1990:BCC

[429] K. Voigt. On boson condensation considering a generalized Casimir example. Journal of Mathematical Physics, 31(12):2996–3002, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p2996_s1.

Osland:1990:CSM

[430] Per Osland and Tai Tsun Wu. Completely solvable models of the nonlinear Boltzmann equation. I. Case of three velocities. Journal of Mathematical Physics, 31(12):3003–3029, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i12/p3003_s1.

Gaveau:1990:EPM

[431] B. Gaveau and L. S. Schulman. Equilibrium properties of a multi-urn model. Journal of Mathematical Physics, 31(12):3030–3040, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i12/p3030_s1.

Ferreira:1990:SHH

[432] L. A. Ferreira and R. M. Londe. Solutions to higher Hamiltonians in the Toda hierarchies. Journal of Mathematical Physics, 31(12):3041– 3046, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i12/p3041_s1.

Verwimp:1990:UPG

[433] Theo Verwimp. Unified prescription for the generation of electroweak and gravitational gauge field Lagrangian on a principal fiber bundle. Journal of Mathematical Physics, 31(12):3047–3052, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i12/p3047_s1.

Naito:1990:CMV

[434] Seichi Naito and Masahiro Nishimoto. Conformal mapping and vertex operators in the light-cone gauge. Journal of Mathematical Physics, 31 (12):3053–3059, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i12/p3053_s1.

Oshima:1990:CRM

[435] Kazuto Oshima. Completeness relations for Maass Laplacians and heat kernels on the super Poincaré upper half-plane. Journal of Mathematical Physics, 31(12):3060–3063, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v31/i12/p3060_s1.

Yurtsever:1990:TFC

[436] Ulvi Yurtsever. Test fields on compact space–times. Journal of Mathematical Physics, 31(12):3064–3078, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v31/i12/p3064_s1.

Ma:1990:NLP

[437] Zhong Qi Ma. New link polynomial obtained from octet representation of quantum sl(3) enveloping algebra. Journal of Mathematical Physics, 31(12):3079–3084, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v31/i12/p3079_s1.

Hillion:1990:GPM

[438] Pierre Hillion. The Goursat problem for Maxwell’s equations. Journal of Mathematical Physics, 31(12):3085–3088, December 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v31/i12/p3085_s1.

Lim:1991:GTS

[439] Chjan C. Lim. Graph theory and a special class of symplectic transformations: The generalized Jacobi variables. Journal of Mathematical Physics, 32(1):1–7, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p1_s1.

Broadbridge:1991:IHN

[440] P. Broadbridge and S. E. Godfrey. Integrable heterogeneous nonlinear Schrödinger equations with dielectric loss: Lie–Bäcklund symmetries. Journal of Mathematical Physics, 32(1):8–18, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i1/p8_s1.

Haba:1991:RKM

[441] Z. Haba. Representations of the SL(2, C) Kac–Moody current algebra in the Fock space. Journal of Mathematical Physics, 32(1):19–23, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p19_s1.

Pestov:1991:ESL

[442] Vladimir Pestov. Even sectors of Lie superalgebras as locally convex Lie algebras. Journal of Mathematical Physics, 32(1):24–32, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p24_s1.

Grigore:1991:CLS

[443] D. R. Grigore and O. T. Popp. On the classification of Lie subalgebras and Lie subgroups. Journal of Mathematical Physics, 32(1):33– 39, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p33_s1.

Konopelchenko:1991:SPS

[444] Boris Konopel′chenko and Walter Strampp. On the structure and properties of the singularity manifold equations of the KP hierarchy. Journal of Mathematical Physics, 32(1):40–49, January 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i1/p40_s1.

Kvitsinsky:1991:TBP

[445] A. A. Kvitsinsky and S. P. Merkuriev. Three-body plane wave at zero angular momentum and some addition theorems. Journal of Mathematical Physics, 32(1):50–56, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p50_s1.

Lawton:1991:NSC

[446] Wayne M. Lawton. Necessary and sufficient conditions for constructing orthonormal wavelet bases. Journal of Mathematical Physics, 32(1):57– 61, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p57_s1.

Morita:1991:EST

[447] T. Morita. An extension of Szegő’s theorem. Journal of Mathematical Physics, 32(1):62–68, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p62_s1.

Brugarino:1991:PAR

[448] Tommaso Brugarino and Antonio M. Greco. Painlevé analysis and reducibility to the canonical form for the generalized Kadomtsev– Petviashvili equation. Journal of Mathematical Physics, 32(1):69–71, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p69_s1.

Calogero:1991:LPG

[449] F. Calogero and M. C. Nucci. Lax pairs galore. Journal of Mathematical Physics, 32(1):72–74, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p72_s1.

Lakshmanan:1991:GLS

[450] M. Lakshmanan and R. Sahadevan. Generalized Lie symmetries and complete integrability of certain nonlinear Hamiltonian systems with three degrees of freedom. Journal of Mathematical Physics, 32(1):75– 83, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p75_s1.

Harnett:1991:MDO

[451] Gerald Harnett. Metrics and dual operators. Journal of Mathematical Physics, 32(1):84–91, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p84_s1.

Kishida:1991:CIS

[452] Kuniharu Kishida. Contraction of information in systems far from equilibrium. Journal of Mathematical Physics, 32(1):92–98, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i1/
p92_s1.

Calogero:1991:BES

[453] F. Calogero and S. De Lillo. The Burgers equation on the semiline with general boundary conditions at the origin. Journal of Mathematical Physics, 32(1):99–105, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i1/p99_s1.

Bombelli:1991:WFW

[454] L. Bombelli, W. E. Couch, and R. J. Torrence. Wake-free waves in one and three dimensions. Journal of Mathematical Physics, 32(1):106– 108, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p106_s1.

Cooke:1991:CRD

[455] David B. Cooke. Classification results and the Darboux theorem for low-order Hamiltonian operators. Journal of Mathematical Physics, 32(1):109–119, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p109_s1.

Matsuno:1991:LNN

[456] Y. Matsuno. Linearization of novel nonlinear diffusion equations with the Hilbert kernel and their exact solutions. Journal of Mathematical Physics, 32(1):120–126, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i1/p120_s1.

Damour:1991:RPV

[457] Thibault Damour and Gerhard Schäfer. Redefinition of position variables and the reduction of higher-order Lagrangians. Journal of Mathematical Physics, 32(1):127–134, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i1/p127_s1.

Ellinas:1991:POG

[458] D. Ellinas. Phase operators via group contraction. Journal of Mathematical Physics, 32(1):135–141, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i1/p135_s1.

Hojman:1991:NLN

[459] Sergio A. Hojman and L. C. Shepley. No Lagrangian? No quantization. Journal of Mathematical Physics, 32(1):142–146, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i1/p142_s1.

Zelenov:1991:API

[460] E. I. Zelenov. p-adic path integrals. Journal of Mathematical Physics, 32(1):147–152, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p147_s1.

Kuperin:1991:TBQ

[461] Yu. A. Kuperin, S. P. Merkuriev, and E. A. Yarevsky. Three-body quantum system on a line: Wave function asymptotics for increasing interactions. Journal of Mathematical Physics, 32(1):153–156, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p153_s1.

Cheng:1991:ISA

[462] Y. Cheng. Integrable systems associated with the Schrödinger spectral problem in the plane. Journal of Mathematical Physics, 32(1):157– 162, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p157_s1.

Klaus:1991:ABJ

[463] Martin Klaus. Asymptotic behavior of Jost functions near resonance points for Wigner–von Neumann type potentials. Journal of Mathematical Physics, 32(1):163–174, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i1/p163_s1.

Svaiter:1991:CED

[464] N. F. Svaiter and B. F. Svaiter. Casimir effect in a D-dimensional flat space–time and the cut-off method. Journal of Mathematical Physics, 32(1):175–180, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p175_s1.

Hall:1991:CCT

[465] G. S. Hall. Covariantly constant tensors and holonomy structure in general relativity. Journal of Mathematical Physics, 32(1):181–187, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p181_s1.

Bonanos:1991:VEE

[466] S. Bonanos. The vacuum Einstein equations as closed ideals of matrixvalued differential forms. Journal of Mathematical Physics, 32(1):188– 192, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p188_s1.

Bonanos:1991:ICI

[467] S. Bonanos. Integrability conditions and the integration of the Bianchi identities in Petrov type-I vacuum space–times. Journal of Mathematical Physics, 32(1):193–197, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i1/p193_s1.

Hauser:1991:IVP

[468] Isidore Hauser and Frederick J. Ernst. Initial value problem for colliding gravitational plane waves. IV. Journal of Mathematical Physics, 32(1):198–209, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p198_s1.

Keszthelyi:1991:SSD

[469] B. Keszthelyi, E. Piard, and P. Ramond. Study of some dual solutions in 2 + 1 dimensions. Journal of Mathematical Physics, 32(1):210–222, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p210_s1.

Sussman:1991:SSS

[470] Roberto A. Sussman. Spherically symmetric solutions admitting a spacelike self-similar motion. Journal of Mathematical Physics, 32(1):223– 230, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p223_s1.

Wils:1991:CES

[471] Patrick Wils. A class of exact solutions of the Einstein–Dirac equations. Journal of Mathematical Physics, 32(1):231–233, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i1/p231_s1.

Hojman:1991:ACR

[472] Sergio A. Hojman and Darío Núñez. Affine collineations in Riemannian spaces. Journal of Mathematical Physics, 32(1):234–238, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p234_s1.

Koehler:1991:ESC

[473] K. Koehler, F. Mansouri, Cenalo Vaz, and L. Witten. Extended supergravity: Chern–Simons theories in 2 + 1 dimensions. Journal of Mathematical Physics, 32(1):239–246, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i1/p239_s1.

Zinoun:1991:PSD

[474] A. Zinoun and J. Cortois. Pertinent statistics for describing a system of particles. Journal of Mathematical Physics, 32(1):247–249, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p247_s1.

Doebner:1991:QLC

[475] H. D. Doebner and W. Lücke. Quantum logic as a consequence of realistic measurements on deterministic systems. Journal of Mathematical Physics, 32(1):250–253, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i1/p250_s1.

Bolsterli:1991:CTM

[476] M. Bolsterli. Canonical transformation method for static-source meson Hamiltonians. Journal of Mathematical Physics, 32(1):254–258, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p254_s1.

Fioravanti:1991:TRY

[477] Mario A. Fioravanti and John R. Porter. Twistor representation of Yang–Mills fields on Minkowski space. Journal of Mathematical Physics, 32(1):259–265, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i1/p259_s1.

Short:1991:PSV

[478] K. M. Short. Polyakov string in variable dimensions. Journal of Mathematical Physics, 32(1):266–279, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i1/p266_s1.

Short:1991:NGF

[479] Kevin M. Short. New geometric formulation of string theory. Journal of Mathematical Physics, 32(1):280–301, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i1/p280_s1.

Bhansali:1991:IRS

[480] Vineer Bhansali. Induced representations for SIM(d − 1, 1) and scaling of massless fields in even-dimensional spaces. Journal of Mathematical Physics, 32(1):302–304, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i1/p302_s1.

Legare:1991:SRW

[481] M. Légaré. Symmetry reduction of the Wess–Zumino model and supersymmetric reduced systems. Journal of Mathematical Physics, 32(1):305– 309, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i1/p305_s1.

Tezcan:1991:OSA

[482] C. Tezcan. The one-speed albedo and the slab albedo problems for the FBIS kernel. Journal of Mathematical Physics, 32(1):310–313, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i1/p310_s1.

Zak:1991:BSE

[483] J. Zak. Biorthogonal sets for energy bands in solids. Journal of Mathematical Physics, 32(1):314–317, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i1/p314_s1.

Inoue:1991:EUL

[484] A. Inoue. Extension of unbounded left Hilbert algebras to partial ^{∗}- algebras. Journal of Mathematical Physics, 32(2):323–331, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i2/p323_s1.

Madore:1991:CLM

[485] J. Madore. The commutative limit of a matrix geometry. Journal of Mathematical Physics, 32(2):332–335, February 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i2/p332_s1.

Guhr:1991:DCF

[486] Thomas Guhr. Dyson’s correlation functions and graded symmetry. Journal of Mathematical Physics, 32(2):336–347, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/p336_s1.

Krause:1991:NAG

[487] J. Krause. Non-Abelian group quantization and quantum kinematic invariants of some noncompact Lie groups. Journal of Mathematical Physics, 32(2):348–355, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p348_s1.

Morrison:1991:IBM

[488] I. Morrison, P. W. Pieruschka, and B. G. Wybourne. The interacting boson model with the exceptional groups G_{2} and E_{6}. Journal of Mathematical Physics, 32(2):356–372, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i2/p356_s1.

Moshinsky:1991:ABR

[489] M. Moshinsky, G. Loyola, and C. Villegas. Anomalous basis for representations of the Poincaré group. Journal of Mathematical Physics, 32(2):373–381, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i2/p373_s1.

Nakamura:1991:SDY

[490] Yoshimasa Nakamura. Self-dual Yang–Mills hierarchy: A nonlinear integrable dynamical system in higher dimensions. Journal of Mathematical Physics, 32(2):382–385, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i2/p382_s1.

Pan:1991:MRB

[491] Feng Pan. The matrix representation of SO_{7} in an (SU_{2})^{3} basis. Journal of Mathematical Physics, 32(2):386–391, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:1991:MRS

[492] Feng Pan. The matrix representation of SO_{7} in an (SU_{2})^{3} basis. Journal of Mathematical Physics, 32(2):386–391, February 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i2/p386_s1.

Sue:1991:ISD

[493] Michael Sué. Involutive systems of differential equations: Einstein’s strength versus Cartan’s degré d’arbitraire. Journal of Mathematical Physics, 32(2):392–399, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p392_s1.

Suzuki:1991:GTF

[494] Masuo Suzuki. General theory of fractal path integrals with applications to many-body theories and statistical physics. Journal of Mathematical Physics, 32(2):400–407, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p400_s1.

Channell:1991:SSM

[495] Paul J. Channell. Stable symplectic maps. Journal of Mathematical Physics, 32(2):408–412, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p408_s1.

Swart:1991:SME

[496] J. H. Swart. The stability of Maxwell’s equations associated with an initial-boundary value problem in quarter space. Journal of Mathematical Physics, 32(2):413–418, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p413_s1.

Oteo:1991:BCH

[497] J. A. Oteo. The Baker–Campbell–Hausdorff formula and nested commutator identities. Journal of Mathematical Physics, 32(2):419–424, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i2/p419_s1.

vanWyk:1991:LOR

[498] C. B. van Wyk. The Lorentz operator revisited. Journal of Mathematical Physics, 32(2):425–430, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p425_s1.

Newton:1991:WIS

[499] Paul K. Newton. Wave interactions in the singular Zakharov system. Journal of Mathematical Physics, 32(2):431–440, February 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/p431_
s1.

Boudjedaa:1991:PIT

[500] T. Boudjedaa, L. Chétouani, L. Guéchi, and T. F. Hammann. Path integral treatment for a screened potential. Journal of Mathematical Physics, 32(2):441–446, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p441_s1.

Adhikari:1991:ESN

[501] R. Adhikari, R. Dutt, and Y. P. Varshni. Exact solutions for nonpolynomial potentials in N-space dimensions using a factorization method and supersymmetry. Journal of Mathematical Physics, 32(2):447–456, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i2/p447_s1.

Nakamura:1991:NRF

[502] Touru Nakamura. A nonstandard representation of Feynman’s path integrals. Journal of Mathematical Physics, 32(2):457–463, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i2/p457_s1.

Chen:1991:CGI

[503] Robert L. W. Chen. Concerning the gauge invariance and the apparent nonlocality of the real form of Schrödinger’s equation. Journal of Mathematical Physics, 32(2):464–465, February 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i2/p464_s1.

Caliceti:1991:QDC

[504] Emanuela Caliceti. Quantum degeneracy and classical perturbation theory. Journal of Mathematical Physics, 32(2):466–469, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/
p466_s1.

Blood:1991:USC

[505] F. A. Blood, Jr. Unitary spin crystals with broken symmetry eigenvectors and spin wave-mediated interactions. Journal of Mathematical Physics, 32(2):470–475, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p470_s1.

Dunn:1991:KNG

[506] K. A. Dunn, T. A. Harriott, and J. G. Williams. Kink number in general relativity. Journal of Mathematical Physics, 32(2):476–479, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i2/p476_s1.

Mendoza:1991:IBT

[507] A. Mendoza and A. Restuccia. Infinitesimal Bäcklund transformations and Kac–Moody algebras in general relativity. Journal of Mathematical Physics, 32(2):480–484, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p480_s1.

Fennelly:1991:ISR

[508] A. J. Fennelly, Jean P. Krisch, John R. Ray, and Larry L. Smalley. Including spin in the Raychaudhuri equation. Journal of Mathematical Physics, 32(2):485–487, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p485_s1.

Tippett:1991:BWT

[509] Michael K. Tippett and Richard W. Ziolkowski. A bidirectional wave transformation of the cold plasma equations. Journal of Mathematical Physics, 32(2):488–492, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p488_s1.

Mason:1991:TVR

[510] D. P. Mason and A. M. Kgathi. Thermodynamic variables of a relativistic gas in collision-dominated equilibrium. Journal of Mathematical Physics, 32(2):493–499, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i2/p493_s1.

Ranada:1991:ETB

[511] Manuel F. Rañada. Extended tangent bundle formalism for timedependent Lagrangian systems. Journal of Mathematical Physics, 32 (2):500–505, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i2/p500_s1.

Kajiwara:1991:CQS

[512] Kenji Kajiwara and Junkichi Satsuma. The conserved quantities and symmetries of the two-dimensional Toda lattice hierarchy. Journal of Mathematical Physics, 32(2):506–514, February 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i2/p506_s1.

Reisz:1991:FMG

[513] T. Reisz. Functional measure for gauge fields on periodic lattices. Journal of Mathematical Physics, 32(2):515–523, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/p515_s1.

Flato:1991:NAS

[514] M. Flato and C. Fronsdal. Non-Abelian singletons. Journal of Mathematical Physics, 32(2):524–531, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i2/p524_s1.

Batalin:1991:CQG

[515] I. A. Batalin, P. M. Lavrov, and I. V. Tyutin. An Sp(2)-covariant quantization of gauge theories with linearly dependent generators. Journal of Mathematical Physics, 32(2):532–539, February 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i2/p532_s1.

Wotzasek:1991:DWZ

[516] C. Wotzasek. Derivation of the Wess–Zumino term for the chiral Schwinger model using Dirac bracket formalism. Journal of Mathematical Physics, 32(2):540–543, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i2/p540_s1.

Anile:1991:TAE

[517] A. M. Anile, S. Pennisi, and M. Sammartino. A thermodynamical approach to Eddington factors. Journal of Mathematical Physics, 32 (2):544–550, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i2/p544_s1.

Newton:1991:BWS

[518] Roger G. Newton. Bloch-wave scattering by crystal defects. Journal of Mathematical Physics, 32(2):551–560, February 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i2/p551_s1.

Audit:1991:PCP

[519] Philippe Audit. Properties of cyclic permutation matrices on Potts spin spaces and associated arithmetical polynomials. Journal of Mathematical Physics, 32(2):561–563, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i2/p561_s1.

Grigore:1991:ECC

[520] D. R. Grigore and O. T. Popp. Erratum: “The complete classification of generalized homogeneous symplectic manifolds”, [J. Math. Phys. 30 (1989), no. 11, 2476–2483; MR 90i:58052]. Journal of Mathematical Physics, 32(2):564, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i2/p564_s2. See [15].

Srivastava:1991:EAP

[521] Sunita Srivastava and Vishwamittar. Erratum: “Application of perturbation theory to the damped sextic oscillator”, [J. Math. Phys. 30, 2815 (1989)]. Journal of Mathematical Physics, 32(2):564, February 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/p564_

s1. See [16].

Alisauskas:1991:OCE

[522] Sigitas Ališauskas. Orthogonalization coefficients for the Elliott–Draayer states of the two parametric irreps of SU(n) ⊃ SO(n). Journal of Mathematical Physics, 32(3):569–575, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i3/p569_s1.

Crawford:1991:CAN

[523] J. P. Crawford. Clifford algebra: notes on the spinor metric and Lorentz, Poincaré, and conformal groups. Journal of Mathematical Physics, 32(3): 576–583, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p576_s1.

Fan:1991:NOB

[524] Hong yi Fan and Hai Guang Weng. Normally ordered Bose–Fermi operator realization of the SU(N/M) group. Journal of Mathematical Physics, 32(3):584–587, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i3/p584_s1.

Minnaert:1991:SBT

[525] Pierre Minnaert and Marek Mozrzymas. On the superalgebras built from SO(3) tensor operators. Journal of Mathematical Physics, 32(3):588– 594, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p588_s1.

Sun:1991:ABA

[526] Chang-Pu Sun and Mo-Lin Ge. The q-analog of the boson algebra, its representation on the Fock space, and applications to the quantum group. Journal of Mathematical Physics, 32(3):595–598, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i3/p595_s1.

Gunaydin:1991:ULW

[527] M. Günaydin and R. J. Scalise. Unitary lowest weight representations of the noncompact supergroup OSp(2m ∗ /2n). Journal of Mathematical Physics, 32(3):599–606, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i3/p599_s1.

Isham:1991:CSD

[528] C. J. Isham and J. R. Klauder. Coherent states for n-dimensional Euclidean groups E_{(}n) and their application. Journal of Mathematical Physics, 32(3):607–620, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i3/p607_s1.

Benamor:1991:ERL

[529] Hédi Benamor and Georges Pinczon. Extensions of representations of Lie superalgebras. Journal of Mathematical Physics, 32(3):621–629, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i3/p621_s1.

vanDiejen:1991:SGP

[530] J. F. van Diejen and A. Tip. Scattering from generalized point interactions using self-adjoint extensions in Pontryagin spaces. Journal of Mathematical Physics, 32(3):630–641, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i3/p630_s1.

Maximon:1991:EIP

[531] Leonard C. Maximon. On the evaluation of the integral over the product of two spherical Bessel functions. Journal of Mathematical Physics, 32(3): 642–648, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p642_s1.

Raynal:1991:CRB

[532] Jacques Raynal. Comments on: “Relations between hyperspherical harmonic transformations and generalized Talmi–Moshinsky transformations”, [J. Math. Phys. 31 (1990), no. 7, 1621–1626; MR 91h:33006] by X. H. Liu. Journal of Mathematical Physics, 32(3):649–650, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i3/p649_s1. See [241].

Robinson:1991:HFA

[533] Sam L. Robinson. High-frequency asymptotic solution of the wave equation in an inhomogeneous medium. Journal of Mathematical Physics, 32(3):651–655, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i3/p651_s1.

Gudder:1991:RQM

[534] S. Gudder and C. Schindler. Regular quantum Markov processes. Journal of Mathematical Physics, 32(3):656–668, March 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i3/p656_s1.

Inayat-Hussain:1991:MIT

[535] A. A. Inayat-Hussain. Mathieu integral transforms. Journal of Mathematical Physics, 32(3):669–675, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i3/p669_s1.

Curtright:1991:QAD

[536] T. L. Curtright, G. I. Ghandour, and C. K. Zachos. Quantum algebra deforming maps, Clebsch–Gordan coefficients, coproducts, R and U matrices. Journal of Mathematical Physics, 32(3):676–688, March 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i3/p676_
s1.

Verbeure:1991:DFF

[537] A. Verbeure. Dynamics of field fluctuations in the Schwinger model. Journal of Mathematical Physics, 32(3):689–693, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i3/p689_s1.

Shahid-Saless:1991:PVC

[538] Bahman Shahid-Saless. Palatini variation of curvature-squared action and gravitational collapse. Journal of Mathematical Physics, 32(3):694– 697, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p694_s1.

Kalnins:1991:CSFa

[539] E. G. Kalnins and W. Miller, Jr. Complete sets of functions for perturbations of Robertson–Walker cosmologies and spin 1 equations in Robertson–Walker-type space–times. Journal of Mathematical Physics, 32(3):698–707, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i3/p698_s1.

GarciaD:1991:SAI

[540] Alberto García D. Stationary axisymmetric interior metric for a charged perfect fluid source. Journal of Mathematical Physics, 32(3):708–713, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p708_s1.

Karas:1991:IMK

[541] V. Karas and D. Vokrouhlický. On interpretation of the magnetized Kerr–Newman black hole. Journal of Mathematical Physics, 32(3):714– 716, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p714_s1.

AitMoussa:1991:GMS

[542] Karim Ait Moussa and Gérard Clément. Geodesic motion and scattering in the field of a Kaluza–Klein dyon. Journal of Mathematical Physics, 32(3):717–722, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i3/p717_s1.

Li:1991:CGP

[543] Wei Li, Isidore Hauser, and Frederick J. Ernst. Colliding gravitational plane waves with noncollinear polarizations. Journal of Mathematical Physics, 32(3):723–727, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i3/p723_s1.

Manoff:1991:IPE

[544] S. Manoff. Invariant projection of the energy-momentum tensors for field theories in spaces with affine connection and metric. Journal of Mathematical Physics, 32(3):728–734, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i3/p728_s1.

Ranganathan:1991:CQW

[545] D. Ranganathan. Charge quantization in a Weyl model. Journal of Mathematical Physics, 32(3):735–738, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i3/p735_s1.

Amann:1991:ISS

[546] Anton Amann. Invariant states of C^{∗}-systems without norm-continuity properties. Journal of Mathematical Physics, 32(3):739–743, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i3/
p739_s1.

Chen:1991:DST

[547] M. Chen. Dynamical stability and thermodynamic stability in irreversible thermodynamics. Journal of Mathematical Physics, 32(3):744– 748, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p744_s1.

Tscheuschner:1991:CVN

[548] Ralf D. Tscheuschner. Coinciding versus noncoinciding: Is the topological spin-statistics theorem already proven in quantum mechanics? Journal of Mathematical Physics, 32(3):749–752, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i3/p749_s1.

Schiff:1991:ICS

[549] Jeremy Schiff. Integrability of Chern–Simons–Higgs and Abelian Higgs vortex equations in a background metric. Journal of Mathematical Physics, 32(3):753–761, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i3/p753_s1.

Tanimura:1991:DEE

[550] N. Tanimura and O. Tanimura. Differential equation for SU(3) extension of Bessel function with two indices. Journal of Mathematical Physics, 32(3):762–766, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i3/p762_s1.

Fu:1991:IDR

[551] Hong Chen Fu. Inhomogeneous differential realization, boson-fermion realization of Lie superalgebras and their indecomposable representations. Journal of Mathematical Physics, 32(3):767–775, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i3/p767_s1.

Li:1991:HOS

[552] Shiwei Li and J. K. Percus. High-order K-space linear response of a noninteracting fermion system in an external potential. Journal of Mathematical Physics, 32(3):776–782, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i3/p776_s1.

Saxena:1991:PSS

[553] G. M. Saxena and C. L. Mehta. Para-squeezed states. Journal of Mathematical Physics, 32(3):783–786, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i3/p783_s1.

Eringen:1991:MDN

[554] A. Cemal Eringen. Memory-dependent nonlocal electromagnetic elastic solids and superconductivity. Journal of Mathematical Physics, 32(3): 787–796, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p787_s1.

McLaughlin:1991:SNP

[555] D. McLaughlin and O. Pironneau. Some notes on periodic Beltrami fields in Cartesian geometry. Journal of Mathematical Physics, 32(3): 797–804, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p797_s1.

Malvagi:1991:IBC

[556] F. Malvagi and G. C. Pomraning. Initial and boundary conditions for diffusive linear transport problems. Journal of Mathematical Physics, 32(3): 805–820, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i3/p805_s1.

King:1991:ACR

[557] R. B. King and E. R. Canfield. An algorithm for calculating the roots of a general quintic equation from its coefficients. Journal of Mathematical Physics, 32(4):823–825, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i4/p823_s1.

Coquereaux:1991:RIM

[558] Robert Coquereaux and Gilles Esposito-Farese. Right-invariant metrics on the Lie group SU(3) and the Gell-Mann–Okubo formula. Journal of Mathematical Physics, 32(4):826–831, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i4/p826_s1.

DHoker:1991:NIS

[559] Eric D’Hoker. A new invariant of superconformal OSp(n|1). Journal of Mathematical Physics, 32(4):832–836, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i4/p832_s1.

Gromov:1991:GTR

[560] N. A. Gromov. The Gel′fand–Tsetlin representations of the unitary Cayley–Klein algebras. Journal of Mathematical Physics, 32(4):837– 844, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i4/p837_s1.

Kim:1991:QMG

[561] J. D. Kim, I. G. Koh, and Zhong Qi Ma. Quantum Ř matrix for E_{7} and F_{4} groups. Journal of Mathematical Physics, 32(4):845–856, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:1991:QMM

[562] J. D. Kim, I. G. Koh, and Zhong-Qi Ma. Quantum R̂ matrix for E_{7} and F_{4} groups. Journal of Mathematical Physics, 32(4):845–856, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p845_s1.

Wolski:1991:APDa

[563] A. Wolski and J. S. Dowker. Area-preserving diffeomorphisms of the tetrahedron. Journal of Mathematical Physics, 32(4):857–863, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p857_s1.

Antoniou:1991:CSS

[564] Ioannis E. Antoniou and Baidyanath Misra. Characterization of semidirect sum Lie algebras. Journal of Mathematical Physics, 32(4):864–868, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p864_s1.

Das:1991:CLO

[565] Ashok Das and Shibaji Roy. Covariant Lax operators and Kac–Moody algebras. Journal of Mathematical Physics, 32(4):869–874, April 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i4/p869_
s1.

Calogero:1991:INPa

[566] F. Calogero and Ji Xiaoda. C-integrable nonlinear partial differentiation equations. I. Journal of Mathematical Physics, 32(4):875–887, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p875_s1.

McInnes:1991:MHT

[567] Brett McInnes. Methods of holonomy theory for Ricci-flat Riemannian manifolds. Journal of Mathematical Physics, 32(4):888–896, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p888_s1.

Tamura:1991:PCC

[568] Hiroshi Tamura. On the possibility of confinement caused by nonlinear electromagnetic interaction. Journal of Mathematical Physics, 32(4):897– 904, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i4/p897_s1.

Chen:1991:NRM

[569] Gen Shun Chen and Anthony W. Leung. Nonlinear reactor multigroup neutron transport system: Existence and stability problems. Journal of Mathematical Physics, 32(4):905–915, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i4/p905_s1.

Alber:1991:AIS

[570] Solomon J. Alber. Associated integrable systems. Journal of Mathematical Physics, 32(4):916–922, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i4/p916_s1.

Labelle:1991:NSK

[571] P. Labelle and P. Mathieu. A new N = 2 supersymmetric Korteweg–de Vries equation. Journal of Mathematical Physics, 32(4):923–927, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p923_s1.

Magnon:1991:SPD

[572] Anne M. R. Magnon. Spin-plane defects and emergence of Planck’s constant in gravity. Journal of Mathematical Physics, 32(4):928–931, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p928_s1.

Khrennikov:1991:AQM

[573] A. Yu. Khrennikov. p-adic quantum mechanics with p-adic valued functions. Journal of Mathematical Physics, 32(4):932–937, April 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i4/p932_
s1.

Saksena:1991:AAQ

[574] A. Saksena, T. A. Osborn, and F. H. Molzahn. An asymptotic analysis of quantum evolution with electromagnetic fields. Journal of Mathematical Physics, 32(4):938–955, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i4/p938_s1.

Herve:1991:MTP

[575] Rose-Marie Hervé and Michel Hervé. A mathematical treatment for a perturbation of the one-dimensional harmonic oscillator. Journal of Mathematical Physics, 32(4):956–958, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i4/p956_s1.

Gompa:1991:TDN

[576] Raghu R. Gompa. Time-dependent normal form theory and Schrödinger initial value problems. Journal of Mathematical Physics, 32(4):959–966, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p959_s1.

Papp:1991:TPS

[577] E. Papp. Transformation properties of the Schrödinger equation with the momentum-dependent potential (x · ∇)_{ord}^{2}. Journal of Mathematical Physics, 32(4):967–970, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i4/p967_s1.

Dashen:1991:NTS

[578] Roger Dashen and Daniel Wurmser. A new theory for scattering from a surface. Journal of Mathematical Physics, 32(4):971–985, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i4/

p971_s1. See erratum [1938].

Dashen:1991:ARS

[579] Roger Dashen and Daniel Wurmser. Approximate representations of the scattering amplitude. Journal of Mathematical Physics, 32(4):986– 996, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i4/p986_s1.

Dashen:1991:ANS

[580] Roger Dashen and Daniel Wurmser. Applications of the new scattering formalism: The Dirichlet boundary condition. Journal of Mathematical Physics, 32(4):997–1003, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i4/p997_s1.

Przanowski:1991:KVF

[581] M. Przanowski. Killing vector fields in self-dual, Euclidean Einstein spaces with Λ ≠ 0. Journal of Mathematical Physics, 32(4):1004–1010, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p1004_s1.

Edgar:1991:SCC

[582] S. Brian Edgar. Sufficient conditions for a curvature tensor to be Riemannian and for determining its metric. Journal of Mathematical Physics, 32(4):1011–1016, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i4/p1011_s1.

Tsoubelis:1991:GIP

[583] Dimitri Tsoubelis and An Zhong Wang. On the gravitational interaction of plane symmetric clouds of null dust. Journal of Mathematical Physics, 32(4):1017–1024, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i4/p1017_s1.

Li:1991:CGW

[584] Wei Li, Isidore Hauser, and Frederick J. Ernst. Colliding gravitational waves with Killing–Cauchy horizons. Journal of Mathematical Physics, 32(4):1025–1029, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i4/p1025_s1.

Li:1991:CWS

[585] Wei Li, Isidore Hauser, and Frederick J. Ernst. Colliding wave solutions of the Einstein–Maxwell field equations. Journal of Mathematical Physics, 32(4):1030–1034, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i4/p1030_s1.

Sukhov:1991:EMT

[586] Andrei M. Sukhov. Energy-momentum tensor and transformation properties of the gauge potentials in gravitation theory. Journal of Mathematical Physics, 32(4):1035–1036, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i4/p1035_s1.

Antoniadis:1991:GFS

[587] I. Antoniadis and E. Mottola. Graviton fluctuations in de Sitter space. Journal of Mathematical Physics, 32(4):1037–1044, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i4/p1037_s1.

Delbourgo:1991:RBU

[588] R. Delbourgo and Brian G. Kenny. Relations between universal scaling constants for the circle map near the golden mean. Journal of Mathematical Physics, 32(4):1045–1051, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i4/p1045_s1.

Davydychev:1991:SER

[589] A. I. Davydychev. Some exact results for N-point massive Feynman integrals. Journal of Mathematical Physics, 32(4):1052–1060, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i4/
p1052_s1.

Liu:1991:SCG

[590] Yu Fen Liu. Spinor connections and geometrization of fermions. Journal of Mathematical Physics, 32(4):1061–1064, April 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i4/p1061_s1.

Keyl:1991:AGS

[591] Michael Keyl. About the geometric structure of symmetry-breaking. Journal of Mathematical Physics, 32(4):1065–1071, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i4/p1065_s1.

Mainland:1991:DCS

[592] G. Bruce Mainland and L. Alayne Parson. Derivation of the coordinate system that separates the Bethe–Salpeter equation for the Wick– Cutkosky model. Journal of Mathematical Physics, 32(4):1072–1075, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p1072_s1.

Cassinelli:1991:HSR

[593] G. Cassinelli, P. Truini, and V. S. Varadarajan. Hilbert space representations of the Poincaré group for the Landau gauge. Journal of Mathematical Physics, 32(4):1076–1090, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i4/p1076_s1.

Barut:1991:CMB

[594] A. O. Barut. The covariant many-body problem in quantumelectrodynamics. Journal of Mathematical Physics, 32(4):1091–1095, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p1091_s1.

Epifanio:1991:PAV

[595] Giuseppina Epifanio and Camillo Trapani. Partial ∗-algebras and Volterra convolution of distribution kernels. Journal of Mathematical Physics, 32(4):1096–1101, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i4/p1096_s1.

Kent:1991:MOM

[596] R. D. Kent and M. Schlesinger. Multibody operator matrix elements and subduction coefficients in U(n). Journal of Mathematical Physics, 32(4):1102–1108, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i4/p1102_s1.

Maia:1991:EMM

[597] Adolfo Maia, Jr., Erasmo Recami, Waldyr A. Rodrigues, Jr., and Marcio A. F. Rosa. Erratum: “Magnetic monopoles without string in the Kähler–Clifford algebra bundle: a geometrical interpretation”, [J. Math. Phys. 31 (1990), no. 2, 502–505; MR 91d:81158]. Journal of Mathematical Physics, 32(4):1109, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i4/p1109_s2. See [90].

Mitra:1991:EFD

[598] Bani Mitra. Erratum: “Finite-dimensional irreducible representations of the SU(3/1) superalgebra”, [J. Math. Phys. 30 (1989), no. 3, 719– 731; MR 90a:17018]. Journal of Mathematical Physics, 32(4):1109, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p1109_s1. See [8].

Tkadlec:1991:VSO

[599] Josef Tkadlec. Vector states on orthoposets and extremality. Journal of Mathematical Physics, 32(5):1117–1120, May 1991. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i5/p1117_s1.

Tsohantjis:1991:STC

[600] I. Tsohantjis and J. F. Cornwell. Sugawara-type constructions of the Virasoro algebra based on the twisted affine Kac–Moody superalgebras. Journal of Mathematical Physics, 32(5):1121–1149, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1121_s1.

Caride:1991:CSF

[601] A. O. Caride, S. I. Zanette, and S. R. A. Nogueira. On the characters of solvable finite groups. Journal of Mathematical Physics, 32(5):1150– 1154, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1150_s1.

Celeghini:1991:QHG

[602] Enrico Celeghini, Riccardo Giachetti, Emanuele Sorace, and Marco Tarlini. The quantum Heisenberg group H(1)_{q}. Journal of Mathematical Physics, 32(5):1155–1158, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1155_s1.

Celeghini:1991:TDE

[603] Enrico Celeghini, Riccardo Giachetti, Emanuele Sorace, and Marco Tarlini. The three-dimensional Euclidean quantum group E(3)_{q} and its Rmatrix. Journal of Mathematical Physics, 32(5):1159–1165, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/
p1159_s1.

Deumens:1991:CSF

[604] E. Deumens, Y. Öhrn, and B. Weiner. Coherent state formulation of multiconfiguration states. Journal of Mathematical Physics, 32(5):1166– 1175, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1166_s1.

Dundarer:1991:ORS

[605] A. Reşit Dündarer and Feza Gürsey. Octonionic representations of SO(8) and its subgroups and cosets. Journal of Mathematical Physics, 32(5):

1176–1181, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1176_s1.

Gadella:1991:SWC

[606] Manuel Gadella, Miguel Á. Martín, Luis M. Nieto, and Mariano A. del Olmo. The Stratonovich–Weyl correspondence for one-dimensional kinematical groups. Journal of Mathematical Physics, 32(5):1182–1192, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i5/p1182_s1.

Maekawa:1991:UIR

[607] Takayoshi Maekawa. On the unitary irreducible representations of SO(n, 1) and U(n, 1) in the scalar product with the nonintertwining operator. Journal of Mathematical Physics, 32(5):1193–1202, May 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1193_
s1.

Ponnapalli:1991:GME

[608] P. S. Ponnapalli, M. Schlesinger, and R. D. Kent. U(n) generator matrix elements using orthogonal Gel′fand basis states. Journal of Mathematical Physics, 32(5):1203–1217, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1203_s1.

Fradkin:1991:IDG

[609] E. S. Fradkin and V. Ya. Linetsky. Infinite-dimensional generalizations of finite-dimensional symmetries. Journal of Mathematical Physics, 32(5): 1218–1226, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1218_s1.

Chakrabarti:1991:A

[610] A. Chakrabarti. q-analogs of IU(n) and U(n, 1). Journal of Mathematical Physics, 32(5):1227–1234, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1227_s1.

Yukalov:1991:MSS

[611] V. I. Yukalov. Method of self-similar approximations. Journal of Mathematical Physics, 32(5):1235–1239, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i5/p1235_s1.

Laurence:1991:WVP

[612] Peter Laurence and Marco Avellaneda. On Woltjer’s variational principle for force-free fields. Journal of Mathematical Physics, 32(5):1240– 1253, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1240_s1.

Lindblom:1991:IAW

[613] Lee Lindblom and R. T. Robiscoe. Improving the accuracy of WKB eigenvalues. Journal of Mathematical Physics, 32(5):1254–1258, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i5/p1254_s1.

Robinson:1991:FDC

[614] D. C. Robinson. Four-dimensional conformal and quaternionic structures. Journal of Mathematical Physics, 32(5):1259–1262, May 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1259_
s1.

Galicki:1991:DYM

[615] Krzysztof Galicki and Yat Sun Poon. Duality and Yang–Mills fields on quaternionic Kähler manifolds. Journal of Mathematical Physics, 32(5): 1263–1268, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1263_s1.

Kopec:1991:PME

[616] S. Kopeć. Properties of maximum entropy approximate solutions to Fredholm integral equations. Journal of Mathematical Physics, 32(5):1269– 1272, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1269_s1.

Torresani:1991:WAR

[617] Bruno Torrésani. Wavelets associated with representations of the affine Weyl–Heisenberg group. Journal of Mathematical Physics, 32(5):1273– 1279, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1273_s1.

Beltrametti:1991:CCN

[618] Enrico G. Beltrametti and Maciej J. Maczyński. On a characterization of classical and nonclassical probabilities. Journal of Mathematical Physics, 32(5):1280–1286, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1280_s1.

Pykacz:1991:HVQ

[619] Jarosław Pykacz and Emilio Santos. Hidden variables in quantum logic approach reexamined. Journal of Mathematical Physics, 32(5):1287– 1292, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1287_s1.

daCosta:1991:RIL

[620] G. A. T. F. da Costa. A rational interaction: Lower bound on the number of I-PI graphs without tadpoles. Journal of Mathematical Physics, 32(5): 1293–1295, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1293_s1.

deLange:1991:SOD

[621] O. L. de Lange. Shift operators for a Dirac oscillator. Journal of Mathematical Physics, 32(5):1296–1300, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i5/p1296_s1.

Ge:1991:NSB

[622] Mo Lin Ge and Kang Xue. New solutions of braid group representations associated with Yang–Baxter equation. Journal of Mathematical Physics, 32(5):1301–1309, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1301_s1.

Richardson:1991:IPC

[623] P. A. Richardson and M. A. Hooshyar. The inverse problem of central, spin-orbit and L^{2}-dependent potentials. Journal of Mathematical Physics, 32(5):1310–1317, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1310_s1.

Kirst:1991:ISO

[624] Th. Kirst. Inverse scattering, output kernels, and Jost solutions. Journal of Mathematical Physics, 32(5):1318–1321, May 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1318_s1.

Taub:1991:IND

[625] A. H. Taub. Interaction of null dust clouds fronted by plane impulsive gravitational waves. II. Journal of Mathematical Physics, 32(5):1322– 1327, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1322_s1.

Li:1991:RHP

[626] Wei Li and Bo Yu Hou. The Riemann–Hilbert problem approach to a representation of the Virasoro group. II. Journal of Mathematical Physics, 32(5):1328–1333, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1328_s1.

Jing:1991:ESE

[627] Ji Liang Jing. Exact solutions of Einstein’s equations for space–time with local rotational symmetry in which the Dirac equation separates. Journal of Mathematical Physics, 32(5):1334–1336, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1334_s1.

Ng:1991:UTG

[628] Y. Jack Ng and H. van Dam. Unimodular theory of gravity and the cosmological constant. Journal of Mathematical Physics, 32(5):1337– 1340, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1337_s1.

Kostin:1991:PDE

[629] M. D. Kostin. Partial differential equations for the probability density and charge density of quantum statistical mechanics. Journal of Mathematical Physics, 32(5):1341–1343, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i5/p1341_s1.

Zhang:1991:ERF

[630] M. Q. Zhang. Exact response functions of a 2-D fermion fluid. Journal of Mathematical Physics, 32(5):1344–1349, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i5/p1344_s1.

SicardiSchifino:1991:GGP

[631] A. C. Sicardi Schifino, A. Costa, and C. Ferro Fontán. Generalization of the Glansdorff–Prigogine stability criterion. Journal of Mathematical Physics, 32(5):1350–1353, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1350_s1.

Audretsch:1991:RPF

[632] Jürgen Audretsch and Claus Lämmerzahl. Reasons for a physical field to obey linear partial differential equations. Journal of Mathematical Physics, 32(5):1354–1359, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1354_s1.

Ewen:1991:STG

[633] H. Ewen, P. Schaller, and Günter Schwarz. Schwinger terms from geometric quantization of field theories. Journal of Mathematical Physics, 32(5): 1360–1367, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1360_s1.

Dodd:1991:ESS

[634] L. R. Dodd and M. A. Lohe. Exact solutions for a soliton crystal. Journal of Mathematical Physics, 32(5):1368–1375, May 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1368_s1.

Balasin:1991:APR

[635] H. Balasin and M. Stierle. The action principle revisited. Journal of Mathematical Physics, 32(5):1376–1382, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i5/p1376_s1.

Marques-Bonham:1991:DEN

[636] Sirley Marques-Bonham. The Dirac equation in a non-Riemannian manifold. III. An analysis using the algebra of quaternions and octonions.

Journal of Mathematical Physics, 32(5):1383–1394, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1383_s1.

Yang:1991:EMV

[637] Yisong Yang. Existence of the massive SO(3) vortices. Journal of Mathematical Physics, 32(5):1395–1399, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i5/p1395_s1.

Willis:1991:EMS

[638] Barton L. Willis. Electromagnetic momenta for systems of charged particles. Journal of Mathematical Physics, 32(5):1400–1402, May 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/p1400_
s1.

Kvitsinsky:1991:AEI

[639] Andrei Kvitsinsky and Seth Putterman. Adiabatic evolution of an irreversible two level system. Journal of Mathematical Physics, 32(5): 1403–1407, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1403_s1.

Wu:1991:TCD

[640] Yihren Wu and K. Y. Szeto. Topological classification of defects in twodimensional quasicrystals. Journal of Mathematical Physics, 32(5):1408– 1414, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i5/p1408_s1.

Kalnins:1991:CSFb

[641] E. G. Kalnins and W. Miller, Jr. Complete sets of functions for perturbations of Robertson Walker cosmologies. Journal of Mathematical Physics, 32(5):1415–1423, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i5/p1415_s1.

DHoker:1991:ORM

[642] Eric D’Hoker, Roberto Floreanini, and Luc Vinet. q-oscillator realizations of the metaplectic representation of quantum osp(3, 2). Journal of Mathematical Physics, 32(6):1427–1429, June 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1427_s1.

Ginocchio:1991:NIS

[643] Joseph N. Ginocchio. A new identity for some six-j symbols. Journal of Mathematical Physics, 32(6):1430–1432, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i6/p1430_s1.

Dowker:1991:CES

[644] J. S. Dowker. Central extensions of sphere groups. Journal of Mathematical Physics, 32(6):1433–1436, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i6/p1433_s1.

Kopp:1991:SPT

[645] A. Kopp and K. Schindler. Singular perturbation theory applied to magnetohydrostatic equilibria: Proof of convergence. Journal of Mathematical Physics, 32(6):1437–1439, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i6/p1437_s1.

Lawton:1991:MPW

[646] Wayne M. Lawton. Multiresolution properties of the wavelet Galerkin operator. Journal of Mathematical Physics, 32(6):1440–1443, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/
p1440_s1.

CohendeLara:1991:NSG

[647] M. Cohen de Lara. A note on the symmetry group and perturbation algebra of a parabolic partial differential equation. Journal of Mathematical Physics, 32(6):1444–1449, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i6/p1444_s1.

Musette:1991:AMD

[648] M. Musette and R. Conte. Algorithmic method for deriving Lax pairs from the invariant Painlevé analysis of nonlinear partial differential equations. Journal of Mathematical Physics, 32(6):1450–1457, June 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1450_
s1.

Chakravarty:1991:CSA

[649] Sarbarish Chakravarty, Lionel Mason, and Ezra T. Newman. Canonical structures on anti-self-dual four-manifolds and the diffeomorphism group. Journal of Mathematical Physics, 32(6):1458–1464, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1458_s1.

Bailey:1991:SLS

[650] T. N. Bailey. The space of leaves of a shear-free congruence, multipole expansions, and Robinson’s theorem. Journal of Mathematical Physics, 32(6):1465–1469, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i6/p1465_s1.

Martin:1991:TPQ

[651] Pablo Martín and George A. Baker, Jr. Two-point quasifractional approximant in physics. Truncation error. Journal of Mathematical Physics, 32(6):1470–1477, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i6/p1470_s1.

Kitada:1991:RPF

[652] Akihiko Kitada and Hiroyuki Umehara. A rapid proof for a flattening property of a classical solution of the Mullins equation. Journal of Mathematical Physics, 32(6):1478–1479, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i6/p1478_s1.

Bouquet:1991:PSO

[653] S. E. Bouquet, M. R. Feix, and P. G. L. Leach. Properties of second-order ordinary differential equations invariant under time translation and selfsimilar transformation. Journal of Mathematical Physics, 32(6):1480– 1490, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i6/p1480_s1.

Hojman:1991:ARG

[654] Sergio A. Hojman, Sergio Chayet, Darío Núñez, and Marco A. Roque. An algorithm to relate general solutions of different bidimensional problems. Journal of Mathematical Physics, 32(6):1491–1497, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1491_s1.

Gu:1991:CCI

[655] Zhu Quan Gu. Complex confocal involutive systems associated with the solutions of the AKNS evolution equations. Journal of Mathematical Physics, 32(6):1498–1504, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i6/p1498_s1.

Hajra:1991:SQD

[656] K. Hajra. Stochastic quantization of a dissipative dynamical system and its hydrodynamical interpretation. Journal of Mathematical Physics, 32(6):1505–1509, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i6/p1505_s1.

Srivastava:1991:QLD

[657] Sunita Srivastava, Vishwamittar, and I. S. Minhas. On the quantization of linearly damped harmonic oscillator. Journal of Mathematical Physics, 32(6):1510–1515, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i6/p1510_s1.

Labelle:1991:SDC

[658] Simon Labelle, Michel Mayrand, and Luc Vinet. Symmetries and degeneracies of a charged oscillator in the field of a magnetic monopole. Journal of Mathematical Physics, 32(6):1516–1521, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1516_s1.

Demuth:1991:LCO

[659] M. Demuth. On large coupling operator norm convergences of resolvent differences. Journal of Mathematical Physics, 32(6):1522–1530, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i6/p1522_s1.

Gu:1991:TFD

[660] Zhu Quan Gu. Two finite-dimensional completely integrable Hamiltonian systems associated with the solutions of the MKDV hierarchy. Journal of Mathematical Physics, 32(6):1531–1536, June 1991. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i6/p1531_s1.

Fan:1991:JCN

[661] Zu Hui Fan, Zhi Quan Kuang, Can Bin Liang, and Yun Qiang Yu. Junction conditions on null hypersurfaces for spherically symmetric metrics. Journal of Mathematical Physics, 32(6):1537–1540, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1537_s1.

Saridakis:1991:SIC

[662] E. Saridakis and M. Tsamparlis. Symmetry inheritance of conformal Killing vectors. Journal of Mathematical Physics, 32(6):1541–1551, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i6/p1541_s1.

Stephani:1991:SRO

[663] H. Stephani. Similarity reduction of the one-Killing-vector vacuum equations in general relativity. Journal of Mathematical Physics, 32(6):1552– 1555, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i6/p1552_s1.

Maluf:1991:APD

[664] José W. Maluf. On the absence of physical degrees of freedom in the Stephenson–Kilmister–Yang theory. Journal of Mathematical Physics, 32(6):1556–1559, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i6/p1556_s1.

Davidson:1991:BBC

[665] W. Davidson. A big-bang cylindrically symmetric radiation universe. Journal of Mathematical Physics, 32(6):1560–1561, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1560_s1.

Schmidt:1991:SIG

[666] Hans-Jürgen Schmidt. Scale-invariant gravity in two dimensions. Journal of Mathematical Physics, 32(6):1562–1566, June 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1562_s1.

Accardi:1991:WCL

[667] L. Accardi, A. Frigerio, and Y. G. Lu. The weak coupling limit for fermions. Journal of Mathematical Physics, 32(6):1567–1581, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/
p1567_s1.

Gibbs:1991:MSF

[668] Richard L. Gibbs. The massive scalar field propagator on finite lattices. Journal of Mathematical Physics, 32(6):1582–1585, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1582_s1.

Heidenreich:1991:UMF

[669] W. F. Heidenreich. Unitarizable massless fields of higher order and higher spin. Journal of Mathematical Physics, 32(6):1586–1590, June 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1586_
s1.

Keller:1991:MGC

[670] Jaime Keller and Suemi Rodríguez-Romo. Multivectorial generalization of the Cartan map. Journal of Mathematical Physics, 32(6):1591–1598, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i6/p1591_s1.

Bandelloni:1991:GTL

[671] Giuseppe Bandelloni. Gauge theories on loop groups. Journal of Mathematical Physics, 32(6):1599–1606, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i6/p1599_s1.

Bremner:1991:SLA

[672] Murray R. Bremner. Structure of the Lie algebra of polynomial vector fields on the Riemann sphere with three punctures. Journal of Mathematical Physics, 32(6):1607–1608, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i6/p1607_s1.

Klauder:1991:WTP

[673] J. R. Klauder and R. F. Streater. A wavelet transform for the Poincaré group. Journal of Mathematical Physics, 32(6):1609–1611, June 1991.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/
p1609_s1.

Nakariki:1991:ESP

[674] Shin ichi Nakariki. An exact solution in Poincaré gauge theory: Monopole solution. Journal of Mathematical Physics, 32(6):1612–1618, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i6/
p1612_s1.

Newton:1991:HEL

[675] Conrad Newton and Tai Tsun Wu. High-energy, large-momentum-transfer processes: The N-rung ladder diagram of φ^{3} theory. Journal of Mathematical Physics, 32(6):1619–1637, June 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i6/p1619_s1.

Temme:1991:SGP

[676] F. P. Temme. The structure and Gel′fand patterns inherent in the Heisenberg supergenerator algebra of dual Liouville-space |kq{K̃.}(k_{1} − k_{n}): [λ̃], S_{n}⟩⟩ tensors and the Cayley algebra of scalar invariants over a (k_{1} − k_{n}) field. Journal of Mathematical Physics, 32(6):1638–1650, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i6/p1638_s1.

Davies:1991:EDR

[677] K. T. R. Davies, R. W. Davies, and G. D. White. Erratum: “Dispersion relations for causal Green’s functions: derivations using the Poincaré– Bertrand theorem and its generalizations”, [J. Math. Phys. 31 (1990), no. 6, 1356–1373; MR 91h:81179]. Journal of Mathematical Physics, 32 (6):1651, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i6/p1651_s1. See [204].

Okubo:1991:RRFa

[678] Susumu Okubo. Real representations of finite Clifford algebras. I. Classification. Journal of Mathematical Physics, 32(7):1657–1668, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/
p1657_s1.

Okubo:1991:RRFb

[679] Susumu Okubo. Real representations of finite Clifford algebras. II. Explicit construction and pseudo-octonion. Journal of Mathematical Physics, 32(7):1669–1673, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1669_s1.

Brown:1991:PRH

[680] R. W. Brown, M. E. Convery, and D. B. Delaney. Product representation for the harmonic series of a unit vector. Journal of Mathematical Physics, 32(7):1674–1682, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1674_s1.

Filippov:1991:EHW

[681] G. F. Filippov and A. L. Blokhin. Effective Hamiltonian within the microscopic unitary nuclear model. I. Coherent-state restriction technique for microscopic Hamiltonian. Journal of Mathematical Physics, 32(7): 1683–1695, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1683_s1.

Avramenko:1991:EHW

[682] V. I. Avramenko and A. L. Blokhin. Effective Hamiltonian within the microscopic unitary nuclear model. II: The study of sd nuclei with two valence particles. Journal of Mathematical Physics, 32(7):1696–1702, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1696_s1.

Fuchs:1991:SGS

[683] J. C. Fuchs. Symmetry groups and similarity solutions of MHD equations. Journal of Mathematical Physics, 32(7):1703–1708, July 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1703_
s1.

vanBemmelen:1991:NSR

[684] Theo van Bemmelen and Paul Kersten. Nonlocal symmetries and recursion operator of the Landau–Lifshitz equation. Journal of Mathematical Physics, 32(7):1709–1716, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1709_s1.

Ciulli:1991:AME

[685] S. Ciulli, M. Mounsif, N. Gorman, and T. D. Spearman. On the application of maximum entropy to the moments problem. Journal of Mathematical Physics, 32(7):1717–1719, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i7/p1717_s1.

Jiang:1991:IDS

[686] Zhuhan Jiang and S. Rauch-Wojciechowski. Integrable dynamical systems with quadratic and cubic nonlinearities related to graded Lie algebras. Journal of Mathematical Physics, 32(7):1720–1732, July 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1720_
s1.

Verosky:1991:NPO

[687] John M. Verosky. Negative powers of Olver recursion operators. Journal of Mathematical Physics, 32(7):1733–1736, July 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1733_s1.

Chau:1991:AEC

[688] Ling-Lie Chau, J. C. Shaw, and H. C. Yen. An alternative explicit construction of N-soliton solutions in 1 + 1 dimensions. Journal of Mathematical Physics, 32(7):1737–1743, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i7/p1737_s1.

Guil:1991:HMM

[689] F. Guil and M. Mañas. Homogeneous manifolds and modified KdV equations. Journal of Mathematical Physics, 32(7):1744–1749, July 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1744_
s1.

Kerbrat:1991:GGF

[690] Yvan Kerbrat, Hélène Kerbrat-Lunc, and Jedrzej Śniatycki. Generalized gauge fields. Journal of Mathematical Physics, 32(7):1750–1754, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1750_s1.

Luckock:1991:MBC

[691] Hugh Luckock. Mixed boundary conditions in quantum field theory. Journal of Mathematical Physics, 32(7):1755–1766, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1755_s1.

Wentworth:1991:ADF

[692] Richard Wentworth. Asymptotics of determinants from functional integration. Journal of Mathematical Physics, 32(7):1767–1773, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/
p1767_s1.

Pfirsch:1991:GFW

[693] D. Pfirsch and R. N. Sudan. Green’s functions in WKB approximation. Journal of Mathematical Physics, 32(7):1774–1779, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1774_s1.

Harnad:1991:GTM

[694] J. Harnad and J. Hurtubise. Generalized tops and moment maps to loop algebras. Journal of Mathematical Physics, 32(7):1780–1787, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1780_s1.

Zerzion:1991:CMS

[695] Daniel Zerzion, L. P. Horwitz, and R. I. Arshansky. Classical mechanics of special relativity in a Riemannian space–time. Journal of Mathematical Physics, 32(7):1788–1795, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1788_s1.

Muraki:1991:RCE

[696] Naohumi Muraki. Remarks on continuity of entropy of states on C^{∗}- dynamical systems. Journal of Mathematical Physics, 32(7):1796–1798, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1796_s1.

Carpio-Bernido:1991:PIQ

[697] M. Victoria Carpio-Bernido. Path integral quantization of certain noncentral systems with dynamical symmetries. Journal of Mathematical Physics, 32(7):1799–1807, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1799_s1.

Beckers:1991:LSPa

[698] J. Beckers and N. Debergh. Lie structures in parasupersymmetric quantum mechanics. I. The standard supersymmetrization procedure. Journal of Mathematical Physics, 32(7):1808–1814, July 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1808_s1.

Beckers:1991:LSPb

[699] J. Beckers and N. Debergh. Lie structures in parasupersymmetric quantum mechanics. II. The spin-orbit coupling supersymmetrization procedure. Journal of Mathematical Physics, 32(7):1815–1821, July 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1815_
s1.

Dammert:1991:EEA

[700] Örjan Dammert. Energy eigenvalues for an arbitrary potential well with N minima. Journal of Mathematical Physics, 32(7):1822–1837, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1822_s1.

Arai:1991:NBT

[701] Asao Arai. Noninvertible Bogoliubov transformations and instability of embedded eigenvalues. Journal of Mathematical Physics, 32(7):1838– 1846, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1838_s1.

Hall:1991:CVF

[702] G. S. Hall and J. D. Steele. Conformal vector fields in general relativity. Journal of Mathematical Physics, 32(7):1847–1853, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1847_s1.

Sharma:1991:PSC

[703] Ramesh Sharma. Proper special conformal Killing vectors and the quadratic theory of gravity. Journal of Mathematical Physics, 32(7):

1854–1856, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1854_s1.

Kramer:1991:CSP

[704] Dietrich Kramer and Jaume Carot. Conformal symmetry of perfect fluids in general relativity. Journal of Mathematical Physics, 32(7):1857– 1860, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1857_s1.

Reboucas:1991:EUM

[705] M. J. Rebouças and A. F. F. Teixeira. Einstein universe as a moving fluid cosmology. Journal of Mathematical Physics, 32(7):1861–1865, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1861_s1.

Xanthopoulos:1991:LPP

[706] Basilis C. Xanthopoulos. Linear perturbations of plane polarized plane waves. II. The Einstein–Maxwell perturbations. Journal of Mathematical Physics, 32(7):1866–1874, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1866_s1.

Xanthopoulos:1991:UBB

[707] Basilis C. Xanthopoulos and Thomas Zannias. The uniqueness of the Bekenstein black hole. Journal of Mathematical Physics, 32(7):1875– 1880, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1875_s1.

Bakas:1991:CAG

[708] Ioannis Bakas and Mark J. Bowick. Curiosities of arithmetic gases. Journal of Mathematical Physics, 32(7):1881–1884, July 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1881_s1.

Narumi:1991:GEN

[709] Hideyuki Narumi, Haruo Hosoya, and Hiroshi Murakami. Generalized expression for the numbers of perfect matching of cylindrical m×n graphs. Journal of Mathematical Physics, 32(7):1885–1889, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1885_s1.

Arcuri:1991:CCE

[710] Regina C. Arcuri. Conformal and critical embeddings, infinite magic square, and a new Clifford product. Journal of Mathematical Physics, 32(7):1890–1899, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i7/p1890_s1.

Tu:1991:TIA

[711] G. Z. Tu, R. I. Andrushkiw, and X. C. Huang. A trace identity and its application to integrable systems of 1 + 2 dimensions. Journal of Mathematical Physics, 32(7):1900–1907, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i7/p1900_s1.

Zhang:1991:SCQ

[712] Hongwei Zhang, Gui Zhang Tu, Walter Oevel, and Benno Fuchssteiner. Symmetries, conserved quantities, and hierarchies for some lattice systems with soliton structure. Journal of Mathematical Physics, 32(7): 1908–1918, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1908_s1.

Kopper:1991:HRR

[713] Christoph Kopper. The ’t Hooft reparametrization respects Borel summability. Journal of Mathematical Physics, 32(7):1919–1920, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1919_s1.

deAzcarraga:1991:SLT

[714] J. A. de Azcárraga, M. S. Rashid, and W. J. Zakrzewski. Skyrmelike and topological terms in sigma models. Journal of Mathematical Physics, 32(7):1921–1927, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1921_s1.

Tanimura:1991:REC

[715] N. Tanimura and O. Tanimura. Recursive equation of character expansion coefficient in lattice gauge theory. Journal of Mathematical Physics, 32(7):1928–1933, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1928_s1.

Khudaverdian:1991:GSE

[716] O. M. Khudaverdian. Geometry of superspace with even and odd brackets. Journal of Mathematical Physics, 32(7):1934–1937, July 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i7/p1934_
s1.

Khudaverdian:1991:CPB

[717] O. M. Khudaverdian and A. P. Nersessian. Canonical Poisson brackets of different gradings and strange superalgebras. Journal of Mathematical Physics, 32(7):1938–1941, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1938_s1.

AnaniasNeto:1991:SCS

[718] Jorge Ananias Neto and Erasmo Ferreira. Solutions for the classical sigma model in the hedgehog ansatz. Journal of Mathematical Physics, 32(7): 1942–1948, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i7/p1942_s1.

AnaniasNeto:1991:PSS

[719] Jorge Ananias Neto, Ramón Méndez Galain, and Erasmo Ferreira. Properties of the Skyrme soliton configuration. Journal of Mathematical Physics, 32(7):1949–1957, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1949_s1.

Helsing:1991:CTP

[720] Johan Helsing, Göran Grimvall, and Ke Da Bao. Conduction in a two-phase plane with diamond-shaped tiling. Journal of Mathematical Physics, 32(7):1958–1960, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i7/p1958_s1.

Freeman:1991:PLR

[721] D. Kent Freeman. The pseudodipole as the limit of a resonant cylindrical dipole. Journal of Mathematical Physics, 32(7):1961–1964, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i7/p1961_s1.

Folkins:1991:FTD

[722] Ian Folkins. Functions of two-dimensional Bravais lattices. Journal of Mathematical Physics, 32(7):1965–1969, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i7/p1965_s1.

Batalin:1991:ECQ

[723] I. A. Batalin, P. M. Lavrov, and I. V. Tyutin. Erratum: “Covariant quantization of gauge theories in the framework of extended BRST symmetry”, [J. Math. Phys. 31 (1990), no. 6, 1487–1493; MR 91e:81103]. Journal of Mathematical Physics, 32(7):1970, July 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i7/p1970_s1. See [224].

Albertin:1991:DGF

[724] Uwe K. Albertin. The diffeomorphism group and flat principal bundles. Journal of Mathematical Physics, 32(8):1975–1980, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p1975_
s1.

Gajdzinski:1991:AM

[725] C. Gajdzinski and R. F. Streater. Anomaly meltdown. Journal of Mathematical Physics, 32(8):1981–1983, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i8/p1981_s1.

Grosche:1991:GPI

[726] Christian Grosche. The SU(u, v)-group path integration. Journal of Mathematical Physics, 32(8):1984–1997, August 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i8/p1984_s1.

Kim:1991:STG

[727] Chang-Ho Kim and Y. S. Kim. Space–time geometry of relativistic particles in four-dimensional phase space. Journal of Mathematical Physics, 32(8):1998–2006, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i8/p1998_s1.

Pilato:1991:STR

[728] Alex Pilato. Super twistor realization of unitary super representations of U(p, q|N). Journal of Mathematical Physics, 32(8):2007–2019, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2007_s1.

RodriguezPlaza:1991:CO

[729] M. J. Rodríguez Plaza. Casimir operators of U_{q}sl(3). Journal of Mathematical Physics, 32(8):2020–2027, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i8/p2020_s1.

Weimar-Woods:1991:TDR

[730] Evelyn Weimar-Woods. The three-dimensional real Lie algebras and their contractions. Journal of Mathematical Physics, 32(8):2028–2033, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2028_s1.

Juriev:1991:PRQ

[731] Denis Juriev. Projective reduction in quantum conformal field theory. Journal of Mathematical Physics, 32(8):2034–2038, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2034_
s1.

Purkait:1991:BHS

[732] S. Purkait and A. Roy Chowdhury. On the bi-Hamiltonian structure, strong and hereditary operator of symmetries for a new nonlinear system. Journal of Mathematical Physics, 32(8):2039–2042, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2039_
s1.

Kersten:1991:GSA

[733] Paul H. M. Kersten. The general symmetry algebra structure of the underdetermined equation u_{x} = (v_{xx})^{2}. Journal of Mathematical Physics, 32(8):2043–2050, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i8/p2043_s1.

Mahomed:1991:CSA

[734] F. M. Mahomed and P. G. L. Leach. Contact symmetry algebras of scalar second-order ordinary differential equations. Journal of Mathematical Physics, 32(8):2051–2055, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i8/p2051_s1.

Gerold:1991:SII

[735] A. Gerold and K. Buchner. Solitons and isometric immersions. Journal of Mathematical Physics, 32(8):2056–2062, August 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2056_s1.

Lopez:1991:EEV

[736] M. C. López, R. J. Noriega, and C. G. Schifini. Equivalence and s-equivalence of vector-tensor Lagrangians. Journal of Mathematical Physics, 32(8):2063–2064, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i8/p2063_s1.

Tratnik:1991:SMOa

[737] M. V. Tratnik. Some multivariable orthogonal polynomials of the Askey tableau—continuous families. Journal of Mathematical Physics, 32(8):2065–2073, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i8/p2065_s1.

Antoine:1991:NFP

[738] J.-P. Antoine and A. Inoue. Normal forms on partial O^{∗}-algebras. Journal of Mathematical Physics, 32(8):2074–2081, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2074_s1.

Arai:1991:TES

[739] Asao Arai. A theorem on essential self-adjointness with application to Hamiltonians in nonrelativistic quantum field theory. Journal of Mathematical Physics, 32(8):2082–2088, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i8/p2082_s1.

Gilkey:1991:HEA

[740] Peter B. Gilkey, Thomas P. Branson, and Stephen A. Fulling. Heat equation asymptotics of “nonminimal” operators on differential forms. Journal of Mathematical Physics, 32(8):2089–2091, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2089_s1.

Rylov:1991:NRM

[741] Yuri A. Rylov. Non-Riemannian model of the space–time responsible for quantum effects. Journal of Mathematical Physics, 32(8):2092–2098, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2092_s1.

Audretsch:1991:ERS

[742] Jürgen Audretsch and Claus Lämmerzahl. Establishing the Riemannian structure of space–time by means of light rays and free matter waves. Journal of Mathematical Physics, 32(8):2099–2105, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2099_
s1.

Gudder:1991:AUI

[743] Stanley Gudder. Amplitudes and the universal influence function. Journal of Mathematical Physics, 32(8):2106–2113, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2106_s1.

Stovicek:1991:KFA

[744] P. Štovíček. Kreĭn’s formula approach to the multisolenoid Aharonov– Bohm effect. Journal of Mathematical Physics, 32(8):2114–2122, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2114_s1.

deFaria:1991:FIH

[745] M. de Faria, J. Potthoff, and L. Streit. The Feynman integrand as a Hida distribution. Journal of Mathematical Physics, 32(8):2123–2127, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2123_s1.

Isar:1991:QDO

[746] Aurelian Isar, Werner Scheid, and Aurel Săndulescu. Quasiprobability distributions for open quantum systems within the Lindblad theory.

Journal of Mathematical Physics, 32(8):2128–2134, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2128_
s1.

Aguilera-Navarro:1991:NSH

[747] V. C. Aguilera-Navarro and R. Guardiola. Nonsingular spiked harmonic oscillator. Journal of Mathematical Physics, 32(8):2135–2141, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2135_s1.

Meyer:1991:BFB

[748] Jutta Meyer. Boson-fermion and baryon mapping of multiquark states. Journal of Mathematical Physics, 32(8):2142–2151, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2142_
s1.

Gatarek:1991:CQJ

[749] Dariusz Gatarek and Nicolas Gisin. Continuous quantum jumps and infinite-dimensional stochastic equations. Journal of Mathematical Physics, 32(8):2152–2157, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i8/p2152_s1.

Lehmann:1991:PAE

[750] Detlef Lehmann. A probabilistic approach to Euclidean Dirac fields. Journal of Mathematical Physics, 32(8):2158–2166, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2158_
s1.

Joshi:1991:SNS

[751] P. S. Joshi and I. H. Dwivedi. Strengths of naked singularities in radiation collapse with nonlinear mass functions. Journal of Mathematical Physics, 32(8):2167–2168, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i8/p2167_s1.

Hehl:1991:CST

[752] Friedrich W. Hehl, Wojciech Kopczyński, J. Dermott McCrea, and Eckehard W. Mielke. Chern–Simons terms in metric-affine space–time:

Bianchi identities as Euler–Lagrange equations. Journal of Mathematical Physics, 32(8):2169–2180, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i8/p2169_s1.

Klein:1991:LTS

[753] David Klein and Wei-Shih Yang. Low-temperature smoothness of the pressure for antiferromagnets and other models. Journal of Mathematical Physics, 32(8):2181–2185, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i8/p2181_s1.

Ferrari:1991:BTA

[754] Franco Ferrari. Bosonization techniques on algebraic curves and conformal field theories. Journal of Mathematical Physics, 32(8):2186– 2199, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i8/p2186_s1.

Kubyshin:1991:HPM

[755] Yu. A. Kubyshin. Higgs potentials from multidimensional Yang–Mills theories. Journal of Mathematical Physics, 32(8):2200–2209, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2200_s1.

Hou:1991:SYB

[756] Bo Yu Hou, Bo Yuan Hou, Zhong Qi Ma, and Yu Dong Yin. Solutions of Yang–Baxter equation in the vertex model and the face model for octet representation. Journal of Mathematical Physics, 32(8):2210– 2218, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i8/p2210_s1.

Farris:1991:SBR

[757] Timothy H. Farris and Thomas W. Kephart. Symmetry breaking and restoration in models with extended Higgs sectors. Journal of Mathematical Physics, 32(8):2219–2223, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i8/p2219_s1.

Arai:1991:LTB

[758] Asao Arai. Long-time behavior of an electron interacting with a quantized radiation field. Journal of Mathematical Physics, 32(8):2224– 2242, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i8/p2224_s1.

Savatier:1991:AHS

[759] F. Savatier. Asymptotic hadron spectrum in the M.I.T. bag model. Journal of Mathematical Physics, 32(8):2243–2249, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2243_s1.

Tu:1991:AST

[760] Khiet Tu. Analytic solution to the Thomas–Fermi and Thomas– Fermi–Dirac–Weizsäcker equations. Journal of Mathematical Physics, 32(8):2250–2253, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i8/p2250_s1.

De:1991:DRH

[761] Mrityunjoy De and A. Roy Chowdhury. On a double Riemann–Hilbert approach to stimulated Raman scattering. Journal of Mathematical Physics, 32(8):2254–2257, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i8/p2254_s1.

Wang:1991:MEA

[762] D. Wang, L. R. Mead, and M. de Llano. Maximum-entropy approach to classical hard-sphere and hard-disk equations of state. Journal of Mathematical Physics, 32(8):2258–2262, August 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2258_s1.

Nesi:1991:MII

[763] V. Nesi. Multiphase interchange inequalities. Journal of Mathematical Physics, 32(8):2263–2275, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i8/p2263_s1.

Buffet:1991:CCT

[764] E. Buffet and R. F. Werner. A counterexample in coagulation theory. Journal of Mathematical Physics, 32(8):2276–2278, August 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i8/p2276_
s1.

Liu:1991:SSD

[765] Hong Ya Liu. Spherically symmetric density perturbation in relativistic cosmology. Journal of Mathematical Physics, 32(8):2279–2282, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i8/p2279_s1.

Hasiewicz:1991:NSA

[766] Z. Hasiewicz, F. Defever, and W. Troost. Nonassociative superconformal algebras. Journal of Mathematical Physics, 32(9):2285–2297, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2285_s1.

Gould:1991:GGI

[767] M. D. Gould, R. B. Zhang, and A. J. Bracken. Generalized Gel′fand invariants and characteristic identities for quantum groups. Journal of Mathematical Physics, 32(9):2298–2303, September 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2298_s1.

Wolski:1991:APDb

[768] A. Wolski and J. S. Dowker. Area-preserving diffeomorphisms of Riemann surfaces. Journal of Mathematical Physics, 32(9):2304–2309, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2304_s1.

Couture:1991:GC

[769] M. Couture, J. Patera, R. T. Sharp, and P. Winternitz. Graded contractions of sl(3, C). Journal of Mathematical Physics, 32(9):2310–2318, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2310_s1.

Bhaskara:1991:QPS

[770] K. H. Bhaskara and K. Rama. Quadratic Poisson structures. Journal of Mathematical Physics, 32(9):2319–2322, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2319_s1.

Cao:1991:NGI

[771] Ce Wen Cao and Xian Guo Geng. A nonconfocal generator of involutive systems and three associated soliton hierarchies. Journal of Mathematical Physics, 32(9):2323–2328, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i9/p2323_s1.

Gielerak:1991:BPG

[772] Roman Gielerak. Bounded perturbations of the generalized random fields. Journal of Mathematical Physics, 32(9):2329–2336, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2329_s1.

Tratnik:1991:SMOb

[773] M. V. Tratnik. Some multivariable orthogonal polynomials of the Askey tableau-discrete families. Journal of Mathematical Physics, 32(9):2337– 2342, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2337_s1.

Sen:1991:IPC

[774] Shibani Sen and A. Roy Chowdhury. On the invariance properties and conservation laws for Landau–Lifshitz equation in three dimensions with Tjon–Wright ansatz. Journal of Mathematical Physics, 32(9):2343–2346, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2343_s1.

Levay:1991:QGF

[775] Péter Lévay. Quaternionic gauge fields and the geometric phase. Journal of Mathematical Physics, 32(9):2347–2357, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2347_s1.

Hasiewicz:1991:BFG

[776] Z. Hasiewicz, J. Kowalski-Glikman, J. Lukierski, and J. W. van Holten. BRST formulation of the Gupta–Bleuler quantization method. Journal of Mathematical Physics, 32(9):2358–2364, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2358_s1.

Jiang:1991:ESI

[777] Wenxin Jiang and Feza Gürsey. Extensions of some infinite algebras in space–time. Journal of Mathematical Physics, 32(9):2365–2369, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2365_s1.

Tikochinsky:1991:VIS

[778] Y. Tikochinsky. Variation-iteration solution of the classical trajectory boundary value problem. Journal of Mathematical Physics, 32(9):2370– 2373, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2370_s1.

Crater:1991:CEN

[779] Horace Crater and Dujiu Yang. A covariant extrapolation of the noncovariant two particle Wheeler–Feynman Hamiltonian from the Todorov equation and Dirac’s constraint mechanics. Journal of Mathematical Physics, 32(9):2374–2394, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i9/p2374_s1.

Cieslinski:1991:EMC

[780] Jan Cieśliński. An effective method to compute N-fold Darboux matrix and N-soliton surfaces. Journal of Mathematical Physics, 32(9):2395– 2399, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2395_s1.

Hu:1991:RCBa

[781] Ji Shan Hu and Martin Kruskal. Reflection coefficient beyond all orders for singular problems. I. Separated critical points on the nearest critical level line. Journal of Mathematical Physics, 32(9):2400–2405, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2400_s1.

Chung:1991:SQY

[782] H. J. Chung and I. G. Koh. Solutions to the quantum Yang–Baxter equation for the exceptional Lie algebras with a spectral parameter. Journal of Mathematical Physics, 32(9):2406–2408, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2406_s1.

Sun:1991:NDB

[783] Chang Pu Sun, Xu Feng Liu, and Mo Lin Ge. A new q-deformed boson realization of quantum algebra sl_{q}(n+1) and nongeneric sl_{q}(2) R-matrices. Journal of Mathematical Physics, 32(9):2409–2412, September 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2409_
s1.

Weiner:1991:SPF

[784] Brian Weiner, Erik Deumens, and Yngve Öhrn. Spin projection of fermion coherent states. Journal of Mathematical Physics, 32(9):2413– 2426, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2413_s1.

Arai:1991:MWZ

[785] Asao Arai and Osamu Ogurisu. Meromorphic N = 2 Wess–Zumino supersymmetric quantum mechanics. Journal of Mathematical Physics, 32(9):2427–2434, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i9/p2427_s1.

Tanimura:1991:SRR

[786] N. Tanimura and O. Tanimura. Sum rules for SU(2) Racah coefficients from group integral. Journal of Mathematical Physics, 32(9):2435–2441, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2435_s1.

Amar:1991:SEC

[787] V. Amar, M. Pauri, and A. Scotti. Schrödinger equation for convex plane polygons: A tiling method for the derivation of eigenvalues and eigenfunctions. Journal of Mathematical Physics, 32(9):2442–2449, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2442_s1.

Coll:1991:SFL

[788] Bartolomé Coll and Juan Antonio Morales. Symmetric frames on Lorentzian spaces. Journal of Mathematical Physics, 32(9):2450–2455, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2450_s1.

Singh:1991:BTC

[789] T. Singh and G. P. Singh. Bianchi type-I cosmological models in Lyra’s geometry. Journal of Mathematical Physics, 32(9):2456–2458, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2456_s1.

Xanthopoulos:1991:GTF

[790] Basilis C. Xanthopoulos and Thomas Zannias. The gravity of threeforms. Journal of Mathematical Physics, 32(9):2459–2467, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2459_s1.

Papacostas:1991:SSS

[791] Taxiarchis Papacostas. Spherically symmetric scalar fields. Journal of Mathematical Physics, 32(9):2468–2472, September 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2468_s1.

Fuchs:1991:FIE

[792] H. Fuchs. First integrals for the equations of motion for test particles with internal structure. Journal of Mathematical Physics, 32(9):2473–2477, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i9/p2473_s1.

Li:1991:NCG

[793] Wei Li, Isidore Hauser, and Frederick J. Ernst. Nonimpulsive colliding gravitational waves with noncollinear polarizations. Journal of Mathematical Physics, 32(9):2478–2482, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i9/p2478_s1.

Katanaev:1991:CIE

[794] M. O. Katanaev. Conformal invariance, extremals, and geodesics in twodimensional gravity with torsion. Journal of Mathematical Physics, 32 (9):2483–2496, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i9/p2483_s1.

Iyer:1991:SVD

[795] B. R. Iyer and N. Kamran. Separation of variables for the Dirac equation in an extended class of Lorentzian metrics with local rotational symmetry. Journal of Mathematical Physics, 32(9):2497–2502, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i9/p2497_s1.

Aldrovandi:1991:EGT

[796] R. Aldrovandi. Extended gauge theories. Journal of Mathematical Physics, 32(9):2503–2512, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i9/p2503_s1.

Batalin:1991:RCL

[797] I. A. Batalin, P. M. Lavrov, and I. V. Tyutin. Remarks on the Sp(2)- covariant Lagrangian quantization of gauge theories. Journal of Mathematical Physics, 32(9):2513–2521, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i9/p2513_s1.

Socolovsky:1991:GTF

[798] M. Socolovsky. Gauge transformations and fiber bundle theory. Journal of Mathematical Physics, 32(9):2522–2531, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2522_s1.

Tchrakian:1991:HOG

[799] D. H. Tchrakian and A. Chakrabarti. How overdetermined are the generalized self-duality relations? Journal of Mathematical Physics, 32 (9):2532–2539, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i9/p2532_s1.

Hu:1991:IPC

[800] Zhi Jin Hu. On the isovectors of the principal chiral model. Journal of Mathematical Physics, 32(9):2540–2542, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2540_s1.

Brankov:1991:FSS

[801] J. G. Brankov and D. M. Danchev. Finite-size scaling for the correlation function of the spherical model with long-range interactions. Journal of Mathematical Physics, 32(9):2543–2560, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i9/p2543_s1.

Zhang:1991:EFA

[802] Guihua Zhang and J. K. Percus. The effective field approximation for Ising spin glasses on simply connected lattices. Journal of Mathematical Physics, 32(9):2561–2562, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i9/p2561_s1.

Tomaschitz:1991:RQC

[803] Roman Tomaschitz. Relativistic quantum chaos in Robertson–Walker cosmologies. Journal of Mathematical Physics, 32(10):2571–2579, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2571_s1.

Carter:1991:SAS

[804] B. Carter and R. N. Henriksen. A systematic approach to self-similarity in Newtonian space–time. Journal of Mathematical Physics, 32(10):2580– 2597, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i10/p2580_s1.

Maekawa:1991:WRC

[805] Takayoshi Maekawa. On the Wigner and the Racah coefficients of su_{q}(2) and su_{q}(1, 1). Journal of Mathematical Physics, 32(10):2598–2604, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2598_s1.

Zhang:1991:GRT

[806] R. B. Zhang. Graded representations of the Temperley–Lieb algebra, quantum supergroups, and the Jones polynomial. Journal of Mathematical Physics, 32(10):2605–2613, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2605_s1.

Rowe:1991:VCS

[807] D. J. Rowe and J. Repka. Vector-coherent-state theory as a theory of induced representations. Journal of Mathematical Physics, 32(10):2614– 2634, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i10/p2614_s1.

Tamizhmani:1991:LSH

[808] K. M. Tamizhmani, A. Ramani, and B. Grammaticos. Lie symmetries of Hirota’s bilinear equations. Journal of Mathematical Physics, 32(10):2635–2659, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2635_s1.

Weimar-Woods:1991:CLA

[809] Evelyn Weimar-Woods. Contraction of Lie algebra representations. Journal of Mathematical Physics, 32(10):2660–2665, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2660_s1.

Savatier:1991:DPT

[810] F. Savatier. Deconfinement phase transition within the MIT bag model. Journal of Mathematical Physics, 32(10):2666–2675, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2666_s1.

Hu:1991:RCBb

[811] Ji Shan Hu and Martin Kruskal. Reflection coefficient beyond all orders for singular problems. II. Closed-spaced critical points on the nearest critical level line. Journal of Mathematical Physics, 32(10):2676– 2678, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i10/p2676_s1.

Saenz:1991:HOA

[812] A. W. Sáenz. Higher-order averaging for nonperiodic systems. Journal of Mathematical Physics, 32(10):2679–2694, October 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2679_s1.

Beals:1991:NIC

[813] Richard Beals and Boris Konopel′chenko. On nontrivial interactions and complete integrability of soliton equations. Journal of Mathematical Physics, 32(10):2695–2702, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2695_s1.

Calogero:1991:INPb

[814] F. Calogero and Ji Xiaoda. C-integrable nonlinear PDEs. II. Journal of Mathematical Physics, 32(10):2703–2717, October 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2703_s1.

Thron:1991:TSE

[815] Christopher P. Thron. Taylor series expansion of the dominating eigenvalue of the Ruelle–Araki transfer operator. Journal of Mathematical Physics, 32(10):2718–2721, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2718_s1.

Maia:1991:HOS

[816] M. D. Maia and W. L. Roque. Higher-order space–time contacts and equivalence. Journal of Mathematical Physics, 32(10):2722–2726, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2722_s1.

Kramer:1991:BTC

[817] D. Kramer, G. Neugebauer, and T. Matos. Bäcklund transforms of chiral fields. Journal of Mathematical Physics, 32(10):2727–2730, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2727_s1.

Nogueira:1991:STN

[818] P. Nogueira. N = 2 superprojective transformations and the nilpotent invariant. Journal of Mathematical Physics, 32(10):2731–2732, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2731_s1.

Das:1991:GFF

[819] Ashok Das, Wen-Jui Huang, and Shibaji Roy. A geometrical formulation of fermionic integrable systems. Journal of Mathematical Physics, 32(10):2733–2738, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2733_s1.

Goldberg:1991:HSF

[820] Joshua Goldberg, Ezra T. Newman, and Carlo Rovelli. On Hamiltonian systems with first-class constraints. Journal of Mathematical Physics, 32(10):2739–2743, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2739_s1.

Gracia:1991:HOL

[821] X. Gràcia, J. M. Pons, and N. Román-Roy. Higher-order Lagrangian systems: Geometric structures, dynamics, and constraints. Journal of Mathematical Physics, 32(10):2744–2763, October 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2744_s1.

Ranada:1991:ALC

[822] Manuel F. Rañada. Alternative Lagrangians for central potentials. Journal of Mathematical Physics, 32(10):2764–2769, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2764_s1.

Lahti:1991:SIC

[823] Pekka J. Lahti, Paul Busch, and Peter Mittelstaedt. Some important classes of quantum measurements and their information gain. Journal of Mathematical Physics, 32(10):2770–2775, October 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2770_s1.

Kvitsinsky:1991:SQC

[824] Andrei A. Kvitsinsky and Vadim V. Kostrykin. Semiclassical quantization for Coulomb systems on a hypersphere. Journal of Mathematical Physics, 32(10):2776–2782, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2776_s1.

Zhao:1991:MZF

[825] Zu Sen Zhao. Modified Zassenhaus formula and Stratonovich stochastic integrals. Journal of Mathematical Physics, 32(10):2783–2785, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2783_s1.

Aktosun:1991:SIS

[826] Tuncay Aktosun and Cornelis van der Mee. Scattering and inverse scattering for the 1-D Schrödinger equation with energy-dependent potentials. Journal of Mathematical Physics, 32(10):2786–2801, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/
p2786_s1.

Kvitsinsky:1991:QTB

[827] Andrei A. Kvitsinsky and Vadim V. Kostrykin. Quantum three-body scattering problem in the adiabatic hyperspherical representation. Journal of Mathematical Physics, 32(10):2802–2812, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2802_s1.

Folacci:1991:QFT

[828] A. Folacci. Quantum field theory of p-forms in curved space–time. Journal of Mathematical Physics, 32(10):2813–2827, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2813_s1.

Folacci:1991:TPF

[829] A. Folacci. Two-point functions and stress-energy tensors of p-forms in de Sitter and anti-de Sitter spaces. Journal of Mathematical Physics, 32(10):2828–2838, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2828_s1. See erratum [1140].

GuzmanSanchez:1991:DAT

[830] A. R. Guzmán Sánchez, J. F. Plebański, and M. Przanowski. Deformations of algebraic types of the energy-momentum tensor. Journal of Mathematical Physics, 32(10):2839–2847, October 1991. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i10/p2839_s1.

Hall:1991:CCGa

[831] G. S. Hall and J. da Costa. Curvature collineations in general relativity. I. Journal of Mathematical Physics, 32(10):2848–2853, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/
p2848_s1.

Hall:1991:CCGb

[832] G. S. Hall and J. da Costa. Curvature collineations in general relativity. II. Journal of Mathematical Physics, 32(10):2854–2862, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/
p2854_s1.

Wang:1991:PWI

[833] An Zhong Wang. Plane walls interacting with gravitational waves and matter fields. Journal of Mathematical Physics, 32(10):2863–2868, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i10/p2863_s1.

Baadhio:1991:GGI

[834] Randy A. Baadhio and Pascal Lee. On the global gravitational instanton and soliton that are homotopy spheres. Journal of Mathematical Physics, 32(10):2869–2874, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2869_s1.

Amann:1991:PTS

[835] Anton Amann. Perturbation theory of C^{∗}-systems without normcontinuity properties. Journal of Mathematical Physics, 32(10):2875– 2879, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i10/p2875_s1.

Capitani:1991:CCL

[836] Stefano Capitani. Convergence of compact lattice scalar field theory to its continuum limit. Journal of Mathematical Physics, 32(10):2880– 2885, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i10/p2880_s1.

Maximov:1991:ARR

[837] S. I. Maximov and V. I. Vovk. Asymptotic representations of the running parameters in Yang–Mills theories. Journal of Mathematical Physics, 32(10):2886–2889, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2886_s1.

Collins:1991:SQR

[838] Graham P. Collins and William I. Weisberger. Surface quadrupling and regularized off-shell amplitudes for unoriented, open, bosonic strings. Journal of Mathematical Physics, 32(10):2890–2902, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/
p2890_s1.

Calzetta:1991:NSC

[839] Esteban Calzetta. A necessary and sufficient condition for convergence to equilibrium in Kolmogorov systems. Journal of Mathematical Physics, 32(10):2903–2906, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2903_s1.

Goldstein:1991:TFT

[840] Jerome A. Goldstein and Gisèle Ruiz Rieder. Thomas–Fermi theory with an external magnetic field. Journal of Mathematical Physics, 32 (10):2907–2917, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i10/p2907_s1.

Evans:1991:AOB

[841] J. W. Evans and H. C. Kang. Analytic observations for the d = 1 + 1 bridge site (or single-step) deposition model. Journal of Mathematical Physics, 32(10):2918–2922, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2918_s1.

Porsezian:1991:DRS

[842] K. Porsezian and M. Lakshmanan. On the dynamics of the radially symmetric Heisenberg ferromagnetic spin system. Journal of Mathematical Physics, 32(10):2923–2928, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2923_s1.

Broidioi:1991:OME

[843] Marino Broidioi, Bruno Nachtergaele, and André Verbeure. The Overhauser model: Equilibrium fluctuation dynamics. Journal of Mathematical Physics, 32(10):2929–2935, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i10/p2929_s1.

Davies:1991:OAD

[844] Brian Davies. Onsager’s algebra and the Dolan–Grady condition in the non-self-dual case. Journal of Mathematical Physics, 32(11):2945–2950, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p2945_s1.

Ekhaguere:1991:CPC

[845] G. O. S. Ekhaguere and P. O. Odiobala. Completely positive conjugatebilinear maps on partial ∗-algebras. Journal of Mathematical Physics, 32(11):2951–2958, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p2951_s1.

Duplij:1991:SNS

[846] Steven Duplij. On the semigroup nature of superconformal symmetry. Journal of Mathematical Physics, 32(11):2959–2966, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/
p2959_s1.

Jumarie:1991:QEN

[847] Guy Jumarie. Quantum entropies of nonprobabilistic matrices. Journal of Mathematical Physics, 32(11):2967–2971, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/p2967_s1.

Uehara:1991:MGS

[848] Shozo Uehara and Yukinori Yasui. On the monodromy group for the super Schwarzian differential equation. Journal of Mathematical Physics, 32(11):2972–2978, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p2972_s1.

Vokos:1991:MDJ

[849] Stamatis P. Vokos. Multiparameter deformations of GL(N) and Jordan– Schwinger constructions. Journal of Mathematical Physics, 32(11):2979– 2985, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p2979_s1.

Maekawa:1991:NUI

[850] T. Maekawa. Note on the unitary irreducible representations of U(n) in the scalar product with the nonintertwining operator. Journal of Mathematical Physics, 32(11):2986–2987, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i11/p2986_s1.

Ibragimov:1991:PGC

[851] N. H. Ibragimov, M. Torrisi, and A. Valenti. Preliminary group classification of equations v_{tt} = f(x, v_{x})v_{xx} + g(x, v_{x}). Journal of Mathematical Physics, 32(11):2988–2995, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i11/p2988_s1.

Dickey:1991:FMH

[852] L. A. Dickey. On the τ-function of matrix hierarchies of integrable equations. Journal of Mathematical Physics, 32(11):2996–3002, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i11/p2996_s1.

Heiss:1991:ALC

[853] W. D. Heiss and W.-H. Steeb. Avoided level crossings and Riemann sheet structure. Journal of Mathematical Physics, 32(11):3003–3007, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p3003_s1.

Kirsten:1991:IME

[854] Klaus Kirsten. Inhomogeneous multidimensional Epstein zeta functions. Journal of Mathematical Physics, 32(11):3008–3014, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/
p3008_s1.

Hinton:1991:ISL

[855] D. B. Hinton, A. K. Jordan, M. Klaus, and J. K. Shaw. Inverse scattering on the line for a Dirac system. Journal of Mathematical Physics, 32 (11):3015–3030, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3015_s1.

Bender:1991:NPA

[856] Carl M. Bender, Stefan Boettcher, and Kimball A. Milton. A new perturbative approach to nonlinear partial differential equations. Journal of Mathematical Physics, 32(11):3031–3038, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/p3031_s1.

Law:1991:NEM

[857] Peter R. Law. Neutral Einstein metrics in four dimensions. Journal of Mathematical Physics, 32(11):3039–3042, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/p3039_s1.

Fusco:1991:CLM

[858] D. Fusco and N. Manganaro. A class of linearizable models and generation of material response functions to nonlinear hyperbolic heat conduction. Journal of Mathematical Physics, 32(11):3043–3046, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i11/p3043_s1.

Mignemi:1991:CQM

[859] S. Mignemi. Classical and quantum motion in an Eguchi–Hanson space. Journal of Mathematical Physics, 32(11):3047–3054, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/
p3047_s1.

Zhang:1991:CAI

[860] Peng Zhang. On a conjecture about the invariance of the collision between solitary waves. Journal of Mathematical Physics, 32(11):3055– 3059, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p3055_s1.

Dorren:1991:MEN

[861] H. J. S. Dorren and A. Tip. Maxwell’s equations for nonsmooth media: fractal-shaped and pointlike objects. Journal of Mathematical Physics, 32(11):3060–3070, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3060_s1.

Cooke:1991:CCH

[862] David B. Cooke. Compatibility conditions for Hamiltonian pairs. Journal of Mathematical Physics, 32(11):3071–3076, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/p3071_s1.

Ban:1991:RNS

[863] Masashi Ban. Relative number state representation and phase operator for physical systems. Journal of Mathematical Physics, 32(11):3077– 3087, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p3077_s1.

Durdevic:1991:SHV

[864] Mićo Durdević, Milan Vujičić, and Fedor Herbut. Symplectic hiddenvariables theories—the missing link in algebraic contextual approaches. Journal of Mathematical Physics, 32(11):3088–3093, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/
p3088_s1.

Beckers:1991:SHO

[865] J. Beckers and N. Debergh. On supersymmetric harmonic oscillators and the Green–Cusson ansätze. Journal of Mathematical Physics, 32 (11):3094–3100, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3094_s1.

deMonvel-Berthier:1991:GAB

[866] Anne Boutet de Monvel-Berthier and Vladimir Georgescu. Graded C^{∗}- algebras in the N-body problem. Journal of Mathematical Physics, 32 (11):3101–3110, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3101_s1.

Common:1991:BKE

[867] A. K. Common. Bounds on the kinetic energy, mean radius, and electric dipole moment of ground states in a class of potential models of quarkonium. Journal of Mathematical Physics, 32(11):3111–3116, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i11/p3111_s1.

Gibson:1991:NMA

[868] A. G. Gibson, A. J. Waters, G. H. Berthold, and C. Chandler. A new Kmatrix approach to N-body scattering. Journal of Mathematical Physics, 32(11):3117–3124, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3117_s1.

Fricke:1991:UAC

[869] S. H. Fricke, A. B. Balantekin, and T. Uzer. Uniform approximation and coherent state path integrals. Journal of Mathematical Physics, 32 (11):3125–3129, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3125_s1.

Newton:1991:USQ

[870] Roger G. Newton. Uniqueness in some quasi-Goursat problems in 3 + 1 dimensions and the inverse scattering problem. Journal of Mathematical Physics, 32(11):3130–3134, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i11/p3130_s1.

Carminati:1991:AIR

[871] J. Carminati and R. G. McLenaghan. Algebraic invariants of the Riemann tensor in a four-dimensional Lorentzian space. Journal of Mathematical Physics, 32(11):3135–3140, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i11/p3135_s1.

Maszczyk:1991:FMO

[872] Roman A. Maszczyk. Fluid of multidimensional objects in cosmology. Journal of Mathematical Physics, 32(11):3141–3147, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/
p3141_s1.

Perlick:1991:BPS

[873] V. Perlick. The brachistochrone problem in a stationary space–time. Journal of Mathematical Physics, 32(11):3148–3157, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/
p3148_s1.

Bengtsson:1991:SDM

[874] Ingemar Bengtsson. Self-duality and the metric in a family of neighbors of Einstein’s equations. Journal of Mathematical Physics, 32(11):3158– 3161, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p3158_s1.

Koch:1991:SLG

[875] Hans Koch. A scaling limit of the Glauber dynamics for a class of neural network models. Journal of Mathematical Physics, 32(11):3162–3167, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p3162_s1.

Kheyfets:1991:BBP

[876] Arkady Kheyfets and Warner A. Miller. The boundary of a boundary principle in field theories and the issue of austerity of the laws of physics. Journal of Mathematical Physics, 32(11):3168–3175, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i11/p3168_s1.

Sophocleous:1991:CBT

[877] C. Sophocleous and J. G. Kingston. A class of Bäcklund transformations for equations of the type u_{xy} = f(u, u_{x}). Journal of Mathematical Physics, 32(11):3176–3183, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i11/p3176_s1.

Shishkin:1991:DEE

[878] German V. Shishkin and William D. Cabos. The Dirac equation in external fields: Variable separation in Cartesian coordinates. Journal of Mathematical Physics, 32(11):3184–3188, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/p3184_s1.

Lopuszanski:1991:APL

[879] Jan Łopuszański. On the action of the Pauli–Lubanski Casimir operator in a relativistic supersymmetric quantum field theory. Journal of Mathematical Physics, 32(11):3189–3194, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i11/p3189_s1.

Bassetto:1991:CAV

[880] A. Bassetto and L. Griguolo. Chiral anomalies for vortex potentials in two dimensions and a decompactification limit. Journal of Mathematical Physics, 32(11):3195–3198, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i11/p3195_s1.

Balaban:1991:PBR

[881] Tadeusz Bałaban, Michael O’Carroll, and Ricardo Schor. Properties of block renormalization group operators for Euclidean fermions in an external gauge field. Journal of Mathematical Physics, 32(11):3199–3208, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i11/p3199_s1.

Avellaneda:1991:EDI

[882] Marco Avellaneda. Enhanced diffusivity and intercell transition layers in 2-D models of passive advection. Journal of Mathematical Physics, 32(11):3209–3212, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i11/p3209_s1.

Melkonian:1991:SPS

[883] S. Melkonian and P. Winternitz. Symmetry properties and solutions of nonlinear dispersive thin-film equations in three dimensions. Journal of Mathematical Physics, 32(11):3213–3222, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i11/p3213_s1.

MonteiroS:1991:SCG

[884] W. Monteiro S., Jr. Stably causal Gödel-type models. Journal of Mathematical Physics, 32(11):3223–3227, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i11/p3223_s1.

Halpern:1991:GGU

[885] M. B. Halpern and N. A. Obers. Generalized graphs and unitary irrational central charge in the superconformal master equation. Journal of Mathematical Physics, 32(12):3231–3240, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3231_s1.

Chang:1991:RQG

[886] Zhe Chang, Jian Xiong Wang, and Hong Yan. The realizations of quantum groups of A_{n−1} and C_{n} types in q-deformed oscillator systems at classical and quantum levels. Journal of Mathematical Physics, 32 (12):3241–3245, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3241_s1.

Majid:1991:EBG

[887] Shahn Majid. Examples of braided groups and braided matrices. Journal of Mathematical Physics, 32(12):3246–3253, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3246_s1.

Wehrhahn:1991:SS

[888] R. F. Wehrhahn, Yu. F. Smirnov, and A. M. Shirokov. Symmetry scattering. Journal of Mathematical Physics, 32(12):3254–3260, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i12/p3254_s1.

Zhang:1991:UMI

[889] R. B. Zhang and M. D. Gould. Universal R matrices and invariants of quantum supergroups. Journal of Mathematical Physics, 32(12):3261– 3267, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3261_s1.

Frappat:1991:DLD

[890] L. Frappat, A. Sciarrino, and P. Sorba. Dynkin-like diagrams and representations of the strange superalgebra P(n). Journal of Mathematical Physics, 32(12):3268–3277, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i12/p3268_s1. See erratum [1250].

Gingold:1991:TSR

[891] Harry Gingold and Ji Shan Hu. Transcendentally small reflection of waves for problems with/without turning points near infinity: a new uniform approach. Journal of Mathematical Physics, 32(12):3278–3284, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3278_s1.

Darkhosh:1991:HIW

[892] Teymour Darkhosh. Hyperbolicity of invariant wave equations. Journal of Mathematical Physics, 32(12):3285–3287, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3285_s1.

Dorfman:1991:LSO

[893] I. Ya. Dorfman and O. I. Mokhov. Local symplectic operators and structures related to them. Journal of Mathematical Physics, 32(12):3288– 3296, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3288_s1.

Inayat-Hussain:1991:SCC

[894] A. A. Inayat-Hussain. Steady currents in composite conductors, eigenfunction expansions, and exactly solvable linear integral equations. Journal of Mathematical Physics, 32(12):3297–3311, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3297_s1.

Kopec:1991:EBM

[895] S. Kopeć. Error bounds for maximum entropy approximate solutions to Fredholm integral equations. Journal of Mathematical Physics, 32 (12):3312–3314, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3312_s1.

Nesterenko:1991:CTW

[896] V. V. Nesterenko. Curvature and torsion of the world curve in the action of the relativistic particle. Journal of Mathematical Physics, 32(12):3315– 3320, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3315_s1.

Claude:1991:NRS

[897] C. Claude, A. Latifi, and J. Leon. Nonlinear resonant scattering and plasma instability: an integrable model. Journal of Mathematical Physics, 32(12):3321–3330, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i12/p3321_s1.

Rollins:1991:PAL

[898] David K. Rollins. Painlevé analysis and Lie group symmetries of the regularized long-wave equation. Journal of Mathematical Physics, 32 (12):3331–3332, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3331_s1.

Okubo:1991:KMN

[899] Susumu Okubo. Kac–Moody and new infinite-dimensional Lie algebras. Journal of Mathematical Physics, 32(12):3333–3338, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i12/p3333_s1.

Vaisman:1991:GQP

[900] Izu Vaisman. On the geometric quantization of Poisson manifolds. Journal of Mathematical Physics, 32(12):3339–3345, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3339_s1.

Aldaya:1991:QDS

[901] V. Aldaya, R. Loll, and J. Navarro-Salas. Quantum dynamics on (super)groups: constraints and the “BRST” supergroup. Journal of Mathematical Physics, 32(12):3346–3360, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v32/i12/p3346_s1.

Colegrave:1991:CMD

[902] R. K. Colegrave and P. Croxson. Constants of the motion and diagonalization of the Hamiltonian for some scattering processes. Journal of Mathematical Physics, 32(12):3361–3368, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3361_s1.

Evans:1991:GTS

[903] N. W. Evans. Group theory of the Smorodinsky–Winternitz system. Journal of Mathematical Physics, 32(12):3369–3375, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/
p3369_s1.

Burzlaff:1991:SVV

[904] J. Burzlaff and P. McCarthy. A study of a 90^{∘} vortex-vortex scattering process. Journal of Mathematical Physics, 32(12):3376–3380, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i12/p3376_s1.

Chaichian:1991:PIS

[905] M. Chaichian, D. Ellinas, and P. Prešnajder. Path integrals and supercoherent states. Journal of Mathematical Physics, 32(12):3381–3391, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3381_s1.

Vinette:1991:ULB

[906] F. Vinette and J. Čížek. Upper and lower bounds of the ground state energy of anharmonic oscillators using renormalized inner projection. Journal of Mathematical Physics, 32(12):3392–3404, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3392_s1.

Steeb:1991:SJC

[907] W.-H. Steeb, N. Euler, and P. Mulser. Semiclassical Jaynes–Cummings model, Painlevé test, and exact solutions. Journal of Mathematical Physics, 32(12):3405–3406, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i12/p3405_s1.

Jiang:1991:NUB

[908] Wenxin Jiang. A note on the universal behavior of Schwinger terms in higher-dimensional conformal field theories. Journal of Mathematical Physics, 32(12):3407–3408, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i12/p3407_s1.

Jiang:1991:AWI

[909] Wenxin Jiang. The anomalous Ward identities in gauge and gravitational theories. Journal of Mathematical Physics, 32(12):3409–3411, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i12/p3409_s1.

Rosquist:1991:KTT

[910] Kjell Rosquist and Claes Uggla. Killing tensors in two-dimensional space– times with applications to cosmology. Journal of Mathematical Physics, 32(12):3412–3422, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3412_s1.

Kozameh:1991:GAL

[911] Carlos Kozameh, Pedro W. Lamberti, and Oscar Reula. Global aspects of light cone cuts. Journal of Mathematical Physics, 32(12):3423–3426, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3423_s1.

Page:1991:MCM

[912] Don N. Page. Minisuperspaces with conformally and minimally coupled scalar fields. Journal of Mathematical Physics, 32(12):3427–3438, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3427_s1.

Cornille:1991:DHM

[913] Henri Cornille. (1 + 1)-dimensional hierarchies of multispeed discrete Boltzmann model equations. Journal of Mathematical Physics, 32 (12):3439–3454, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3439_s1.

deSouza:1991:GAF

[914] S. M. de Souza and M. T. Thomaz. Grassmann algebra and fermions in the lattice: A simple example. Journal of Mathematical Physics, 32 (12):3455–3462, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3455_s1.

Haba:1991:EIM

[915] Z. Haba. Ergodicity and invariant measures of some randomly perturbed classical fields. Journal of Mathematical Physics, 32(12):3463– 3472, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3463_s1.

Chowdhury:1991:CDO

[916] A. Roy Chowdhury and N. Das Gupta. Commutative differential operator and a class of solutions for the BKP equation on the singular rational curve. Journal of Mathematical Physics, 32(12):3473–3475, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i12/p3473_s1.

Manolessou:1991:ZDP

[917] M. Manolessou. The zero-dimensional problem in Φ^{4} field theory. Journal of Mathematical Physics, 32(12):3476–3487, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3476_s1.

Fushchich:1991:NLR

[918] W. I. Fushchich and R. Z. Zhdanov. On the non-Lie reduction of the nonlinear Dirac equation. Journal of Mathematical Physics, 32(12):3488– 3490, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3488_s1.

Bandelloni:1991:RGC

[919] Giuseppe Bandelloni. Renormalization of gauge coset models. Journal of Mathematical Physics, 32(12):3491–3499, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3491_s1.

deAzcarraga:1991:NLS

[920] José A. de Azcárraga and Damián Ginestar. Nonrelativistic limit of supersymmetric theories. Journal of Mathematical Physics, 32(12):3500– 3508, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3500_s1.

Motovilov:1991:RED

[921] A. K. Motovilov. The removal of an energy dependence from the interaction in two-body systems. Journal of Mathematical Physics, 32 (12):3509–3518, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i12/p3509_s1.

Moshinsky:1991:RIM

[922] M. Moshinsky and G. López Laurrabaquio. Relativistic interactions by means of boundary conditions: The Breit–Wigner formula. Journal of Mathematical Physics, 32(12):3519–3528, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3519_s1.

He:1991:ESP

[923] Sailing He and Staffan Ström. The electromagnetic scattering problem in the time domain for a dissipative slab and a point source using invariant imbedding. Journal of Mathematical Physics, 32(12):3529–3539, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v32/i12/p3529_s1.

Greenberg:1991:RHP

[924] W. Greenberg, P. F. Zweifel, and S. Paveri-Fontana. The Riemann– Hilbert problem for nonsymmetric systems. Journal of Mathematical Physics, 32(12):3540–3545, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i12/p3540_s1. See erratum [1098].

PoncedeLeon:1991:CMD

[925] J. Ponce de León. Cosmological models that describe particle creation in the early universe and evolve into the “present-day” universe. Journal of Mathematical Physics, 32(12):3546–3552, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v32/i12/p3546_s1.

Johannesson:1992:LVA

[926] Henrik Johannesson. Lattice Virasoro algebras, unitarity, and the eightvertex model. Journal of Mathematical Physics, 33(1):1–3, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p1_s1.

King:1992:RTH

[927] R. C. King and B. G. Wybourne. Representations and traces of the Hecke algebras H_{n}(q) of type A_{n−1}. Journal of Mathematical Physics, 33(1): 4–14, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p4_s1.

Rao:1992:ACT

[928] K. Srinivasa Rao, J. Van der Jeugt, and G. Vanden Berghe. On the algebra of coupled SO(3) tensors. Journal of Mathematical Physics, 33(1): 15–18, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p15_s1.

Brzezinski:1992:QDC

[929] T. Brzeziński, H. Dabrowski, and J. Rembieliński. On the quantum differential calculus and the quantum holomorphicity. Journal of Mathematical Physics, 33(1):19–24, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i1/p19_s1.

Dattoli:1992:GFM

[930] G. Dattoli, C. Chiccoli, S. Lorenzutta, G. Maino, M. Richetta, and A. Torre. Generating functions of multivariable generalized Bessel functions and Jacobi-elliptic functions. Journal of Mathematical Physics, 33(1):25–36, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p25_s1.

Hu:1992:RCB

[931] Ji Shan Hu. Reflection coefficient beyond all orders for singular problems. III. Essential singularities on the real axis. Journal of Mathematical Physics, 33(1):37–43, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p37_s1.

Newton:1992:RSH

[932] Roger G. Newton. A representation of solutions of the Helmholtz equation with application to crystal Green’s functions. Journal of Mathematical Physics, 33(1):44–55, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i1/p44_s1.

Lu:1992:GAR

[933] Jing Fa Lü, Mo Lin Ge, and Chao Liu. On the group aspects of Riemann– Hilbert transformations. Journal of Mathematical Physics, 33(1):56– 63, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p56_s1.

Buchdahl:1992:GIP

[934] H. A. Buchdahl. On gauge-invariant and phase-invariant spinor analysis. II. Journal of Mathematical Physics, 33(1):64–67, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i1/p64_s1.

Gielerak:1992:PDC

[935] Roman Gielerak. On the phase diagram for a class of continuous systems. Journal of Mathematical Physics, 33(1):68–83, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i1/p68_s1.

Ungar:1992:NLT

[936] Abraham A. Ungar. A note on the Lorentz transformations linking initial and final four-vectors. Journal of Mathematical Physics, 33(1):84– 85, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p84_s1.

Moses:1992:GST

[937] Harry E. Moses, Raymond J. Nagem, and Guido v. H. Sandri. The general solution of the three-dimensional acoustic equation and of Maxwell’s equations in the infinite domain in terms of the asymptotic solution in the wave zone. Journal of Mathematical Physics, 33(1):86–101, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p86_s1.

Strachan:1992:LVS

[938] I. A. B. Strachan. Low-velocity scattering of vortices in a modified Abelian Higgs model. Journal of Mathematical Physics, 33(1):102– 110, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p102_s1.

Kostin:1992:SFQ

[939] M. D. Kostin. Schrödinger-Fuerth quantum diffusion theory: generalized complex diffusion equation. Journal of Mathematical Physics, 33(1):111– 113, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p111_s1.

Basu:1992:BGC

[940] Debabrata Basu. The Barut–Girardello coherent states. Journal of Mathematical Physics, 33(1):114–121, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i1/p114_s1.

Ragiadakos:1992:QFD

[941] C. N. Ragiadakos. Quantization of a four-dimensional generally covariant conformal model. Journal of Mathematical Physics, 33(1):122–130, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p122_s1.

Singer:1992:PSR

[942] Matthias Singer and Werner Stulpe. Phase-space representations of general statistical physical theories. Journal of Mathematical Physics, 33(1):131–142, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p131_s1.

Hertle:1992:LGS

[943] Jochen Hertle and Reinhard Honegger. Limiting Gibbs states and dynamics for thermal photons. Journal of Mathematical Physics, 33(1):143– 151, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p143_s1.

Beckers:1992:SNQ

[944] J. Beckers, N. Debergh, and A. G. Nikitin. On supersymmetries in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 33(1):152–160, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p152_s1.

Keller:1992:GSD

[945] Jaime Keller and Adán Rodríguez. Geometric superalgebra and the Dirac equation. Journal of Mathematical Physics, 33(1):161–170, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p161_s1.

Marchiafava:1992:QQ

[946] Stefano Marchiafava and Jakub Rembieliński. Quantum quaternions. Journal of Mathematical Physics, 33(1):171–173, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i1/p171_s1.

Wehrhahn:1992:LBS

[947] R. F. Wehrhahn and Ian L. Cooper. Local behavior in supersymmetric quantum mechanics. Journal of Mathematical Physics, 33(1):174– 177, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p174_s1.

Zelenov:1992:RCR

[948] E. I. Zelenov. Representations of commutations relations for p-adic systems of infinitely many degrees of freedom. Journal of Mathematical Physics, 33(1):178–188, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i1/p178_s1.

Ahia:1992:FDS

[949] Francis Ahia. Finiteness of the discrete spectrum of (N ≥ 3) N-body Schrödinger operators, which have some determinate subsystems that are virtual at the bottom of the continuum. Journal of Mathematical Physics, 33(1):189–202, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i1/p189_s1.

Mayrand:1992:HST

[950] Michel Mayrand and Luc Vinet. Hidden symmetries of two-dimensional systems with local degeneracies. Journal of Mathematical Physics, 33(1):203–212, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p203_s1.

Znojil:1992:AAO

[951] Miloslav Znojil. Asymmetric anharmonic oscillators in the Hilldeterminant picture. Journal of Mathematical Physics, 33(1):213–221, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p213_s1.

Cognola:1992:RDV

[952] Guido Cognola, Luciano Vanzo, and Sergio Zerbini. Regularization dependence of vacuum energy in arbitrarily shaped cavities. Journal of Mathematical Physics, 33(1):222–228, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i1/p222_s1.

Ozsvath:1992:EPI

[953] István Ozsváth. Embedding problem. II. Journal of Mathematical Physics, 33(1):229–247, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i1/p229_s1.

Wald:1992:IEY

[954] Robert M. Wald. On the instability of the n = 1 Einstein–Yang–Mills black holes and mathematically related systems. Journal of Mathematical Physics, 33(1):248–255, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i1/p248_s1.

Williams:1992:KMD

[955] J. G. Williams and P. Zvengrowski. Kink metrics in (2 + 1)-dimensional space–time. Journal of Mathematical Physics, 33(1):256–266, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p256_s1.

Bona:1992:IGT

[956] Carles Bona and Bartolomé Coll. Isometry groups of three-dimensional Riemannian metrics. Journal of Mathematical Physics, 33(1):267–272, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p267_s1.

Galtsov:1992:GTE

[957] Dmitry V. Gal′tsov and Basilis C. Xanthopoulos. A generating technique for Einstein gravity conformally coupled to a scalar field with Higgs potential. Journal of Mathematical Physics, 33(1):273–277, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p273_s1.

Gao:1992:NDS

[958] Ya Jun Gao and Zai Zhe Zhong. Nondiagonal double seed solutions and double soliton solution family of the Einstein equations. Journal of Mathematical Physics, 33(1):278–285, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i1/p278_s1.

Kalnins:1992:SSD

[959] E. G. Kalnins and W. Miller, Jr. Series solutions for the Dirac equation in Kerr–Newman space–time. Journal of Mathematical Physics, 33(1):286– 296, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p286_s1.

Shishkin:1992:DEEa

[960] German V. Shishkin and William D. Cabos. Dirac equation in external fields: Separation of variables in nondiagonal metrics. Journal of Mathematical Physics, 33(1):297–303, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i1/p297_s1.

Wahlquist:1992:PEI

[961] H. D. Wahlquist. The problem of exact interior solutions for rotating rigid bodies in general relativity. Journal of Mathematical Physics, 33(1):304– 335, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p304_s1. See erratum [1282].

Song:1992:DCS

[962] Jong June Song, Sam Nyung Yi, and Sang Don Choi. Derivation of correlation spectra by the operator algebra technique. Journal of Mathematical Physics, 33(1):336–342, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i1/p336_s1.

Hertle:1992:RDP

[963] Jochen Hertle and Reinhard Honegger. Rigorous derivation of Planck’s law in the thermodynamic limit. Journal of Mathematical Physics, 33(1):343–348, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p343_s1.

Sarkar:1992:SCD

[964] Biplab Sarkar and Kamal Bhattacharyya. Solution of convergence difficulties in the Madelung-sum problem: an extrapolation scheme for sawtooth sequences. Journal of Mathematical Physics, 33(1):349–357, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i1/p349_s1.

Davydychev:1992:GRM

[965] A. I. Davydychev. General results for massive N-point Feynman diagrams with different masses. Journal of Mathematical Physics, 33 (1):358–369, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p358_s1.

Evans:1992:HIR

[966] T. S. Evans, I. Hardman, H. Umezawa, and Y. Yamanaka. Heisenberg and interaction representations in thermo field dynamics. Journal of Mathematical Physics, 33(1):370–378, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i1/p370_s1.

Ahmed:1992:IDF

[967] E. Ahmed and A. S. Hegazi. Infinitesimally deformed field and string theories. Journal of Mathematical Physics, 33(1):379–381, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i1/
p379_s1.

Chakravarty:1992:DAN

[968] Sarbarish Chakravarty, Steven Kent, and Ezra T. Newman. Diffeomorphism algebras and the Nahm and Ward equations. Journal of Mathematical Physics, 33(1):382–387, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i1/p382_s1.

Cederwall:1992:OPS

[969] Martin Cederwall. Octonionic particles and the S^{7} symmetry. Journal of Mathematical Physics, 33(1):388–393, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i1/p388_s1.

Debergh:1992:PSC

[970] N. Debergh. Parasupercoherent states and color supergroups. Journal of Mathematical Physics, 33(1):394–398, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i1/p394_s1.

Jarvis:1992:UMGa

[971] P. D. Jarvis and D. S. McAnally. Unified models from gauged supergroups. II. Exceptional series D(2, 1; α), G(3), F (4). Journal of Mathematical Physics, 33(1):399–402, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i1/p399_s1.

Kerner:1992:GAC

[972] Richard Kerner. Z_{3}-graded algebras and the cubic root of the supersymmetry translations. Journal of Mathematical Physics, 33(1):403– 411, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p403_s1.

Matsuno:1992:TDD

[973] Y. Matsuno. Two-dimensional dynamical system associated with Abel’s nonlinear differential equation. Journal of Mathematical Physics, 33 (1):412–421, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i1/p412_s1.

Fu:1992:BRP

[974] Hong Chen Fu and Mo Lin Ge. The q-boson realization of parametrized cyclic representations of quantum algebras at q^{p} = 1. Journal of Mathematical Physics, 33(2):427–435, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i2/p427_s1.

Rittenberg:1992:TOQ

[975] V. Rittenberg and M. Scheunert. Tensor operators for quantum groups and applications. Journal of Mathematical Physics, 33(2):436–445, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p436_s1.

Lu:1992:QAP

[976] Zhi Cheng Lu. Quantum algebras on phase space. Journal of Mathematical Physics, 33(2):446–453, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i2/p446_s1.

Shin:1992:POC

[977] H. J. Shin. Primary operators of coset conformal field theories. Journal of Mathematical Physics, 33(2):454–461, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p454_s1.

Pan:1992:MRB

[978] Feng Pan and Yu Fang Cao. Matrix representations of SO_{n+2} in an SO_{n}×SO_{2} basis and some isoscalar factors for SO_{n+2} ⊃ SO_{n}×SO_{2}. Journal of Mathematical Physics, 33(2):462–469, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i2/p462_s1.

Hughes:1992:CFK

[979] J. W. B. Hughes, R. C. King, and J. Van der Jeugt. On the composition factors of Kac modules for the Lie superalgebras sl(m/n). Journal of Mathematical Physics, 33(2):470–491, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p470_s1.

Juriev:1992:ROA

[980] D. V. Juriev. On the representation of operator algebras of the quantum conformal field theory. Journal of Mathematical Physics, 33(2):492– 496, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p492_s1.

Koca:1992:SOR

[981] Mehmet Koca. Symmetries of the octonionic root system of E_{8}. Journal of Mathematical Physics, 33(2):497–510, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p497_s1.

Negro:1992:BRI

[982] Javier Negro and Mariano A. del Olmo. Bundle realizations and invariant connections in an Abelian principal bundle. Journal of Mathematical Physics, 33(2):511–523, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i2/p511_s1.

Exton:1992:GRT

[983] Harold Exton. Generating relations for Tratnik’s multivariable biorthogonal continuous Hahn polynomials. Journal of Mathematical Physics, 33(2):524–527, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i2/p524_s1.

Calogero:1992:SNI

[984] F. Calogero. Solvable nonlinear integrodifferential equations of Boltzmann type. I. Journal of Mathematical Physics, 33(2):528–536, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i2/p528_s1.

Rabelo:1992:CPS

[985] Mauro L. Rabelo and Keti Tenenblat. A classification of pseudospherical surface equations of type u_{t} = u_{xxx} + G(u, u_{x}, u_{xx}). Journal of Mathematical Physics, 33(2):537–549, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p537_s1.

Kopczynski:1992:SSR

[986] Wojciech Kopczyński and Andrzej Trautman. Simple spinors and real structures. Journal of Mathematical Physics, 33(2):550–559, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i2/p550_s1.

Yamaguchi:1992:FTS

[987] Yoshihiro Yamaguchi and Kiyotaka Tanikawa. On the first tangency of stable and unstable manifolds in C^{r} planar diffeomorphisms. Journal of Mathematical Physics, 33(2):560–569, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p560_s1.

Garcia-Pelayo:1992:NF

[988] Ricardo García-Pelayo and William C. Schieve. Noisy fractals. Journal of Mathematical Physics, 33(2):570–577, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p570_s1.

Arnoldus:1992:AEE

[989] Henk F. Arnoldus and Thomas F. George. Analytical evaluation of elastic Coulomb integrals. Journal of Mathematical Physics, 33(2):578– 583, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p578_s1.

Hojman:1992:LDE

[990] Sergio Hojman, Francisco Pardo, Luis Aulestia, and Francisco de Lisa. Lagrangians for differential equations of any order. Journal of Mathematical Physics, 33(2):584–590, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i2/p584_s1.

Lewis:1992:ROD

[991] H. Ralph Lewis, P. G. L. Leach, Serge Bouquet, and Marc R. Feix. Representations of one-dimensional Hamiltonians in terms of their invariants. Journal of Mathematical Physics, 33(2):591–598, February 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i2/p591_
s1.

Perlick:1992:HPG

[992] Volker Perlick. The Hamiltonization problem from a global viewpoint. Journal of Mathematical Physics, 33(2):599–606, February 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i2/p599_
s1.

Zakharov:1992:HFN

[993] O. A. Zakharov. Hamiltonian formalism for nonregular Lagrangian theories in fibered manifolds. Journal of Mathematical Physics, 33(2):607– 611, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p607_s1.

Stephas:1992:ASW

[994] Paul Stephas. Analytic solutions for Wheeler–Feynman interaction: Two bodies in straight-line motion. Journal of Mathematical Physics, 33(2):612–624, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i2/p612_s1.

Sonego:1992:HPC

[995] S. Sonego and V. Faraoni. Huygens’ principle and characteristic propagation property for waves in curved space–times. Journal of Mathematical Physics, 33(2):625–632, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i2/p625_s1.

Kulshreshtha:1992:HFT

[996] Usha Kulshreshtha. Hamiltonian formulation of a theory with constraints. Journal of Mathematical Physics, 33(2):633–638, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i2/

p633_s1. See comment [1372].

Khalil:1992:RBT

[997] A. B. Khalil. A relation between two perturbation methods generating Padé upper and lower bounds to ground state eigenvalues. Journal of Mathematical Physics, 33(2):639–642, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p639_s1.

Kleinert:1992:SSR

[998] H. Kleinert and I. Mustapic. Summing the spectral representations of Pöschl–Teller and Rosen–Morse fixed-energy amplitudes. Journal of Mathematical Physics, 33(2):643–662, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p643_s1.

Hooshyar:1992:IST

[999] M. A. Hooshyar and Lakshman S. Tamil. Inverse scattering theory at fixed energy and the design of circular optical waveguides. Journal of Mathematical Physics, 33(2):663–669, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p663_s1.

Bona:1992:ICS

[1000] Carles Bona, Bartolomé Coll, and Juan Antonio Morales. Intrinsic characterization of space–time symmetric tensors. Journal of Mathematical Physics, 33(2):670–681, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i2/p670_s1.

Gursey:1992:EST

[1001] Feza Gürsey and Wenxin Jiang. Euclidean space–time diffeomorphisms and their Fueter subgroups. Journal of Mathematical Physics, 33(2):682– 700, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p682_s1.

Krolak:1992:SCS

[1002] Andrzej Królak. Strong curvature singularities and causal simplicity. Journal of Mathematical Physics, 33(2):701–704, February 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i2/p701_
s1.

Bona:1992:DMC

[1003] C. Bona and P. Palou. Dust metrics with comoving constant curvature slices. Journal of Mathematical Physics, 33(2):705–711, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i2/p705_s1.

deFelice:1992:SIM

[1004] F. de Felice, S. Usseglio-Tomasset, and M. Bradley. Singularities inside matter: The static case. Journal of Mathematical Physics, 33(2):712– 720, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p712_s1.

Baadhio:1992:GGI

[1005] Randy A. Baadhio. Global gravitational instantons and their degrees of symmetry. Journal of Mathematical Physics, 33(2):721–724, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i2/p721_s1.

Kachkachi:1992:CSS

[1006] M. Kachkachi and T. Lhallabi. N = 2 conformal supergravity and superconformal anomaly. Journal of Mathematical Physics, 33(2):725– 734, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p725_s1.

Merkulov:1992:SGT

[1007] S. A. Merkulov. Superconformal geometry in three dimensions. Journal of Mathematical Physics, 33(2):735–757, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p735_s1.

Warner:1992:MPD

[1008] Roland C. Warner and R. Delbourgo. Morphology of the period-doubling universal function. Journal of Mathematical Physics, 33(2):758–770, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p758_s1.

Daniel:1992:SSA

[1009] M. Daniel, M. D. Kruskal, M. Lakshmanan, and K. Nakamura. Singularity structure analysis of the continuum Heisenberg spin chain with anisotropy and transverse field: Nonintegrability and chaos. Journal of Mathematical Physics, 33(2):771–776, February 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i2/p771_s1.

Morchio:1992:IVS

[1010] G. Morchio, D. Pierotti, and F. Strocchi. Infrared and vacuum structure in two-dimensional models of local quantum field theory. II. Fermion bosonization. Journal of Mathematical Physics, 33(2):777–790, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i2/p777_s1.

Dito:1992:SPN

[1011] Joseph Dito. Star-products and nonstandard quantization for Klein– Gordon equation. Journal of Mathematical Physics, 33(2):791–801, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p791_s1.

Nurowski:1992:SNS

[1012] Paweł Nurowski. Some null solutions of the Yang–Mills equations and Cauchy–Riemann structures. Journal of Mathematical Physics, 33(2):802–807, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i2/p802_s1.

Strube:1992:TIG

[1013] Dieter Strube. Translationally invariant global charges in a local collision theory for massless particles. Journal of Mathematical Physics, 33(2):808–813, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i2/p808_s1.

Dimock:1992:RGA

[1014] J. Dimock and T. R. Hurd. A renormalization group analysis of infrared QED. Journal of Mathematical Physics, 33(2):814–821, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i2/p814_s1.

Martinov:1992:NCS

[1015] N. Martinov, D. Ouroushev, and A. Grigorov. New class solutions of the 3-D Laplace equation. Journal of Mathematical Physics, 33(2):822– 825, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p822_s1.

Rebolledo:1992:DPA

[1016] Rolando Rebolledo, Juan C. Retamal, and Carlos Saavedra. Diffusion processes associated to a laser model. Journal of Mathematical Physics, 33(2):826–831, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i2/p826_s1.

Figueroa-OFarrill:1992:EUU

[1017] José M. Figueroa-O’Farrill and Eduardo Ramos. Existence and uniqueness of the universal W algebra. Journal of Mathematical Physics, 33(3):

833–839, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p833_s1. See erratum [1468].

Dzieciol:1992:SCA

[1018] Aleksander Dzieciol. The Stokes constants associated with the differential equation d^{2}ψ/dz^{2} − 4A^{2}z^{n−2}(z^{n} + c)ψ = 0. Journal of Mathematical Physics, 33(3):840–853, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i3/p840_s1.

Gannon:1992:LFIa

[1019] Terry Gannon and C. S. Lam. Lattices and θ-function identities. I. Theta constants. Journal of Mathematical Physics, 33(3):854–870, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p854_s1.

Gannon:1992:LFIb

[1020] Terry Gannon and C. S. Lam. Lattices and Θ-function identities. II: Theta series. Journal of Mathematical Physics, 33(3):871–887, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p871_s1.

Hlavaty:1992:PCP

[1021] Ladislav Hlavatý. The Painlevé classification of partial differential equations. III. Rational equations with several poles. Journal of Mathematical Physics, 33(3):888–894, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p888_s1.

Okubo:1992:NBI

[1022] Susumu Okubo. Nijenhuis–Bianchi identity and BRST-like operators. Journal of Mathematical Physics, 33(3):895–900, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/p895_s1.

Penna:1992:COR

[1023] Vittorio Penna and Mauro Spera. On coadjoint orbits of rotational perfect fluids. Journal of Mathematical Physics, 33(3):901–909, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p901_s1.

Nester:1992:SOF

[1024] James M. Nester. Special orthonormal frames. Journal of Mathematical Physics, 33(3):910–913, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p910_s1.

Shishkin:1992:DEEb

[1025] German V. Shishkin and William D. Cabos. Dirac equation in external fields: Separation of variables in curvilinear coordinates. Journal of Mathematical Physics, 33(3):914–925, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i3/p914_s1.

Fecko:1992:GFN

[1026] Marián Fecko. On a geometrical formulation of the Nambu dynamics. Journal of Mathematical Physics, 33(3):926–929, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/p926_s1.

Fecko:1992:VPN

[1027] Marián Fecko. On a variational principle for the Nambu dynamics. Journal of Mathematical Physics, 33(3):930–933, March 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i3/p930_s1.

Kamath:1992:RT

[1028] S. G. Kamath. Relativistic tautochrone. Journal of Mathematical Physics, 33(3):934–940, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i3/p934_s1.

Morosi:1992:MTR

[1029] C. Morosi. The R-matrix theory and the reduction of Poisson manifolds. Journal of Mathematical Physics, 33(3):941–952, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/p941_s1.

He:1992:FDW

[1030] Sailing He. Factorization of a dissipative wave equation and the Green functions technique for axially symmetric fields in a stratified slab. Journal of Mathematical Physics, 33(3):953–966, March 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i3/p953_s1.

Laedke:1992:SCC

[1031] E. W. Laedke, R. Blaha, K. H. Spatschek, and E. A. Kuznetsov. On the stability of collapse in the critical case. Journal of Mathematical Physics, 33(3):967–973, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i3/p967_s1.

Kerrick:1992:CGK

[1032] D. M. Kerrick. A comment on the generalized Kirchhoff–d’Adhémar formula for massless free fields in Minkowski space–time. Journal of Mathematical Physics, 33(3):974–976, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i3/p974_s1.

Bohm:1992:BCB

[1033] A. Bohm, B. Kendrick, Mark E. Loewe, and L. J. Boya. The Berry connection and Born–Oppenheimer method. Journal of Mathematical Physics, 33(3):977–989, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i3/p977_s1.

Banerjee:1992:TAF

[1034] Dipti Banerjee and Pratul Bandyopadhyay. Topological aspects of a fermion, chiral anomaly, and Berry phase. Journal of Mathematical Physics, 33(3):990–997, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i3/p990_s1.

Berceanu:1992:EMC

[1035] S. Berceanu and A. Gheorghe. On equations of motion on compact Hermitian symmetric spaces. Journal of Mathematical Physics, 33(3):998– 1007, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p998_s1.

Gould:1992:MEW

[1036] M. D. Gould, J. Links, and A. J. Bracken. Matrix elements and Wigner coefficients for U_{q}[gl(n)]. Journal of Mathematical Physics, 33(3):1008– 1022, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p1008_s1.

Gould:1992:RWC

[1037] M. D. Gould. Reduced Wigner coefficients for U_{q}[gl(n)]. Journal of Mathematical Physics, 33(3):1023–1031, March 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1023_s1.

Grosse:1992:SQS

[1038] H. Grosse and E. Langmann. A superversion of quasifree second quantization. I. Charged particles. Journal of Mathematical Physics, 33(3):1032–1046, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p1032_s1.

Rendall:1992:IVP

[1039] A. D. Rendall. The initial value problem for a class of general relativistic fluid bodies. Journal of Mathematical Physics, 33(3):1047–1053, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p1047_s1.

Tsoubelis:1992:HCG

[1040] Dimitri Tsoubelis and An Zhong Wang. Head-on collision of gravitational plane waves with noncollinear polarization: A new class of analytic models. Journal of Mathematical Physics, 33(3):1054–1064, March 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1054_
s1.

Wang:1992:GIP

[1041] An Zhong Wang. Gravitational interaction of plane gravitational waves and matter shells. Journal of Mathematical Physics, 33(3):1065–1072, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/

jmapaq/v33/i3/p1065_s1. See comments [2320, 2321].

Smalley:1992:MCE

[1042] Larry L. Smalley and Jean P. Krisch. Minimal coupling of electromagnetic fields in Riemann–Cartan space–times for perfect fluids with spin density. Journal of Mathematical Physics, 33(3):1073–1081, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p1073_s1.

Grandati:1992:EPF

[1043] Y. Grandati, A. Bérard, and P. Grangé. Exact partition functions of interacting lattice fermions. Journal of Mathematical Physics, 33(3):1082– 1089, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p1082_s1.

Goswami:1992:FSS

[1044] Gautam Goswami and Pratul Bandyopadhyay. Fractional statistics, Z_{p} spin system, and XY model. Journal of Mathematical Physics, 33(3):1090–1096, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p1090_s1.

Castagnino:1992:CTR

[1045] M. Castagnino, G. Domenech, V. Frolov, R. J. Noriega, and G. Schifini. Concomitants theory and renormalization in even dimensions. Journal of Mathematical Physics, 33(3):1097–1102, March 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1097_s1.

Rcheulishvili:1992:CAK

[1046] G. Rcheulishvili. The curvature and the algebra of Killing vectors in five-dimensional space. Journal of Mathematical Physics, 33(3):1103– 1108, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p1103_s1.

Cardoso:1992:BFE

[1047] J. G. Cardoso. Bessel function expressions for Dirac fields. Journal of Mathematical Physics, 33(3):1109–1113, March 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1109_s1.

Dixit:1992:SDO

[1048] V. V. Dixit, T. S. Santhanam, and W. D. Thacker. A supersymmetric Dirac oscillator with scalar coupling. Journal of Mathematical Physics, 33(3):1114–1117, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p1114_s1.

Kovalyov:1992:CSS

[1049] M. Kovalyov and M. Légaré. A complete set of solutions to the massive Dirac equation in static open Robertson–Walker space. Journal of Mathematical Physics, 33(3):1118–1126, March 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1118_s1.

Wang:1992:HMQ

[1050] Zheng Dong Wang. von Hove model of quantum field theory—from the point of view of infinite-dimensional stationary Markov process theory. Journal of Mathematical Physics, 33(3):1127–1131, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1127_s1.

Hanckowiak:1992:SEE

[1051] J. Hańćkowiak. In search of evolution equations in quantum and statistical field theory. Journal of Mathematical Physics, 33(3):1132– 1140, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p1132_s1.

Finkelstein:1992:MIP

[1052] Robert J. Finkelstein. Modular invariance of the propagator of an extended particle. Journal of Mathematical Physics, 33(3):1141–1147, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p1141_s1.

Freund:1992:AFD

[1053] Peter G. O. Freund. Arithmetic features of 2-D strings. Journal of Mathematical Physics, 33(3):1148–1152, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i3/p1148_s1.

Juriev:1992:ASQa

[1054] D. Juriev. Algebraic structures of the quantum projective [sl(2, C)- invariant] field theory: local projective field algebra DSCR [sl(2, C)]. Journal of Mathematical Physics, 33(3):1153–1157, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i3/p1153_s1.

Oshima:1992:EHK

[1055] Kazuto Oshima. Eigenfunctions and heat kernels of super Maass Laplacians on the super Poincaré upper half-plane. Journal of Mathematical Physics, 33(3):1158–1177, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i3/p1158_s1.

Roy:1992:TAL

[1056] Ashim Roy and Pratul Bandyopadhyay. Topological aspects of Liouville field theory and conformal anomaly. Journal of Mathematical Physics, 33(3):1178–1184, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p1178_s1.

Ferreira:1992:MSS

[1057] Erasmo Ferreira and Jorge Ananias Neto. Movable singularities in Sigma and Skyrme models. Journal of Mathematical Physics, 33(3):1185– 1199, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i3/p1185_s1.

Dattoli:1992:LUB

[1058] G. Dattoli, L. Giannessi, C. Mari, M. Richetta, and A. Torre. Linear undulator brightness: Exact analytical treatment. Journal of Mathematical Physics, 33(3):1200–1207, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i3/p1200_s1.

Tasgal:1992:SSC

[1059] R. S. Tasgal and M. J. Potasek. Soliton solutions to coupled higherorder nonlinear Schrödinger equations. Journal of Mathematical Physics, 33(3):1208–1215, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i3/p1208_s1.

Icha:1992:AFI

[1060] Andrzej Icha. An application of first integrals method in magnetic diffusion problem. Journal of Mathematical Physics, 33(3):1216–1220, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p1216_s1.

Esebbag:1992:CPO

[1061] C. Esebbag, J. M. Getino, M. de Llano, S. A. Moszkowski, U. Oseguera, A. Plastino, and H. Rubio. Cooper pairing in one, two, and three dimensions. Journal of Mathematical Physics, 33(3):1221–1223, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p1221_s1.

Brauer:1992:ETP

[1062] Uwe Brauer. An existence theorem for perturbed Newtonian cosmological models. Journal of Mathematical Physics, 33(3):1224–1233, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i3/p1224_s1.

Kim:1992:STM

[1063] Y. S. Kim and Leehwa Yeh. E(2)-symmetric two-mode sheared states. Journal of Mathematical Physics, 33(4):1237–1246, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i4/p1237_s1.

Rangarajan:1992:CR

[1064] Govindan Rangarajan and Filippo Neri. Canonical representations of sp(2n, R). Journal of Mathematical Physics, 33(4):1247–1251, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1247_s1.

Yoshida:1992:EEA

[1065] Zensho Yoshida. Eigenfunction expansions associated with the curl derivatives in cylindrical geometries: Completeness of Chandrasekhar– Kendall eigenfunctions. Journal of Mathematical Physics, 33(4):1252– 1256, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i4/p1252_s1.

Calogero:1992:INP

[1066] F. Calogero. C-integrable nonlinear partial differential equations in N +1 dimensions. Journal of Mathematical Physics, 33(4):1257–1271, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1257_s1.

deAzcarraga:1992:MHT

[1067] J. A. de Azcárraga, J. M. Izquierdo, and W. J. Zakrzewski. Models with Hopf terms. Journal of Mathematical Physics, 33(4):1272–1280, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1272_s1.

Lin:1992:SRT

[1068] F. J. Lin and J. E. Marsden. Symplectic reduction and topology for applications in classical molecular dynamics. Journal of Mathematical Physics, 33(4):1281–1294, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i4/p1281_s1.

Yurtsever:1992:SPG

[1069] Ulvi Yurtsever. A simple proof of geodesical completeness for compact space–times of zero curvature. Journal of Mathematical Physics, 33(4):1295–1300, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i4/p1295_s1.

Kolsrud:1992:ISL

[1070] Torbjörn Kolsrud and Jean-Claude Zambrini. An introduction to the semiclassical limit of Euclidean quantum mechanics. Journal of Mathematical Physics, 33(4):1301–1334, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i4/p1301_s1.

Abraham-Shrauner:1992:TSP

[1071] B. Abraham-Shrauner, Carl M. Bender, and R. N. Zitter. Taylor series and δ-perturbation expansions for a nonlinear semiconductor transport equation. Journal of Mathematical Physics, 33(4):1335–1340, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1335_s1.

deAzeredoCampos:1992:CQS

[1072] R. de Azeredo Campos. Canonical quantization of supersymmetric second-order systems. Journal of Mathematical Physics, 33(4):1341– 1347, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i4/p1341_s1.

Houard:1992:SQQ

[1073] Jean-Claude Houard and Michèle Irac-Astaud. Stochastic quantization for quantum mechanics and conditional probabilities. Journal of Mathematical Physics, 33(4):1348–1357, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i4/p1348_s1.

Floreanini:1992:MRG

[1074] Roberto Floreanini and Luc Vinet. The metaplectic representation of su_{q}(1, 1) and the q-Gegenbauer polynomials. Journal of Mathematical Physics, 33(4):1358–1362, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i4/p1358_s1.

Gromov:1992:GTR

[1075] N. A. Gromov. The Gel′fand–Tsetlin representations of the orthogonal Cayley–Klein algebras. Journal of Mathematical Physics, 33(4):1363– 1373, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i4/p1363_s1.

Gromov:1992:CIR

[1076] N. A. Gromov and V. I. Man′ko. Contractions of the irreducible representations of the quantum algebras su_{q}(2) and so_{q}(3). Journal of Mathematical Physics, 33(4):1374–1378, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i4/p1374_s1.

Hadjiivanov:1992:COV

[1077] L. K. Hadjiivanov, R. R. Paunov, and I. T. Todorov. U_{q} covariant oscillators and vertex operators. Journal of Mathematical Physics, 33(4):1379– 1394, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i4/p1379_s1.

Aktosun:1992:ISD

[1078] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Inverse scattering in 1-D nonhomogeneous media and recovery of the wave speed. Journal of Mathematical Physics, 33(4):1395–1402, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i4/p1395_s1.

Bera:1992:IRC

[1079] P. K. Bera, U. Das, C. Bhattacharyya, and B. Talukdar. On integral representations of the Coulomb functions. Journal of Mathematical Physics, 33(4):1403–1407, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i4/p1403_s1.

Hall:1992:SCS

[1080] G. S. Hall, A. D. Hossack, and J. R. Pulham. Sectional curvature, symmetries, and conformally flat plane waves. Journal of Mathematical Physics, 33(4):1408–1414, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i4/p1408_s1.

Xanthopoulos:1992:KSMa

[1081] Basilis C. Xanthopoulos and Thomas Zannias. Kantowski–Sachs metrics with source: A massless scalar field. Journal of Mathematical Physics, 33(4):1415–1419, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i4/p1415_s1.

Xanthopoulos:1992:KSMb

[1082] Basilis C. Xanthopoulos and Thomas Zannias. Kantowski–Sachs metrics with source: A conformally invariant scalar field. Journal of Mathematical Physics, 33(4):1420–1430, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i4/p1420_s1.

Verwimp:1992:BTL

[1083] Theo Verwimp. Boundary terms in Lovelock gravity from dimensionally continued Chern–Simons forms. Journal of Mathematical Physics, 33(4):1431–1436, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i4/p1431_s1.

Dunn:1992:TMG

[1084] K. A. Dunn, T. A. Harriott, and J. G. Williams. Toy model for gravitational kinks. Journal of Mathematical Physics, 33(4):1437–1444, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1437_s1.

Gomez:1992:AGC

[1085] R. Gómez and J. Winicour. Asymptotics of gravitational collapse of scalar waves. Journal of Mathematical Physics, 33(4):1445–1457, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1445_s1.

Kramer:1992:SSR

[1086] D. Kramer. Spherically symmetric radiating solution with heat flow in general relativity. Journal of Mathematical Physics, 33(4):1458–1462, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1458_s1.

Xanthopoulos:1992:EGC

[1087] Basilis C. Xanthopoulos and Thanassis E. Dialynas. Einstein gravity coupled to a massless conformal scalar field in arbitrary space–time dimensions. Journal of Mathematical Physics, 33(4):1463–1471, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1463_s1.

Tsamparlis:1992:GGC

[1088] M. Tsamparlis. Geometrization of a general collineation. Journal of Mathematical Physics, 33(4):1472–1479, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i4/p1472_s1.

Sammartino:1992:DLP

[1089] M. Sammartino, F. Malvagi, and G. C. Pomraning. Diffusive limits for particle transport in stochastic mixtures. Journal of Mathematical Physics, 33(4):1480–1501, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i4/p1480_s1.

Narnhofer:1992:QAC

[1090] Heide Narnhofer. Quantized Arnol′d cat maps can be entropic Ksystems. Journal of Mathematical Physics, 33(4):1502–1510, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i4/
p1502_s1.

Broda:1992:TFT

[1091] B. Broda. A topological-field-theory approach to the non-Abelian Stokes theorem. Journal of Mathematical Physics, 33(4):1511–1514, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i4/
p1511_s1.

Leo:1992:GET

[1092] R. A. Leo, L. Martina, and G. Soliani. Gauge equivalence theory of the noncompact Ishimori model and the Davey–Stewartson equation. Journal of Mathematical Physics, 33(4):1515–1522, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i4/p1515_s1.

Modanese:1992:PRG

[1093] Giovanni Modanese. The propagator in the radial gauge. Journal of Mathematical Physics, 33(4):1523–1528, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i4/p1523_s1.

MacKay:1992:FMU

[1094] N. J. MacKay. The fusion of R-matrices using the Birman–Wenzl– Murakami algebra. Journal of Mathematical Physics, 33(4):1529–1537, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1529_s1.

DellaSelva:1992:RES

[1095] A. Della Selva and A. Sciarrino. Realization of exceptional superalgebras in terms of fermion-boson creation-annihilation operators. Journal of Mathematical Physics, 33(4):1538–1545, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i4/p1538_s1.

Sardanashvily:1992:GSS

[1096] Gennadi Sardanashvily. On the geometry of spontaneous symmetry breaking. Journal of Mathematical Physics, 33(4):1546–1549, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i4/
p1546_s1.

Vanagas:1992:ECM

[1097] V. Vanagas and J. A. Castilho Alcarás. On the evaluation of the U(3) content of the matrix elements of one- and two-body operators. Journal of Mathematical Physics, 33(4):1550–1565, April 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i4/p1550_s1.

Greenberg:1992:ERH

[1098] W. Greenberg, P. F. Zweifel, and S. Paveri-Fontana. Erratum: “The Riemann–Hilbert problem for nonsymmetric systems”, [J. Math. Phys. 32, 3540 (1991)]. Journal of Mathematical Physics, 33(4):1566, April 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i4/p1566_s1. See [924].

Aguirre:1992:GST

[1099] M. Aguirre, C. Friedli, and J. Krause. SL(3, R) as the group of symmetry transformations for all one-dimensional linear systems. III. Equivalent Lagrangian formalisms. Journal of Mathematical Physics, 33(5):1571– 1578, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1571_s1.

Brihaye:1992:RRC

[1100] Y. Brihaye, P. Maślanka, S. Giler, and P. Kosiński. Real representations of Clifford algebras. Journal of Mathematical Physics, 33(5):1579– 1581, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1579_s1.

Minnaert:1992:RWC

[1101] Pierre Minnaert and Marek Mozrzymas. Racah–Wigner calculus for the super-rotation algebra. I. Journal of Mathematical Physics, 33(5):1582– 1593, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1582_s1.

Minnaert:1992:AST

[1102] Pierre Minnaert and Marek Mozrzymas. Algebraic structure of tensor superoperators for the super-rotation algebra. II. Journal of Mathematical Physics, 33(5):1594–1600, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i5/p1594_s1.

Wijnands:1992:MDG

[1103] Frank Wijnands and Ad Thiers. Methods to determine generators and defining relations for an n-dimensional point group. Journal of Mathematical Physics, 33(5):1601–1611, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i5/p1601_s1.

Wu:1992:TPP

[1104] Yihren Wu. On the topological phases in a planar N-point vortex system. Journal of Mathematical Physics, 33(5):1612–1617, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1612_s1.

Gingras:1992:OOB

[1105] F. Gingras, J. Patera, and R. T. Sharp. Orbit-orbit branching rules between simple low-rank algebras and equal-rank subalgebras. Journal of Mathematical Physics, 33(5):1618–1626, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i5/p1618_s1.

MacDowell:1992:RDS

[1106] Samuel W. MacDowell. Realization of the discrete series of unitary representations of Sl(2, R) in terms of creation and annihilation operators. Journal of Mathematical Physics, 33(5):1627–1630, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1627_s1.

Marcinek:1992:GAG

[1107] W. Marcinek. Graded algebras and geometry based on Yang–Baxter operators. Journal of Mathematical Physics, 33(5):1631–1635, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/
p1631_s1.

Khrennikov:1992:AASa

[1108] A. Yu. Khrennikov. Analysis on the p-adic superspace. I. Generalized functions: Gaussian distribution. Journal of Mathematical Physics, 33(5):1636–1642, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i5/p1636_s1.

Khrennikov:1992:AASb

[1109] A. Yu. Khrennikov. Analysis on the p-adic superspace. II. Differential equations on p-adic superspace. Journal of Mathematical Physics, 33(5): 1643–1647, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1643_s1.

Boenkost:1992:FCB

[1110] W. Boenkost and F. Constantinescu. Functional calculus of the Borel transform. Journal of Mathematical Physics, 33(5):1648–1651, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/
p1648_s1.

Giachetta:1992:JMN

[1111] Giovanni Giachetta. Jet methods in nonholonomic mechanics. Journal of Mathematical Physics, 33(5):1652–1665, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i5/p1652_s1.

Smith:1992:NAF

[1112] A. P. Smith. A new asymptotic form for the Laguerre polynomials. Journal of Mathematical Physics, 33(5):1666–1671, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1666_s1.

Ullah:1992:VDE

[1113] Nazakat Ullah. Variational density of electrons for large closed shell atoms. Journal of Mathematical Physics, 33(5):1672–1674, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/
p1672_s1.

Kerner:1992:NRM

[1114] Edward H. Kerner. Note on the representation of many-particle dynamics as higher-order single-particle dynamics, with a means to relativistic elevation of Newtonian dynamics. Journal of Mathematical Physics, 33(5): 1675–1684, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1675_s1.

Popov:1992:ETO

[1115] I. Yu. Popov. The extension theory and the opening in semitransparent surface. Journal of Mathematical Physics, 33(5):1685–1689, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i5/p1685_s1.

Abe:1992:SWF

[1116] Sumiyoshi Abe. Super Wigner function. Journal of Mathematical Physics, 33(5):1690–1694, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i5/p1690_s1.

Jaekel:1992:SL

[1117] M. T. Jaekel. Stochastic lengths. Journal of Mathematical Physics, 33(5): 1695–1704, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1695_s1.

Fradkin:1992:NSO

[1118] David M. Fradkin. Novel solution of the one-dimensional quantummechanical harmonic oscillator. Journal of Mathematical Physics, 33(5): 1705–1709, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1705_s1.

Hall:1992:MPQ

[1119] Richard L. Hall. Mixtures of potentials in quantum mechanics. Journal of Mathematical Physics, 33(5):1710–1716, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i5/p1710_s1.

Aktosun:1992:SIS

[1120] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Scattering and inverse scattering in one-dimensional nonhomogeneous media. Journal of Mathematical Physics, 33(5):1717–1744, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1717_s1.

Radosz:1992:RTC

[1121] A. Radosz and W. Magierski. Real trajectories in complex-time method and barrier penetrationlike phenomena. Journal of Mathematical Physics, 33(5):1745–1753, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i5/p1745_s1.

Coley:1992:ACV

[1122] A. A. Coley and B. O. J. Tupper. Affine conformal vectors in space–time. Journal of Mathematical Physics, 33(5):1754–1764, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1754_s1.

Galicki:1992:NSC

[1123] Krzysztof Galicki and Takashi Nitta. Nonzero scalar curvature generalizations of the ALE hyper-Kähler metrics. Journal of Mathematical Physics, 33(5):1765–1771, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i5/p1765_s1.

Coley:1992:QAC

[1124] A. A. Coley and K. A. Dunn. Qualitative analysis of a class of Bianchi V imperfect fluid cosmologies. Journal of Mathematical Physics, 33(5): 1772–1779, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1772_s1.

Krori:1992:MSO

[1125] K. D. Krori and D. Goswami. A multipolar stationary object embedded in a gravitational field. Journal of Mathematical Physics, 33(5):1780– 1781, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1780_s1.

Choquet-Bruhat:1992:CYM

[1126] Yvonne Choquet-Bruhat. Cosmological Yang–Mills hydrodynamics. Journal of Mathematical Physics, 33(5):1782–1785, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1782_s1.

Wiesen:1992:PGB

[1127] Bernd Wiesen. On a phenomenological generalized Boltzmann equation. Journal of Mathematical Physics, 33(5):1786–1798, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1786_s1.

Long:1992:CST

[1128] Ming Long and M. Howard Lee. The classical susceptibility and the temperature-dependent susceptibility of a free electron gas. Journal of Mathematical Physics, 33(5):1799–1806, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i5/p1799_s1.

Porsezian:1992:IAO

[1129] K. Porsezian, M. Daniel, and M. Lakshmanan. On the integrability aspects of the one-dimensional classical continuum isotropic biquadratic Heisenberg spin chain. Journal of Mathematical Physics, 33(5):1807– 1816, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i5/p1807_s1.

Hillion:1992:NSK

[1130] Pierre Hillion. Nondispersive solutions of the Klein–Gordon equation. Journal of Mathematical Physics, 33(5):1817–1821, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1817_s1.

Hillion:1992:NSD

[1131] Pierre Hillion. Nondispersive solutions of the Dirac equation. Journal of Mathematical Physics, 33(5):1822–1830, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i5/p1822_s1.

MartinezyRomero:1992:CID

[1132] R. P. Martínez y Romero and A. L. Salas-Brito. Conformal invariance in a Dirac oscillator. Journal of Mathematical Physics, 33(5):1831–1836, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i5/p1831_s1.

Wiesbrock:1992:NCA

[1133] Hans-Werner Wiesbrock. A note on the construction of the C^{∗}-algebra of bosonic strings. Journal of Mathematical Physics, 33(5):1837–1840, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i5/p1837_s1.

Ky:1992:TSI

[1134] Nguyen Anh Ky, Tchavdar D. Palev, and Nedjalka I. Stoilova. Transformations of some induced osp(3/2) modules in an so(3) ⊕ sp(2) basis. Journal of Mathematical Physics, 33(5):1841–1863, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1841_s1.

Talmadge:1992:GFH

[1135] Andy Talmadge. A geometric formulation of the Higgs mechanism. Journal of Mathematical Physics, 33(5):1864–1868, May 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1864_s1.

Bloore:1992:HSD

[1136] F. Bloore and P. A. Horváthy. Helicity-supersymmetry of dyons. Journal of Mathematical Physics, 33(5):1869–1877, May 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1869_s1.

Moses:1992:BEM

[1137] Harry E. Moses and Reese T. Prosser. Bounds for the exact matrix elements for radiation processes in hydrogenic atoms. Journal of Mathematical Physics, 33(5):1878–1886, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i5/p1878_s1.

Fishman:1992:EOR

[1138] Louis Fishman. Exact and operator rational approximate solutions of the Helmholtz, Weyl composition equation in underwater acoustics—the quadratic profile. Journal of Mathematical Physics, 33(5):1887–1914, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i5/p1887_s1.

Petters:1992:MTG

[1139] A. O. Petters. Morse theory and gravitational microlensing. Journal of Mathematical Physics, 33(5):1915–1931, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i5/p1915_s1.

Folacci:1992:ETP

[1140] A. Folacci. Erratum: “Two-point functions and stress-energy tensors of p-forms in de Sitter and anti-de Sitter spaces”, [J. Math. Phys. 32, 2828 (1991)]. Journal of Mathematical Physics, 33(5):1932, May 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i5/p1932_

s1. See [829].

Klimyk:1992:RQA

[1141] A. U. Klimyk and S. Z. Pakuliak. Representations of the quantum algebras U_{q}(u_{r,s}) and U_{q}(u_{r+s}) related to the quantum hyperboloid and sphere. Journal of Mathematical Physics, 33(6):1937–1947, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/
p1937_s1.

Healy:1992:CEE

[1142] Liam M. Healy, Alex J. Dragt, and Ivan M. Gjaja. Computation of error effects in nonlinear Hamiltonian systems using Lie algebraic methods. Journal of Mathematical Physics, 33(6):1948–1963, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/p1948_s1.

Sullivan:1992:COD

[1143] J. J. Sullivan and T. H. Siddall, III. Casimir operators, duality, and the point groups. Journal of Mathematical Physics, 33(6):1964–1969, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i6/p1964_s1.

Zhang:1992:UOI

[1144] R. B. Zhang. Universal L operator and invariants of the quantum supergroup U_{q}(gl(m/n)). Journal of Mathematical Physics, 33(6):1970– 1979, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p1970_s1.

Alisauskas:1992:CMR

[1145] S. Ališauskas. Comment on: “The matrix representation of U_{4} in the U_{2} × U_{2} basis and some isoscalar factors for U_{p+q} ⊃ U_{p} × U_{q}”, [J. Math. Phys. 31 (1990), no. 6, 1333–1339; MR 91d:81040] by F. Pan. Journal of Mathematical Physics, 33(6):1980–1982, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i6/p1980_s1. See [200].

Alisauskas:1992:ECT

[1146] Sigitas Ališauskas. Explicit canonical tensor operators and orthonormal coupling coefficients of SU(3). Journal of Mathematical Physics, 33(6): 1983–2004, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p1983_s1.

Maximon:1992:ESB

[1147] Leonard C. Maximon. Evaluation of the spherical Bessel transform of a Whittaker function: An application of a difference equation method. Journal of Mathematical Physics, 33(6):2005–2012, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/p2005_s1.

Yordanov:1992:CPL

[1148] Russi G. Yordanov. Cauchy problem for the linearized version of the generalized polynomial KdV equation. Journal of Mathematical Physics, 33(6):2013–2022, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i6/p2013_s1.

Leach:1992:FIC

[1149] P. G. L. Leach and S. D. Maharaj. A first integral for a class of timedependent anharmonic oscillators with multiple anharmonicities. Journal of Mathematical Physics, 33(6):2023–2030, June 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i6/p2023_s1.

Mugan:1992:STP

[1150] U. Muğan and A. S. Fokas. Schlesinger transformations of Painlevé II–V. Journal of Mathematical Physics, 33(6):2031–2045, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/p2031_s1.

Barraza:1992:DBV

[1151] O. Barraza, H. Falomir, R. E. Gamboa Saraví, and E. M. Santangelo. P determinants and boundary values. Journal of Mathematical Physics, 33(6):2046–2052, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2046_s1.

Bleecker:1992:SSA

[1152] David D. Bleecker. The supertrace of the steady asymptotic of the spinorial heat kernel. Journal of Mathematical Physics, 33(6):2053–2070, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i6/p2053_s1.

Kelnhofer:1992:RBG

[1153] Gerald Kelnhofer. A relation between group- and Čech-cohomology in principal fiber bundles and anomalies. Journal of Mathematical Physics, 33(6):2071–2079, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2071_s1.

Okubo:1992:CIS

[1154] Susumu Okubo. Continuous integrable systems with multi-Poisson brackets. Journal of Mathematical Physics, 33(6):2080–2092, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/
p2080_s1.

Shishkin:1992:NPGa

[1155] German V. Shishkin and Víctor M. Villalba. Neutrino in the presence of gravitational fields: Separation of variables. Journal of Mathematical Physics, 33(6):2093–2110, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2093_s1.

Thiran:1992:DSS

[1156] E. Thiran and J. Weyers. Dixon–Selberg summation over local number fields. Journal of Mathematical Physics, 33(6):2111–2114, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/
p2111_s1.

Antonowicz:1992:HCF

[1157] Marek Antonowicz and Stefan Rauch-Wojciechowski. How to construct finite-dimensional bi-Hamiltonian systems from soliton equations: Jacobi integrable potentials. Journal of Mathematical Physics, 33(6):2115– 2125, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2115_s1.

Fawcett:1992:CCS

[1158] R. J. B. Fawcett. Compact classical systems: so(n) systems. Journal of Mathematical Physics, 33(6):2126–2137, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i6/p2126_s1.

Lacomba:1992:IIM

[1159] E. A. Lacomba and J. Llibre. Integrals, invariant manifolds, and degeneracy for central force problems in R^{n}. Journal of Mathematical Physics, 33(6):2138–2147, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2138_s1.

Okubo:1992:EGF

[1160] Susumu Okubo. Existence of gauge field in any partially integrable systems. Journal of Mathematical Physics, 33(6):2148–2157, June 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/p2148_
s1.

Bette:1992:HLF

[1161] Andreas Bette. Hamiltonization of the Lorentz force equation for a massive particle with spin: A twistorial approach. Journal of Mathematical Physics, 33(6):2158–2163, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2158_s1.

Blaive:1992:SMP

[1162] B. Blaive and A. O. Barut. On some mathematical problems in the evaluation of fields produced by currents which are generalized functions. Journal of Mathematical Physics, 33(6):2164–2168, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/p2164_s1.

Cardoso:1992:RWE

[1163] J. G. Cardoso. Real wave equations for positive-frequency electromagnetic fields and their Green’s function solutions. Journal of Mathematical Physics, 33(6):2169–2172, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2169_s1.

Pulmannova:1992:BIQ

[1164] Sylvia Pulmannová and Vladimir Majernik. Bell inequalities on quantum logics. Journal of Mathematical Physics, 33(6):2173–2178, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i6/p2173_s1.

Fesseha:1992:EPQ

[1165] K. Fesseha. Evaluation of the propagator for a quadratic Hamiltonian with arbitrary ordering. Journal of Mathematical Physics, 33(6):2179– 2184, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2179_s1.

Roelofs:1992:SEN

[1166] G. H. M. Roelofs and P. H. M. Kersten. Supersymmetric extensions of the nonlinear Schrödinger equation: Symmetries and coverings. Journal of Mathematical Physics, 33(6):2185–2206, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i6/p2185_s1.

Dittrich:1992:DHC

[1167] J. Dittrich, P. Exner, and P. Šeba. Dirac Hamiltonian with Coulomb potential and spherically symmetric shell contact interaction. Journal of Mathematical Physics, 33(6):2207–2214, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i6/p2207_s1.

Freeman:1992:GPI

[1168] D. Kent Freeman. Generalization of the pseudopotential to an interaction with nonzero range. Journal of Mathematical Physics, 33(6):2215– 2227, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2215_s1.

Folacci:1992:GFR

[1169] A. Folacci. Green functions of the de rham Laplacian in maximally symmetric spaces. Journal of Mathematical Physics, 33(6):2228–2231, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i6/p2228_s1.

Fullwood:1992:NTS

[1170] D. T. Fullwood. A new topology on space–time. Journal of Mathematical Physics, 33(6):2232–2241, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2232_s1.

Kheyfets:1992:CMR

[1171] Arkady Kheyfets and Warner A. Miller. É. Cartan moment of rotation in Ashtekar’s self-dual representation of gravitation. Journal of Mathematical Physics, 33(6):2242–2248, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i6/p2242_s1.

Beem:1992:CCP

[1172] John K. Beem and Andrzej Królak. Cosmic censorship and pseudoconvexity. Journal of Mathematical Physics, 33(6):2249–2253, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/
p2249_s1.

Garcia:1992:SAC

[1173] Alberto García and Jorge Téllez. Stationary axisymmetric charged perfect fluid type D solutions. Journal of Mathematical Physics, 33(6): 2254–2257, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2254_s1.

Melfo:1992:CGT

[1174] A. Melfo, L. Núñez, U. Percoco, and V. M. Villalba. Collineations of Gödel-type space–times. Journal of Mathematical Physics, 33(6):2258– 2261, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2258_s1.

Qadir:1992:ESS

[1175] Asghar Qadir. Essential singularities of spherically symmetric space– times. Journal of Mathematical Physics, 33(6):2262–2264, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i6/
p2262_s1.

Kremer:1992:RT

[1176] G. M. Kremer and Ingo Müller. Radiation thermodynamics. Journal of Mathematical Physics, 33(6):2265–2268, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i6/p2265_s1.

Sutcliffe:1992:NSS

[1177] P. M. Sutcliffe. Nontrivial soliton scattering in an integrable chiral model in (2 + 1)-dimensions. Journal of Mathematical Physics, 33(6):2269– 2278, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2269_s1.

Bagrov:1992:NMI

[1178] V. G. Bagrov and V. V. Obukhov. New method of integration for the Dirac equation on a curved space–time. Journal of Mathematical Physics, 33(6):2279–2289, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2279_s1.

Marx:1992:SER

[1179] Egon Marx. Spinor equations in relativistic quantum mechanics. Journal of Mathematical Physics, 33(6):2290–2302, June 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i6/p2290_s1.

Ohta:1992:DMB

[1180] Tadayuki Ohta and Toshiei Kimura. Derivation of many-body potential among charged particles in the S-matrix method. Journal of Mathematical Physics, 33(6):2303–2322, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i6/p2303_s1.

Mihalache:1992:ESN

[1181] D. Mihalache and N. C. Panoiu. Exact solutions of nonlinear Schrödinger equation for positive group velocity dispersion. Journal of Mathematical Physics, 33(6):2323–2328, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i6/p2323_s1.

Baldo:1992:TDR

[1182] M. Baldo, A. Grassi, and A. Raudino. Time-dependent rototranslational diffusion equation with multiple mixed boundary conditions: A formally exact analytical solution. Journal of Mathematical Physics, 33(6):2329– 2335, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i6/p2329_s1.

Chakraborty:1992:SCS

[1183] Subenoy Chakraborty and Ashok Kr. Chakraborty. String cosmology in spherically symmetric space–time. Journal of Mathematical Physics, 33(6):2336–2338, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i6/p2336_s1.

Begin:1992:WOT

[1184] F. Bégin and R. T. Sharp. Weyl orbits and their expansions in irreducible representations for affine Kac–Moody algebras. Journal of Mathematical Physics, 33(7):2343–2356, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i7/p2343_s1.

deAzcarraga:1992:SBS

[1185] J. A. de Azcárraga, J. M. Izquierdo, and W. J. Zakrzewski. A supergroup based supersigma model. Journal of Mathematical Physics, 33(7):2357– 2364, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2357_s1.

Kalnins:1992:MAR

[1186] E. G. Kalnins, H. L. Manocha, and Willard Miller, Jr. Models of qalgebra representations: Tensor products of special unitary and oscillator algebras. Journal of Mathematical Physics, 33(7):2365–2383, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i7/p2365_s1.

Wehrhahn:1992:SSH

[1187] R. F. Wehrhahn, Yu. F. Smirnov, and A. M. Shirokov. Symmetry scattering on the hyperboloid SO(2, 1)/SO(2) in different coordinate systems. Journal of Mathematical Physics, 33(7):2384–2389, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i7/p2384_s1.

Wolf:1992:ERS

[1188] Kurt Bernardo Wolf. The Euclidean root of Snell’s law. I. Geometric polarization optics. Journal of Mathematical Physics, 33(7):2390–2408, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i7/p2390_s1.

Zhang:1992:QAZ

[1189] H. Zhang. Quantum aspects of zero modes, topological vacua and dynamical supersymmetry breaking in nonabelian gauge theories. Journal of Mathematical Physics, 33(7):2409–2416, July 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i7/p2409_s1.

VanderJeugt:1992:TPG

[1190] J. Van der Jeugt. Tensor product of group representations and structural zeros of Racah coefficients. Journal of Mathematical Physics, 33(7):2417– 2421, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2417_s1.

Su:1992:SSR

[1191] Yu Cai Su. Some simple representations of the affine Lie superalgebra A^{(1)}(1, 0). Journal of Mathematical Physics, 33(7):2422–2427, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i7/p2422_s1.

Miyazawa:1992:DAL

[1192] Toru Miyazawa. Diffusion in asymptotically logarithmic potentials. Journal of Mathematical Physics, 33(7):2428–2433, July 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i7/p2428_s1.

Worden:1992:RST

[1193] K. Worden. Renormalized stress tensors on ellipses. Journal of Mathematical Physics, 33(7):2434–2439, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2434_s1.

Cairo:1992:FIM

[1194] L. Cairó and M. R. Feix. Families of invariants of the motion for the Lotka–Volterra equations: The linear polynomials family. Journal of Mathematical Physics, 33(7):2440–2455, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i7/p2440_s1.

Glockle:1992:PTE

[1195] W. G. Glöckle, G. Baumann, and T. F. Nonnenmacher. Painlevé test and exact similarity solutions of a class of nonlinear diffusion equations. Journal of Mathematical Physics, 33(7):2456–2463, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i7/p2456_s1.

Ma:1992:LRL

[1196] Wen Xiu Ma. Lax representations and Lax operator algebras of isospectral and nonisospectral hierarchies of evolution equations. Journal of Mathematical Physics, 33(7):2464–2476, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i7/p2464_s1.

Strachan:1992:NFI

[1197] I. A. B. Strachan. A new family of integrable models in (2+1) dimensions associated with Hermitian symmetric spaces. Journal of Mathematical Physics, 33(7):2477–2482, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i7/p2477_s1.

Chalbaud:1992:TPQ

[1198] Eduardo Chalbaud and Pablo Martín. Two-point quasifractional approximant in physics: Method improvement and application to J_{ν}(x). Journal of Mathematical Physics, 33(7):2483–2486, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i7/p2483_s1.

Das:1992:HSA

[1199] Ashok Das and Wen-Jui Huang. The Hamiltonian structures associated with a generalized Lax operator. Journal of Mathematical Physics, 33(7): 2487–2497, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2487_s1.

Grundland:1992:FNK

[1200] A. M. Grundland and E. Infeld. A family of nonlinear Klein–Gordon equations and their solutions. Journal of Mathematical Physics, 33(7):

2498–2503, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2498_s1.

Dorfman:1992:HTN

[1201] I. Ya. Dorfman and A. S. Fokas. Hamiltonian theory over noncommutative rings and integrability in multidimensions. Journal of Mathematical Physics, 33(7):2504–2514, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i7/p2504_s1.

Bertrand:1992:CAW

[1202] J. Bertrand and P. Bertrand. A class of affine Wigner functions with extended covariance properties. Journal of Mathematical Physics, 33(7): 2515–2527, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2515_s1. See erratum [1466].

Bohm:1992:UBC

[1203] Arno Bohm, Luis J. Boya, and Brian Kendrick. On the uniqueness of the Berry connection. Journal of Mathematical Physics, 33(7):2528– 2532, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2528_s1.

Rydzynska:1992:QAB

[1204] Ewa Rydzyńska. Quantum aspects of Bellert’s theory. Journal of Mathematical Physics, 33(7):2533–2540, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2533_s1.

Ge:1992:RQM

[1205] Mo Lin Ge, Xu Feng Liu, and Chang Pu Sun. Representations of quantum matrix algebra m_{q}(2) and its q-boson realization. Journal of Mathematical Physics, 33(7):2541–2545, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2541_s1.

Tsukada:1992:SVO

[1206] Haruo Tsukada. Shifted vertex operator algebras and G-elliptic systems. Journal of Mathematical Physics, 33(7):2546–2556, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i7/p2546_s1.

Darewych:1992:BRV

[1207] J. W. Darewych and I. Schlifer. Bounding the remainder in variational calculations of potential scattering. Journal of Mathematical Physics, 33(7):2557–2561, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i7/p2557_s1.

Bandyopadhyay:1992:PFQ

[1208] N. Bandyopadhyay. Partition functions for quantized fields in curved space–times. Journal of Mathematical Physics, 33(7):2562–2566, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i7/p2562_s1.

Aliev:1992:ESV

[1209] B. N. Aliev and A. N. Leznov. Exact solutions of the vacuum Einstein’s equations allowing for two noncommutative Killing vectors (type G_{2}II of Petrov classification). Journal of Mathematical Physics, 33(7):2567– 2573, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2567_s1.

Fonseca-Neto:1992:EPT

[1210] J. B. Fonseca-Neto, M. J. Rebouças, and A. F. F. Teixeira. The equivalence problem in torsion theories of gravitation. Journal of Mathematical Physics, 33(7):2574–2577, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2574_s1.

Bobylev:1992:GBT

[1211] A. V. Bobylëv and G. Toscani. On the generalization of the Boltzmann H-theorem for a spatially homogeneous Maxwell gas. Journal of Mathematical Physics, 33(7):2578–2586, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2578_s1.

Cornille:1992:ESH

[1212] H. Cornille and T. Płatkowski. Exact solutions of a hierarchy of mixing speeds models. Journal of Mathematical Physics, 33(7):2587–2606, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i7/p2587_s1.

Rudnicki:1992:LDL

[1213] Sławomir Rudnicki, Robert Alicki, and Sławomir Sadowski. The lowdensity limit in terms of collective squeezed vectors. Journal of Mathematical Physics, 33(7):2607–2617, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2607_s1.

Galeao:1992:GMR

[1214] A. P. Galeão and P. Leal Ferreira. General method for reducing the two-body Dirac equation. Journal of Mathematical Physics, 33(7):2618– 2625, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2618_s1.

Ferreira:1992:SSM

[1215] Erasmo Ferreira and Jorge Ananias Neto. Singularities in the solutions of the massive Skyrme model. Journal of Mathematical Physics, 33(7): 2626–2632, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i7/p2626_s1.

Hall:1992:WMC

[1216] G. S. Hall. Weyl manifolds and connections. Journal of Mathematical Physics, 33(7):2633–2638, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i7/p2633_s1.

Malvagi:1992:AED

[1217] F. Malvagi, M. Sammartino, and G. C. Pomraning. Asymptotically exact diffusive boundary conditions in linear kinetic theory. Journal of Mathematical Physics, 33(7):2639–2647, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i7/p2639_s1.

Su:1992:RCA

[1218] Yu Cai Su. Realization of a class of affine Lie superalgebras. Journal of Mathematical Physics, 33(8):2653–2655, August 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2653_s1.

Lemke:1992:WAC

[1219] Jürgen Lemke, Yuval Ne’eman, and Jose Pecina-Cruz. Wigner analysis and Casimir operators of \overline{SA}(4, R). Journal of Mathematical Physics, 33(8):2656–2659, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2656_s1.

Tong:1992:IRB

[1220] Dian Min Tong, Cheng Jiu Zhu, and Zhong Qi Ma. Irreducible representations of braid groups. Journal of Mathematical Physics, 33(8):2660– 2663, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i8/p2660_s1.

Carbon:1992:WDC

[1221] Steven L. Carbon and Eric J. Piard. Weight diagram construction of Lax operators. Journal of Mathematical Physics, 33(8):2664–2674, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2664_s1.

Jha:1992:SRG

[1222] Pradip K. Jha and Kishor C. Tripathy. Symmetries of the renormalization group equations. Journal of Mathematical Physics, 33(8):2675– 2681, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i8/p2675_s1.

Pereira:1992:ODN

[1223] L. G. Pereira and J. Goedert. One-dimensional nonautonomous dynamical systems with exact transcendental invariants. Journal of Mathematical Physics, 33(8):2682–2687, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i8/p2682_s1.

Chan:1992:GLF

[1224] Fuk-Lok Chan, Kim-Ming Lau, and Robert J. Finkelstein. Grassmannian Legendre functions. Journal of Mathematical Physics, 33(8):2688– 2696, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i8/p2688_s1.

Zhu:1992:SCC

[1225] Chaoyuan Zhu and Hiroki Nakamura. Stokes constants for a certain class of second-order ordinary differential equations. Journal of Mathematical Physics, 33(8):2697–2717, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i8/p2697_s1.

Martina:1992:PIS

[1226] L. Martina and P. Winternitz. Partially invariant solutions of nonlinear Klein–Gordon and Laplace equations. Journal of Mathematical Physics, 33(8):2718–2727, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2718_s1.

Goriely:1992:IPP

[1227] Alain Goriely. Investigation of Painlevé property under time singularities transformations. Journal of Mathematical Physics, 33(8):2728–2742, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2728_s1.

Barcelos-Neto:1992:BHS

[1228] J. Barcelos-Neto and Ashok Das. On the bi-Hamiltonian structures of the sKdV. Journal of Mathematical Physics, 33(8):2743–2748, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2743_s1.

Hillion:1992:CHS

[1229] Pierre Hillion. The Courant–Hilbert solutions of the wave equation. Journal of Mathematical Physics, 33(8):2749–2753, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2749_s1.

Matsuno:1992:DSD

[1230] Y. Matsuno. Dynamics of solitons in a damped sine-Hilbert equation. Journal of Mathematical Physics, 33(8):2754–2764, August 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2754_
s1.

Benn:1992:DC

[1231] I. M. Benn and W. P. Wood. Dilation currents. Journal of Mathematical Physics, 33(8):2765–2767, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i8/p2765_s1.

Lenczewski:1992:AQC

[1232] Romuald Lenczewski and Krzysztof Podgórski. A q-analog of the quantum central limit theorem for SU_{q}(2). Journal of Mathematical Physics, 33(8):2768–2778, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2768_s1.

Witwit:1992:ELTa

[1233] M. R. M. Witwit. Energy levels of a three-dimensional anharmonic oscillator with sextic perturbation. Journal of Mathematical Physics, 33(8):2779–2784, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2779_s1.

Znojil:1992:ESB

[1234] M. Znojil and P. G. L. Leach. On the elementary Schrödinger bound states and their multiplets. Journal of Mathematical Physics, 33(8):2785– 2794, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i8/p2785_s1.

Kuperin:1992:TBR

[1235] Yu. A. Kuperin and Yu. B. Mel′nikov. Two-body resonance scattering and annihilation of composite charged particles. Journal of Mathematical Physics, 33(8):2795–2807, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v33/i8/p2795_s1.

Albeverio:1992:FSF

[1236] Sergio Albeverio, Erling G. B. Hohler, and Bogusław Zegarliński. Ferromagneticity of simplicial fields on two-dimensional compact manifolds. Journal of Mathematical Physics, 33(8):2808–2818, August 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2808_
s1.

Juriev:1992:NGC

[1237] Denis Juriev. Noncommutative geometry, chiral anomaly in the quantum projective [sl(2, C)-invariant] field theory and jl(2, C)-invariance. Journal of Mathematical Physics, 33(8):2819–2822, August 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2819_s1. See er-

ratum [1523].

Horuzhy:1992:BQS

[1238] S. S. Horuzhy and A. V. Voronin. BRST quantization of the Schwinger model. Journal of Mathematical Physics, 33(8):2823–2841, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i8/
p2823_s1.

Kannar:1992:SST

[1239] János Kánnár and István Rácz. On the strength of space–time singularities. Journal of Mathematical Physics, 33(8):2842–2848, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2842_s1.

Maluf:1992:SDC

[1240] José W. Maluf. Self-dual connections, torsion and Ashtekar’s variables. Journal of Mathematical Physics, 33(8):2849–2854, August 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2849_
s1.

Reboucas:1992:HST

[1241] M. J. Rebouças and A. F. F. Teixeira. Homogeneous space–times with seven isometries. Journal of Mathematical Physics, 33(8):2855–2862, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2855_s1.

Christodoulakis:1992:QMG

[1242] T. Christodoulakis and E. Korfiatis. Quantum mechanics of the general spatially homogeneous geometry coupled to a scalar field. Journal of Mathematical Physics, 33(8):2863–2876, August 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2863_s1.

McQuistan:1992:CNN

[1243] R. B. McQuistan and J. L. Hock. Composite nearest-neighbor degeneracies for several kinds of simple particles distributed on two-dimensional, square-cell lattices. Journal of Mathematical Physics, 33(8):2877–2883, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i8/p2877_s1.

Jakobczyk:1992:IRS

[1244] Lech Jakóbczyk. Irreducibility and reducing subspaces for Kreĭn space representations of field algebra. Journal of Mathematical Physics, 33 (8):2884–2888, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2884_s1.

Cho:1992:ATF

[1245] H. S. Cho, J. D. Kim, and I. G. Koh. D_{5} affine Toda field theory. Journal of Mathematical Physics, 33(8):2889–2897, August 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i8/p2889_s1.

Brihaye:1992:PSS

[1246] Yves Brihaye and Jutta Kunz. Periodic solutions in spontaneously broken theories. Journal of Mathematical Physics, 33(8):2898–2902, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2898_s1.

Budreck:1992:NME

[1247] David E. Budreck and James H. Rose. A Newton–Marchenko equation and generalized optical theorem for elastodynamics. Journal of Mathematical Physics, 33(8):2903–2915, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i8/p2903_s1.

Pavlov:1992:ETQ

[1248] B. S. Pavlov and M. A. Pankratov. Electron transport in quasi-one-dimensional random structure. Journal of Mathematical Physics, 33 (8):2916–2922, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2916_s1.

Makhankov:1992:CCH

[1249] V. G. Makhan′kov and O. K. Pashaev. Continual classical Heisenberg models defined on graded su(2, 1) and su(3) algebras. Journal of Mathematical Physics, 33(8):2923–2936, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i8/p2923_s1.

Frappat:1992:EDL

[1250] L. Frappat, A. Sciarrino, and P. Sorba. Erratum: “Dynkin-like diagrams and representations of the strange superalgebra P(n)”, [J. Math. Phys. 32, 3268 (1991)]. Journal of Mathematical Physics, 33(8):2937, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i8/p2937_s1. See [890].

Li:1992:EPN

[1251] Fu Bin Li. Erratum: “Proof of the noninteraction theorem for a system of N relativistic particles in direct interaction”, [J. Math. Phys. 31 (1990), no. 6, 1395–1399; MR 92a:70013]. Journal of Mathematical Physics, 33(8):2938, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i8/p2938_s1. See [208].

MartinezAlonso:1992:LCS

[1252] L. Martínez Alonso and E. Medina Reus. Localized coherent structures of the Davey–Stewartson equation in the bilinear formalism. Journal of Mathematical Physics, 33(9):2947–2957, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p2947_s1.

Bonatsos:1992:WEP

[1253] Dennis Bonatsos, C. Daskaloyannis, and K. Kokkotas. WKB equivalent potentials for q-deformed harmonic and anharmonic oscillators. Journal of Mathematical Physics, 33(9):2958–2965, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p2958_s1.

Gat:1992:SAT

[1254] Omri Gat. Symmetry algebras of third-order ordinary differential equations. Journal of Mathematical Physics, 33(9):2966–2971, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i9/p2966_s1.

Marshalek:1992:NBR

[1255] Eugene R. Marshalek. A new boson realization for the isotropic threedimensional oscillator. Journal of Mathematical Physics, 33(9):2972– 2983, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p2972_s1.

Floreanini:1992:AFB

[1256] Roberto Floreanini and Luc Vinet. Addition formulas for q-Bessel functions. Journal of Mathematical Physics, 33(9):2984–2988, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i9/p2984_s1.

Duggal:1992:CIS

[1257] K. L. Duggal. Curvature inheritance symmetry in Riemannian spaces with applications to fluid space times. Journal of Mathematical Physics, 33(9):2989–2997, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i9/p2989_s1.

Godoy:1992:RPB

[1258] S. Godoy and S. Fujita. Reflection principles for biased correlated walks. Simple applications. Journal of Mathematical Physics, 33(9):2998–3003, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p2998_s1.

McRae:1992:CPE

[1259] S. M. McRae and E. R. Vrscay. Canonical perturbation expansions to large order from classical hypervirial and Hellmann–Feynman theorems. Journal of Mathematical Physics, 33(9):3004–3024, September 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3004_
s1.

Bilge:1992:ISO

[1260] Ayşe Hümeyra Bilge. Integrability of seventh-order KdV-like evolution equations using formal symmetries and perturbations of conserved densities. Journal of Mathematical Physics, 33(9):3025–3038, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i9/p3025_s1.

Matsuno:1992:PFD

[1261] Y. Matsuno. Pulse formation in a dissipative nonlinear system. Journal of Mathematical Physics, 33(9):3039–3045, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3039_s1.

Bezerra:1992:SSS

[1262] Valdir B. Bezerra and Ivonete B. dos Santos. Scalar and spinor solutions in the space–time of a domain wall. Journal of Mathematical Physics, 33(9):3046–3049, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i9/p3046_s1.

Djehiche:1992:BPS

[1263] Boualem Djehiche. Bernstein processes and spin-1/2 particles. Journal of Mathematical Physics, 33(9):3050–3059, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3050_s1.

Lu:1992:WCL

[1264] Y. G. Lu. The weak coupling limit for quadratic interaction. Journal of Mathematical Physics, 33(9):3060–3086, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3060_s1.

Aldaya:1992:SQH

[1265] V. Aldaya, J. Navarro-Salas, J. Bisquert, and R. Loll. Symmetry and quantization: Higher-order polarization and anomalies. Journal of Mathematical Physics, 33(9):3087–3097, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i9/p3087_s1.

Razon:1992:USP

[1266] A. Razon and L. P. Horwitz. Uniqueness of the scalar product in the tensor product of quaternion Hilbert modules. Journal of Mathematical Physics, 33(9):3098–3104, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i9/p3098_s1.

Latypov:1992:CSK

[1267] D. M. Latypov and A. M. Mukhamedzhanov. On the Coulomb singular kernel of Lippmann–Schwinger-type equation. Journal of Mathematical Physics, 33(9):3105–3107, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i9/p3105_s1.

Bytsenko:1992:VED

[1268] Andrei A. Bytsenko, Guido Cognola, and Luciano Vanzo. Vacuum energy for (3+1)-dimensional space–time with compact hyperbolic spatial part.

Journal of Mathematical Physics, 33(9):3108–3111, September 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3108_

s1. See erratum [1522].

Juriev:1992:ASQb

[1269] D. Juriev. Algebraic structures of the quantum projective [sl(2, C)- invariant] field theory: the commutative version of Élie Cartan’s exterior differential calculus. Journal of Mathematical Physics, 33(9):3112–3116, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3112_s1.

Srivastava:1992:DTC

[1270] S. K. Srivastava. D-dimensional torus as compact manifold and Kaluza– Klein cosmological model. Journal of Mathematical Physics, 33(9):3117– 3127, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3117_s1.

Sainty:1992:CCV

[1271] Philippe Sainty. Construction of a complex-valued fractional Brownian motion of order N. Journal of Mathematical Physics, 33(9):3128–3149, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3128_s1.

Cohen:1992:RPF

[1272] E. Cohen, H. Kluberg-Stern, and H. Navelet. Regulated propagator on the flat torus and its Weyl properties. Journal of Mathematical Physics, 33(9):3150–3158, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i9/p3150_s1.

Merkulov:1992:GQS

[1273] Sergey A. Merkulov. Geometric quantization of N = 2 superstring. Journal of Mathematical Physics, 33(9):3159–3171, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3159_s1.

Chang:1992:GHP

[1274] Zhe Chang. Generalized Holstein–Primakoff realizations and quantum group-theoretic coherent states. Journal of Mathematical Physics, 33 (9):3172–3179, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i9/p3172_s1. See erratum [1715].

Yu:1992:BHS

[1275] Feng Yu. Bi-Hamiltonian structure of super KP hierarchy. Journal of Mathematical Physics, 33(9):3180–3189, September 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3180_s1.

Kim:1992:LLG

[1276] Y. S. Kim and Leehwa Yeh. O(2, 1)-like little group for spacelike fourmomenta and localized light waves in continuous media. Journal of Mathematical Physics, 33(9):3190–3196, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i9/p3190_s1.

Harnad:1992:IGT

[1277] J. Harnad and S. Shnider. Isotropic geometry and twistors in higher dimensions. I. The generalized Klein correspondence and spinor flags in even dimensions. Journal of Mathematical Physics, 33(9):3197–3208, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3197_s1.

Chan:1992:ASQ

[1278] A. H. Chan and C. K. Chew. Asymptotic solution to the QCD-parton branching equation. Journal of Mathematical Physics, 33(9):3209–3212, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3209_s1.

Ban:1992:LAA

[1279] Masashi Ban. SU(1, 1) Lie algebraic approach to linear dissipative processes in quantum optics. Journal of Mathematical Physics, 33(9):3213– 3228, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3213_s1.

Pennisi:1992:HML

[1280] S. Pennisi and M. Sammartino. Hyperbolic models for linear transport theory. Journal of Mathematical Physics, 33(9):3229–3239, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i9/p3229_s1.

Kuznetsov:1992:QRR

[1281] Vadim B. Kuznetsov. Quadrics on real Riemannian spaces of constant curvature: Separation of variables and connection with Gaudin magnet. Journal of Mathematical Physics, 33(9):3240–3254, September 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i9/p3240_
s1.

Wahlquist:1992:EPE

[1282] H. D. Wahlquist. Erratum: “The problem of exact interior solutions for rotating rigid bodies in general relativity”, [J. Math. Phys. 33 (1992), no. 1, 304–335; MR 92m:83017]. Journal of Mathematical Physics, 33(9): 3255, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i9/p3255_s1. See [961].

Newton:1992:EPI

[1283] Roger G. Newton. Editorial. Journal of Mathematical Physics, 33(9): iii–iv, September 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pestov:1992:CNS

[1284] Vladimir G. Pestov. A contribution to nonstandard superanalysis. Journal of Mathematical Physics, 33(10):3263–3273, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3263_s1.

Halpern:1992:GGT

[1285] M. B. Halpern and N. A. Obers. Generalized graph theory on Lie g and applications in conformal field theory. Journal of Mathematical Physics, 33(10):3274–3295, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i10/p3274_s1.

Alisauskas:1992:EOS

[1286] Sigitas Ališauskas. Explicit orthogonalization of some biorthogonal bases for SU(n) ⊃ SO(n) and Sp(4) ⊃ U(2). Journal of Mathematical Physics, 33(10):3296–3312, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i10/p3296_s1.

Sun:1992:BDC

[1287] Xiao Dong Sun and Shi Kun Wang. Bicovariant differential calculus on the two-parameter quantum group GL_{p,q}(2). Journal of Mathematical Physics, 33(10):3313–3329, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i10/p3313_s1.

Antone:1992:ASC

[1288] T. A. Antone and A. A. Al-Maaitah. Analytical solutions to classes of linear oscillator equations with time varying frequencies. Journal of Mathematical Physics, 33(10):3330–3339, October 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3330_s1.

Baird:1992:RTH

[1289] Paul Baird. Riemannian twistors and Hermitian structures on lowdimensional space forms. Journal of Mathematical Physics, 33(10):3340– 3355, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3340_s1.

Fleury:1992:LP

[1290] N. Fleury and M. Rausch de Traubenberg. Linearization of polynomials. Journal of Mathematical Physics, 33(10):3356–3366, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/
p3356_s1.

Thomas:1992:SHS

[1291] Lawrence E. Thomas and Stephen R. Wassell. Stability of Hamiltonian systems at high energy. Journal of Mathematical Physics, 33(10):3367– 3373, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3367_s1.

Arai:1992:MOG

[1292] Asao Arai. Momentum operators with gauge potentials, local quantization of magnetic flux, and representation of canonical commutation relations. Journal of Mathematical Physics, 33(10):3374–3378, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3374_s1.

Ballesteros:1992:MQD

[1293] A. Ballesteros, M. Gadella, and M. A. Del Olmo. Moyal quantization of (2 + 1)-dimensional Galilean systems. Journal of Mathematical Physics, 33(10):3379–3386, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i10/p3379_s1.

Beckers:1992:PQO

[1294] J. Beckers, N. Debergh, and A. G. Nikitin. On pararelativistic quantum oscillators. Journal of Mathematical Physics, 33(10):3387–3392, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3387_s1.

Garbaczewski:1992:RDS

[1295] Piotr Garbaczewski. Rotational diffusions as seen by relativistic observers. Journal of Mathematical Physics, 33(10):3393–3401, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3393_s1.

Nieto:1992:GFC

[1296] Luis M. Nieto. Green’s function for crossed time-dependent electric and magnetic fields. Phase-space quantum mechanics approach. Journal of Mathematical Physics, 33(10):3402–3409, October 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3402_s1.

Chetouani:1992:ATG

[1297] L. Chetouani, L. Guechi, and T. F. Hammann. Algebraic treatment of a general noncentral potential. Journal of Mathematical Physics, 33 (10):3410–3418, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i10/p3410_s1.

Dobrev:1992:DMQ

[1298] V. K. Dobrev. Duality for the matrix quantum group GL_{p,q}(2, C). Journal of Mathematical Physics, 33(10):3419–3430, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3419_s1.

Majid:1992:SQG

[1299] Shahn Majid. C-statistical quantum groups and Weyl algebras. Journal of Mathematical Physics, 33(10):3431–3444, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3431_s1.

Meyer:1992:BFBa

[1300] Jutta Meyer. Boson-fermion and baryon mapping: construction of collective subspaces. I. Theory. Journal of Mathematical Physics, 33(10):3445– 3453, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3445_s1.

Meyer:1992:BFBb

[1301] Jutta Meyer. Boson-fermion and baryon mapping: Construction of collective subspaces. II. Example: A simple quark shell model. Journal of Mathematical Physics, 33(10):3454–3462, October 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3454_s1.

Meyer:1992:BFBc

[1302] Jutta Meyer. Boson-fermion and baryon mapping: construction of collective subspaces. III. Application of the baryon mapping to many-baryon states. Journal of Mathematical Physics, 33(10):3463–3467, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3463_s1.

Rabello:1992:ESW

[1303] S. J. Rabello and A. N. Vaidya. Evaluation of Schwinger–de Witt kernels for spin-1/2 fields through the SO(2, 1) Lie algebra. Journal of Mathematical Physics, 33(10):3468–3471, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i10/p3468_s1.

Hall:1992:SWR

[1304] Richard L. Hall. Square-well representations for potentials in quantum mechanics. Journal of Mathematical Physics, 33(10):3472–3476, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3472_s1.

Chandler:1992:BQS

[1305] Colston Chandler and Archie G. Gibson. N-body quantum scattering theory in two Hilbert spaces. VI. Compactness conditions. Journal of Mathematical Physics, 33(10):3477–3492, October 1992. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i10/p3477_s1.

Prosser:1992:FSI

[1306] Reese T. Prosser. Formal solutions of inverse scattering problems. V. Journal of Mathematical Physics, 33(10):3493–3496, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/
p3493_s1.

Heusler:1992:NHT

[1307] Markus Heusler. A no-hair theorem for self-gravitating nonlinear sigma models. Journal of Mathematical Physics, 33(10):3497–3502, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3497_s1.

Jiang:1992:GSU

[1308] Sheng Jiang. Generalized Stephani universes. Journal of Mathematical Physics, 33(10):3503–3505, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i10/p3503_s1.

Li:1992:STP

[1309] Jian Zeng Li. Space–times with plane-symmetric scalar waves. Journal of Mathematical Physics, 33(10):3506–3508, October 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/

p3506_s1. See comment [1870].

Khodunov:1992:GES

[1310] A. V. Khodunov and V. V. Zhytnikov. Gravitational equations in space–time with torsion. Journal of Mathematical Physics, 33(10):3509– 3520, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3509_s1.

Matos:1992:ESI

[1311] Tonatiuh Matos, Guadalupe Rodríguez, and Ricardo Becerril. Exact solutions of SL(N, R)-invariant chiral equations one- and two-dimensional subspaces. Journal of Mathematical Physics, 33(10):3521–3535, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3521_s1.

Jumarie:1992:FPE

[1312] Guy Jumarie. A Fokker–Planck equation of fractional order with respect to time. Journal of Mathematical Physics, 33(10):3536–3542, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i10/p3536_s1.

Fairlie:1992:EHU

[1313] D. B. Fairlie and J. Govaerts. Euler hierarchies and universal equations. Journal of Mathematical Physics, 33(10):3543–3566, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i10/
p3543_s1.

Catto:1992:ILS

[1314] Sultan Catto and Yusuf Gürsey. Instanton-like solutions in conformal chiral superfield theories. Journal of Mathematical Physics, 33(10):3567– 3573, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3567_s1.

vanderMee:1992:SMT

[1315] C. V. M. van der Mee and J. W. Hovenier. Structure of matrices transforming Stokes parameters. Journal of Mathematical Physics, 33 (10):3574–3584, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i10/p3574_s1.

Coggeshall:1992:GIS

[1316] S. V. Coggeshall and J. Meyer ter Vehn. Group-invariant solutions and optimal systems for multidimensional hydrodynamics. Journal of Mathematical Physics, 33(10):3585–3601, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i10/p3585_s1.

Demys:1992:ECH

[1317] K. Demys. Erratum: “Comment: ‘On the hyperbolic complex linear symmetry groups and their local gauge transformation actions’ by Z. Z. Zhong”, [J. Math. Phys. 26, 404 (1985)]; [J. Math. Phys. 28 (1987), no. 2, 339; MR 88a:83062]. Journal of Mathematical Physics, 33(10): 3602, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3602_s1. See [1, 3].

Gould:1992:PCT

[1318] Mark Gould and L. C. Biedenharn. The pattern calculus for tensor operators in quantum groups. Journal of Mathematical Physics, 33(11):3613– 3635, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3613_s1.

Nomura:1992:SRQ

[1319] Masao Nomura and L. C. Biedenharn. On the q-symplecton realization of the quantum group SU_{q}(2). Journal of Mathematical Physics, 33 (11):3636–3648, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3636_s1.

Parks:1992:FBC

[1320] A. D. Parks. The Fermi and Bose congruences for free semigroups on two generators. Journal of Mathematical Physics, 33(11):3649–3652, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3649_s1.

Ohnuki:1992:SAO

[1321] Yoshio Ohnuki and Shuji Watanabe. Self-adjointness of the operators in Wigner’s commutation relations. Journal of Mathematical Physics, 33 (11):3653–3665, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3653_s1.

Zubelli:1992:ZCP

[1322] Jorge P. Zubelli. On a zero curvature problem related to the ZS-AKNS operator. Journal of Mathematical Physics, 33(11):3666–3675, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3666_s1.

Konopelchenko:1992:NRK

[1323] Boris Konopel′chenko and Walter Strampp. New reductions of the Kadomtsev–Petviashvili and two-dimensional Toda lattice hierarchies via symmetry constraints. Journal of Mathematical Physics, 33(11):3676– 3686, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3676_s1.

Sophocleous:1992:SRS

[1324] Christodoulos Sophocleous. On symmetries of radially symmetric nonlinear diffusion equations. Journal of Mathematical Physics, 33(11):3687– 3693, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3687_s1.

Wisse:1992:QSM

[1325] M. A. Wisse. Quasiperiodic solutions for matrix nonlinear Schrödinger equations. Journal of Mathematical Physics, 33(11):3694–3699, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3694_s1.

Klauder:1992:UPA

[1326] John R. Klauder and Wolfgang A. Tomé. The universal propagator for affine [or SU(1, 1)] coherent states. Journal of Mathematical Physics, 33 (11):3700–3709, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3700_s1.

Chang:1992:MAQ

[1327] D. Chang, I. Phillips, and L. Rozansky. R-matrix approach to quantum superalgebras su_{q}(m|n). Journal of Mathematical Physics, 33(11):3710– 3715, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3710_s1.

Edgar:1992:CSC

[1328] S. Brian Edgar. Conditions for a symmetric connection to be a metric connection. Journal of Mathematical Physics, 33(11):3716–3722, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3716_s1.

Swift:1992:NBM

[1329] S. Timothy Swift. Natural bundles. I. A minimal resolution of superspace. Journal of Mathematical Physics, 33(11):3723–3730, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3723_s1.

Hinsen:1992:RDE

[1330] K. Hinsen and B. U. Felderhof. Reduced description of electric multipole potential in Cartesian coordinates. Journal of Mathematical Physics, 33 (11):3731–3735, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3731_s1.

Iguchi:1992:QSC

[1331] Kazumoto Iguchi. Quasiperiodic systems without Cantor-set-like energy bands. Journal of Mathematical Physics, 33(11):3736–3739, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3736_s1.

Linde:1992:AAS

[1332] Johan Linde and Håkan Snellman. Approximate analytic solutions to the Skyrme equation. Journal of Mathematical Physics, 33(11):3740– 3745, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3740_s1.

Scherer:1992:BHS

[1333] W. Scherer. Bi-Hamiltonian systems and Maslov indices. Journal of Mathematical Physics, 33(11):3746–3756, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3746_s1.

deOliveira:1992:SHG

[1334] E. Capelas de Oliveira. On the solution of homogeneous generalized wave equation. Journal of Mathematical Physics, 33(11):3757–3758, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3757_s1.

Chan:1992:LSI

[1335] W. L. Chan, Kam-Shun Li, and Yi Shen Li. Line soliton interactions of a nonisospectral and variable-coefficient Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 33(11):3759–3773, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3759_s1.

Cheng:1992:CKP

[1336] Yi Cheng. Constraints of the Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 33(11):3774–3782, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3774_s1.

Degasperis:1992:GIL

[1337] A. Degasperis, D. Lebedev, M. Olshanetsky, S. Pakuliak, A. Perelomov, and P. M. Santini. Generalized intermediate long-wave hierarchy in zero-curvature representation with noncommutative spectral parameter. Journal of Mathematical Physics, 33(11):3783–3793, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/
p3783_s1.

Popov:1992:RNS

[1338] I. Yu. Popov. The resonator with narrow slit and the model based on the operator extensions theory. Journal of Mathematical Physics, 33 (11):3794–3801, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3794_s1.

Torre:1992:CPS

[1339] C. G. Torre. Covariant phase space formulation of parametrized field theories. Journal of Mathematical Physics, 33(11):3802–3812, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3802_s1.

Klibanov:1992:PIS

[1340] Michael V. Klibanov and Paul E. Sacks. Phaseless inverse scattering and the phase problem in optics. Journal of Mathematical Physics, 33 (11):3813–3821, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3813_s1.

Fukumi:1992:WNC

[1341] Toshio Fukumi. White-noise calculation of annihilation and creation operators in stochastic mechanics. Journal of Mathematical Physics, 33 (11):3822–3825, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3822_s1.

Basu:1992:CGP

[1342] Debabrata Basu. The Clebsch–Gordan problem of the SL(2, R) coherent states. Journal of Mathematical Physics, 33(11):3826–3835, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3826_s1.

Tanimura:1992:MGI

[1343] N. Tanimura and O. Tanimura. Multifold group integral for SU(2) character and sum rule among 6-j symbols. Journal of Mathematical Physics, 33(11):3836–3847, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3836_s1.

Tsukada:1992:FRQ

[1344] Haruo Tsukada. Functional realization of quantum affine algebras. Journal of Mathematical Physics, 33(11):3848–3864, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/
p3848_s1.

Aktosun:1992:FSM

[1345] Tuncay Aktosun. A factorization of the scattering matrix for the Schrödinger equation and for the wave equation in one dimension. Journal of Mathematical Physics, 33(11):3865–3869, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/
p3865_s1.

Combescure:1992:SSA

[1346] M. Combescure. The squeezed state approach of the semiclassical limit of the time-dependent Schrödinger equation. Journal of Mathematical Physics, 33(11):3870–3880, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i11/p3870_s1.

Frampton:1992:PSS

[1347] Paul H. Frampton, Y. Jack Ng, and Hendrik Van Dam. Possible solution of strong CP problem in generalized unimodular gravity. Journal of Mathematical Physics, 33(11):3881–3882, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3881_s1.

Wesson:1992:KKE

[1348] Paul S. Wesson and J. Ponce de León. Kaluza–Klein equations, Einstein’s equations, and an effective energy-momentum tensor. Journal of Mathematical Physics, 33(11):3883–3887, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3883_s1.

Liu:1992:ESG

[1349] Hong Ya Liu and P. S. Wesson. Exact solutions of general relativity derived from 5-D “black hole” solutions of Kaluza–Klein theory. Journal of Mathematical Physics, 33(11):3888–3891, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3888_s1.

Grosse:1992:NSN

[1350] Harald Grosse, Wolfgang Kummer, Peter Prešnajder, and Dominik J. Schwarz. Novel symmetry of non-Einsteinian gravity in two dimensions. Journal of Mathematical Physics, 33(11):3892–3900, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3892_s1.

You:1992:PTA

[1351] J. Q. You, J. R. Yan, and J. X. Zhong. Phase transition in aperiodic quantum Ising models. Journal of Mathematical Physics, 33(11):3901– 3906, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3901_s1.

Kaup:1992:CFS

[1352] D. J. Kaup and Gregory H. Paine. A closed form solution of the onedimensional Born–Green–Yvon equation for a hard-rod fluid. Journal of Mathematical Physics, 33(11):3907–3910, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3907_s1.

Frappat:1992:ORS

[1353] L. Frappat and A. Sciarrino. Oscillator realization of the strange superalgebras P(n) and Q(n). Journal of Mathematical Physics, 33(11):3911– 3917, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i11/p3911_s1.

Zhang:1992:BGR

[1354] R. B. Zhang. Braid group representations arising from quantum supergroups with arbitrary q and link polynomials. Journal of Mathematical Physics, 33(11):3918–3930, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i11/p3918_s1.

Ibragimov:1992:SMG

[1355] N. H. Ibragimov and M. Torrisi. A simple method for group analysis and its application to a model of detonation. Journal of Mathematical Physics, 33(11):3931–3937, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i11/p3931_s1.

Iguchi:1992:EWF

[1356] Kazumoto Iguchi. Exact wave functions of an electron on a quasiperiodic lattice: Definition of an infinite-dimensional Riemann theta function. Journal of Mathematical Physics, 33(11):3938–3947, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i11/p3938_s1.

Martin:1992:TPG

[1357] Paul Martin. On a two-parameter generalization of Hecke algebras and the XY chain in a magnetic field. Journal of Mathematical Physics, 33 (11):3948–3953, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i11/p3948_s1.

Magnon:1992:CSP

[1358] Anne Magnon. Cosmological singularity at the Planck scale. Journal of Mathematical Physics, 33(11):3954–3956, November 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i11/p3954_s1.

Cornwell:1992:MDU

[1359] J. F. Cornwell. Multiparameter deformations of the universal enveloping algebras of the simple Lie algebras A_{l} for all l ≥ 2 and the Yang– Baxter equation. Journal of Mathematical Physics, 33(12):3963–3977, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p3963_s1.

Przanowski:1992:DRL

[1360] M. Przanowski and A. Maciołek-Niedżwiecki. Duality rotation as a Lie– Bäcklund transformation. Journal of Mathematical Physics, 33(12):3978– 3982, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p3978_s1.

Herlt:1992:ITC

[1361] E. Herlt and H. Stephani. Invariance transformations of the class y′′ = F (x)y^{n} of differential equations. Journal of Mathematical Physics, 33 (12):3983–3988, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p3983_s1.

Turbiner:1992:PSD

[1362] A. Turbiner. On polynomial solutions of differential equations. Journal of Mathematical Physics, 33(12):3989–3993, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p3989_s1.

Yukalov:1992:SCM

[1363] V. I. Yukalov. Stability conditions for method of self-similar approximations. Journal of Mathematical Physics, 33(12):3994–4001, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i12/p3994_s1.

Tariq:1992:CSI

[1364] N. Tariq and B. O. J. Tupper. Conformal symmetry inheritance with cosmological constant. Journal of Mathematical Physics, 33(12):4002– 4007, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4002_s1.

Thompson:1992:NFM

[1365] G. Thompson. Normal form of a metric admitting a parallel field of planes. Journal of Mathematical Physics, 33(12):4008–4010, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i12/p4008_s1.

Trautman:1992:SDO

[1366] Andrzej Trautman. Spinors and the Dirac operator on hypersurfaces. I. General theory. Journal of Mathematical Physics, 33(12):4011–4019, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4011_s1.

Nagatomo:1992:RSD

[1367] Yasuyuki Nagatomo. Rigidity of c_{1}-self-dual connections on quaternionic Kähler manifolds. Journal of Mathematical Physics, 33(12):4020–4025, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4020_s1.

Liu:1992:ASG

[1368] J. W. Liu. Analytical solutions to the generalized spheroidal wave equation and the Green’s function of one-electron diatomic molecules. Journal of Mathematical Physics, 33(12):4026–4036, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4026_s1.

Shishkin:1992:NPGb

[1369] German V. Shishkin and Víctor M. Villalba. Neutrino in the presence of gravitational fields: Exact solutions. Journal of Mathematical Physics, 33(12):4037–4045, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4037_s1.

Awane:1992:SS

[1370] A. Awane. k-symplectic structures. Journal of Mathematical Physics, 33(12):4046–4052, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4046_s1.

Chiang:1992:GJD

[1371] R. C. Chiang and S. Y. Shieh. Generalization of Jacobi’s decomposition theorem to the rotation and translation of a solid in a fluid. Journal of Mathematical Physics, 33(12):4053–4065, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4053_s1.

Corichi:1992:CHF

[1372] Alejandro Corichi. Comment on: “Hamiltonian formulation of a theory with constraints”, [J. Math. Phys. 33 (1992), no. 2, 633–638; MR 92m:81116] by U. Kulshreshta. Journal of Mathematical Physics, 33 (12):4066–4067, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4066_s1. See [996].

Lakshmanan:1992:DIG

[1373] M. Lakshmanan and M. Senthil Velan. Direct integration of generalized Lie or dynamical symmetries of three degrees of freedom nonlinear Hamiltonian systems: Integrability and separability. Journal of Mathematical Physics, 33(12):4068–4077, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i12/p4068_s1.

Eringen:1992:CTM

[1374] A. Cemal Eringen. Continuum theory of microstretch liquid crystals. Journal of Mathematical Physics, 33(12):4078–4086, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/
p4078_s1.

Gazeau:1992:SVC

[1375] J. P. Gazeau and P. Winternitz. Symmetries of variable coefficient Korteweg–de Vries equations. Journal of Mathematical Physics, 33 (12):4087–4102, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4087_s1.

He:1992:THG

[1376] Sailing He. A time-harmonic Green’s function technique and wave propagation in a stratified nonreciprocal chiral slab with multiple discontinuities. Journal of Mathematical Physics, 33(12):4103–4110, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i12/p4103_s1.

Vaklev:1992:SSS

[1377] Y. S. Vaklev. Some soliton solutions for the quadratic bundle. Journal of Mathematical Physics, 33(12):4111–4115, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4111_s1.

Bosco:1992:CTI

[1378] B. Bosco, P. Ricci, and M. T. Sacchi. The Casimir theorem and the inversion problem in electrodynamics at first quantized level. Journal of Mathematical Physics, 33(12):4116–4132, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4116_s1.

Lahti:1992:OQT

[1379] Pekka J. Lahti and Maciej J. Maczyński. Orthomodularity and quadratic transformations in probabilistic theories of physics. Journal of Mathematical Physics, 33(12):4133–4138, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i12/p4133_s1.

Carreau:1992:FIF

[1380] Michel Carreau. The functional integral for a free particle on a halfplane. Journal of Mathematical Physics, 33(12):4139–4147, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i12/p4139_s1.

Dittmann:1992:CCG

[1381] Jochen Dittmann and Gerd Rudolph. A class of connections governing parallel transport along density matrices. Journal of Mathematical Physics, 33(12):4148–4154, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i12/p4148_s1.

Anzaldo-Meneses:1992:KNA

[1382] A. Anzaldo-Meneses. Krichever–Novikov algebras on Riemann surfaces of genus zero and one with N punctures. Journal of Mathematical Physics, 33(12):4155–4163, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4155_s1.

Harnad:1992:MNS

[1383] J. Harnad and M.-A. Wisse. Matrix nonlinear Schrödinger equations and moment maps into loop algebras. Journal of Mathematical Physics, 33 (12):4164–4176, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4164_s1.

Meister:1992:AOC

[1384] Andreas Meister. The C^{∗}-algebra obtained by completing the quantum group Fun_{q}(SU(2)). Journal of Mathematical Physics, 33(12):4177– 4189, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4177_s1.

Stochel:1992:CCR

[1385] Jan Stochel and Todor. S. Todorov. Characterizations of cyclic ∗- representations with equal weak and strong commutants. Journal of Mathematical Physics, 33(12):4190–4195, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4190_s1.

Witwit:1992:ELTb

[1386] M. R. M. Witwit. Energy levels of two-dimensional anharmonic oscillators with sextic and octic perturbations. Journal of Mathematical Physics, 33(12):4196–4205, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i12/p4196_s1.

Charalambopoulos:1992:ISL

[1387] Antonios Charalambopoulos and George Dassios. Inverse scattering via low-frequency moments. Journal of Mathematical Physics, 33(12):4206– 4216, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4206_s1.

Modanese:1992:MTP

[1388] Giovanni Modanese. On the motion of test particles in a fluctuating gravitational field. Journal of Mathematical Physics, 33(12):4217–4219, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4217_s1.

Rylov:1992:DRS

[1389] Yuri A. Rylov. Distorted Riemannian space and technique of differential geometry. Journal of Mathematical Physics, 33(12):4220–4224, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4220_s1.

Halilsoy:1992:NMS

[1390] Mustafa Halilsoy. New metrics for spinning spheroids in general relativity. Journal of Mathematical Physics, 33(12):4225–4230, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i12/p4225_s1.

Tresguerres:1992:TGT

[1391] Romualdo Tresguerres. Topological gravity in a three-dimensional metric-affine space–time. Journal of Mathematical Physics, 33(12):4231– 4238, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4231_s1.

Wang:1992:SGB

[1392] Sheng Wang and Yi Song Yang. Solutions of the generalized Bogomol′nyĭequations via monotone iterations. Journal of Mathematical Physics, 33(12):4239–4249, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i12/p4239_s1.

Schimming:1992:CPS

[1393] R. Schimming and E. Mundt. Constant potential solutions of the Yang– Mills equation. Journal of Mathematical Physics, 33(12):4250–4254, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i12/p4250_s1.

Furlan:1992:QDL

[1394] Paolo Furlan, Ludmil K. Hadjiivanov, and Ivan T. Todorov. Quantum deformation of the ladder representations of U(1, 1) for |q| = 1. Journal of Mathematical Physics, 33(12):4255–4258, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4255_s1.

Finkelstein:1992:DO

[1395] R. J. Finkelstein. q deformations of the oscillator. Journal of Mathematical Physics, 33(12):4259–4266, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i12/p4259_s1.

Jarvis:1992:UMGb

[1396] P. D. Jarvis and D. S. McAnally. Unified models from gauged supergroups. III. The “super-2-sphere” OSp(3|2)/OSp(2|2). Journal of Mathematical Physics, 33(12):4267–4273, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v33/i12/p4267_s1.

Nagamachi:1992:HS

[1397] Shigeaki Nagamachi and Yuji Kobayashi. Hilbert superspace. Journal of Mathematical Physics, 33(12):4274–4282, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v33/i12/p4274_s1.

Kim:1992:MCS

[1398] Won-Tae Kim, Jae Kwan Kim, and Young-Jai Park. Minimal chiral Schwinger model in the matter gauge. Journal of Mathematical Physics, 33(12):4283–4287, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4283_s1.

Suzuki:1992:SCL

[1399] Noboru Suzuki. Simple calculation of Löwdin’s alpha function. IV. Procedure for evaluating the coefficients in expansion of a Slater-type atomic orbital. Journal of Mathematical Physics, 33(12):4288–4299, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v33/i12/p4288_s1.

Lou:1992:NSR

[1400] Sen Yue Lou. Nonclassical symmetry reductions for the dispersive wave equations in shallow water. Journal of Mathematical Physics, 33 (12):4300–4305, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v33/i12/p4300_s1.

Burrows:1993:LBQ

[1401] B. L. Burrows, M. Cohen, and Tova Feldmann. Lower bounds for quartic anharmonic and double-well potentials. Journal of Mathematical Physics, 34(1):1–11, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i1/p1_s1.

Cohen:1993:ELD

[1402] Jeffrey M. Cohen and Boris Kuharetz. Energy level degeneracy. Journal of Mathematical Physics, 34(1):12–22, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i1/p12_s1.

Exner:1993:BSQ

[1403] Pavel Exner. Bound states in quantum waveguides of a slowly decaying curvature. Journal of Mathematical Physics, 34(1):23–28, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i1/p23_s1.

Mekideche:1993:SBP

[1404] F. Mekideche and B. G. Giraud. Symmetry breaking by propagation. Journal of Mathematical Physics, 34(1):29–49, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i1/p29_s1.

Nakamura:1993:QDS

[1405] M. Nakamura and H. Minowa. Quantization of dynamical systems subject to second class constraints. Journal of Mathematical Physics, 34(1): 50–58, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p50_s1.

Pereyra:1993:PRF

[1406] P. Pereyra and R. Quezada. Probabilistic representation formulas for the time evolution of quantum systems. Journal of Mathematical Physics, 34(1):59–68, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i1/p59_s1.

Swanson:1993:TRC

[1407] D. G. Swanson and V. F. Shvets. Transmission and reflection coefficients from the tunneling equation with absorption in the complex z plane.

Journal of Mathematical Physics, 34(1):69–88, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i1/p69_s1.

Tanimura:1993:RRIa

[1408] N. Tanimura and O. Tanimura. Recursion relations in U(N) integral in lattice QCD. Journal of Mathematical Physics, 34(1):89–99, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanimura:1993:RRI

[1409] N. Tanimura and O. Tanimura. Recursion relations in U(N) integral in lattice QCD. Journal of Mathematical Physics, 34(1):89–99, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i1/p89_s1.

Trifonov:1993:CGS

[1410] D. A. Trifonov. Completeness and geometry of Schrödinger minimum uncertainty states. Journal of Mathematical Physics, 34(1):100–110, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i1/p100_s1.

Berman:1993:SHI

[1411] M. S. Berman and M. M. Som. Spatially homogeneous isotropic models with nonminimally coupled electromagnetic field. Journal of Mathematical Physics, 34(1):111–114, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i1/p111_s1.

Braham:1993:GEV

[1412] Stephen P. Braham. Generalized embedding variables for geometrodynamics and space–time diffeomorphisms: Ultralocal coordinate conditions. Journal of Mathematical Physics, 34(1):115–129, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i1/
p115_s1.

Capovilla:1993:SES

[1413] Riccardo Capovilla and Jerzy Plebański. Some exact solutions of the Einstein field equations in terms of the self-dual spin connection. Journal of Mathematical Physics, 34(1):130–138, January 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i1/p130_s1.

Falbo-Kenkel:1993:NTB

[1414] M. K. Falbo-Kenkel and F. Mansouri. Nonperturbative two-body dynamics in (2 + 1)-dimensional gravity. Journal of Mathematical Physics, 34(1):139–153, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i1/p139_s1.

Hogan:1993:SHF

[1415] P. A. Hogan and T. Futamase. Some high-frequency spherical gravity waves. Journal of Mathematical Physics, 34(1):154–169, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i1/p154_s1.

Manko:1993:KNM

[1416] V. S. Man′ko and N. R. Sibgatullin. Kerr–Newman metric endowed with magnetic dipole moment. Journal of Mathematical Physics, 34(1):170– 177, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p170_s1.

Nolan:1993:SMM

[1417] Brien Nolan. Sources for McVittie’s mass particle in an expanding universe. Journal of Mathematical Physics, 34(1):178–185, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i1/
p178_s1.

Santos:1993:CSF

[1418] J. Santos, M. J. Rebouças, and A. F. F. Teixeira. Classification of scalar fields in general relativity. Journal of Mathematical Physics, 34(1):186– 192, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p186_s1.

Guthrie:1993:NSK

[1419] Graeme A. Guthrie and Mark S. Hickman. Nonlocal symmetries of the KdV equation. Journal of Mathematical Physics, 34(1):193–205, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i1/p193_s1.

Kharik:1993:GIW

[1420] V. M. Kharik. Generation of internal waves in a rotating, stratified fluid: Some symmetry properties. Journal of Mathematical Physics, 34(1):206– 213, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p206_s1.

Konopelchenko:1993:GLS

[1421] B. G. Konopel′chenko and C. Rogers. On generalized Loewner systems: Novel integrable equations in 2 + 1 dimensions. Journal of Mathematical Physics, 34(1):214–242, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i1/p214_s1.

Strachan:1993:SIH

[1422] I. A. B. Strachan. Some integrable hierarchies in (2 + 1) dimensions and their twistor description. Journal of Mathematical Physics, 34(1):243– 259, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p243_s1.

Hock:1993:ECT

[1423] J. L. Hock and R. B. McQuistan. Exact, composite ternary nearestneighbor degeneracies. Journal of Mathematical Physics, 34(1):260– 269, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p260_s1.

Kolsrud:1993:MRB

[1424] Marius Kolsrud. Maximal reductions in the Baker–Hausdorff formula. Journal of Mathematical Physics, 34(1):270–285, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i1/p270_s1.

Kulish:1993:CSR

[1425] P. P. Kulish, R. Sasaki, and C. Schwiebert. Constant solutions of reflection equations and quantum groups. Journal of Mathematical Physics, 34(1):286–304, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i1/p286_s1.

Lesniewski:1993:TPR

[1426] Andrzej Lesniewski and Maurizio Rinaldi. Tensor products of representations of C(SU_{q}(2)). Journal of Mathematical Physics, 34(1):305– 314, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i1/p305_s1.

Low:1993:CSL

[1427] R. J. Low. Celestial spheres, light cones, and cuts. Journal of Mathematical Physics, 34(1):315–319, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i1/p315_s1.

Scheunert:1993:ASO

[1428] M. Scheunert. The antipode of and star operations in a Hopf algebra. Journal of Mathematical Physics, 34(1):320–325, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i1/p320_s1.

Tagliani:1993:AME

[1429] Aldo Tagliani. On the application of maximum entropy to the moments problem. Journal of Mathematical Physics, 34(1):326–337, January 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i1/p326_s1.

Baadhio:1993:VCI

[1430] Randy A. Baadhio. Vacuum configuration for inflationary superstring. Journal of Mathematical Physics, 34(2):345–357, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p345_
s1.

Baadhio:1993:GGI

[1431] Randy A. Baadhio. On global gravitational instantons in superstring theory. Journal of Mathematical Physics, 34(2):358–368, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/
p358_s1.

Batalin:1993:GSA

[1432] I. A. Batalin and I. V. Tyutin. General solution to the associativity equation in the Euclidean phase space. Journal of Mathematical Physics, 34(2):369–380, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p369_s1.

Charpentier:1993:BBM

[1433] Eric Charpentier and Krzysztof Gawedzki. Bosonization in background of metric and gauge. Journal of Mathematical Physics, 34(2):381–436, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p381_s1.

Estevez-Breton:1993:NSH

[1434] Erika Sabrina Estévez-Bretón and Gentil A. Estévez-Bretón. The nonsingular spiked harmonic oscillator. Journal of Mathematical Physics, 34(2):437–440, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p437_s1. See comment [1755].

Hadjiivanov:1993:EPF

[1435] Ludmil K. Hadjiivanov. Existence of primary fields as a generalization of the Lüscher–Mack theorem. Journal of Mathematical Physics, 34(2):441– 453, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p441_s1.

Joye:1993:SEA

[1436] Alain Joye and Charles-Edouard Pfister. Superadiabatic evolution and adiabatic transition probability between two nondegenerate levels isolated in the spectrum. Journal of Mathematical Physics, 34(2):454– 479, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p454_s1.

Lenczewski:1993:AQC

[1437] Romuald Lenczewski. A q-analog of the quantum central limit theorem for SU_{q}(2), q complex. Journal of Mathematical Physics, 34(2):480– 489, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p480_s1.

Lev:1993:FPP

[1438] Felix Lev. Finiteness of physics and its possible consequences. Journal of Mathematical Physics, 34(2):490–527, February 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i2/p490_s1.

Lynch:1993:CGS

[1439] R. Lynch and H. A. Mavromatis. Consideration of the general state which minimizes the Nth order minimum uncertainty product. Journal of Mathematical Physics, 34(2):528–534, February 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i2/p528_s1.

Moschella:1993:WOE

[1440] U. Moschella. The Wick-ordered exponential of the dipole field as a field of type S. Journal of Mathematical Physics, 34(2):535–548, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i2/p535_s1.

Pfalzner:1993:TFT

[1441] S. Pfalzner and N. H. March. Thomas–Fermi theory in magnetic fields of arbitrary strength. Journal of Mathematical Physics, 34(2):549–557, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p549_s1.

Rideau:1993:EEI

[1442] G. Rideau and P. Winternitz. Evolution equations invariant under twodimensional space–time Schrödinger group. Journal of Mathematical Physics, 34(2):558–570, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i2/p558_s1.

Weniger:1993:SOR

[1443] Ernst Joachim Weniger, Jiří Čížek, and Francine Vinette. The summation of the ordinary and renormalized perturbation series for the ground state energy of the quartic, sextic, and octic anharmonic oscillators using nonlinear sequence transformations. Journal of Mathematical Physics, 34(2):571–609, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p571_s1.

Bollini:1993:APD

[1444] C. G. Bollini and J. J. Giambiagi. Arbitrary powers of D’Alembertians and the Huygens’ principle. Journal of Mathematical Physics, 34(2):610– 621, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p610_s1.

deLeon:1993:CTD

[1445] Manuel de León and Juan C. Marrero. Constrained time-dependent Lagrangian systems and Lagrangian submanifolds. Journal of Mathematical Physics, 34(2):622–644, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i2/p622_s1.

Horwitz:1993:DGR

[1446] L. P. Horwitz. Dynamical group of the relativistic Kepler problem. Journal of Mathematical Physics, 34(2):645–648, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p645_s1.

Inamoto:1993:QIL

[1447] Tadashi Inamoto. Quasi-invariant Lagrangians on Lie groups and the method of coadjoint orbits. Journal of Mathematical Physics, 34(2):649– 673, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p649_s1.

Ivanova:1993:SSY

[1448] T. A. Ivanova and A. D. Popov. Some solutions of the Yang–Mills equations in dimension greater than four. Journal of Mathematical Physics, 34(2):674–680, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p674_s1.

Castro:1993:SGT

[1449] Carlos Castro. SU(∞) (super)gauge theories and self-dual (super)gravity. Journal of Mathematical Physics, 34(2):681–689, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p681_
s1.

Hortacsu:1993:FSG

[1450] M. Hortaçsu. Fluctuations of a spherical gravitational impulsive wave. Journal of Mathematical Physics, 34(2):690–699, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p690_

s1. See errata [2318].

Katanaev:1993:AUC

[1451] M. O. Katanaev. All universal coverings of two-dimensional gravity with torsion. Journal of Mathematical Physics, 34(2):700–736, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i2/p700_s1.

Basu:1993:GTF

[1452] B. Basu. A gauge theoretical formulation of a frustrated Ising system. Journal of Mathematical Physics, 34(2):737–748, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p737_
s1.

Goswami:1993:SSG

[1453] Gautam Goswami and Pratul Bandyopadhyay. Spin system, gauge theory, and renormalization group equation. Journal of Mathematical Physics, 34(2):749–756, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i2/p749_s1.

Li:1993:YB

[1454] You Quan Li. Yang–Baxterization. Journal of Mathematical Physics, 34(2):757–767, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p757_s1.

Li:1993:MSY

[1455] You Quan Li. Multiparameter solutions of Yang–Baxter equation from braid group representations. Journal of Mathematical Physics, 34(2):768–774, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p768_s1.

Misawa:1993:SDI

[1456] Tetsuya Misawa. Stochastic dynamical interpolation formula to timereversible diffusion processes. Journal of Mathematical Physics, 34 (2):775–784, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p775_s1.

Verbeure:1993:ALM

[1457] A. Verbeure and V. A. Zagrebnov. About the Luttinger model. Journal of Mathematical Physics, 34(2):785–800, February 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i2/p785_s1.

Gupta:1993:IML

[1458] Neelam Gupta. Integrals of motion for the Lorenz system. Journal of Mathematical Physics, 34(2):801–804, February 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i2/p801_s1. See com-

ment [1714].

Geng:1993:FDD

[1459] Xian Guo Geng. Finite-dimensional discrete systems and integrable systems through nonlinearization of the discrete eigenvalue problem. Journal of Mathematical Physics, 34(2):805–817, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p805_s1.

Akemann:1993:ZCC

[1460] G. Akemann and R. Grimm. Zero curvature conditions and conformal covariance. Journal of Mathematical Physics, 34(2):818–835, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i2/p818_s1.

Gieres:1993:MCF

[1461] François Gieres. Maurer–Cartan forms and equations for two-dimensional superdiffeomorphisms. Journal of Mathematical Physics, 34(2):836–843, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p836_s1.

Kachkachi:1993:SBP

[1462] M. Kachkachi and T. Lhallabi. (2, 0) super-Beltrami parametrization and superconformal structures. Journal of Mathematical Physics, 34(2):844– 863, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p844_s1.

Ludde:1993:YBD

[1463] Mirko Lüdde. Yang–Baxter deformations of semisimple Lie algebras. Journal of Mathematical Physics, 34(2):864–874, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p864_
s1.

Millane:1993:MTP

[1464] R. P. Millane. Möbius transform pairs. Journal of Mathematical Physics, 34(2):875–877, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i2/p875_s1.

Wang:1993:STD

[1465] Alan P. Wang and Sueo Ueno. Searchlight on a target with diffuse background. I. Journal of Mathematical Physics, 34(2):878–884, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i2/p878_s1.

Bertrand:1993:ECA

[1466] J. Bertrand and P. Bertrand. Erratum: “A class of affine Wigner functions with extended covariance properties [J. Math. Phys. 33, 2515 (1992)]. Journal of Mathematical Physics, 34(2):885, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i2/p885_

s1. See [1202].

Chandler:1993:EBQ

[1467] Colston Chandler and A. G. Gibson. Erratum: “N-body quantum scattering theory in two Hilbert spaces. V. Computation strategy”, [J. Math. Phys. 30 (1989), no. 7, 1533–1544; MR 90h:81198]. Journal of Mathematical Physics, 34(2):886, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i2/p886_s1. See [12].

Figueroa-OFarrill:1993:EEU

[1468] José M. Figueroa-O’Farrill and Eduardo Ramos. Erratum: “Existence and uniqueness of the universal W algebra”, [J. Math. Phys. 33 (1992), no. 3, 833–839; MR 92m:17042]. Journal of Mathematical Physics, 34(2): 887, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i2/p887_s1. See [1017].

Lieb:1993:UDT

[1469] Elliott H. Lieb, Michael Loss, and Robert J. McCann. Uniform density theorem for the Hubbard model. Journal of Mathematical Physics, 34(3): 891–898, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i3/p891_s1.

Acerbi:1993:ISF

[1470] F. Acerbi, G. Morchio, and F. Strocchi. Infrared singular fields and nonregular representations of canonical commutation relation algebras. Journal of Mathematical Physics, 34(3):899–914, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i3/p899_s1.

Arai:1993:PDW

[1471] Asao Arai. Properties of the Dirac–Weyl operator with a strongly singular gauge potential. Journal of Mathematical Physics, 34(3):915– 935, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i3/p915_s1.

Galloonov:1993:ENS

[1472] G. V. Galloonov, V. L. Oleĭnik, and B. S. Pavlov. Estimations for negative spectral bands of three-dimensional periodical Schrödinger operator. Journal of Mathematical Physics, 34(3):936–942, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i3/p936_s1.

Gerisch:1993:LDG

[1473] Thomas Gerisch, Reinhard Honegger, and Alfred Rieckers. Limiting dynamics of generalized Bardeen–Cooper–Schrieffer models beyond the pair algebra. Journal of Mathematical Physics, 34(3):943–968, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i3/p943_s1.

Kempf:1993:QGS

[1474] Achim Kempf. Quantum group symmetric Bargmann–Fock space: Integral kernels, Green functions, driving forces. Journal of Mathematical Physics, 34(3):969–987, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i3/p969_s1.

Yu:1993:FCQ

[1475] Yue Yu, Zhong Yuan Zhu, and H. C. Lee. Flat connections on quantum bundles and fractional statistics in geometric quantization. Journal of Mathematical Physics, 34(3):988–996, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i3/p988_s1.

Dewisme:1993:FIS

[1476] Alain Dewisme and Serge Bouquet. First integrals and symmetries of time-dependent Hamiltonian systems. Journal of Mathematical Physics, 34(3):997–1006, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i3/p997_s1.

Hadley:1993:NGG

[1477] R. H. Hadley and R. L. Mallett. Null geodesics in the gravitational field of a rotating, radiating body. Journal of Mathematical Physics, 34(3):1007– 1015, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i3/p1007_s1.

Kuang:1993:AST

[1478] Zhi Quan Kuang and Can Bin Liang. All space–times admitting strongly synchronizable reference frames are static. Journal of Mathematical Physics, 34(3):1016–1021, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i3/p1016_s1.

Tomaschitz:1993:CQD

[1479] Roman Tomaschitz. Classical and quantum dispersion in Robertson– Walker cosmologies. Journal of Mathematical Physics, 34(3):1022– 1042, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i3/p1022_s1.

Narumi:1993:GEP

[1480] Hideyuki Narumi and Haruo Hosoya. Generalized expression of the perfect matching number for 2 × 3 × n lattices. Journal of Mathematical Physics, 34(3):1043–1051, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i3/p1043_s1.

Kundu:1993:CQI

[1481] A. Kundu and B. Basu Mallick. Classical and quantum integrability of a derivative nonlinear Schrödinger model related to quantum group. Journal of Mathematical Physics, 34(3):1052–1062, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i3/p1052_s1.

Pang:1993:CLQ

[1482] Gen Di Pang and Bao Heng Zhao. Conservation laws of the quantized Davey–Stewartson I system in (2 + 1) dimensions. Journal of Mathematical Physics, 34(3):1063–1075, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i3/p1063_s1. See addendum [1789].

Baker:1993:TPD

[1483] T. H. Baker. Tensor product decompositions and branching rules of the N = 2 superconformal algebra. Journal of Mathematical Physics, 34(3):1076–1094, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i3/p1076_s1.

Battle:1993:KGP

[1484] Guy Battle. Klein–Gordon propagation of Daubechies wavelets. Journal of Mathematical Physics, 34(3):1095–1109, March 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i3/p1095_s1.

Cougo-Pinto:1993:AIP

[1485] M. V. Cougo-Pinto. The algebras of inhibited parafermionic violation of the Pauli principle. Journal of Mathematical Physics, 34(3):1110– 1124, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i3/p1110_s1.

Cougo-Pinto:1993:RKM

[1486] M. V. Cougo-Pinto. Representations of Kac–Moody and Virasoro algebras on modules of Greenberg–Mohapatra algebras. Journal of Mathematical Physics, 34(3):1125–1142, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i3/p1125_s1.

Dragovic:1993:PSE

[1487] Branko G. Dragović. Power series everywhere convergent on R and all Q_{p}. Journal of Mathematical Physics, 34(3):1143–1148, March 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i3/p1143_
s1.

Duan:1993:TSG

[1488] Yi Shi Duan and Xin He Meng. Topological structure of Gauss–Bonnet– Chern density and its topological current. Journal of Mathematical Physics, 34(3):1149–1161, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i3/p1149_s1.

Dundarer:1993:VYM

[1489] A. Reşit Dündarer. On SO(8)_{±} valued Yang–Mills fields and multiinstanton solutions in eight-dimensional curved space. Journal of Mathematical Physics, 34(3):1162–1168, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i3/p1162_s1.

Goyal:1993:AEC

[1490] I. C. Goyal, R. L. Gallawa, and A. K. Ghatak. Accuracy of eigenvalues: A comparison of two methods. Journal of Mathematical Physics, 34(3):1169–1175, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i3/p1169_s1.

Majid:1993:QBL

[1491] S. Majid. Quantum and braided linear algebra. Journal of Mathematical Physics, 34(3):1176–1196, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i3/p1176_s1.

Micu:1993:CHP

[1492] M. Micu. Continuous Hahn polynomials. Journal of Mathematical Physics, 34(3):1197–1205, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i3/p1197_s1.

Negro:1993:LEU

[1493] Javier Negro and Mariano A. del Olmo. On the local equivalence in a unidimensional world. Journal of Mathematical Physics, 34(3):1206– 1217, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i3/p1206_s1.

Sun:1993:NQH

[1494] Chang Pu Sun, Xu Feng Liu, and Mo Lin Ge. A new quasitriangular Hopf algebra as the nontrivial quantum double of a simplest C-algebra and its universal R matrix for Yang–Baxter equation. Journal of Mathematical Physics, 34(3):1218–1222, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i3/p1218_s1.

Vourdas:1993:CSS

[1495] A. Vourdas. Coherent states on the m-sheeted covering group of SU(1, 1). Journal of Mathematical Physics, 34(3):1223–1235, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i3/p1223_s1.

Zhang:1993:FDI

[1496] R. B. Zhang. Finite dimensional irreducible representations of the quantum supergroup U_{q}(gl(m/n)). Journal of Mathematical Physics, 34(3):1236–1254, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i3/p1236_s1.

Chetouani:1993:PIT

[1497] L. Chetouani, L. Guechi, A. Lecheheb, and T. F. Hammann. Path integral treatment for the one-dimensional Natanzon potentials. Journal of Mathematical Physics, 34(4):1257–1269, April 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1257_s1.

Debergh:1993:PDS

[1498] N. Debergh. Positive discrete series of osp(2|2, R) and (para)supersymmetric quantum mechanics. Journal of Mathematical Physics, 34(4):1270–1276, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i4/p1270_s1.

Khare:1993:PQM

[1499] Avinash Khare. Parasupersymmetry in quantum mechanics. Journal of Mathematical Physics, 34(4):1277–1294, April 1993. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i4/p1277_s1.

Rupertsberger:1993:BSS

[1500] H. Rupertsberger. Bound states for square well potentials extending to infinity in D ≥ 2. Journal of Mathematical Physics, 34(4):1295– 1299, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i4/p1295_s1.

Zaikov:1993:CEO

[1501] R. P. Zaĭkov and S. I. Zlatev. Casimir energy for open membranes. Journal of Mathematical Physics, 34(4):1300–1308, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1300_s1.

Brouzet:1993:AER

[1502] R. Brouzet. About the existence of recursion operators for completely integrable Hamiltonian systems near a Liouville torus. Journal of Mathematical Physics, 34(4):1309–1313, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i4/p1309_s1.

Doyle:1993:DGP

[1503] Philip W. Doyle. Differential geometric Poisson bivectors in one space variable. Journal of Mathematical Physics, 34(4):1314–1338, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i4/
p1314_s1.

Kristensson:1993:TWP

[1504] Gerhard Kristensson and Sten Rikte. Transient wave propagation in reciprocal bi-isotropic media at oblique incidence. Journal of Mathematical Physics, 34(4):1339–1359, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i4/p1339_s1.

Manganaro:1993:GAL

[1505] N. Manganaro and G. Valenti. Group analysis and linearization procedure for a nonautonomous model describing rate-type materials. Journal of Mathematical Physics, 34(4):1360–1369, April 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1360_s1.

Weston:1993:TDW

[1506] V. H. Weston. Time-domain wave splitting of Maxwell’s equations. Journal of Mathematical Physics, 34(4):1370–1392, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1370_s1.

Yuan:1993:GSN

[1507] Miao En Yuan and Achilaltu Wang. General solutions of nonlinear Ianiro–Lebowitz model equation. Journal of Mathematical Physics, 34(4):1393–1420, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i4/p1393_s1.

Leble:1993:KVM

[1508] S. B. Leble and N. V. Ustinov. Korteweg–de Vries–modified Korteweg– de Vries systems and Darboux transforms in 1 + 1 and 2 + 1 dimensions. Journal of Mathematical Physics, 34(4):1421–1428, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1421_s1.

Sidorenko:1993:MIR

[1509] Jurij Sidorenko and Walter Strampp. Multicomponent integrable reductions in the Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 34(4):1429–1446, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i4/p1429_s1.

Dabrowski:1993:IES

[1510] Mariusz P. Dabrowski. Isometric embedding of the spherically symmetric Stephani universe: Some explicit examples. Journal of Mathematical Physics, 34(4):1447–1479, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i4/p1447_s1.

Shiraishi:1993:MSM

[1511] Kiyoshi Shiraishi. Multicentered solution for maximally charged dilaton black holes in arbitrary dimensions. Journal of Mathematical Physics, 34(4):1480–1486, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i4/p1480_s1.

Zardecki:1993:FSB

[1512] Andrew Zardecki. A formulation of supergravity based on Cartan’s connection. Journal of Mathematical Physics, 34(4):1487–1496, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i4/p1487_s1.

Allen:1993:AEC

[1513] Scott Allen and R. K. Pathria. Analytical evaluation of a class of phasemodulated lattice sums. Journal of Mathematical Physics, 34(4):1497– 1507, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i4/p1497_s1.

Belavkin:1993:UIF

[1514] V. P. Belavkin. The unified Itô formula has the pseudo-Poisson structure df(x) = [f(x + b) − f(x)]_{ν}^{μ}da_{μ}^{ν}. Journal of Mathematical Physics, 34(4):1508–1518, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i4/p1508_s1.

Daumens:1993:SDO

[1515] Michel Daumens and Yves Noirot. A simple differential operator realization of the super-rotation algebra. Journal of Mathematical Physics, 34(4):1519–1522, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i4/p1519_s1.

Flath:1993:TOI

[1516] Dan Flath and Jacob Towber. Tensor operators. III. Some fundamental tensor operator identities. Journal of Mathematical Physics, 34(4):1523– 1547, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i4/p1523_s1.

Kopsky:1993:TNEa

[1517] Vojtěch Kopský. Translation normalizers of Euclidean groups. I. Elementary theory. Journal of Mathematical Physics, 34(4):1548–1556, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i4/p1548_s1.

Kopsky:1993:TNEb

[1518] Vojtěch Kopský. Translation normalizers of Euclidean groups. II. Systematic calculation. Journal of Mathematical Physics, 34(4):1557–1576, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i4/p1557_s1.

Links:1993:RLO

[1519] J. Links and M. D. Gould. Raising and lowering operators for U_{q}(gl(n)). Journal of Mathematical Physics, 34(4):1577–1586, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1577_s1.

Miyazawa:1993:FEM

[1520] Toru Miyazawa. Frequency expansion method for the one-dimensional Fokker–Planck equation. Journal of Mathematical Physics, 34(4):1587– 1605, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i4/p1587_s1.

Stovicek:1993:QLB

[1521] Pavel Štovíček. Quantum line bundles on S^{2} and method of orbits for SU_{q}(2). Journal of Mathematical Physics, 34(4):1606–1613, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i4/p1606_s1.

Bystsenko:1993:EVE

[1522] Andrei A. Bystsenko, Guido Cognola, and Luciano Vanzo. Erratum: “Vacuum energy for 3 + 1 dimensional space–time with compact hyperbolic spatial part”, [J. Math. Phys. 33, 3108 (1992)]. Journal of Mathematical Physics, 34(4):1614, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i4/p1614_s1. See [1268].

Juriev:1993:ENG

[1523] Denis Juriev. Erratum: “Noncommutative geometry, chiral anomaly in the quantum projective [sl(2, C)-invariant] field theory and jl(2, C)- invariance”, [J. Math. Phys. 33 (1992), no. 8, 2819–2822; MR 93e:81042]. Journal of Mathematical Physics, 34(4):1615, April 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i4/p1615_

s1. See [1237].

Aktosun:1993:SIS

[1524] Tuncay Aktosun. Scattering and inverse scattering for a second-order differential equation. Journal of Mathematical Physics, 34(5):1619–1634, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1619_s1.

Angelidis:1993:PLE

[1525] Thomas D. Angelidis. On the problem of a local extension of the quantum formalism. Journal of Mathematical Physics, 34(5):1635–1653, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1635_s1.

Bracken:1993:BFM

[1526] A. J. Bracken, M. D. Gould, and I. Tsohantjis. Boson-fermion models for osp(1, 2) and U_{q}[osp(1, 2)]. Journal of Mathematical Physics, 34(5): 1654–1664, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i5/p1654_s1.

Burzlaff:1993:ASI

[1527] J. Burzlaff, A. Chakrabarti, and D. H. Tchrakian. Axially symmetric instantons in generalized Yang–Mills theory in 4p dimensions. Journal of Mathematical Physics, 34(5):1665–1680, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i5/p1665_s1.

Duclos:1993:SLQ

[1528] P. Duclos and H. Hogreve. On the semiclassical localization of the quantum probability. Journal of Mathematical Physics, 34(5):1681–1691, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1681_s1.

Fittler:1993:ANQ

[1529] Robert Fittler. Asymptotic nonstandard quantum electrodynamics. Journal of Mathematical Physics, 34(5):1692–1724, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/p1692_s1.

Hietarinta:1993:STD

[1530] Jarmo Hietarinta. Solving the two-dimensional constant quantum Yang– Baxter equation. Journal of Mathematical Physics, 34(5):1725–1756, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1725_s1.

Nagamachi:1993:LCS

[1531] S. Nagamachi and T. Nishimura. Linear canonical supertransformation. Journal of Mathematical Physics, 34(5):1757–1772, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/p1757_s1.

Ogurisu:1993:ESI

[1532] Osamu Ogurisu. Existence and structure of infinitely degenerate zeroenergy ground states of a Wess–Zumino-type model in supersymmetric quantum mechanics. Journal of Mathematical Physics, 34(5):1773–1790, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1773_s1.

Semay:1993:VTT

[1533] Claude Semay. Virial theorem for two-body Dirac equation. Journal of Mathematical Physics, 34(5):1791–1793, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i5/p1791_s1.

Slater:1993:FIP

[1534] Paul B. Slater. Fisher information, prior probabilities, and the state determination of spin-1/2 and spin-1 systems. Journal of Mathematical Physics, 34(5):1794–1798, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i5/p1794_s1.

VanderJeugt:1993:DAD

[1535] J. Van der Jeugt. Dynamical algebra of the q-deformed three-dimensional oscillator. Journal of Mathematical Physics, 34(5):1799–1806, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/
p1799_s1.

Wotzasek:1993:HGS

[1536] Clovis Wotzasek and Clifford Neves. On the hidden U(1) gauge symmetry in the nonlinear sigma model. Journal of Mathematical Physics, 34(5): 1807–1817, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i5/p1807_s1.

Chmielowski:1993:LCC

[1537] Piotr Chmielowski. Local classification of constraints. Journal of Mathematical Physics, 34(5):1818–1823, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i5/p1818_s1.

Corones:1993:SRM

[1538] James Corones and Zhiming Sun. Simultaneous reconstruction of material and transient source parameters using the invariant imbedding method. Journal of Mathematical Physics, 34(5):1824–1845, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/
p1824_s1.

Levy:1993:DKR

[1539] Bernard C. Levy and Arthur J. Krener. Dynamics and kinematics of reciprocal diffusions. Journal of Mathematical Physics, 34(5):1846–1875, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1846_s1.

Annamalai:1993:ITL

[1540] A. Annamalai and K. M. Tamizhmani. Integrability of Toda lattice by generalized variational symmetry approach. Journal of Mathematical Physics, 34(5):1876–1883, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i5/p1876_s1.

Sen:1993:AKS

[1541] Dolan Chapa Sen and A. Roy Chowdhury. Adler–Kostant–Syms [Symes] theorem and supersymmetric integrable systems. Journal of Mathematical Physics, 34(5):1884–1891, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i5/p1884_s1.

Tafel:1993:TDR

[1542] Jacek Tafel. Two-dimensional reductions of the self-dual Yang–Mills equations in self-dual spaces. Journal of Mathematical Physics, 34(5): 1892–1907, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i5/p1892_s1.

Davidson:1993:CSS

[1543] W. Davidson. A cylindrically symmetric solution of Einstein’s equations describing gravitational collapse of stiff fluid. Journal of Mathematical Physics, 34(5):1908–1913, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i5/p1908_s1.

Klimcik:1993:SCS

[1544] Ctirad Klimčík. Search for the conformal scalar hair at arbitrary D. Journal of Mathematical Physics, 34(5):1914–1926, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/p1914_s1.

Patel:1993:CSV

[1545] L. K. Patel and Naresh Dadhich. Cylindrically symmetric viscous universes. Journal of Mathematical Physics, 34(5):1927–1936, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/
p1927_s1.

Aizawa:1993:TOC

[1546] N. Aizawa. Tensor operators and Clebsch–Gordan coefficients of the quantum algebra su_{q}(1, 1). Journal of Mathematical Physics, 34(5):1937– 1963, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i5/p1937_s1.

Chakrabarti:1993:CSR

[1547] A. Chakrabarti. Canonical structure in SO(4)_{q} and relations to E(3)_{q} and SO(3, 1)_{q}. Journal of Mathematical Physics, 34(5):1964–1985, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p1964_s1.

Cianci:1993:ASIa

[1548] Roberto Cianci and Andrew Khrennikov. p-adic superanalysis. I. Infinitely generated Banach superalgebras and algebraic groups. Journal of Mathematical Physics, 34(5):1986–1994, May 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i5/p1986_s1.

Cianci:1993:ASIb

[1549] Roberto Cianci and Andrew Khrennikov. p-adic superanalysis. II. Supermanifolds and differential operators. Journal of Mathematical Physics, 34(5):1995–2003, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i5/p1995_s1.

Foulis:1993:OME

[1550] D. L. Foulis. Optimized multicenter expansions. Journal of Mathematical Physics, 34(5):2004–2027, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i5/p2004_s1.

Garcia:1993:GSR

[1551] Alberto García. Generalization of the Szekeres results on perfect fluids with u ∧ k ∧ l = 0 to u ∧ k ∧ l ≠ 0: Bianchi identities. Journal of Mathematical Physics, 34(5):2028–2044, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i5/p2028_s1.

Majid:1993:BMP

[1552] Shahn Majid. Braided momentum in the q-Poincaré group. Journal of Mathematical Physics, 34(5):2045–2058, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i5/p2045_s1.

Majumdar:1993:IWC

[1553] Parthasarathi Majumdar. Inönü-Wigner contraction of Kac–Moody algebras. Journal of Mathematical Physics, 34(5):2059–2065, May 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i5/p2059_
s1.

Mignaco:1993:LAD

[1554] Juan Alberto Mignaco and Cesar Augusto Linhares. Lie algebras for the Dirac–Clifford ring. Journal of Mathematical Physics, 34(5):2066– 2074, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i5/p2066_s1.

Nagao:1993:NCR

[1555] Taro Nagao and Keith Slevin. Nonuniversal correlations for random matrix ensembles. Journal of Mathematical Physics, 34(5):2075–2085, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i5/p2075_s1.

Barut:1993:NAD

[1556] A. O. Barut, A. J. Bracken, S. Komy, and N. Ünal. New approach to the determination of eigenvalues and eigenfunctions for a relativistic two-fermion equation. Journal of Mathematical Physics, 34(6):2089– 2106, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2089_s1.

Braibant:1993:QES

[1557] S. Braibant and Y. Brihaye. Quasi-exactly-solvable system and sphaleron stability. Journal of Mathematical Physics, 34(6):2107–2114, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/
p2107_s1.

Fitzsimmons:1993:OMM

[1558] Sean Fitzsimmons. Obtaining magnetic monopole solutions by use of the Laplace transform. Journal of Mathematical Physics, 34(6):2115– 2118, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2115_s1.

Gomez:1993:QHO

[1559] César Gómez and Germán Sierra. Quantum harmonic oscillator algebra and link invariants. Journal of Mathematical Physics, 34(6):2119– 2131, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2119_s1.

Gomis:1993:PTL

[1560] Joaquim Gomis and Jordi París. Perturbation theory and locality in the field-antifield formalism. Journal of Mathematical Physics, 34(6):2132– 2152, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2132_s1.

Kowalski:1993:CSQ

[1561] K. Kowalski and J. Rembieliński. Coherent states for the quantum complex plane. Journal of Mathematical Physics, 34(6):2153–2165, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i6/p2153_s1.

Naito:1993:VOG

[1562] Seichi Naito. Vertex operators and generating functional for one loop scattering amplitudes. Journal of Mathematical Physics, 34(6):2166– 2184, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2166_s1.

Robinson:1993:SMT

[1563] Sam L. Robinson. Semiclassical mechanics for time-dependent Wigner functions. Journal of Mathematical Physics, 34(6):2185–2205, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/
p2185_s1.

Romeo:1993:FMG

[1564] August Romeo. Fermion mass generation at finite temperature. Journal of Mathematical Physics, 34(6):2206–2214, June 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i6/p2206_s1.

Semay:1993:TBD

[1565] C. Semay, R. Ceuleneer, and B. Silvestre-Brac. Two-body Dirac equation with diagonal central potentials. Journal of Mathematical Physics, 34(6): 2215–2225, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2215_s1.

Andronov:1993:AZR

[1566] Ivan V. Andronov. Application of zero-range models to the problem of diffraction by a small hole in elastic plate. Journal of Mathematical Physics, 34(6):2226–2241, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2226_s1.

Chang:1993:EBS

[1567] Shau-Jin Chang, Bruno Crespi, and Kang Jie Shi. Elliptical billiard systems and the full Poncelet’s theorem in n dimensions. Journal of Mathematical Physics, 34(6):2242–2256, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i6/p2242_s1.

Crespi:1993:EBH

[1568] Bruno Crespi, Shau-Jin Chang, and Kang Jie Shi. Elliptical billiards and hyperelliptic functions. Journal of Mathematical Physics, 34(6):2257– 2289, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2257_s1.

Grmela:1993:EKT

[1569] Miroslav Grmela and David Jou. Extended kinetic theory. Journal of Mathematical Physics, 34(6):2290–2316, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i6/p2290_s1.

Nagao:1993:LER

[1570] Taro Nagao and Keith Slevin. Laguerre ensembles of random matrices: Nonuniversal correlation functions. Journal of Mathematical Physics, 34(6):2317–2330, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2317_s1.

Samuel:1993:SIM

[1571] Stuart Samuel. The solution of the Ising model on the truncated icosahedron lattice of C_{60}. Journal of Mathematical Physics, 34(6):2331– 2341, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2331_s1.

Wu:1993:QPP

[1572] Yihren Wu. On the quantization of the planar N-point vortex problem. Journal of Mathematical Physics, 34(6):2342–2352, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/p2342_s1.

Berceanu:1993:LDS

[1573] S. Berceanu and L. Boutet de Monvel. Linear dynamical systems, coherent state manifolds, flows, and matrix Riccati equation. Journal of Mathematical Physics, 34(6):2353–2371, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i6/p2353_s1.

Cieslinski:1993:GIS

[1574] Jan Cieśliński. Group interpretation of the spectral parameter in the case of nonhomogeneous, nonlinear Schrödinger system. Journal of Mathematical Physics, 34(6):2372–2384, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i6/p2372_s1.

Ravoson:1993:SLP

[1575] V. Ravoson, L. Gavrilov, and R. Caboz. Separability and Lax pairs for Hénon–Heiles system. Journal of Mathematical Physics, 34(6):2385– 2393, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2385_s1.

Roelofs:1993:PSL

[1576] G. H. M. Roelofs and R. Martini. Prolongation structure of the Landau– Lifshitz equation. Journal of Mathematical Physics, 34(6):2394–2399, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i6/p2394_s1.

Tajiri:1993:LSS

[1577] Masayoshi Tajiri and Youichi Murakami. Lattice soliton: Solutions to the Kadomtsev–Petviashvili equation with positive dispersion. Journal of Mathematical Physics, 34(6):2400–2411, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i6/p2400_s1.

Brodbeck:1993:GBT

[1578] Othmar Brodbeck and Norbert Straumann. A generalized Birkhoff theorem for the Einstein–Yang–Mills system. Journal of Mathematical Physics, 34(6):2412–2423, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2412_s1.

Kachkachi:1993:SKE

[1579] M. Kachkachi and T. Lhallabi. N = 2 superconformal Killing equations and the harmonic differentials. Journal of Mathematical Physics, 34(6): 2424–2434, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i6/p2424_s1.

Kim:1993:TVS

[1580] Jong Chul Kim and Jin Hwan Kim. Totally vicious space–times. Journal of Mathematical Physics, 34(6):2435–2439, June 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i6/p2435_s1.

Pant:1993:NCE

[1581] D. N. Pant and Neeraj Pant. A new class of exact solutions in general relativity representing perfect fluid balls. Journal of Mathematical Physics, 34(6):2440–2447, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2440_s1.

Racz:1993:STE

[1582] István Rácz. Space–time extensions. I. Journal of Mathematical Physics, 34(6):2448–2464, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i6/p2448_s1.

Tiwari:1993:NUT

[1583] S. C. Tiwari. A note on the unimodular theory of gravitation. Journal of Mathematical Physics, 34(6):2465–2467, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i6/p2465_s1.

Coll:1993:CST

[1584] Bartolomé Coll and Juan Antonio Morales. Comments on space–time signature. Journal of Mathematical Physics, 34(6):2468–2474, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/
p2468_s1.

Daumens:1993:SRR

[1585] M. Daumens, P. Minnaert, M. Mozrzymas, and S. Toshev. The superrotation Racah–Wigner calculus revisited. Journal of Mathematical Physics, 34(6):2475–2507, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2475_s1.

Daumens:1993:SAM

[1586] Michel Daumens. Super-angular momentum and super-spherical harmonics. Journal of Mathematical Physics, 34(6):2508–2522, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/
p2508_s1.

Guhr:1993:FBA

[1587] Thomas Guhr. Fourier–Bessel analysis for ordinary and graded 2 × 2 Hermitian matrices. Journal of Mathematical Physics, 34(6):2523–2540, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i6/p2523_s1.

Guhr:1993:GG

[1588] Thomas Guhr. On the graded group U(1/1). Journal of Mathematical Physics, 34(6):2541–2553, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2541_s1.

Hakobyan:1993:SNS

[1589] T. S. Hakobyan and A. G. Sedrakyan. Some new spinor representations of quantum groups B_{q}(n), C_{q}(n), G_{q}(2). Journal of Mathematical Physics, 34(6):2554–2560, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2554_s1.

Karimipour:1993:MQS

[1590] V. Karimipour and A. Aghamohammadi. Multiparametric quantization of the special linear superalgebra. Journal of Mathematical Physics, 34(6):2561–2571, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i6/p2561_s1.

Sasiela:1993:MTM

[1591] Richard J. Sasiela and John D. Shelton. Mellin transform methods applied to integral evaluation: Taylor series and asymptotic approximations. Journal of Mathematical Physics, 34(6):2572–2617, June 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i6/p2572_
s1.

Yan:1993:RLW

[1592] Chun Tao Yan. Regularized long wave equation and inverse scattering transform. Journal of Mathematical Physics, 34(6):2618–2630, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i6/p2618_s1.

Zhedanov:1993:ROT

[1593] A. S. Zhedanov. Q rotations and other Q transformations as unitary nonlinear automorphisms of quantum algebras. Journal of Mathematical Physics, 34(6):2631–2647, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i6/p2631_s1.

Aktosun:1993:RHP

[1594] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. On the Riemann–Hilbert problem for the one-dimensional Schrödinger equation. Journal of Mathematical Physics, 34(7):2651–2690, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p2651_s1.

Borthwick:1993:EMF

[1595] David Borthwick. Euclidean Majorana fermions, fermionic integrals, and relative Pfaffians. Journal of Mathematical Physics, 34(7):2691–2712, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i7/p2691_s1.

Cedillo:1993:PAT

[1596] Andrés Cedillo. A perturbative approach to the Thomas–Fermi equation in terms of the density. Journal of Mathematical Physics, 34(7):2713– 2717, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2713_s1.

Chhajlany:1993:GSD

[1597] S. C. Chhajlany. Generalized Schwinger duality in bound state theory. Journal of Mathematical Physics, 34(7):2718–2722, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p2718_s1.

Corns:1993:PIS

[1598] R. Corns. Path integrals in N = 2 supersymmetric quantum mechanics. Journal of Mathematical Physics, 34(7):2723–2741, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p2723_s1.

Dodonov:1993:QPR

[1599] V. V. Dodonov, A. B. Klimov, and D. E. Nikonov. Quantum phenomena in resonators with moving walls. Journal of Mathematical Physics, 34(7): 2742–2756, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2742_s1.

Ferraro:1993:QRG

[1600] Rafael Ferraro, Marc Henneaux, and Marcel Puchin. On the quantization of reducible gauge systems. Journal of Mathematical Physics, 34(7): 2757–2778, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2757_s1.

Hall:1993:ETS

[1601] Richard L. Hall. Envelope theory in spectral geometry. Journal of Mathematical Physics, 34(7):2779–2788, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p2779_s1.

Hirokane:1993:WCS

[1602] Tohru Hirokane, Mitsuharu Miyajima, and Yukinori Yasui. Witten complex and supersymmetric quantum mechanics on SO(3). Journal of Mathematical Physics, 34(7):2789–2806, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p2789_s1.

Jaksic:1993:LZF

[1603] Vojkan Jakšić and Jan Segert. On the Landau–Zener formula for twolevel systems. Journal of Mathematical Physics, 34(7):2807–2820, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i7/p2807_s1.

Ktitarev:1993:FPI

[1604] D. V. Ktitarev and R. S. Yegikian. Feynman path integral for Dirac system with analytic potential. Journal of Mathematical Physics, 34(7):

2821–2826, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2821_s1.

Ohnuki:1993:FAQ

[1605] Y. Ohnuki and S. Kitakado. Fundamental algebra for quantum mechanics on S^{D} and gauge potentials. Journal of Mathematical Physics, 34(7): 2827–2851, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2827_s1.

Polyzou:1993:TLF

[1606] Wayne N. Polyzou. A theorem on light-front quantum models. Journal of Mathematical Physics, 34(7):2852–2865, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i7/p2852_s1.

Saar:1993:SLR

[1607] R. Saar, M. Kõiv, I. Ots, and R. K. Loide. A second look at the Rarita– Schwinger theory. Journal of Mathematical Physics, 34(7):2866–2874, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i7/p2866_s1.

Saxena:1993:EBP

[1608] G. M. Saxena, C. L. Mehta, and B. S. Mathur. Eigenstates of bilinears in paraboson operators and their inverses. Journal of Mathematical Physics, 34(7):2875–2892, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p2875_s1.

Tome:1993:UPS

[1609] Wolfgang A. Tomé and John R. Klauder. The universal propagator for spin [or SU(2)] coherent states. Journal of Mathematical Physics, 34(7): 2893–2913, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2893_s1.

Wehrhahn:1993:QMI

[1610] R. F. Wehrhahn and Yu. B. Mel′nikov. Quantum mechanics inversion for symmetry scattering. Journal of Mathematical Physics, 34(7):2914– 2925, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2914_s1.

Zarzo:1993:QRH

[1611] A. Zarzo and A. Martínez. The quantum relativistic harmonic oscillator: Spectrum of zeros of its wave functions. Journal of Mathematical Physics, 34(7):2926–2935, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p2926_s1.

Benn:1993:SIE

[1612] I. M. Benn and W. P. Wood. Surface integrals for electromagnetic energy. Journal of Mathematical Physics, 34(7):2936–2949, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p2936_s1.

Degasperis:1993:SRF

[1613] A. Degasperis and F. Tinebra. 1 + 1 solvable relativistic field models. Journal of Mathematical Physics, 34(7):2950–2964, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p2950_s1.

Dworkin:1993:STX

[1614] Steven Dworkin. Spectral theory and X-ray diffraction. Journal of Mathematical Physics, 34(7):2965–2967, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p2965_s1.

Hojman:1993:SON

[1615] Sergio A. Hojman. Small oscillations: A new solution based on non-Noetherian symmetries. Journal of Mathematical Physics, 34(7):2968– 2974, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p2968_s1.

Overfelt:1993:RQF

[1616] P. L. Overfelt. Rings, quadratic forms, and complete degeneracy for a subclass of highly overmoded rectangular waveguides. Journal of Mathematical Physics, 34(7):2975–2989, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p2975_s1.

Skorupski:1993:WPC

[1617] Andrzej A. Skorupski. Wave propagation in complex systems of cutoffs and resonances. Journal of Mathematical Physics, 34(7):2990–3011, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i7/p2990_s1.

Chen:1993:IDF

[1618] M. Chen and Byung Chan Eu. On the integrability of differential forms related to nonequilibrium entropy and irreversible thermodynamics. Journal of Mathematical Physics, 34(7):3012–3029, July 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p3012_
s1.

Maes:1993:IPS

[1619] Christian Maes and Koen Vande Velde. The interaction potential of the stationary measure of a high-noise spinflip process. Journal of Mathematical Physics, 34(7):3030–3038, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p3030_s1.

Barashenkov:1993:UAIa

[1620] I. V. Barashenkov, B. S. Getmanov, and V. E. Kovtun. The unified approach to integrable relativistic equations: Soliton solutions over nonvanishing backgrounds. I. Journal of Mathematical Physics, 34(7):3039– 3053, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p3039_s1.

Barashenkov:1993:UAIb

[1621] I. V. Barashenkov and B. S. Getmanov. The unified approach to integrable relativistic equations: Soliton solutions over nonvanishing backgrounds. II. Journal of Mathematical Physics, 34(7):3054–3072, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i7/p3054_s1.

Bauhardt:1993:ZSI

[1622] Wolfgang Bauhardt and Christoph Pöppe. The Zakharov–Shabat inverse spectral problem for operators. Journal of Mathematical Physics, 34(7):

3073–3086, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p3073_s1.

Bellucci:1993:SKV

[1623] Stefano Bellucci, Evgenyi Ivanov, and Sergey Krivonos. On N = 3 super Korteweg–de Vries equation. Journal of Mathematical Physics, 34(7): 3087–3097, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p3087_s1.

Beutler:1993:PSS

[1624] R. Beutler. Positon solutions of the sine–Gordon equation. Journal of Mathematical Physics, 34(7):3098–3109, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i7/p3098_s1.

Qiao:1993:NCI

[1625] Zhi Jun Qiao. A new completely integrable Liouville’s system produced by the Kaup–Newell eigenvalue problem. Journal of Mathematical Physics, 34(7):3110–3120, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p3110_s1.

Figueiredo:1993:SSW

[1626] B. D. B. Figueiredo and M. Novello. Solutions to a spheroidal wave equation. Journal of Mathematical Physics, 34(7):3121–3132, July 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p3121_
s1.

Tomaschitz:1993:ERM

[1627] Roman Tomaschitz. Electromagnetic radiation in multiply connected Robertson–Walker cosmologies. Journal of Mathematical Physics, 34(7): 3133–3150, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p3133_s1.

Auberson:1993:MCV

[1628] G. Auberson. Monogenic continuation for vector fields. Journal of Mathematical Physics, 34(7):3151–3161, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i7/p3151_s1.

Borzdov:1993:ITT

[1629] George N. Borzdov. An intrinsic tensor technique in Minkowski space with applications to boundary value problems. Journal of Mathematical Physics, 34(7):3162–3196, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p3162_s1.

Calogero:1993:UIN

[1630] F. Calogero. Universal C-integrable nonlinear partial differential equation in N + 1 dimensions. Journal of Mathematical Physics, 34(7):3197– 3209, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p3197_s1.

Collins:1993:SFD

[1631] R. Eugene Collins. On some features of differentiable probabilities: a new viewpoint in physics. Journal of Mathematical Physics, 34(7):3210– 3221, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i7/p3210_s1.

Elizalde:1993:SRH

[1632] E. Elizalde. A simple recurrence for the higher derivatives of the Hurwitz zeta function. Journal of Mathematical Physics, 34(7):3222–3226, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i7/p3222_s1.

Holschneider:1993:LPW

[1633] Matthias Holschneider. Localization properties of wavelet transforms. Journal of Mathematical Physics, 34(7):3227–3244, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p3227_s1.

Kovalyov:1993:RIS

[1634] M. Kovalyov, M. Légaré, and L. Gagnon. Reductions by isometries of the self-dual Yang–Mills equations in four-dimensional Euclidean space. Journal of Mathematical Physics, 34(7):3245–3268, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i7/p3245_s1.

Nash:1993:CES

[1635] Charles Nash, Denjoe O’Connor, and Siddhartha Sen. Central extensions of sphere groups and their Lie algebras. Journal of Mathematical Physics, 34(7):3269–3272, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p3269_s1.

Okubo:1993:TPYa

[1636] Susumu Okubo. Triple products and Yang–Baxter equation. I. Octonionic and quaternionic triple systems. Journal of Mathematical Physics, 34(7):3273–3291, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p3273_s1.

Okubo:1993:TPYb

[1637] Susumu Okubo. Triple products and Yang–Baxter equation. II. Orthogonal and symplectic ternary systems. Journal of Mathematical Physics, 34(7):3292–3315, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i7/p3292_s1.

Pestov:1993:SES

[1638] Vladimir G. Pestov. Soul expansion of G^{∞} superfunctions. Journal of Mathematical Physics, 34(7):3316–3323, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i7/p3316_s1.

Albeverio:1993:CLA

[1639] S. Albeverio, K. Iwata, and M. Schmidt. A convergent lattice approximation for nonlinear electromagnetic fields in four dimensions. Journal of Mathematical Physics, 34(8):3327–3342, August 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3327_s1.

Amar:1993:SEC

[1640] V. Amar, M. Pauri, and A. Scotti. Schrödinger equation for convex plane polygons. II. A no-go theorem for plane waves representation of solutions. Journal of Mathematical Physics, 34(8):3343–3350, August 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3343_
s1.

Aquilanti:1993:ASE

[1641] Vincenzo Aquilanti, Simonetta Cavalli, and Mikhail B. Sevryuk. Asymptotic (semiclassical) equivalence for Schrödinger equations with singular potentials and for related systems of two first-order equations. Journal of Mathematical Physics, 34(8):3351–3377, August 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3351_s1.

Beckers:1993:FSP

[1642] J. Beckers and N. Debergh. On a family of supersymmetrization procedures. Journal of Mathematical Physics, 34(8):3378–3390, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3378_s1.

Dodonov:1993:QPB

[1643] V. V. Dodonov, A. B. Klimov, and D. E. Nikonov. Quantum particle in a box with moving walls. Journal of Mathematical Physics, 34(8):3391– 3404, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i8/p3391_s1.

Horwitz:1993:SSP

[1644] L. P. Horwitz. Some spectral properties of anti-self-adjoint operators on a quaternionic Hilbert space. Journal of Mathematical Physics, 34(8):3405–3419, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3405_s1.

Kochubei:1993:STE

[1645] Anatoly N. Kochubei. A Schrödinger-type equation over the field of p-adic numbers. Journal of Mathematical Physics, 34(8):3420–3428, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3420_s1.

Pinson:1993:ACU

[1646] Haru T. Pinson. Alternative characterizations of unitarity of conformal models with 0 < c < 1. Journal of Mathematical Physics, 34(8):3429– 3439, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i8/p3429_s1.

Sun:1993:EQD

[1647] Chang Pu Sun. Exotic quantum double, its universal R matrix, and their representations. Journal of Mathematical Physics, 34(8):3440–3452, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3440_s1.

Suzuki:1993:DPA

[1648] Takashi Suzuki. Deformations of Poisson algebra in terms of star products and q-deformed mechanics. Journal of Mathematical Physics, 34 (8):3453–3462, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3453_s1.

Dhara:1993:DCS

[1649] H. C. Dhara and A. Roy Chowdhury. On a different class of solutions for Minkowskian σ model on Grassmanian [Grassmannian] manifold. Journal of Mathematical Physics, 34(8):3463–3470, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3463_s1.

Fan:1993:PSA

[1650] Qi Hong Fan. Phase space analysis of the identity decompositions. Journal of Mathematical Physics, 34(8):3471–3477, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3471_s1.

Helfer:1993:NID

[1651] Adam D. Helfer. Null infinity does not carry massive fields. Journal of Mathematical Physics, 34(8):3478–3480, August 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3478_s1.

Iguchi:1993:EBN

[1652] Kazumoto Iguchi. Equivalence between the Nielsen and the scaling transformations in one-dimensional quasiperiodic systems. Journal of Mathematical Physics, 34(8):3481–3490, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i8/p3481_s1.

Andreev:1993:HLK

[1653] V. A. Andreev and M. V. Shmakova. Hierarchy of lower Korteweg–de Vries equations and supersymmetry structure of Miura transformations.

Journal of Mathematical Physics, 34(8):3491–3506, August 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3491_
s1.

Athorne:1993:HSD

[1654] C. Athorne and I. Ya. Dorfman. The Hamiltonian structure of the (2+1)- dimensional Ablowitz–Kaup–Newell–Segur hierarchy. Journal of Mathematical Physics, 34(8):3507–3517, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i8/p3507_s1.

Harnad:1993:IFL

[1655] J. Harnad and M.-A. Wisse. Isospectral flow in loop algebras and quasiperiodic solutions of the sine–Gordon equation. Journal of Mathematical Physics, 34(8):3518–3526, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i8/p3518_s1.

Tenenblat:1993:SGI

[1656] K. Tenenblat and P. Winternitz. On the symmetry groups of the intrinsic generalized wave and sine–Gordon equations. Journal of Mathematical Physics, 34(8):3527–3542, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i8/p3527_s1.

Bokhari:1993:CRT

[1657] Ashfaque H. Bokhari and Asghar Qadir. Collineations of the Ricci tensor. Journal of Mathematical Physics, 34(8):3543–3552, August 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3543_
s1.

Halilsoy:1993:NLF

[1658] M. Halilsoy. Note on: “Large family of colliding waves in the Einstein– Maxwell theory”, [J. Math. Phys. 31 (1990), no. 11, 2694–2698; MR 91m:83033]. Journal of Mathematical Physics, 34(8):3553–3554, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3553_s1.

Petters:1993:AST

[1659] A. O. Petters. Arnol′d’s singularity theory and gravitational lensing. Journal of Mathematical Physics, 34(8):3555–3581, August 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3555_
s1.

Sharma:1993:PCS

[1660] Ramesh Sharma. Proper conformal symmetries of space–times with divergence-free Weyl conformal tensor. Journal of Mathematical Physics, 34(8):3582–3587, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3582_s1.

Baskerville:1993:BHG

[1661] W. K. Baskerville and S. Majid. The braided Heisenberg group. Journal of Mathematical Physics, 34(8):3588–3606, August 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3588_s1.

Boenkost:1993:VOH

[1662] W. Boenkost and F. Constantinescu. Vertex operators in Hilbert space. Journal of Mathematical Physics, 34(8):3607–3615, August 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/p3607_
s1.

Delbourgo:1993:GCL

[1663] R. Delbourgo, P. D. Jarvis, and Roland C. Warner. Grassmann coordinates and Lie algebras for unified models. Journal of Mathematical Physics, 34(8):3616–3641, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v34/i8/p3616_s1.

Doran:1993:LGS

[1664] C. Doran, D. Hestenes, F. Sommen, and N. Van Acker. Lie groups as spin groups. Journal of Mathematical Physics, 34(8):3642–3669, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3642_s1.

Fernandez:1993:FMA

[1665] Francisco M. Fernández and E. A. Castro. A factorization method with application to perturbation theory. Journal of Mathematical Physics, 34(8):3670–3675, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3670_s1.

Harin:1993:CDS

[1666] A. O. Harin. On a countable-dimensional subalgebra of the equivalence algebra for equations v_{tt} = f(x, v_{x})v_{xx} + g(x, v_{x}). Journal of Mathematical Physics, 34(8):3676–3682, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i8/p3676_s1.

Lasenby:1993:GCP

[1667] Anthony Lasenby, Chris Doran, and Stephen Gull. Grassmann calculus, pseudoclassical mechanics, and geometric algebra. Journal of Mathematical Physics, 34(8):3683–3712, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i8/p3683_s1.

Levi:1993:SCS

[1668] D. Levi and P. Winternitz. Symmetries and conditional symmetries of differential-difference equations. Journal of Mathematical Physics, 34(8):3713–3730, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3713_s1.

Mankoc-Borstnik:1993:SVR

[1669] Norma Mankoč-Borštnik. Spinor and vector representations in fourdimensional Grassmann space. Journal of Mathematical Physics, 34 (8):3731–3745, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3731_s1.

Manogue:1993:FLT

[1670] Corinne A. Manogue and Jörg Schray. Finite Lorentz transformations, automorphisms, and division algebras. Journal of Mathematical Physics, 34(8):3746–3767, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3746_s1.

Romero:1993:CGO

[1671] Alfonso Romero and Miguel Sánchez. On the completeness of geodesics obtained as a limit. Journal of Mathematical Physics, 34(8):3768– 3774, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i8/p3768_s1.

Saito:1993:ADF

[1672] Yoshihiko Saito, Reiji Sugano, Tadayuki Ohta, and Toshiei Kimura. Addendum to: “A dynamical formalism of singular Lagrangian system with higher derivatives”, [J. Math. Phys. 30 (1989), no. 5, 1122–1132; MR 90d:70047]. Journal of Mathematical Physics, 34(8):3775–3779, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3775_s1. See [9].

Scheunert:1993:PDS

[1673] M. Scheunert. The presentation and q deformation of special linear Lie superalgebras. Journal of Mathematical Physics, 34(8):3780–3808, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3780_s1.

Schlichenmaier:1993:DGK

[1674] Martin Schlichenmaier. Degenerations of generalized Krichever–Novikov algebras on tori. Journal of Mathematical Physics, 34(8):3809–3824, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i8/p3809_s1.

Swift:1993:NBIa

[1675] S. Timothy Swift. Natural bundles. II. Spin and the diffeomorphism group. Journal of Mathematical Physics, 34(8):3825–3840, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i8/
p3825_s1.

Swift:1993:NBIb

[1676] S. Timothy Swift. Natural bundles. III. Resolving the singularities in the space of immersed submanifolds. Journal of Mathematical Physics, 34(8):3841–3855, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3841_s1.

TorresdelCastillo:1993:SWC

[1677] G. F. Torres del Castillo. Spin-weighted cylindrical harmonics and the Euclidean group of the plane. Journal of Mathematical Physics, 34(8):3856–3862, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i8/p3856_s1.

Alicki:1993:WPP

[1678] Robert Alicki and John R. Klauder. Wiener and Poisson process regularization for coherent-state path integrals. Journal of Mathematical Physics, 34(9):3867–3877, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i9/p3867_s1.

deOliveira:1993:SPS

[1679] César R. de Oliveira. Spectral properties of a simple Hamiltonian model. Journal of Mathematical Physics, 34(9):3878–3886, September 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p3878_
s1.

Isar:1993:DMD

[1680] Aurelian Isar, Aurel Săndulescu, and Werner Scheid. Density matrix for the damped harmonic oscillator within the Lindblad theory. Journal of Mathematical Physics, 34(9):3887–3900, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p3887_s1.

Kelnhofer:1993:GSC

[1681] Gerald Kelnhofer. On the geometrical structure of covariant anomalies in Yang–Mills theory. Journal of Mathematical Physics, 34(9):3901–3917, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p3901_s1.

Lammerzahl:1993:POS

[1682] Claus Lämmerzahl. The pseudodifferential operator square root of the Klein–Gordon equation. Journal of Mathematical Physics, 34(9):3918– 3932, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p3918_s1. See addendum [2112].

Nirov:1993:EBL

[1683] Kh. S. Nirov and A. V. Razumov. Equivalence between Lagrangian and Hamiltonian BRST formalisms. Journal of Mathematical Physics, 34 (9):3933–3953, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p3933_s1.

Plyushchay:1993:QCS

[1684] Mikhail S. Plyushchay. Quantization of the classical SL(2, R) system and representations of SL(2, R) group. Journal of Mathematical Physics, 34 (9):3954–3963, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p3954_s1.

Salmistraro:1993:ILA

[1685] Franco Salmistraro and Riccardo Rosso. Invariants and Lie algebraic solutions of Schrödinger equation. Journal of Mathematical Physics, 34 (9):3964–3979, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p3964_s1.

vanderWalt:1993:ASB

[1686] D. M. van der Walt, R. M. Quick, and M. de Llano. Analytic solution of the Bardeen–Cooper–Schrieffer gap equation in one, two, and three dimensions. Journal of Mathematical Physics, 34(9):3980–3985, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p3980_s1. See erratum [1871].

Das:1993:GSM

[1687] A. Das. General solutions of Maxwell–Dirac equations in (1 + 1)- dimensional space–time and a spatially confined solution. Journal of Mathematical Physics, 34(9):3986–3999, September 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i9/p3986_s1.

Ghaboussi:1993:GCM

[1688] F. Ghaboussi. Geometrization of classical mechanics. Journal of Mathematical Physics, 34(9):4000–4006, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i9/p4000_s1.

Palmer:1993:FWS

[1689] Matthew R. Palmer and Rod Donnelly. Focused waves and the scalar wave equation. Journal of Mathematical Physics, 34(9):4007–4013, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p4007_s1.

Chung:1993:SAS

[1690] S. K. Chung and H. F. Chau. On the structure of absolute steady states in sandpile type of cellular automata: The geometrical aspect. Journal of Mathematical Physics, 34(9):4014–4024, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p4014_s1.

Sachdev:1993:ISS

[1691] P. L. Sachdev and Sharadha Ramanan. Integrability and singularity structure of predator-prey system. Journal of Mathematical Physics, 34 (9):4025–4044, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4025_s1.

Yordanov:1993:WDS

[1692] Russi G. Yordanov. Why do soliton equations come in hierarchies? Journal of Mathematical Physics, 34(9):4045–4049, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p4045_s1.

Yung:1993:SAA

[1693] C. M. Yung and Roland C. Warner. The N = 2 super W_{4} algebra and its associated generalized Korteweg–de Vries hierarchies. Journal of Mathematical Physics, 34(9):4050–4063, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p4050_s1.

Griffiths:1993:CGW

[1694] J. B. Griffiths. The collision of gravitational waves in a stiff perfect fluid. Journal of Mathematical Physics, 34(9):4064–4069, September 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p4064_
s1.

Liu:1993:TDK

[1695] Hong Ya Liu, Paul Wesson, and J. Ponce de León. Time-dependent Kaluza–Klein soliton solutions. Journal of Mathematical Physics, 34 (9):4070–4079, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4070_s1.

PoncedeLeon:1993:ESE

[1696] J. Ponce de León and P. S. Wesson. Exact solutions and the effective equation of state in Kaluza–Klein theory. Journal of Mathematical Physics, 34(9):4080–4092, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i9/p4080_s1.

Witt:1993:SOT

[1697] Hans J. Witt and Arlie O. Petters. Singularities of the one- and two-point mass gravitational lens. Journal of Mathematical Physics, 34(9):4093– 4111, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p4093_s1.

Bacry:1993:BRD

[1698] Henri Bacry. Band representations and duals of the nonsymmorphic crystallographic plane groups pg and p4gm. Journal of Mathematical Physics, 34(9):4112–4137, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i9/p4112_s1.

Brownstein:1993:POT

[1699] K. R. Brownstein. Projection operator technique for imposing the boundary conditions on a trial wave function. Journal of Mathematical Physics, 34(9):4138–4143, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4138_s1.

Chen:1993:FSF

[1700] William Y. C. Chen and James D. Louck. The factorial Schur function. Journal of Mathematical Physics, 34(9):4144–4160, September 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p4144_
s1.

Daumens:1993:CDR

[1701] M. Daumens. Canonical and differential realizations of OSP(3|2) superalgebra. Journal of Mathematical Physics, 34(9):4161–4171, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i9/p4161_s1.

Grigore:1993:PUI

[1702] D. R. Grigore. The projective unitary irreducible representations of the Poincaré group in 1 + 2 dimensions. Journal of Mathematical Physics, 34(9):4172–4189, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4172_s1.

Holschneider:1993:GIF

[1703] Matthias Holschneider. General inversion formulas for wavelet transforms. Journal of Mathematical Physics, 34(9):4190–4198, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i9/p4190_s1.

Hussin:1993:SFT

[1704] V. Hussin and L. M. Nieto. Supergroups factorizations through matrix realization. Journal of Mathematical Physics, 34(9):4199–4220, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i9/p4199_s1.

Kauderer:1993:MDF

[1705] Mark Kauderer. Modes in n-dimensional first-order systems. Journal of Mathematical Physics, 34(9):4221–4250, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i9/p4221_s1.

Lesimple:1993:DLG

[1706] Marc Lesimple and Georges Pinczon. Deformations of Lie group and Lie algebra representations. Journal of Mathematical Physics, 34(9):4251– 4272, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p4251_s1.

McInnes:1993:HGC

[1707] Brett McInnes. Holonomy groups of compact Riemannian manifolds: A classification in dimensions up to ten. Journal of Mathematical Physics, 34(9):4273–4286, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4273_s1.

McInnes:1993:EEM

[1708] Brett McInnes. Examples of Einstein manifolds with all possible holonomy groups in dimensions less than seven. Journal of Mathematical Physics, 34(9):4287–4304, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i9/p4287_s1.

Pan:1993:ICHa

[1709] Feng Pan and Jin Quan Chen. The induction coefficients of the Hecke algebra and the Clebsch–Gordan coefficients of the quantum group SU_{q}(N). I. Special Gel′fand basis. Journal of Mathematical Physics, 34(9):4305–4315, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4305_s1.

Pan:1993:ICHb

[1710] Feng Pan and Jin Quan Chen. The induction coefficients of the Hecke algebra and the Clebsch–Gordan coefficients of the quantum group SU_{q}(N). II. General Gel′fand basis. Journal of Mathematical Physics, 34(9):4316–4329, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i9/p4316_s1.

Rezende:1993:NBS

[1711] Jorge Rezende. A note on Borel summability. Journal of Mathematical Physics, 34(9):4330–4339, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i9/p4330_s1.

Schnizer:1993:RUR

[1712] W. A. Schnizer. Roots of unity: Representations for symplectic and orthogonal quantum groups. Journal of Mathematical Physics, 34(9):4340– 4363, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p4340_s1.

Shawagfeh:1993:NAS

[1713] N. T. Shawagfeh. Nonperturbative approximate solution for Lane– Emden equation. Journal of Mathematical Physics, 34(9):4364–4369, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p4364_s1.

Cairo:1993:CIM

[1714] L. Cairó and D. Hua. Comments on: “Integrals of motion for the Lorenz system”, [J. Math. Phys. 34 (1993), no. 2, 801–804; MR 93j:34071] by N. Gupta. Journal of Mathematical Physics, 34(9):4370–4371, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i9/p4370_s1. See [1458].

Chang:1993:EGH

[1715] Zhe Chang. Erratum: “Generalized Holstein–Primakoff realizations and quantum group-theoretic coherent states”, [J. Math. Phys. 33 (1992), no. 9, 3172–3179; MR 93h:81063]. Journal of Mathematical Physics, 34(9): 4372, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i9/p4372_s1. See [1274].

Blokhin:1993:CNP

[1716] A. L. Blokhin. Collective nuclear potential within the microscopic Sp(6, R) model. Journal of Mathematical Physics, 34(10):4377–4420, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4377_s1.

Chantelau:1993:RSO

[1717] Klaus Chantelau. On the regularization of Schrödinger operators with Coulombic singularities. Journal of Mathematical Physics, 34(10):4421– 4427, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4421_s1.

Crawford:1993:DOL

[1718] James P. Crawford. The Dirac oscillator and local automorphism invariance. Journal of Mathematical Physics, 34(10):4428–4435, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4428_s1.

Dalarsson:1993:PIA

[1719] Nils Dalarsson. Phase-integral approach to quantum-mechanical tunneling. Journal of Mathematical Physics, 34(10):4436–4440, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i10/
p4436_s1.

Fairlie:1993:SQQ

[1720] David Fairlie and Jean Nuyts. Study of quommutators of quantum variables and generalized derivatives. Journal of Mathematical Physics, 34 (10):4441–4454, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4441_s1.

Fukui:1993:FPI

[1721] Takahiro Fukui. Fluctuation path integrals in the generalized coherent states representation. Journal of Mathematical Physics, 34(10):4455– 4468, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4455_s1.

Grosse:1993:STC

[1722] Harald Grosse, Walter Maderner, and Christian Reitberger. Schwinger terms and cyclic cohomology for massive (1 + 1)-dimensional fermions and Virasoro algebras. Journal of Mathematical Physics, 34(10):4469– 4477, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4469_s1.

Hiroshima:1993:SLM

[1723] Fumio Hiroshima. Scaling limit of a model of quantum electrodynamics. Journal of Mathematical Physics, 34(10):4478–4518, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i10/
p4478_s1.

Kay:1993:SCQ

[1724] Bernard S. Kay. Sufficient conditions for quasifree states and an improved uniqueness theorem for quantum fields on space–times with horizons. Journal of Mathematical Physics, 34(10):4519–4539, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i10/
p4519_s1.

Lammerzahl:1993:DST

[1725] Claus Lämmerzahl and Alfredo Macías. On the dimensionality of space– time. Journal of Mathematical Physics, 34(10):4540–4553, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i10/

p4540_s1. See addendum [2309].

Latypov:1993:TTK

[1726] D. M. Latypov and A. M. Mukhamedzhanov. Two theorems on the kernel of Faddeev equations for the scattering of three charged particles and some applications. Journal of Mathematical Physics, 34(10):4554– 4561, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4554_s1.

Nag:1993:OSD

[1727] Subhashis Nag and Parameswaran Sankaran. Open string diagrams. I. Topological type. Journal of Mathematical Physics, 34(10):4562– 4574, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4562_s1.

Oseguera:1993:TSP

[1728] U. Oseguera and M. de Llano. Two singular potentials: The spacesplitting effect. Journal of Mathematical Physics, 34(10):4575–4589, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4575_s1.

Rodriguez-Romo:1993:ASD

[1729] Suemi Rodríguez-Romo. Algebraic spinors in Dirac propagators as a directed random walk. Journal of Mathematical Physics, 34(10):4590– 4600, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4590_s1.

Tanimura:1993:RRIb

[1730] N. Tanimura and O. Tanimura. Recursion relation in SU(N) integral in quantum chromodynamics. Journal of Mathematical Physics, 34(10):4601–4616, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4601_s1.

Bette:1993:TPS

[1731] Andreas Bette. Twistor phase space dynamics and the Lorentz force equation. Journal of Mathematical Physics, 34(10):4617–4627, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4617_s1.

He:1993:CGF

[1732] Sailing He. A “compact Green function” approach to the time-domain direct and inverse problems for a stratified dissipative slab. Journal of Mathematical Physics, 34(10):4628–4645, October 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i10/p4628_s1.

Lewandowski:1993:VPP

[1733] Jerzy Lewandowski, Ezra T. Newman, and Carlo Rovelli. Variations of the parallel propagator and holonomy operator and the Gauss law constraint. Journal of Mathematical Physics, 34(10):4646–4654, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4646_s1.

Ohta:1993:CFA

[1734] Tadayuki Ohta and Toshiei Kimura. Canonical formalism of action-at-a-distance electrodynamics and many-particle potential among charged particles. Journal of Mathematical Physics, 34(10):4655–4667, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4655_s1.

Busoni:1993:ASO

[1735] Giorgio Busoni and Giovanni Frosali. Asymptotic scattering operators and drift velocity in charged particle transport problems. Journal of Mathematical Physics, 34(10):4668–4691, October 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i10/p4668_s1.

Arrigo:1993:NSS

[1736] D. J. Arrigo, P. Broadbridge, and J. M. Hill. Nonclassical symmetry solutions and the methods of Bluman–Cole and Clarkson–Kruskal. Journal of Mathematical Physics, 34(10):4692–4703, October 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i10/p4692_s1.

Calogero:1993:IMC

[1737] F. Calogero. Interpolation in multidimensions, a convenient finitedimensional matrix representation of the (partial) differential operators, and some applications. Journal of Mathematical Physics, 34(10):4704– 4724, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4704_s1.

Levermore:1993:MCP

[1738] C. D. Levermore and G. B. Zimmerman. Modeling charged particle loss in a fuel/shell mixture. Journal of Mathematical Physics, 34(10):4725– 4729, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4725_s1.

Tafel:1993:LAA

[1739] J. Tafel and A. Wisse. Loop algebra approach to generalized sine–Gordon equations. Journal of Mathematical Physics, 34(10):4730–4741, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4730_s1.

Zeng:1993:BHS

[1740] Yun Bo Zeng. The bi-Hamiltonian structure for the restricted flows of the Boussinesq and the modified Boussinesq equation. Journal of Mathematical Physics, 34(10):4742–4754, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i10/p4742_s1.

GarciaDiaz:1993:APP

[1741] A. García Díaz, J. F. Plebański, and N. R. Sibgatullin. On the absence of the Painlevé property in Petrov type III Robinson–Trautman solutions. Journal of Mathematical Physics, 34(10):4755–4759, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4755_s1.

Hernandez-Pastora:1993:SAF

[1742] J. L. Hernández-Pastora, V. S. Man′ko, and J. Martin. Some asymptotically flat generalizations of the Curzon metric. Journal of Mathematical Physics, 34(10):4760–4774, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i10/p4760_s1.

Junevicus:1993:ESD

[1743] G. J. G. Junevicus. An exact solution for a dielectric atmosphere in general relativity. Journal of Mathematical Physics, 34(10):4775–4780, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i10/p4775_s1.

Levine:1993:AST

[1744] H. I. Levine, A. O. Petters, and J. Wambsganss. Applications of singularity theory to gravitational lensing. I. Multiple lens planes. Journal of Mathematical Physics, 34(10):4781–4808, October 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i10/p4781_s1.

Abraham-Shrauner:1993:HSL

[1745] Barbara Abraham-Shrauner. Hidden symmetries and linearization of the modified Painlevé-Ince equation. Journal of Mathematical Physics, 34(10):4809–4816, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4809_s1.

Borthwick:1993:HSS

[1746] David Borthwick, Andrzej Lesniewski, and Maurizio Rinaldi. Hermitian symmetric superspaces of type IV. Journal of Mathematical Physics, 34(10):4817–4833, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4817_s1.

Davtyan:1993:HTG

[1747] Levon S. Davtyan. The Hurwitz transformation and generalized Cayley– Klein parameterization. Journal of Mathematical Physics, 34(10):4834– 4839, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4834_s1.

Low:1993:KNH

[1748] Robert J. Low. Kinking number and the Hopf index theorem. Journal of Mathematical Physics, 34(10):4840–4842, October 1993. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v34/i10/p4840_s1.

Majid:1993:FBD

[1749] S. Majid. Free braided differential calculus, braided binomial theorem, and the braided exponential map. Journal of Mathematical Physics, 34(10):4843–4856, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4843_s1.

McInnes:1993:CSG

[1750] Brett McInnes. Complex symplectic geometry and compact locally hyper-Kählerian manifolds. Journal of Mathematical Physics, 34(10): 4857–4871, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4857_s1.

Palev:1993:SGD

[1751] Tchavdar D. Palev. A superalgebra U_{q}[osp(3/2)] generated by deformed paraoperators and its morphism onto a W_{q}(1/1) Clifford–Weyl algebra. Journal of Mathematical Physics, 34(10):4872–4883, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i10/
p4872_s1.

Ricci:1993:SRZ

[1752] Paolo Emilio Ricci. Sum rules for zeros of polynomials and generalized Lucas polynomials. Journal of Mathematical Physics, 34(10):4884– 4891, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i10/p4884_s1.

Suzuki:1993:CUC

[1753] Masuo Suzuki and Takashi Yamauchi. Convergence of unitary and complex decompositions of exponential operators. Journal of Mathematical Physics, 34(10):4892–4897, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i10/p4892_s1.

Vargas:1993:FSC

[1754] José G. Vargas and Douglas G. Torr. Finslerian structures: The Cartan– Clifton method of the moving frame. Journal of Mathematical Physics, 34(10):4898–4913, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4898_s1.

Znojil:1993:CNS

[1755] Miloslav Znojil. Comment on: “The nonsingular spiked harmonic oscillator”, [J. Math. Phys. 34 (1993), no. 2, 437–440; MR 93j:33002] by E. S. Estévez-Bretón and G. A. Estévez-Bretón. Journal of Mathematical Physics, 34(10):4914, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i10/p4914_s1. See [1434].

Beltrametti:1993:CPG

[1756] Enrico G. Beltrametti and Maciej J. Maczyński. On the characterization of probabilities: A generalization of Bell’s inequalities. Journal of Mathematical Physics, 34(11):4919–4929, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i11/p4919_s1.

Carimalo:1993:SSP

[1757] C. Carimalo. On the spinor structure of the proton wave function. Journal of Mathematical Physics, 34(11):4930–4963, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/
p4930_s1.

Cohen:1993:RHA

[1758] Jeffrey M. Cohen and Boris Kuharetz. Relativistic hydrogen atom: Wave equation in Whittaker form. Journal of Mathematical Physics, 34(11):4964–4974, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i11/p4964_s1.

Jensen:1993:GAB

[1759] Bjørn Jensen and Jaromír Kučera. On a gravitational Aharonov–Bohm effect. Journal of Mathematical Physics, 34(11):4975–4985, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p4975_s1.

Karasev:1993:QSR

[1760] Mikhail Karasev and Mikhail Kozlov. Quantum and semiclassical representations over Lagrangian submanifolds in su(2)^{∗}, so(4)^{∗}, and su(1, 1)^{∗}. Journal of Mathematical Physics, 34(11):4986–5006, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/
p4986_s1.

Majewski:1993:QCE

[1761] W. A. Majewski and M. Kuna. On quantum characteristic exponents. Journal of Mathematical Physics, 34(11):5007–5015, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/
p5007_s1.

Maldonado:1993:BTQ

[1762] Osvaldo Maldonado. On the Bogoliubov transformation for quadratic boson observables. Journal of Mathematical Physics, 34(11):5016–5027, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5016_s1.

Posilicano:1993:CND

[1763] Andrea Posilicano and Stefania Ugolini. Convergence of Nelson diffusions with time-dependent electromagnetic potentials. Journal of Mathematical Physics, 34(11):5028–5036, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i11/p5028_s1.

Shishkin:1993:END

[1764] German V. Shishkin and Víctor M. Villalba. Electrically neutral Dirac particles in the presence of external fields: Exact solutions. Journal of Mathematical Physics, 34(11):5037–5049, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/p5037_s1.

Witwit:1993:QTA

[1765] M. R. M. Witwit. Quantum theory of anharmonic oscillator: Energy levels of a three-dimensional oscillator with quartic anisotropic perturbation. Journal of Mathematical Physics, 34(11):5050–5061, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5050_s1.

Dattoli:1993:QMP

[1766] G. Dattoli and A. Torre. A quantum mechanical point of view to perturbative problems in classical mechanics. Journal of Mathematical Physics, 34(11):5062–5071, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i11/p5062_s1.

vanderMee:1993:ECM

[1767] C. V. M. van der Mee. An eigenvalue criterion for matrices transforming Stokes parameters. Journal of Mathematical Physics, 34(11):5072– 5088, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5072_s1.

Hertle:1993:LDI

[1768] Jochen Hertle. The limiting dynamics of an imperfect Bose gas. Journal of Mathematical Physics, 34(11):5089–5108, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/p5089_s1.

Chau:1993:GAS

[1769] H. F. Chau and K. S. Cheng. Generalized Abelian sandpile model. Journal of Mathematical Physics, 34(11):5109–5117, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/
p5109_s1.

delOlmo:1993:ISB

[1770] M. A. del Olmo, M. A. Rodríguez, and P. Winternitz. Integrable systems based on SU(p, q) homogeneous manifolds. Journal of Mathematical Physics, 34(11):5118–5139, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i11/p5118_s1.

Fuchssteiner:1993:INE

[1771] Benno Fuchssteiner. Integrable nonlinear evolution equations with timedependent coefficients. Journal of Mathematical Physics, 34(11):5140– 5158, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5140_s1.

McIntosh:1993:GEL

[1772] Ian McIntosh. Groups of equivariant loops in SL(n + 1, C) and soliton equations. Journal of Mathematical Physics, 34(11):5159–5180, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5159_s1.

Nahmad-Achar:1993:VDE

[1773] E. Nahmad-Achar. Variational derivation of the equations of motion for a perfect fluid in a rotating cylinder. Journal of Mathematical Physics, 34(11):5181–5189, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i11/p5181_s1.

Ohta:1993:CDS

[1774] Yasuhiro Ohta, Kenji Kajiwara, Junta Matsukidaira, and Junkichi Satsuma. Casorati determinant solution for the relativistic Toda lattice equation. Journal of Mathematical Physics, 34(11):5190–5204, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5190_s1.

Ottolenghi:1993:CNF

[1775] A. Ottolenghi. On convergence of normal forms for complex frequencies. Journal of Mathematical Physics, 34(11):5205–5220, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5205_s1.

Roepstorff:1993:LES

[1776] G. Roepstorff. Liapunov exponents of simple smooth flows. Journal of Mathematical Physics, 34(11):5221–5237, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/p5221_s1.

Rondoni:1993:ARD

[1777] Lamberto Rondoni. Autocatalytic reactions as dynamical systems on the interval. Journal of Mathematical Physics, 34(11):5238–5251, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5238_s1.

Ernst:1993:GSE

[1778] Frederick J. Ernst and Isidore Hauser. On Gürses’s symmetries of the Einstein equations. Journal of Mathematical Physics, 34(11):5252–5267, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5252_s1.

Kubyshin:1993:ICH

[1779] Yu. A. Kubyshin, O. Richter, and G. Rudolph. Invariant connections on homogeneous spaces. Journal of Mathematical Physics, 34(11):5268– 5282, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5268_s1.

Brihaye:1993:QES

[1780] Y. Brihaye and M. Godart. Quasi-exactly solvable extensions of the Lamé equation. Journal of Mathematical Physics, 34(11):5283–5291, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5283_s1.

Clercx:1993:AEA

[1781] H. J. H. Clercx and P. P. J. M. Schram. An alternative expression for the addition theorems of spherical wave solutions of the Helmholtz equation. Journal of Mathematical Physics, 34(11):5292–5302, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5292_s1.

Daumens:1993:SS

[1782] M. Daumens, P. Minnaert, and Y. Noirot. Supersymmetries of the superspace E(3|2). Journal of Mathematical Physics, 34(11):5303–5319, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i11/p5303_s1.

Havlicek:1993:EPL

[1783] M. Havlíček and P. Moylan. An embedding of the Poincaré Lie algebra into an extension of the Lie field of SO_{0}(1, 4). Journal of Mathematical Physics, 34(11):5320–5332, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i11/p5320_s1.

Kalnins:1993:MAR

[1784] E. G. Kalnins, Willard Miller, Jr., and Sanchita Mukherjee. Models of q-algebra representations: Matrix elements of the q-oscillator algebra. Journal of Mathematical Physics, 34(11):5333–5356, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/
p5333_s1.

Luczka:1993:ESE

[1785] Jerzy Łuczka, Mariusz Niemiec, and Edward Piotrowski. Exact solution of evolution equation for randomly interrupted diffusion. Journal of Mathematical Physics, 34(11):5357–5366, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/p5357_s1.

Ludu:1993:GDA

[1786] Andrei Ludu and Raj K. Gupta. A generalization of the deformed algebra of quantum group SU(2)_{q} for Hopf algebra. Journal of Mathematical Physics, 34(11):5367–5375, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i11/p5367_s1.

Ren:1993:ESC

[1787] Wei Ren. Exact solutions of coupled-wave equations in piezoelectric solids. Journal of Mathematical Physics, 34(11):5376–5390, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i11/p5376_s1.

Robinson:1993:CGC

[1788] Ivor Robinson and Andrzej Trautman. The conformal geometry of complex quadrics and the fractional-linear form of Möbius transformations. Journal of Mathematical Physics, 34(11):5391–5406, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/
p5391_s1.

Pang:1993:ACL

[1789] Gen Di Pang and Bao Heng Zhao. Addendum to: “Conservation laws of the quantized Davey–Stewartson I system in (2 + 1) dimensions”, [J. Math. Phys. 34 (1993), no. 3, 1063–1075; MR 94h:81070]. Journal of Mathematical Physics, 34(11):5407–5408, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i11/

p5407_s1. See [1482].

Bandelloni:1993:DCB

[1790] Giuseppe Bandelloni and Serge Lazzarini. Diffeomorphism cohomology in Beltrami parametrization. Journal of Mathematical Physics, 34 (12):5413–5440, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i12/p5413_s1.

Breuer:1993:RQM

[1791] Thomas Breuer, Anton Amann, and Nicolaas P. Landsman. Robustness in quantum measurements. Journal of Mathematical Physics, 34 (12):5441–5450, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i12/p5441_s1.

Caliceti:1993:SAS

[1792] E. Caliceti and A. M. Cherubini. Spectral analysis and stability properties of a relativistic deformation of the harmonic oscillator. Journal of Mathematical Physics, 34(12):5451–5467, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/p5451_s1.

Cassinelli:1993:SPO

[1793] Gianni Cassinelli and Pekka J. Lahti. Spectral properties of observables and convex mappings in quantum mechanics. Journal of Mathematical Physics, 34(12):5468–5475, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p5468_s1.

Hadjiivanov:1993:QDB

[1794] Ludmil K. Hadjiivanov. Quantum deformation of Bose parastatistics. Journal of Mathematical Physics, 34(12):5476–5492, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/
p5476_s1.

Han:1993:STM

[1795] D. Han, Y. S. Kim, Marilyn E. Noz, and Leehwa Yeh. Symmetries of twomode squeezed states. Journal of Mathematical Physics, 34(12):5493– 5508, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5493_s1.

Hostler:1993:CES

[1796] Levere Hostler. Commutator expansion of the self-energy operator for an electron in an external potential. Journal of Mathematical Physics, 34(12):5509–5532, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i12/p5509_s1.

Khudaverdian:1993:EOS

[1797] O. M. Khudaverdian and A. P. Nersessian. Even and odd symplectic and Kählerian structures on projective superspaces. Journal of Mathematical Physics, 34(12):5533–5548, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p5533_s1.

Momont:1993:RCC

[1798] B. Momont. Rigorous calculation of the collective excitation spectrum in a mean field model. Journal of Mathematical Physics, 34(12):5549– 5559, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5549_s1.

Muller:1993:NAS

[1799] V. F. Müller. Non-Abelian solitons in two-dimensional lattice field theories. Journal of Mathematical Physics, 34(12):5560–5588, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5560_s1.

Nesterenko:1993:MRP

[1800] V. V. Nesterenko. On a model of a relativistic particle with curvature and torsion. Journal of Mathematical Physics, 34(12):5589–5595, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5589_s1.

Ozawa:1993:CAQ

[1801] Masanao Ozawa. Canonical approximate quantum measurements. Journal of Mathematical Physics, 34(12):5596–5624, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/p5596_s1.

Vanzo:1993:FIF

[1802] Luciano Vanzo. The functional integral for fields in a cavity. Journal of Mathematical Physics, 34(12):5625–5638, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/p5625_s1.

Zhang:1993:QTT

[1803] Yong De Zhang and Zhong Tang. Quantum transformation theory in fermion Fock space. Journal of Mathematical Physics, 34(12):5639–5645, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5639_s1.

Cabo:1993:GIS

[1804] Alejandro Cabo, Masud Chaichian, and Domingo Louis Martinez. Gauge invariance of systems with first-class constraints. Journal of Mathematical Physics, 34(12):5646–5658, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p5646_s1.

Calogero:1993:SNC

[1805] F. Calogero and Xiao Da Ji. Solvable (nonrelativistic, classical) n-body problems on the line. I. Journal of Mathematical Physics, 34(12):5659– 5670, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5659_s1.

Diacu:1993:PIP

[1806] Florin N. Diacu. The planar isosceles problem for Maneff’s gravitational law. Journal of Mathematical Physics, 34(12):5671–5690, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5671_s1.

Gumral:1993:PSD

[1807] Hasan Gümral and Yavuz Nutku. Poisson structure of dynamical systems with three degrees of freedom. Journal of Mathematical Physics, 34 (12):5691–5723, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i12/p5691_s1.

He:1993:AGF

[1808] Sailing He, Sergei I. Kabanikhin, Vladimir G. Romanov, and Staffan Ström. Analysis of the Green’s function approach to one-dimensional inverse problems. I. One parameter reconstruction. Journal of Mathematical Physics, 34(12):5724–5746, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i12/p5724_s1.

Soloviev:1993:BVH

[1809] Vladimir O. Soloviev. Boundary values as Hamiltonian variables. I. New Poisson brackets. Journal of Mathematical Physics, 34(12):5747–5769, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5747_s1.

Pereira:1993:CST

[1810] Emmanuel A. Pereira. The critical surface for the tridimensional massless Gaussian fixed point: An unstable manifold with two relevant directions. Journal of Mathematical Physics, 34(12):5770–5780, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5770_s1.

Scalas:1993:HMS

[1811] E. Scalas and G. A. Viano. The Hausdorff moments in statistical mechanics. Journal of Mathematical Physics, 34(12):5781–5800, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5781_s1.

Boldin:1993:OAT

[1812] A. Yu. Boldin, S. S. Safin, and R. A. Sharipov. On an old article of Tzitzeica and the inverse scattering method. Journal of Mathematical Physics, 34(12):5801–5809, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p5801_s1.

Calogero:1993:INP

[1813] F. Calogero and Xiao Da Ji. C-integrable nonlinear partial differential equations. III. Journal of Mathematical Physics, 34(12):5810–5831, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5810_s1.

Freund:1993:HAI

[1814] Peter G. O. Freund and Anton V. Zabrodin. A hierarchic array of integrable models. Journal of Mathematical Physics, 34(12):5832–5842, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5832_s1.

Kaushal:1993:DAA

[1815] R. S. Kaushal and S. C. Mishra. Dynamical algebraic approach and invariants for time-dependent Hamiltonian systems in two dimensions. Journal of Mathematical Physics, 34(12):5843–5850, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/
p5843_s1.

Yu:1993:KPHa

[1816] Feng Yu and Yong-Shi Wu. On the Kadomtsev–Petviashvili hierarchy, Ŵ_{∞} algebra, and conformal SL(2, R)/U(1) model. I. The classical case. Journal of Mathematical Physics, 34(12):5851–5871, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5851_s1.

Yu:1993:KPHb

[1817] Feng Yu and Yong-Shi Wu. On the Kadomtsev–Petviashvili hierarchy, Ŵ_{∞} algebra, and conformal SL(2, R)/U(1) model. II. The quantum case. Journal of Mathematical Physics, 34(12):5872–5896, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5872_s1.

Hall:1993:SRS

[1818] G. S. Hall and A. D. Hossack. Some remarks on sectional curvature and tidal accelerations. Journal of Mathematical Physics, 34(12):5897– 5899, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5897_s1.

Lake:1993:CIK

[1819] Kayll Lake. Curvature invariants for Kantowski–Sachs metrics with source: A conformally invariant scalar field. Journal of Mathematical Physics, 34(12):5900–5907, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p5900_s1.

SanMiguel:1993:SPI

[1820] A. San Miguel and F. Vicente. Some properties of the inertia tensor in Dixon’s theory. Journal of Mathematical Physics, 34(12):5908–5915, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p5908_s1.

Wood:1993:OTS

[1821] W. R. Wood, A. Feoli, and G. Papini. Oscillating thin shells in Weyl space. Journal of Mathematical Physics, 34(12):5916–5923, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5916_s1.

Bergdolt:1993:CRC

[1822] G. Bergdolt. The complete reduction of Clifford algebras. Journal of Mathematical Physics, 34(12):5924–5934, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/p5924_s1.

Bincer:1993:COE

[1823] A. M. Bincer and K. Riesselmann. Casimir operators of the exceptional group G_{2}. Journal of Mathematical Physics, 34(12):5935–5941, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v34/i12/p5935_s1.

Capps:1993:RQ

[1824] Richard H. Capps. Representations of quantum SU(3). Journal of Mathematical Physics, 34(12):5942–5950, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i12/p5942_s1.

ElGradechi:1993:SHS

[1825] Amine M. El Gradechi. On the supersymplectic homogeneous superspace underlying the OSp(1/2) coherent states. Journal of Mathematical Physics, 34(12):5951–5963, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p5951_s1.

Gieres:1993:SCO

[1826] François Gieres and Stefan Theisen. Superconformally covariant operators and super-W-algebras. Journal of Mathematical Physics, 34 (12):5964–5985, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v34/i12/p5964_s1.

Hucks:1993:HCS

[1827] Joseph Hucks. Hyperbolic complex structures in physics. Journal of Mathematical Physics, 34(12):5986–6008, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v34/i12/p5986_s1.

Leaseburg:1993:EBM

[1828] Michael J. Leaseburg and Lawrence R. Mead. Error bounds in maximum entropy approximations. Journal of Mathematical Physics, 34(12):6009– 6015, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p6009_s1.

Links:1993:ECI

[1829] J. R. Links and R. B. Zhang. Eigenvalues of Casimir invariants of U_{q}(gl(m/n)). Journal of Mathematical Physics, 34(12):6016–6024, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p6016_s1.

Maslanka:1993:DMH

[1830] Paweł Maślanka. Deformation map and Hermitian representations of k-Poincaré algebra. Journal of Mathematical Physics, 34(12):6025–6029, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i12/p6025_s1.

Rideau:1993:RQA

[1831] G. Rideau and P. Winternitz. Representations of the quantum algebra su_{q}(2) on a real two-dimensional sphere. Journal of Mathematical Physics, 34(12):6030–6044, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i12/p6030_s1.

Zhang:1993:UCR

[1832] Y.-Z. Zhang and M. D. Gould. Unitarity and complete reducibility of certain modules over quantized affine Lie algebras. Journal of Mathematical Physics, 34(12):6045–6059, December 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v34/i12/p6045_s1.

Fagnola:1994:QFA

[1833] Franco Fagnola, Rolando Rebolledo, and Carlos Saavedra. Quantum flows associated to master equations in quantum optics. Journal of Mathematical Physics, 35(1):1–12, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i1/p1_s1.

Foda:1994:VOS

[1834] Omar Foda, Michio Jimbo, Tetsuji Miwa, Kei Miki, and Atsushi Nakayashiki. Vertex operators in solvable lattice models. Journal of Mathematical Physics, 35(1):13–46, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i1/p13_s1.

Kostrykin:1994:EAC

[1835] Vadim V. Kostrykin and Andrei A. Kvitsinsky. Exact and avoided crossings of adiabatic hyperspherical potential curves. Journal of Mathematical Physics, 35(1):47–70, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i1/p47_s1.

Kuperin:1994:GSF

[1836] Yu. A. Kuperin, S. B. Levin, and Yu. B. Mel′nikov. Generalized stringflip model for quantum cluster scattering. Journal of Mathematical Physics, 35(1):71–95, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p71_s1.

Langmann:1994:CBF

[1837] Edwin Langmann. Cocycles for boson and fermion Bogoliubov transformations. Journal of Mathematical Physics, 35(1):96–112, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p96_s1.

Tip:1994:RTE

[1838] A. Tip. Random time evolution and direct integrals: Constants of the motion and the mass operator. Journal of Mathematical Physics, 35(1):113–126, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p113_s1.

Cherkaev:1994:VPC

[1839] A. V. Cherkaev and L. V. Gibiansky. Variational principles for complex conductivity, viscoelasticity, and similar problems in media with complex moduli. Journal of Mathematical Physics, 35(1):127–145, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p127_s1.

Gonera:1994:NLA

[1840] C. Gonera. A non-Lagrangian approach to gauge symmetry in three dimensions. Journal of Mathematical Physics, 35(1):146–172, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p146_s1.

Hortacsu:1994:RST

[1841] M. Hortaçsu and N. Özdemir. Reduction of space–time by a constraint for C^{(0)} metrics. Journal of Mathematical Physics, 35(1):173–180, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p173_s1.

Martin:1994:SFI

[1842] J. Martin and S. Bouquet. Search for first integrals in relativistic time-dependent Hamiltonian systems. Journal of Mathematical Physics, 35(1):181–189, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p181_s1.

Gabetta:1994:CEK

[1843] E. Gabetta and G. Toscani. On convergence to equilibrium for Kac’s caricature of a Maxwell gas. Journal of Mathematical Physics, 35(1):190– 208, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p190_s1.

Gumbs:1994:SBT

[1844] Godfrey Gumbs and Miroslav Kolář. Supercritical bifurcation theory and stability of point attractors. Journal of Mathematical Physics, 35(1):209– 227, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p209_s1.

Knauf:1994:FSC

[1845] Andreas Knauf. On a ferromagnetic spin chain. II. Thermodynamic limit. Journal of Mathematical Physics, 35(1):228–236, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i1/p228_s1.

Shirzad:1994:MIP

[1846] A. Shirzad and H. Arfaei. Modular invariant partition functions and method of shift vector. Journal of Mathematical Physics, 35(1):237– 254, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p237_s1.

Clarkson:1994:NSR

[1847] Peter A. Clarkson and Simon Hood. New symmetry reductions and exact solutions of the Davey–Stewartson system. I. Reductions to ordinary differential equations. Journal of Mathematical Physics, 35(1):255– 283, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p255_s1.

Kiso:1994:SAK

[1848] Kazuhiro Kiso. Symmetries of the Ablowitz–Kaup–Newell–Segur hierarchy. Journal of Mathematical Physics, 35(1):284–293, January 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i1/p284_
s1.

Cardoso:1994:TDP

[1849] J. G. Cardoso. A twistorial description of the processes of mass scattering of Dirac fields. Journal of Mathematical Physics, 35(1):294–309, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p294_s1.

Kubyshin:1994:ICT

[1850] Yu. A. Kubyshin, V. O. Malyshenko, and D. Marín Ricoy. Invariant connections with torsion on group manifolds and their application in Kaluza–Klein theories. Journal of Mathematical Physics, 35(1):310– 320, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p310_s1.

Kuusk:1994:GMT

[1851] Piret Kuusk, Jüri Örd, and Eugen Paal. Geodesic multiplication and the theory of gravity. Journal of Mathematical Physics, 35(1):321–334, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p321_s1.

Maluf:1994:HFT

[1852] José W. Maluf. Hamiltonian formulation of the teleparallel description of general relativity. Journal of Mathematical Physics, 35(1):335–343, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p335_s1.

Alisauskas:1994:PBO

[1853] S. Ališauskas and W. Berej. On the projected bases for Sp(4) ⊃ U(2) and the orthogonalization problem. Journal of Mathematical Physics, 35(1):344–358, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p344_s1.

Bampi:1994:CHD

[1854] Franco Bampi and Clara Zordan. Constrained hyperbolicity and discontinuity waves. Journal of Mathematical Physics, 35(1):359–363, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p359_s1.

Bender:1994:CFD

[1855] Carl M. Bender and Kimball A. Milton. Continued fraction as a discrete nonlinear transform. Journal of Mathematical Physics, 35(1):364– 367, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p364_s1.

Bender:1994:RWN

[1856] Carl M. Bender, Stefan Boettcher, and Lawrence R. Mead. Random walks in noninteger dimension. Journal of Mathematical Physics, 35(1):368–388, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p368_s1.

Brambilla:1994:CBC

[1857] Nora Brambilla and Emilio Montaldi. On the connection between continuous Hahn polynomials and the Heisenberg algebra. Journal of Mathematical Physics, 35(1):389–392, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i1/p389_s1.

Brownstein:1994:IIO

[1858] K. R. Brownstein. Improved inequalities for one-electron moments of r. Journal of Mathematical Physics, 35(1):393–404, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i1/p393_s1.

deMontigny:1994:GCK

[1859] M. de Montigny, J. Patera, and J. Tolar. Graded contractions and kinematical groups of space–time. Journal of Mathematical Physics, 35(1):405–425, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p405_s1.

HernandezEncinas:1994:SSG

[1860] L. Hernández Encinas and J. Muñoz Masqué. Symplectic structure and gauge invariance on the cotangent bundle. Journal of Mathematical Physics, 35(1):426–434, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i1/p426_s1.

Inoue:1994:MSI

[1861] Atsushi Inoue. Modular systems induced by trace functionals on algebras of unbounded operators. Journal of Mathematical Physics, 35(1):435– 442, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i1/p435_s1.

Kantor:1994:CTD

[1862] I. L. Kantor and J. Patera. A construction of transitively differential groups of degree 3. Journal of Mathematical Physics, 35(1):443–458, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p443_s1.

Kirsten:1994:GME

[1863] Klaus Kirsten. Generalized multidimensional Epstein zeta functions. Journal of Mathematical Physics, 35(1):459–470, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i1/p459_s1.

Klauder:1994:WPH

[1864] J. R. Klauder and R. F. Streater. Wavelets and the Poincaré half-plane. Journal of Mathematical Physics, 35(1):471–478, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i1/p471_s1.

Malbouisson:1994:BLE

[1865] A. P. C. Malbouisson. A Borel-like estimate for the rest of the heat kernel expansion. Journal of Mathematical Physics, 35(1):479–485, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p479_s1.

Seiler:1994:AGS

[1866] Werner M. Seiler. On the arbitrariness of the general solution of an involutive partial differential equation. Journal of Mathematical Physics, 35(1):486–498, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i1/p486_s1.

TorresdelCastillo:1994:ECO

[1867] G. F. Torres del Castillo. Eigenfunctions of the curl operator in spherical coordinates. Journal of Mathematical Physics, 35(1):499–507, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p499_s1.

VandenBerghe:1994:SZR

[1868] G. Vanden Berghe. Structural zeros of Racah coefficients and exceptional Lie algebras. Journal of Mathematical Physics, 35(1):508–516, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p508_s1.

Zha:1994:DHO

[1869] Chao Zheng Zha. q-deformation of high-order Virasoro algebra. Journal of Mathematical Physics, 35(1):517–526, January 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i1/p517_s1.

Krasinski:1994:CST

[1870] Andrzej Krasiński. Comment on: “Space–times with plane-symmetric scalar waves”, [J. Math. Phys. 33 (1992), no. 10, 3506–3508; MR 93j:83022] by J. Z. Li. Journal of Mathematical Physics, 35(1):527, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i1/p527_s1. See [1309].

vanderWalt:1994:EAS

[1871] D. M. van der Walt, R. M. Quick, and M. de Llano. Erratum: “Analytic solution of the Bardeen–Cooper–Schrieffer gap equation in one, two, and three dimensions”, [J. Math. Phys. 34, 3980 (1993)]. Journal of Mathematical Physics, 35(1):528, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i1/p528_s1. See [1686].

Adler:1994:PJI

[1872] Stephen L. Adler, Gyan V. Bhanot, and John D. Weckel. Proof of Jacobi identity in generalized quantum dynamics. Journal of Mathematical Physics, 35(2):531–535, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i2/p531_s1.

Aquilanti:1994:APA

[1873] Vincenzo Aquilanti, Simonetta Cavalli, and Mikhail B. Sevryuk. Adiabatic and post-adiabatic representations for multichannel Schrödinger equations. Journal of Mathematical Physics, 35(2):536–559, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p536_s1.

Arai:1994:DEP

[1874] Takahiro Arai. Dynamical evolution of the projector induced from the generalized squeezed state. Journal of Mathematical Physics, 35(2):560– 567, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i2/p560_s1.

Awobode:1994:GTD

[1875] A. M. Awobode. A generalized time-dependent relativistic wave equation. Journal of Mathematical Physics, 35(2):568–576, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/
p568_s1.

Holas:1994:RDF

[1876] A. Holas and N. H. March. Relativistic density-functional theory of a single-electron problem. Journal of Mathematical Physics, 35(2):577– 586, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i2/p577_s1.

Kalogeropoulos:1994:CGA

[1877] Nikolaos Kalogeropoulos. Comments on the Gribov ambiguity. Journal of Mathematical Physics, 35(2):587–595, February 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i2/p587_s1.

Liboff:1994:PQB

[1878] Richard L. Liboff. The polygon quantum-billiard problem. Journal of Mathematical Physics, 35(2):596–607, February 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i2/p596_s1.

Ma:1994:LRI

[1879] Zhong Shui Ma and Xun Xue. Long-range interaction of electromagnetic field induced by Chern–Simons field. Journal of Mathematical Physics, 35(2):608–627, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p608_s1.

Nerurkar:1994:SOD

[1880] M. G. Nerurkar and H. R. Jauslin. Stability of oscillators driven by ergodic processes. Journal of Mathematical Physics, 35(2):628–645, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p628_s1.

Rainer:1994:RSS

[1881] Martin Rainer. Resolution of simple singularities yielding particle symmetries in a space–time. Journal of Mathematical Physics, 35(2):646– 655, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i2/p646_s1.

Shubov:1994:HEA

[1882] Marianna A. Shubov. High-energy asymptotics of resonances for threedimensional Schrödinger operator with screened Coulomb potential. Journal of Mathematical Physics, 35(2):656–679, February 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/p656_
s1.

Tsay:1994:WSS

[1883] W. J. Tsay. Wave scattering in a strip. Journal of Mathematical Physics, 35(2):680–692, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p680_s1.

Aktosun:1994:NIA

[1884] Tuncay Aktosun and Cornelis van der Mee. Nonuniqueness in inverse acoustic scattering on the line. Journal of Mathematical Physics, 35(2):693–709, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p693_s1.

Calogero:1994:SNC

[1885] F. Calogero and Xiao Da Ji. Solvable (nonrelativistic, classical) nbody problems in multidimensions. I. Journal of Mathematical Physics, 35(2):710–733, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p710_s1.

deFreitas:1994:NTS

[1886] Luciane R. de Freitas, M. Novello, and N. Pinto-Neto. Nonlinear theories of spin-2 field in terms of Fierz variables. Journal of Mathematical Physics, 35(2):734–747, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i2/p734_s1.

Ranada:1994:TDL

[1887] Manuel F. Rañada. Time-dependent Lagrangian systems: A geometric approach to the theory of systems with constraints. Journal of Mathematical Physics, 35(2):748–758, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i2/p748_s1.

Temple-Raston:1994:RFD

[1888] M. Temple-Raston. The reduction of five-dimensional Chern–Simons theories. Journal of Mathematical Physics, 35(2):759–768, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p759_s1.

Meurice:1994:DLR

[1889] Yannick Meurice. Duality in long-range Ising ferromagnets. Journal of Mathematical Physics, 35(2):769–779, February 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i2/p769_s1.

Petz:1994:GCC

[1890] Dénes Petz. Geometry of canonical correlation on the state space of a quantum system. Journal of Mathematical Physics, 35(2):780–795, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i2/p780_s1.

Vlad:1994:SRG

[1891] Marcel Ovidiu Vlad. Stochastic renormalization group approach to random point processes. A fractal generalization of Poisson statistics. Journal of Mathematical Physics, 35(2):796–807, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/p796_s1.

Landi:1994:ROM

[1892] G. Landi, G. Marmo, and G. Vilasi. Recursion operators: Meaning and existence for completely integrable systems. Journal of Mathematical Physics, 35(2):808–815, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i2/p808_s1.

Liu:1994:NIS

[1893] Q. P. Liu. New integrable systems related to a Kupershmidt’s equation by Miura maps. Journal of Mathematical Physics, 35(2):816–820, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p816_s1.

Pickering:1994:WTC

[1894] Andrew Pickering. The Weiss–Tabor–Carnevale Painlevé test and Burgers’ hierarchy. Journal of Mathematical Physics, 35(2):821–833, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p821_s1.

Ragnisco:1994:IMS

[1895] O. Ragnisco and S. Rauch-Wojciechowski. Integrable mechanical systems related to the Harry–Dym hierarchy. Journal of Mathematical Physics, 35(2):834–847, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p834_s1.

Rourke:1994:HSSa

[1896] David E. Rourke and John K. Saunders. Half-solitons as solutions to the Zakharov–Shabat eigenvalue problem for rational reflection coefficient. II. Potentials on infinite support. Journal of Mathematical Physics, 35(2):848–872, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p848_s1.

Bona:1994:IGT

[1897] C. Bona and B. Coll. Isometry groups of three-dimensional Lorentzian metrics. Journal of Mathematical Physics, 35(2):873–884, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/
p873_s1.

Bonanno:1994:CFL

[1898] Alfio Bonanno and Vittorio Romano. Covariant flux-limited diffusion theory in Bianchi cosmologies. Journal of Mathematical Physics, 35(2): 885–898, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rourke:1994:HSSb

[1899] David E. Rourke and John K. Saunders. Half-solitons as solutions to the Zakharov–Shabat eigenvalue problem for rational reflection coefficient. II. Potentials on infinite support. Journal of Mathematical Physics, 35(2):885–898, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p885_s1.

Brodbeck:1994:SGY

[1900] Othmar Brodbeck and Norbert Straumann. Self-gravitating Yang–Mills solitons and their Chern–Simons numbers. Journal of Mathematical Physics, 35(2):899–919, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i2/p899_s1.

Castro:1994:SWZ

[1901] Carlos Castro. The N = 2 super-Wess–Zumino–Novikov–Witten model valued in superdiffeomorphism (SDIFF) M^{2} is self-dual supergravity in four dimensions. Journal of Mathematical Physics, 35(2):920–928, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p920_s1.

Foussats:1994:SCC

[1902] A. Foussats, C. Repetto, O. P. Zandron, and O. S. Zandron. Supermultiplets coupling in complex manifolds. Exterior canonical formalism. Journal of Mathematical Physics, 35(2):929–950, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/p929_s1.

Khorrami:1994:CST

[1903] M. Khorrami and R. Mansouri. Cylindrically symmetric thin walls in general relativity. Journal of Mathematical Physics, 35(2):951–958, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i2/p951_s1.

Bongaarts:1994:ACA

[1904] P. J. M. Bongaarts and H. G. J. Pijls. Almost commutative algebra and differential calculus on the quantum hyperplane. Journal of Mathematical Physics, 35(2):959–970, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i2/p959_s1.

Dabrowski:1994:DIO

[1905] L. Da̧browski, V. K. Dobrev, and R. Floreanini. q-difference intertwining operators for a Lorentz quantum algebra. Journal of Mathematical Physics, 35(2):971–985, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i2/p971_s1.

Daumens:1994:BFR

[1906] Michel Daumens. Boson fermion realizations for the super-rotation algebra. Journal of Mathematical Physics, 35(2):986–992, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/
p986_s1.

Huang:1994:CPO

[1907] Wen Jui Huang. On diff(S^{1}) covariantization of pseudodifferential operator. Journal of Mathematical Physics, 35(2):993–1007, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i2/
p993_s1.

Iguchi:1994:CNI

[1908] Kazumoto Iguchi. A class of new invariant surfaces under the trace maps for nary Fibonacci lattices. Journal of Mathematical Physics, 35(2):1008– 1019, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i2/p1008_s1.

Kaushik:1994:AST

[1909] Sumanth Kaushik. An analytical solution of the two-dimensional excitonphonon matrix element. Journal of Mathematical Physics, 35(2):1020– 1031, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i2/p1020_s1.

Krori:1994:ESS

[1910] K. D. Krori, P. Borgohain, and Dipali Das (Kar). Exact scalar and spinor solutions in the field of a stationary cosmic string. Journal of Mathematical Physics, 35(2):1032–1036, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i2/p1032_s1.

Yang:1994:FDI

[1911] Ya Ping Yang and Zu Rong Yu. Finite dimensional irreducible representations of the quantum algebra U_{q}(C_{2}). Journal of Mathematical Physics, 35(2):1037–1043, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p1037_s1.

Gabadadze:1994:NTA

[1912] G. T. Gabadadze and A. A. Pivovarov. Nonrenormalization theorem for the Abelian anomaly within dimensional regularization. Journal of Mathematical Physics, 35(3):1045–1053, March 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1045_s1.

Macfarlane:1994:ASP

[1913] A. J. Macfarlane. Algebraic structure of para-Bose Fock space. I. The Green’s ansatz revisited. Journal of Mathematical Physics, 35(3):1054– 1065, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i3/p1054_s1.

Markowich:1994:WFA

[1914] P. A. Markowich, N. J. Mauser, and F. Poupaud. A Wigner-function approach to (semi)classical limits: electrons in a periodic potential. Journal of Mathematical Physics, 35(3):1066–1094, March 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1066_s1.

Mostafazadeh:1994:SASa

[1915] Ali Mostafazadeh. Supersymmetry and the Atiyah–Singer index theorem. I. Peierls brackets, Green’s functions, and a proof of the index theorem via Gaussian superdeterminants. Journal of Mathematical Physics, 35(3):1095–1124, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i3/p1095_s1.

Mostafazadeh:1994:SASb

[1916] Ali Mostafazadeh. Supersymmetry and the Atiyah–Singer index theorem. II. The scalar curvature factor in the Schrödinger equation. Journal of Mathematical Physics, 35(3):1125–1138, March 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1125_s1.

Weiner:1994:CSA

[1917] B. Weiner, E. Deumens, and Y. Öhrn. Coherent state approach to electron nuclear dynamics with an antisymmetrized geminal power state. Journal of Mathematical Physics, 35(3):1139–1170, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1139_s1.

Shah:1994:TVP

[1918] P. A. Shah and N. S. Manton. Thermodynamics of vortices in the plane. Journal of Mathematical Physics, 35(3):1171–1184, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1171_s1.

Chetouani:1994:PID

[1919] L. Chetouani, L. Guechi, T. F. Hammann, and M. Letlout. Path integral for the damped harmonic oscillator coupled to its dual. Journal of Mathematical Physics, 35(3):1185–1191, March 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1185_s1.

Eilbeck:1994:EBA

[1920] J. C. Eilbeck and V. Z. Ènol′skiĭ. Elliptic Baker–Akhiezer functions and an application to an integrable dynamical system. Journal of Mathematical Physics, 35(3):1192–1201, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i3/p1192_s1.

Kaper:1994:GCA

[1921] Tasso J. Kaper and Gregor Kovačič. A geometric criterion for adiabatic chaos. Journal of Mathematical Physics, 35(3):1202–1218, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i3/p1202_s1.

Ranada:1994:ITP

[1922] Manuel F. Rañada. An integrable three-particle system. Journal of Mathematical Physics, 35(3):1219–1232, March 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1219_s1.

vanderLende:1994:STL

[1923] Epco van der Lende. Super-Toda lattices. Journal of Mathematical Physics, 35(3):1233–1251, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i3/p1233_s1.

Cianci:1994:ASI

[1924] Roberto Cianci and Andrew Khrennikov. p-adic superanalysis. III. The structure of p-adic super-Lie groups. Journal of Mathematical Physics, 35(3):1252–1259, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i3/p1252_s1.

Elizalde:1994:TSB

[1925] Emilio Elizalde and Klaus Kirsten. Topological symmetry breaking in self-interacting theories on toroidal space–time. Journal of Mathematical Physics, 35(3):1260–1273, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i3/p1260_s1.

Gowdy:1994:APT

[1926] Robert H. Gowdy. Affine projection tensor geometry: Decomposing the curvature tensor when the connection is arbitrary and the projection is tilted. Journal of Mathematical Physics, 35(3):1274–1301, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i3/p1274_s1.

Matos:1994:FDA

[1927] Tonatiuh Matos. Five-dimensional axisymmetric stationary solutions as harmonic maps. Journal of Mathematical Physics, 35(3):1302–1321, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i3/p1302_s1.

deAzcarraga:1994:CAR

[1928] J. A. de Azcárraga and D. Ellinas. Complex analytic realizations for quantum algebras. Journal of Mathematical Physics, 35(3):1322– 1333, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i3/p1322_s1.

Dresse:1994:BRS

[1929] Alain Dresse and Marc Henneaux. Becchi–Rouet–Stora–Tyutin structure of polynomial Poisson algebras. Journal of Mathematical Physics, 35(3):1334–1347, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i3/p1334_s1.

Elbert:1994:ILP

[1930] Árpád Elbert and Andrea Laforgia. An inequality for Legendre polynomials. Journal of Mathematical Physics, 35(3):1348–1360, March 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1348_
s1.

Gjaja:1994:CIP

[1931] I. Gjaja. Convergence of an infinite product of Lie transformations. Journal of Mathematical Physics, 35(3):1361–1371, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1361_s1.

Kaiser:1994:DGF

[1932] Gerald Kaiser. Deformations of Gabor frames. Journal of Mathematical Physics, 35(3):1372–1376, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i3/p1372_s1.

Links:1994:MLI

[1933] J. R. Links and R. B. Zhang. Multiparameter link invariants from quantum supergroups. Journal of Mathematical Physics, 35(3):1377– 1386, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i3/p1377_s1.

Maltsev:1994:NFA

[1934] N. E. Mal′tsev. New family of asymptotic solutions of Helmholtz equation. Journal of Mathematical Physics, 35(3):1387–1398, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/
p1387_s1.

Talebaoui:1994:ASS

[1935] Wathek Talebaoui. Algebraic spinor solution to the Kähler equation. Journal of Mathematical Physics, 35(3):1399–1407, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1399_s1.

Ungar:1994:ACLa

[1936] Abraham A. Ungar. The abstract complex Lorentz transformation group with real metric. I. Special relativity formalism to deal with the holomorphic automorphism group of the unit ball in any complex Hilbert space. Journal of Mathematical Physics, 35(3):1408–1426, March 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/p1408_

s1. See erratum [2113].

vandenHijligenberg:1994:DNP

[1937] N. van den Hijligenberg. Deformations of nilpotent parts of the Neveu– Schwarz and Ramond superalgebras. Journal of Mathematical Physics, 35(3):1427–1440, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i3/p1427_s1.

Dashen:1994:ENT

[1938] Roger Dashen and Daniel Wurmser. Erratum: “A new theory for scattering from a surface”, [J. Math. Phys. 32 (1991), no. 4, 971–985; MR 92e:78012]. Journal of Mathematical Physics, 35(3):1441, March 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i3/

p1441_s1. See [578].

Behncke:1994:FHW

[1939] H. Behncke. The m-function for Hamiltonians with Wigner–von Neumann potentials. Journal of Mathematical Physics, 35(4):1445–1462, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1445_s1.

Bohm:1994:RBB

[1940] A. Bohm and A. Mostafazadeh. The relation between the Berry and the Anandan–Aharonov connections for U(N) bundles. Journal of Mathematical Physics, 35(4):1463–1470, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1463_s1.

Campbell:1994:GSE

[1941] Joel Campbell. Ground state energy for the Hartree–Fock equations with Dirichlet boundary conditions. Journal of Mathematical Physics, 35(4):1471–1486, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i4/p1471_s1.

Chandler:1994:BQS

[1942] Colston Chandler and Archie G. Gibson. N-body quantum scattering theory in two Hilbert spaces. VII. Real-energy limits. Journal of Mathematical Physics, 35(4):1487–1512, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1487_s1.

Lange:1994:DWP

[1943] Horst Lange and P. F. Zweifel. Dissipation in Wigner–Poisson systems. Journal of Mathematical Physics, 35(4):1513–1521, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/p1513_s1.

Makarov:1994:QSC

[1944] Konstantin A. Makarov and Boris S. Pavlov. Quantum scattering on a Cantor bar. Journal of Mathematical Physics, 35(4):1522–1531, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1522_s1.

Mykytiuk:1994:GQN

[1945] I. V. Mykytiuk, A. K. Prykarpatsky, R. I. Andrushkiw, and V. Hr. Samoilenko. Geometric quantization of Neumann-type completely integrable Hamiltonian systems. Journal of Mathematical Physics, 35 (4):1532–1548, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i4/p1532_s1.

Nagel:1994:RHP

[1946] Bengt Nagel. The relativistic Hermite polynomial is a Gegenbauer polynomial. Journal of Mathematical Physics, 35(4):1549–1554, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/
p1549_s1.

Pulmannova:1994:QMQ

[1947] Sylvia Pulmannová. Quantum measurements and quantum logics. Journal of Mathematical Physics, 35(4):1555–1572, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/p1555_s1.

Tomaschitz:1994:CCV

[1948] Roman Tomaschitz. Cosmological CP violation. Journal of Mathematical Physics, 35(4):1573–1596, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i4/p1573_s1.

Tomaschitz:1994:CV

[1949] Roman Tomaschitz. Cosmological CP violation. Journal of Mathematical Physics, 35(4):1573–1596, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhevandrov:1994:SAB

[1950] Peter N. Zhevandrov. Semiclassical approximation for bound states of the Schrödinger equation with a Coulomb-like potential. Journal of Mathematical Physics, 35(4):1597–1621, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1597_s1.

Bao:1994:SHG

[1951] Gang Bao and David C. Dobson. Second harmonic generation in nonlinear optical films. Journal of Mathematical Physics, 35(4):1622–1633, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1622_s1.

Elshamy:1994:NRW

[1952] Maged Elshamy. Nonlinear random wave equations. Journal of Mathematical Physics, 35(4):1634–1643, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1634_s1.

Fatibene:1994:NFC

[1953] L. Fatibene, M. Ferraris, and M. Francaviglia. Nöther formalism for conserved quantities in classical gauge field theories. Journal of Mathematical Physics, 35(4):1644–1657, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1644_s1. See erratum [2310].

Shnerb:1994:GTP

[1954] N. Shnerb and L. P. Horwitz. On the group theory of the polarization states of a massless field. Journal of Mathematical Physics, 35(4):1658– 1667, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i4/p1658_s1.

vandenHeuvel:1994:EMC

[1955] B. M. van den Heuvel. Energy-momentum conservation in gauge theories. Journal of Mathematical Physics, 35(4):1668–1687, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/p1668_s1.

Helsing:1994:TOB

[1956] Johan Helsing. Third-order bounds on the conductivity of a random stacking of cubes. Journal of Mathematical Physics, 35(4):1688–1692, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1688_s1.

Blaszak:1994:GHH

[1957] Maciej Błaszak and Stefan Rauch-Wojciechowski. A generalized Hénon– Heiles system and related integrable Newton equations. Journal of Mathematical Physics, 35(4):1693–1709, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1693_s1.

Kalnins:1994:QCR

[1958] E. G. Kalnins, V. B. Kuznetsov, and Willard Miller, Jr. Quadrics on complex Riemannian spaces of constant curvature, separation of variables, and the Gaudin magnet. Journal of Mathematical Physics, 35(4):1710– 1731, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i4/p1710_s1.

Langer:1994:LGI

[1959] Joel Langer and Ron Perline. Local geometric invariants of integrable evolution equations. Journal of Mathematical Physics, 35(4):1732–1737, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1732_s1.

Leznov:1994:CSI

[1960] A. N. Leznov and A. V. Razumov. The canonical symmetry for integrable systems. Journal of Mathematical Physics, 35(4):1738–1754, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1738_s1.

Lou:1994:SAP

[1961] Sen Yue Lou. Symmetry algebras of the potential Nizhnik–Novikov– Veselov model. Journal of Mathematical Physics, 35(4):1755–1762, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1755_s1.

Prykarpatsky:1994:ASG

[1962] A. K. Prykarpatsky, V. Hr. Samoilenko, R. I. Andrushkiw, Yu. O. Mitropolsky, and M. M. Prytula. Algebraic structure of the gradientholonomic algorithm for Lax integrable nonlinear dynamical systems. I. Journal of Mathematical Physics, 35(4):1763–1777, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/p1763_s1.

Rondoni:1994:LTB

[1963] Lamberto Rondoni. Long time behavior of “X-led” reactions. Journal of Mathematical Physics, 35(4):1778–1795, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i4/p1778_s1.

Benn:1994:UDN

[1964] I. M. Benn. A unified description of null and non-null shear-free congruences. Journal of Mathematical Physics, 35(4):1796–1802, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/
p1796_s1.

Coleman:1994:WFS

[1965] Robert Alan Coleman and Herbert Korté. Why fundamental structures are first or second order. Journal of Mathematical Physics, 35(4):1803– 1818, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i4/p1803_s1.

Erler:1994:ACS

[1966] Jens Erler. Anomaly cancellation in six dimensions. Journal of Mathematical Physics, 35(4):1819–1833, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1819_s1.

Glass:1994:TNE

[1967] E. N. Glass and Mark Naber. Taub numbers extended to Einstein– Maxwell space–time. Journal of Mathematical Physics, 35(4):1834–1839, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1834_s1.

Svaiter:1994:STC

[1968] B. F. Svaiter and N. F. Svaiter. The stress tensor conformal anomaly and analytic regularizations. Journal of Mathematical Physics, 35(4):1840– 1849, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i4/p1840_s1.

Szydlowski:1994:CGB

[1969] Marek Szydłowski. Curvature of gravitationally bound mechanical systems. Journal of Mathematical Physics, 35(4):1850–1880, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/
p1850_s1.

Ungar:1994:ACLb

[1970] Abraham A. Ungar. The abstract complex Lorentz transformation group with real metric. II. The invariance of the form |t|^{2} − |x|^{2}. Journal of Mathematical Physics, 35(4):1881–1913, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i4/p1881_s1.

Bender:1994:DAS

[1971] Carl M. Bender and Stefan Boettcher. Determination of f(∞) from the asymptotic series for f(x) about x = 0. Journal of Mathematical Physics, 35(4):1914–1921, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i4/p1914_s1.

Berenstein:1994:EBM

[1972] D. E. Berenstein and L. F. Urrutia. The equivalence between the Mandelstam and the Cayley–Hamilton identities. Journal of Mathematical Physics, 35(4):1922–1930, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i4/p1922_s1.

Engeldinger:1994:SHA

[1973] Ralf A. Engeldinger and Achim Kempf. Sub-Hopf-algebra-induced twists of quantum enveloping algebras. Journal of Mathematical Physics, 35(4):1931–1938, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i4/p1931_s1.

Falomir:1994:CBW

[1974] H. Falomir, M. A. Muschietti, E. M. Santangelo, and J. Solomin. A calculation with a biorthogonal wavelet transformation. Journal of Mathematical Physics, 35(4):1939–1950, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i4/p1939_s1.

Kalnins:1994:MAR

[1975] E. G. Kalnins and Willard Miller, Jr. Models of q-algebra representations: q-integral transforms and “addition theorems”. Journal of Mathematical Physics, 35(4):1951–1975, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i4/p1951_s1.

Maslanka:1994:GDQ

[1976] Paweł Maślanka. The E_{q}(2) group via direct quantization of the Lie– Poisson structure and its Lie algebra. Journal of Mathematical Physics, 35(4):1976–1983, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i4/p1976_s1.

Parashar:1994:NDG

[1977] Preeti Parashar, V. S. Bhasin, and S. K. Soni. Noncommutative differential geometry on deformed quantum mechanical phase spaces. Journal of Mathematical Physics, 35(4):1984–1991, April 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i4/p1984_s1.

Romero:1994:NPE

[1978] Alfonso Romero and Miguel Sánchez. New properties and examples of incomplete Lorentzian tori. Journal of Mathematical Physics, 35(4):1992– 1997, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i4/p1992_s1.

Rosenhaus:1994:SCL

[1979] V. Rosenhaus and Gerald H. Katzin. On symmetries, conservation laws, and variational problems for partial differential equations. Journal of Mathematical Physics, 35(4):1998–2012, April 1994. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i4/p1998_s1.

Su:1994:HCM

[1980] Yu Cai Su. Harish-Chandra modules of the intermediate series over the high rank Virasoro algebras and high rank super-Virasoro algebras. Journal of Mathematical Physics, 35(4):2013–2023, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i4/p2013_s1.

Thron:1994:TSE

[1981] Chris Thron. Taylor series expansions for eigenvalues and eigenfunctions of parametrized composition operators. Journal of Mathematical Physics, 35(4):2024–2035, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i4/p2024_s1.

Chu:1994:PFB

[1982] Wen Chang Chu. Partial fractions and bilateral summations. Journal of Mathematical Physics, 35(4):2036–2042, April 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i4/p2036_s1. See er-

ratum [2313].

Braunss:1994:RHS

[1983] G. Braunss. On the regular Hilbert space representation of a Moyal quantization. Journal of Mathematical Physics, 35(5):2045–2056, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i5/p2045_s1.

Bytsenko:1994:ALS

[1984] A. A. Bytsenko, S. D. Odintsov, and S. Zerbini. Asymptotic level state density for parabosonic strings. Journal of Mathematical Physics, 35(5): 2057–2063, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2057_s1.

Clarkson:1994:EED

[1985] Eric Clarkson. Eigenvalues and eigenfunctions of the Dirac operator on spheres and pseudospheres. Journal of Mathematical Physics, 35(5):

2064–2073, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2064_s1.

Dolgallo:1994:GSQ

[1986] A. D. Dolgallo and K. N. Ilinski. Generalized supersymmetric quantum mechanics on Riemann surfaces with meromorphic superpotentials. Journal of Mathematical Physics, 35(5):2074–2082, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2074_s1.

FernandezC:1994:CQM

[1987] David J. Fernández C. and Bogdan Mielnik. Controlling quantum motion. Journal of Mathematical Physics, 35(5):2083–2104, May 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2083_
s1.

Fonarev:1994:WFQ

[1988] Oleg A. Fonarev. Wigner function and quantum kinetic theory in curved space–time and external fields. Journal of Mathematical Physics, 35(5): 2105–2129, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2105_s1.

Furlan:1994:ERC

[1989] Paolo Furlan, Yassen S. Stanev, and Ivan T. Todorov. Exchange relations and correlation functions for a quantum particle on the SU_{2} group manifold. Journal of Mathematical Physics, 35(5):2130–2141, May 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2130_
s1.

Grant:1994:SQM

[1990] Aaron K. Grant and Jonathan L. Rosner. Supersymmetric quantum mechanics and the Korteweg–de Vries hierarchy. Journal of Mathematical Physics, 35(5):2142–2156, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2142_s1.

Isham:1994:QLH

[1991] C. J. Isham. Quantum logic and the histories approach to quantum theory. Journal of Mathematical Physics, 35(5):2157–2185, May 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2157_
s1.

LaChapelle:1994:EGG

[1992] J. LaChapelle. U(2) as the electroweak gauge group. Journal of Mathematical Physics, 35(5):2186–2198, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i5/p2186_s1. See erratum [2311].

LaChapelle:1994:GMH

[1993] J. LaChapelle. Generating mass without the Higgs particle. Journal of Mathematical Physics, 35(5):2199–2209, May 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2199_s1. See er-

ratum [2312].

Lee:1994:NVP

[1994] Kuo Cheng Lee and Su Long Nyeo. A nonprincipal value prescription for the temporal gauge. Journal of Mathematical Physics, 35(5):2210– 2217, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2210_s1.

Liboff:1994:CSQ

[1995] Richard L. Liboff. Circular-sector quantum-billiard and allied configurations. Journal of Mathematical Physics, 35(5):2218–2228, May 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2218_
s1.

Miller:1994:NTS

[1996] H. G. Miller. Nonperturbative treatment of singular potentials. Journal of Mathematical Physics, 35(5):2229–2232, May 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2229_s1.

Overfelt:1994:PLW

[1997] P. L. Overfelt. On the possibility of localized wave propagation using a classical two-fluid model of superconductors. Journal of Mathematical Physics, 35(5):2233–2258, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2233_s1.

Rajeev:1994:SMC

[1998] S. G. Rajeev, S. Kalyana Rama, and Siddhartha Sen. Symplectic manifolds, coherent states, and semiclassical approximation. Journal of Mathematical Physics, 35(5):2259–2269, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i5/p2259_s1.

Sen:1994:GFA

[1999] Krishna Sen and Pratul Bandyopadhyay. A geometrical formulation of Abelian gauge structure in non-Abelian gauge theories and disconnected gauge group. Journal of Mathematical Physics, 35(5):2270–2281, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i5/p2270_s1.

Sogo:1994:ECS

[2000] Kiyoshi Sogo. Eigenstates of Calogero–Sutherland–Moser model and generalized Schur functions. Journal of Mathematical Physics, 35(5):2282– 2296, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2282_s1.

Trifonov:1994:GIS

[2001] D. A. Trifonov. Generalized intelligent states and squeezing. Journal of Mathematical Physics, 35(5):2297–2308, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i5/p2297_s1.

Vrbik:1994:DEC

[2002] Jan Vrbik. Dirac equation and Clifford algebra. Journal of Mathematical Physics, 35(5):2309–2314, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2309_s1.

He:1994:AGF

[2003] Sailing He, Vladimir G. Romanov, and Staffan Ström. Analysis of the Green’s function approach to one-dimensional inverse problems. II. Simultaneous reconstruction of two parameters. Journal of Mathematical Physics, 35(5):2315–2335, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2315_s1.

Lou:1994:ASD

[2004] Sen Yue Lou. Abundant symmetries for the (1 + 1)-dimensional classical Liouville field theory. Journal of Mathematical Physics, 35(5):2336– 2348, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2336_s1.

Turner:1994:COF

[2005] Leaf Turner, T. N. Tiouririne, and A. W. C. Lau. Compressional oscillation frequency of an anharmonic oscillator: The spherical non-neutral plasma. Journal of Mathematical Physics, 35(5):2349–2356, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/
p2349_s1.

Broidioi:1994:CEI

[2006] M. Broidioi. Critical exponents in the imperfect and the free Bose gas. Journal of Mathematical Physics, 35(5):2357–2370, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2357_s1.

Kostin:1994:DEE

[2007] M. D. Kostin. Differential equations for the entropy density, energy density, and number density of statistical mechanics. Journal of Mathematical Physics, 35(5):2371–2377, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i5/p2371_s1.

Haak:1994:GTS

[2008] Guido Haak. Group theoretical symmetries and generalized Bäcklund transformations for integrable systems. Journal of Mathematical Physics, 35(5):2378–2389, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i5/p2378_s1.

Lou:1994:SKE

[2009] Sen Yue Lou. Symmetries of the KdV equation and four hierarchies of the integrodifferential KdV equations. Journal of Mathematical Physics, 35(5):2390–2396, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i5/p2390_s1.

Morosi:1994:ETS

[2010] Carlo Morosi and Livio Pizzocchero. On the equivalence of two super Korteweg–de Vries theories: A bi-Hamiltonian viewpoint. Journal of Mathematical Physics, 35(5):2397–2407, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i5/p2397_s1.

Tanikawa:1994:SUM

[2011] Kiyotaka Tanikawa and Yoshihiro Yamaguchi. Stable and unstable manifolds in a zone of instability. Journal of Mathematical Physics, 35(5): 2408–2412, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2408_s1.

Wu:1994:OSS

[2012] Ya Ping Wu. Orbital stability of solitary waves of Zakharov system. Journal of Mathematical Physics, 35(5):2413–2422, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2413_s1.

Banerjee:1994:MPF

[2013] A. Banerjee, A. Sil, and S. Chatterjee. Motion of particles in fivedimensional Kaluza–Klein space–time. Journal of Mathematical Physics, 35(5):2423–2429, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i5/p2423_s1.

Christodoulakis:1994:ESD

[2014] T. Christodoulakis, T. Grammenos, E. Korfiatis, and V. C. Spanos. Exact solutions to the Dirac equation for neutrinos propagating in a particular Vaidya background. Journal of Mathematical Physics, 35(5):2430– 2437, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2430_s1.

Gergely:1994:KSMa

[2015] László Á. Gergely and Zoltán Perjés. Kerr–Schild metrics revisited. I. The ground state. Journal of Mathematical Physics, 35(5):2438–2447, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i5/p2438_s1.

Gergely:1994:KSMb

[2016] László Á. Gergely and Zoltán Perjés. Kerr–Schild metrics revisited. II. The complete vacuum solution. Journal of Mathematical Physics, 35(5): 2448–2462, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2448_s1.

Brandt:1994:BTZ

[2017] Friedemann Brandt. Bäcklund transformations and zero-curvature representations of systems of partial differential equations. Journal of Mathematical Physics, 35(5):2463–2484, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i5/p2463_s1.

Chung:1994:CSG

[2018] Won-Sang Chung. Comment on the solutions of the graded Yang–Baxter equation. Journal of Mathematical Physics, 35(5):2485–2489, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/
p2485_s1.

Chung:1994:TPD

[2019] Won-Sang Chung. Two parameter deformation of Virasoro algebra. Journal of Mathematical Physics, 35(5):2490–2496, May 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2490_s1.

Chung:1994:QGS

[2020] Won-Sang Chung. Quantum Z_{3}-graded spaces. Journal of Mathematical Physics, 35(5):2497–2504, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2497_s1.

Daumens:1994:RBF

[2021] Michel Daumens. Reducible boson fermion realizations for osp(1|2) and osp(3|2) superalgebras. Journal of Mathematical Physics, 35(5):2505– 2515, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2505_s1.

Duval:1994:SS

[2022] C. Duval and P. A. Horváthy. On Schrödinger superalgebras. Journal of Mathematical Physics, 35(5):2516–2538, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i5/p2516_s1.

Forrester:1994:ERU

[2023] P. J. Forrester. Exact results and universal asymptotics in the Laguerre random matrix ensemble. Journal of Mathematical Physics, 35(5):2539– 2551, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2539_s1.

Hakobyan:1994:UMQ

[2024] T. S. Hakobyan and A. G. Sedrakyan. Universal R matrix of U_{q}sl(n, m) quantum superalgebras. Journal of Mathematical Physics, 35(5):2552– 2559, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2552_s1.

Hlavaty:1994:QBG

[2025] Ladislav Hlavatý. Quantized braided groups. Journal of Mathematical Physics, 35(5):2560–2569, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2560_s1.

Huang:1994:SCS

[2026] Wen-Jui Huang. Superconformal covariantization of superdifferential operator on (1|1) superspace and classical N = 2 W superalgebras. Journal of Mathematical Physics, 35(5):2570–2582, May 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2570_s1.

Ky:1994:FDR

[2027] Nguyen Anh Ky. Finite-dimensional representations of the quantum superalgebra U_{q}[gl(2/2)]. I. Typical representations at generic q. Journal of Mathematical Physics, 35(5):2583–2606, May 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2583_s1.

Lukierski:1994:QDN

[2028] Jerzy Lukierski, Henri Ruegg, and Anatol Nowicki. Quantum deformations of nonsemisimple algebras: The example of D = 4 inhomogeneous rotations. Journal of Mathematical Physics, 35(5):2607–2616, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i5/p2607_s1.

Majid:1994:AQM

[2029] S. Majid. On the addition of quantum matrices. Journal of Mathematical Physics, 35(5):2617–2632, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i5/p2617_s1.

Marcinek:1994:NGA

[2030] Władysław Marcinek. Noncommutative geometry for arbitrary braidings. Journal of Mathematical Physics, 35(5):2633–2647, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i5/p2633_s1.

Paveri-Fontana:1994:HHH

[2031] Stefano L. Paveri-Fontana and Paul F. Zweifel. The half-Hartley and the half-Hilbert transform. Journal of Mathematical Physics, 35(5):2648– 2656, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2648_s1. See erratum [2265].

Sun:1994:RCA

[2032] Xiao Dong Sun and Shi Kun Wang. de Rham complexes of q-analog of general linear group GL_{q}(N). Journal of Mathematical Physics, 35(5): 2657–2686, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i5/p2657_s1.

Vourdas:1994:CSS

[2033] A. Vourdas. Coherent states on the m-sheeted complex plane. Journal of Mathematical Physics, 35(5):2687–2697, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i5/p2687_s1.

Crawford:1994:LAI

[2034] James P. Crawford. Local automorphism invariance: Gauge boson mass without a Higgs particle. Journal of Mathematical Physics, 35(6):2701– 2718, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p2701_s1.

SassolideBianchi:1994:LTZ

[2035] M. Sassoli de Bianchi. Levinson’s theorem, zero-energy resonances, and time delay in one-dimensional scattering systems. Journal of Mathematical Physics, 35(6):2719–2733, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i6/p2719_s1.

DeJonghe:1994:SDB

[2036] Frank De Jonghe. Schwinger–Dyson Becchi–Rouet–Stora–Tyutin symmetry and the Batalin–Vilkovisky Lagrangian quantization of gauge theories with open or reducible gauge algebras. Journal of Mathematical Physics, 35(6):2734–2742, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p2734_s1.

Horwitz:1994:SMSa

[2037] L. P. Horwitz. A soluble model for scattering and decay in quaternionic quantum mechanics. I. Decay. Journal of Mathematical Physics, 35(6): 2743–2760, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p2743_s1.

Horwitz:1994:SMSb

[2038] L. P. Horwitz. A soluble model for scattering and decay in quaternionic quantum mechanics. II. Scattering. Journal of Mathematical Physics, 35(6):2761–2771, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p2761_s1.

Kuznetsov:1994:DMR

[2039] Yu. A. Kuznetsov and M. S. Plyushchay. (2 + 1)-dimensional models of relativistic particles with curvature and torsion. Journal of Mathematical Physics, 35(6):2772–2784, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p2772_s1.

Okubo:1994:PLS

[2040] Susumu Okubo. Parastatistics as Lie-supertriple systems. Journal of Mathematical Physics, 35(6):2785–2803, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i6/p2785_s1.

Pillin:1994:DRW

[2041] Mathias Pillin. q-deformed relativistic wave equations. Journal of Mathematical Physics, 35(6):2804–2817, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i6/p2804_s1.

Roy:1994:TAW

[2042] Ashim Roy and Pratul Bandyopadhyay. Topological aspects of SU(2) Weyl fermion and global anomaly. Journal of Mathematical Physics, 35(6):2818–2827, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i6/p2818_s1.

Rybkin:1994:RCR

[2043] G. N. Rybkin and Yu. S. Vernov. Representations of commutation relations in BRST quantization. Journal of Mathematical Physics, 35(6): 2828–2832, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p2828_s1.

Wang:1994:SOD

[2044] Yu Peng Wang and Duo Jin. Study on the one dimensional bosonfermion model. Journal of Mathematical Physics, 35(6):2833–2837, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i6/p2833_s1.

Wehrhahn:1994:SSA

[2045] R. F. Wehrhahn and A. O. Barut. Symmetry scattering for SU(2, 2) with applications. Journal of Mathematical Physics, 35(6):2838–2855, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i6/p2838_s1.

Llosa:1994:HFN

[2046] J. Llosa and J. Vives. Hamiltonian formalism for nonlocal Lagrangians. Journal of Mathematical Physics, 35(6):2856–2877, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p2856_s1.

Ali:1994:JCR

[2047] G. Alì and V. Romano. Jump conditions for a radiating relativistic gas. Journal of Mathematical Physics, 35(6):2878–2901, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p2878_s1.

Guil:1994:TDI

[2048] F. Guil and M. Mañas. Two-dimensional integrable systems and selfdual Yang–Mills equations. Journal of Mathematical Physics, 35(6): 2902–2913, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p2902_s1.

Iwai:1994:TCD

[2049] Toshihiro Iwai and Noriaki Katayama. Two classes of dynamical systems all of whose bounded trajectories are closed. Journal of Mathematical Physics, 35(6):2914–2933, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p2914_s1.

Kitaev:1994:CIS

[2050] A. V. Kitaev. Caustics in 1 + 1 integrable systems. Journal of Mathematical Physics, 35(6):2934–2954, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i6/p2934_s1.

Matveev:1994:AMS

[2051] Vladimir B. Matveev. Asymptotics of the multipositon-soliton τ function of the Korteweg–de Vries equation and the supertransparency. Journal of Mathematical Physics, 35(6):2955–2970, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i6/p2955_s1.

Qiao:1994:HNE

[2052] Zhi Jun Qiao. A hierarchy of nonlinear evolution equations and finitedimensional involutive systems. Journal of Mathematical Physics, 35(6): 2971–2977, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p2971_s1.

Qiao:1994:ISI

[2053] Zhi Jun Qiao. An involutive system and integrable C. Neumann system associated with the modified Korteweg–de Vries hierarchy. Journal of Mathematical Physics, 35(6):2978–2982, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i6/p2978_s1.

vanDiejen:1994:IDC

[2054] J. F. van Diejen. Integrability of difference Calogero–Moser systems. Journal of Mathematical Physics, 35(6):2983–3004, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p2983_s1.

Amir:1994:RCS

[2055] M. Jamil Amir, Ashfaque H. Bokhari, and Asghar Qadir. Ricci collineations of static spherically symmetric spacetimes. Journal of Mathematical Physics, 35(6):3005–3012, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i6/p3005_s1. See erratum [2317]

and comment [2316].

Castro:1994:GSY

[2056] Carlos Castro. W gravity, N = 2 strings, and 2 + 2 SU ∗ (∞) Yang– Mills instantons. Journal of Mathematical Physics, 35(6):3013–3024, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i6/p3013_s1.

Chee:1994:NSA

[2057] Gou Ying Chee. Note on a solution of the algebraically special perturbation of the Schwarzschild black hole. Journal of Mathematical Physics, 35(6):3025–3028, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p3025_s1.

Goldblatt:1994:STG

[2058] E. Goldblatt. Symmetries of type D^{+}D^{−} gravitational instantons. Journal of Mathematical Physics, 35(6):3029–3042, June 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i6/p3029_s1.

Hortacsu:1994:FRI

[2059] M. Hortaçsu and R. Kaya. Fluctuations of a rotating impulsive wave. Journal of Mathematical Physics, 35(6):3043–3050, June 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p3043_

s1. See errata [2318].

TorresdelCastillo:1994:GPT

[2060] G. F. Torres del Castillo. Gravitational perturbations of type-D vacuum space–times with cosmological constant. Journal of Mathematical Physics, 35(6):3051–3058, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p3051_s1.

Antoine:1994:STA

[2061] J.-P. Antoine, A. Inoue, and C. Trapani. Spatial theory of ^{∗}- automorphisms of partial O^{∗}-algebras. Journal of Mathematical Physics, 35(6):3059–3073, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i6/p3059_s1.

Arik:1994:BGR

[2062] M. Arik, F. Aydin, E. Hizel, J. Kornfilt, and A. Yildiz. Braid group related algebras, their representations and generalized hydrogenlike spectra. Journal of Mathematical Physics, 35(6):3074–3088, June 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p3074_
s1.

Brihaye:1994:QSM

[2063] Y. Brihaye and P. Kosiński. Quasiexactly solvable 2×2 matrix equations. Journal of Mathematical Physics, 35(6):3089–3098, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p3089_s1.

Chen:1994:NMI

[2064] Nan Xian Chen. A new method to introduce the Möbius function and the Möbius transform. Journal of Mathematical Physics, 35(6):3099– 3108, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p3099_s1.

Cognola:1994:THK

[2065] Guido Cognola and Luciano Vanzo. The trace of the heat kernel on a compact hyperbolic 3-orbifold. Journal of Mathematical Physics, 35(6): 3109–3116, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p3109_s1.

Fokas:1994:QPA

[2066] A. S. Fokas and I. M. Gel′fand. Quadratic Poisson algebras and their infinite-dimensional extensions. Journal of Mathematical Physics, 35(6): 3117–3131, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p3117_s1.

Minnaert:1994:CGC

[2067] Pierre Minnaert and Marek Mozrzymas. Clebsch–Gordan coefficients for the quantum superalgebra U_{q}(osp(1|2)). Journal of Mathematical Physics, 35(6):3132–3145, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p3132_s1.

Pecina-Cruz:1994:ACI

[2068] J. N. Pecina-Cruz. An algorithm to calculate the invariants of any Lie algebra. Journal of Mathematical Physics, 35(6):3146–3162, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i6/p3146_s1.

Rowe:1994:VCS

[2069] D. J. Rowe. Vector-coherent-state representations of so(5) in an so(3) basis. Journal of Mathematical Physics, 35(6):3163–3177, June 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/p3163_
s1.

Rowe:1994:RFS

[2070] D. J. Rowe. Representation of the five- and six-dimensional harmonic oscillators in a u(5) ⊃ so(5) ⊃ so(3) basis. Journal of Mathematical Physics, 35(6):3178–3189, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p3178_s1.

Saidi:1994:GDA

[2071] E. H. Saidi and M. B. Sedra. On the Gel′fand–Dickey algebra GD(SL_{n}) and the W_{n}-symmetry. I. The bosonic case. Journal of Mathematical Physics, 35(6):3190–3210, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p3190_s1.

Schmudgen:1994:ORR

[2072] Konrad Schmüdgen. Operator representations of the real twisted canonical commutation relations. Journal of Mathematical Physics, 35(6):3211– 3229, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i6/p3211_s1.

Singh:1994:GTS

[2073] S. N. Singh and G. S. Singh. Group-theoretic solution of the scalar wave equation in a multiply connected domain. Journal of Mathematical Physics, 35(6):3230–3249, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i6/p3230_s1.

Atakishiyev:1994:SDR

[2074] Natig M. Atakishiyev, Alejandro Frank, and Kurt Bernardo Wolf. A simple difference realization of the Heisenberg q-algebra. Journal of Mathematical Physics, 35(7):3253–3260, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i7/p3253_s1.

Brzezinski:1994:IMR

[2075] T. Brzeziński and A. J. Macfarlane. On integrable models related to the osp(1, 2) Gaudin algebra. Journal of Mathematical Physics, 35(7): 3261–3272, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3261_s1.

Chan:1994:DCP

[2076] F. L. Chan and R. J. Finkelstein. q-deformation of the Coulomb problem. Journal of Mathematical Physics, 35(7):3273–3284, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i7/p3273_s1.

Clemence:1994:CMP

[2077] Dominic P. Clemence and Martin Klaus. Continuity of the S matrix for the perturbed Hill’s equation. Journal of Mathematical Physics, 35(7): 3285–3300, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3285_s1.

Cordero:1994:ASN

[2078] Patricio Cordero and Sebastián Salamó. Algebraic solution for the Natanzon hypergeometric potentials. Journal of Mathematical Physics, 35(7): 3301–3307, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3301_s1.

Elizalde:1994:VED

[2079] E. Elizalde. The vacuum energy density for spherical and cylindrical universes. Journal of Mathematical Physics, 35(7):3308–3321, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i7/
p3308_s1.

Goldin:1994:NSE

[2080] Gerald A. Goldin and George Svetlichny. Nonlinear Schrödinger equations and the separation property. Journal of Mathematical Physics, 35(7):3322–3332, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i7/p3322_s1.

Good:1994:MSF

[2081] R. H. Good, Jr. Mass spectra from field equations. Journal of Mathematical Physics, 35(7):3333–3339, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i7/p3333_s1. See comments [2315].

Gracia-Bondia:1994:QEE

[2082] José M. Gracia-Bondía and Joseph C. Várilly. Quantum electrodynamics in external fields from the spin representation. Journal of Mathematical Physics, 35(7):3340–3367, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i7/p3340_s1.

Juriev:1994:ASQ

[2083] Denis Juriev. Algebraic structures of quantum projective field theory related to fusion and braiding. Hidden additive weight. Journal of Mathematical Physics, 35(7):3368–3379, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i7/p3368_s1.

Rivas:1994:QGS

[2084] Martín Rivas. Quantization of generalized spinning particles: New derivation of Dirac’s equation. Journal of Mathematical Physics, 35(7): 3380–3399, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3380_s1.

Zhu:1994:CLU

[2085] Chengjun Zhu and John R. Klauder. The classical limit of ultralocal scalar fields. Journal of Mathematical Physics, 35(7):3400–3409, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3400_s1.

Chinea:1994:CAT

[2086] Domingo Chinea, Manuel de León, and Juan C. Marrero. The constraint algorithm for time-dependent Lagrangians. Journal of Mathematical Physics, 35(7):3410–3447, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i7/p3410_s1.

Christov:1994:KCS

[2087] Ognyan Christov. On the Kolmogorov’s condition for a special case of the Kirchhoff top. Journal of Mathematical Physics, 35(7):3448–3463, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3448_s1.

Gadylshin:1994:AHR

[2088] Rustem R. Gadyl′shin. On acoustic Helmholtz resonator and on its electromagnetic analog. Journal of Mathematical Physics, 35(7):3464–3481, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3464_s1.

Gaftoi:1994:LDE

[2089] Violeta Gaftoi, José L. López Bonilla, Jesús Morales, and Marco A. Rosales. The Lorentz–Dirac equation in Minkowski space. Journal of Mathematical Physics, 35(7):3482–3489, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i7/p3482_s1.

Zayed:1994:HSR

[2090] E. M. E. Zayed and A. I. Younis. On hearing the shape of rectilinear regions. Journal of Mathematical Physics, 35(7):3490–3496, July 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i7/p3490_
s1.

Barcelos-Neto:1994:SAD

[2091] J. Barcelos-Neto and N. R. F. Braga. Symplectic analysis of a Dirac constrained theory. Journal of Mathematical Physics, 35(7):3497–3503, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3497_s1.

Leon:1994:NSR

[2092] Jérôme Leon. Nonlinear systems related to an arbitrary space–time dependence of the spectral transform. Journal of Mathematical Physics, 35(7):3504–3524, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i7/p3504_s1.

Barvinsky:1994:BNC

[2093] A. O. Barvinsky, Yu. V. Gusev, G. A. Vilkovisky, and V. V. Zhytnikov. The basis of nonlocal curvature invariants in quantum gravity theory. Third order. Journal of Mathematical Physics, 35(7):3525–3542, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3525_s1.

Barvinsky:1994:ABH

[2094] A. O. Barvinsky, Yu. V. Gusev, G. A. Vilkovisky, and V. V. Zhytnikov. Asymptotic behaviors of the heat kernel in covariant perturbation theory. Journal of Mathematical Physics, 35(7):3543–3559, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i7/p3543_s1.

Carot:1994:STA

[2095] J. Carot, L. Mas, and A. M. Sintes. Space–times admitting a threeparameter similarity group. Journal of Mathematical Physics, 35(7): 3560–3570, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3560_s1.

Drechsler:1994:RIM

[2096] W. Drechsler and D. Hartley. The role of the internal metric in generalized Kaluza–Klein theories. Journal of Mathematical Physics, 35(7): 3571–3586, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3571_s1.

Haddow:1994:CRT

[2097] Barry M. Haddow. Characterization of Riemann tensors using Riccitype equations. Journal of Mathematical Physics, 35(7):3587–3593, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3587_s1.

Huang:1994:TIB

[2098] Wung Hong Huang. Thermodynamics of ideal boson and fermion gases in the static Taub universe. Journal of Mathematical Physics, 35(7): 3594–3603, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3594_s1.

Lemos:1994:PAW

[2099] José P. S. Lemos and Ozemar S. Ventura. Planar and axisymmetric walls in general relativity, a comparison. Journal of Mathematical Physics, 35(7):3604–3611, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i7/p3604_s1.

Basu:1994:HSA

[2100] Debabrata Basu. Hilbert spaces of analytic functions for the SL(2, R) coherent states and a class of associated integral transforms. Journal of Mathematical Physics, 35(7):3612–3623, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i7/p3612_s1.

Borisov:1994:SSW

[2101] Victor V. Borisov and Andrei B. Utkin. Some solutions of the wave and Maxwell’s equations. Journal of Mathematical Physics, 35(7):3624– 3630, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3624_s1.

Chung:1994:NDB

[2102] Won-Sang Chung. New deformed boson algebra. Journal of Mathematical Physics, 35(7):3631–3635, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i7/p3631_s1.

Dattoli:1994:TTI

[2103] G. Dattoli, S. Lorenzutta, G. Maino, A. Torre, G. Voykov, and C. Chiccoli. Theory of two-index Bessel functions and applications to physical problems. Journal of Mathematical Physics, 35(7):3636–3649, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3636_s1.

Daumens:1994:SEG

[2104] Michel Daumens. The superalgebra embedding of osp(1|2) in sl(1|2) and the grade star representations of sl(1|2). Journal of Mathematical Physics, 35(7):3650–3658, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i7/p3650_s1.

Fusco:1994:MDE

[2105] D. Fusco and N. Manganaro. A method for determining exact solutions to a class of nonlinear models based on introduction of differential constraints. Journal of Mathematical Physics, 35(7):3659–3669, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i7/p3659_s1.

Gavrilik:1994:RDA

[2106] A. M. Gavrilik and A. U. Klimyk. Representations of the q-deformed algebras U_{q}(so_{2,1}) and U_{q}(so_{3,1}). Journal of Mathematical Physics, 35(7): 3670–3686, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3670_s1.

Gerdjikov:1994:CEC

[2107] V. S. Gerdjikov and A. B. Yanovski. Completeness of the eigenfunctions for the Caudrey–Beals–Coifman system. Journal of Mathematical Physics, 35(7):3687–3725, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i7/p3687_s1.

Gielerak:1994:RKM

[2108] R. Gielerak, L. Jakóbczyk, and R. Olkiewicz. Reconstruction of Kubo– Martin–Schwinger structure from Euclidean Green functions. Journal of Mathematical Physics, 35(7):3726–3744, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i7/p3726_s1.

Holschneider:1994:FCU

[2109] Matthias Holschneider. Functional calculus using wavelet transforms. Journal of Mathematical Physics, 35(7):3745–3752, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i7/p3745_s1.

Majid:1994:RWH

[2110] S. Majid and M. J. Rodríguez-Plaza. Random walk and the heat equation on superspace and anyspace. Journal of Mathematical Physics, 35(7): 3753–3760, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3753_s1.

Zakkari:1994:AAT

[2111] M. Zakkari. Affine algebras on the torus. Journal of Mathematical Physics, 35(7):3761–3768, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i7/p3761_s1.

Lammerzahl:1994:APO

[2112] Claus Lämmerzahl. Addendum to: “The pseudodifferential operator square root of the Klein–Gordon equation”, [J. Math. Phys. 34 (1993), no. 9, 3918–3932; MR 94g:81127]. Journal of Mathematical Physics, 35 (7):3769, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i7/p3769_s1. See [1682].

Ungar:1994:EAC

[2113] Abraham A. Ungar. Erratum: “The abstract complex Lorentz transformation group with real metric. I. Special relativity formalism to deal with the holomorphic automorphism group of the unit ball in any complex Hilbert space”, [J. Math. Phys. 35, 1408-1426 (1994)]. Journal of Mathematical Physics, 35(7):3770, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i7/p3770_s1. See [1936].

DeBievre:1994:QNO

[2114] Stephan De Bièvre and Jacques Renaud. Quantization of the nilpotent orbits in so(1, 2)∗ and massless particles on (anti-)de Sitter space–time. Journal of Mathematical Physics, 35(8):3775–3793, August 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i8/p3775_
s1.

Fairlie:1994:CQD

[2115] David B. Fairlie and Jean Nuyts. On certain quantum deformations of gl(N, R). Journal of Mathematical Physics, 35(8):3794–3804, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i8/p3794_s1.

Grosse:1994:OSE

[2116] H. Grosse, A. Martin, and J. Stubbe. Order and spacings of energy levels for the Klein–Gordon equation. Journal of Mathematical Physics, 35(8):3805–3816, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p3805_s1.

Gudder:1994:NSQ

[2117] Stanley Gudder. Nonstandard scalar quantum fields. Journal of Mathematical Physics, 35(8):3817–3844, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i8/p3817_s1.

Gupta:1994:YMT

[2118] K. S. Gupta, R. J. Henderson, S. G. Rajeev, and O. T. Turgut. Yang– Mills theory on a cylinder coupled to point particles. Journal of Mathematical Physics, 35(8):3845–3865, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i8/p3845_s1.

Holovatch:1994:CTP

[2119] Yu. Holovatch and T. Krokhmal′skiĭ. Compilation of two-point and fourpoint graphs in field theory in noninteger dimensions. Journal of Mathematical Physics, 35(8):3866–3880, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i8/p3866_s1.

Liboff:1994:NSC

[2120] Richard L. Liboff. Nodal-surface conjectures for the convex quantum billiard. Journal of Mathematical Physics, 35(8):3881–3888, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i8/
p3881_s1.

Martinez:1994:PEE

[2121] André Martinez. Precise exponential estimates in adiabatic theory. Journal of Mathematical Physics, 35(8):3889–3915, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i8/p3889_s1.

Morales:1994:DIE

[2122] Daniel A. Morales. An n-dimensional interpretation of the δ-expansion for the Thomas–Fermi equation. Journal of Mathematical Physics, 35 (8):3916–3921, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p3916_s1.

Rylov:1994:SCQ

[2123] Yuri Rylov. Statistical conception of quantum field theory. Journal of Mathematical Physics, 35(8):3922–3935, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i8/p3922_s1.

Selesnick:1994:DEQ

[2124] S. A. Selesnick. Dirac’s equation on the quantum net. Journal of Mathematical Physics, 35(8):3936–3958, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i8/p3936_s1.

Kim:1994:PSS

[2125] Myung Ho Kim and Phillial Oh. Phase space structure of non-Abelian Chern–Simmons particles. Journal of Mathematical Physics, 35(8):3959– 3968, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i8/p3959_s1.

Ola:1994:BIE

[2126] Petri Ola. Boundary integral equations for the scattering of electromagnetic waves by a homogeneous chiral obstacle. Journal of Mathematical Physics, 35(8):3969–3980, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i8/p3969_s1.

Yu:1994:RTE

[2127] Shou Mian Yu, Tian Yu, and Yinbin Liu. Resonance theory of elastic head waves in layered media. I. A fluid layer bottomed by an infinite solid. Journal of Mathematical Physics, 35(8):3981–3997, August 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i8/p3981_
s1.

Curely:1994:ICS

[2128] Jacques Curély and Roland Georges. Isotropic classical spin chains showing nearest neighbor correlated cationic distributions. Journal of Mathematical Physics, 35(8):3998–4004, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i8/p3998_s1.

Albeverio:1994:LTB

[2129] S. Albeverio and A. Klar. Long time behavior of nonlinear stochastic oscillators: The one-dimensional Hamiltonian case. Journal of Mathematical Physics, 35(8):4005–4027, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i8/p4005_s1.

Auberson:1994:CHN

[2130] G. Auberson and P. C. Sabatier. On a class of homogeneous nonlinear Schrödinger equations. Journal of Mathematical Physics, 35(8):4028– 4040, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i8/p4028_s1.

Berest:1994:DPH

[2131] Yuri Yu. Berest. Deformations preserving Huygens’ principle. Journal of Mathematical Physics, 35(8):4041–4056, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i8/p4041_s1.

Chow:1994:LSD

[2132] Kwok W. Chow. Lattice solitons directly by the bilinear method. Journal of Mathematical Physics, 35(8):4057–4066, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i8/p4057_s1.

Leznov:1994:HPC

[2133] A. N. Leznov and A. V. Razumov. Hamiltonian properties of the canonical symmetry. Journal of Mathematical Physics, 35(8):4067–4087, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i8/p4067_s1.

Prykarpatskyj:1994:ASG

[2134] A. K. Prykarpatskyj, V. Hr. Samoilenko, and R. I. Andrushkiw. Algebraic structure of the gradient-holonomic algorithm for Lax integrable nonlinear dynamical systems. II. The reduction via Dirac and canonical quantization procedure. Journal of Mathematical Physics, 35(8):4088– 4116, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i8/p4088_s1.

Coley:1994:QAD

[2135] A. A. Coley and R. J. van den Hoogen. Qualitative analysis of diagonal Bianchi type V imperfect fluid cosmological models. Journal of Mathematical Physics, 35(8):4117–4144, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i8/p4117_s1.

Dunn:1994:ESK

[2136] K. A. Dunn, Tina A. Harriott, and J. G. Williams. Extension of the de Sitter kink. Journal of Mathematical Physics, 35(8):4145–4156, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i8/p4145_s1.

Geroch:1994:TMM

[2137] Robert Geroch and Shyan-Ming Perng. Total mass-momentum of arbitrary initial-data sets in general relativity. Journal of Mathematical Physics, 35(8):4157–4177, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v35/i8/p4157_s1.

Glass:1994:GMA

[2138] E. N. Glass and M. G. Naber. Gravitational mass anomaly. Journal of Mathematical Physics, 35(8):4178–4183, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i8/p4178_s1.

Gomez:1994:NCE

[2139] R. Gómez, P. Papadopoulos, and J. Winicour. Null cone evolution of axisymmetric vacuum space–times. Journal of Mathematical Physics, 35(8):4184–4204, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p4184_s1.

Ali:1994:CFW

[2140] S. Twareque Ali and L. V. Denisov. A composite Fourier-wavelet transform and square-integrable group representations. Journal of Mathematical Physics, 35(8):4205–4216, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i8/p4205_s1.

Camporesi:1994:SFZ

[2141] Roberto Camporesi and Atsushi Higuchi. Spectral functions and zeta functions in hyperbolic spaces. Journal of Mathematical Physics, 35 (8):4217–4246, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p4217_s1.

Chakrabarti:1994:CCB

[2142] A. Chakrabarti. SO(5)_{q} and contraction: Chevalley basis representations for q-generic and root of unity. Journal of Mathematical Physics, 35(8):4247–4267, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p4247_s1.

Chung:1994:GBF

[2143] Won-Sang Chung and Hye-Jung Kang. The q-gamma, q-beta functions, and q-multiplication formula. Journal of Mathematical Physics, 35(8):4268–4276, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p4268_s1.

Dodonov:1994:NRT

[2144] V. V. Dodonov and V. I. Man′ko. New relations for two-dimensional Hermite polynomials. Journal of Mathematical Physics, 35(8):4277– 4294, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i8/p4277_s1.

Fan:1994:HAT

[2145] Qi Hong Fan. Harmonic analysis on tube-type affine homogeneous phase spaces. Journal of Mathematical Physics, 35(8):4295–4302, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i8/
p4295_s1.

Hudetz:1994:TEA

[2146] Thomas Hudetz. Topological entropy for appropriately approximated C^{∗}-algebras. Journal of Mathematical Physics, 35(8):4303–4333, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i8/p4303_s1.

Nagy:1994:BHH

[2147] Gabriel B. Nagy, Omar E. Ortiz, and Oscar A. Reula. The behavior of hyperbolic heat equations’ solutions near their parabolic limits. Journal of Mathematical Physics, 35(8):4334–4356, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i8/p4334_s1.

Rezende:1994:FIF

[2148] Jorge Rezende. Feynman integrals and Fredholm determinants. Journal of Mathematical Physics, 35(8):4357–4371, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i8/p4357_s1.

Ugrinovsky:1994:MEA

[2149] Roman Ugrinovsky. Maximum entropy algorithm for spectral estimation problem with a priori information. Journal of Mathematical Physics, 35(8):4372–4382, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p4372_s1.

VanderJeugt:1994:ABR

[2150] J. Van der Jeugt. Adjoint boson realization of SU(N) and a family of structural zeros of 6j coefficients. Journal of Mathematical Physics, 35(8):4383–4390, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i8/p4383_s1.

Yang:1994:RCQ

[2151] Ya Ping Yang and Zu Rong Yu. Reduced coefficients of quantum algebra SU_{q}(4). Journal of Mathematical Physics, 35(8):4391–4401, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i8/p4391_s1.

Ali:1994:QCS

[2152] S. Twareque Ali and Ute A. Mueller. Quantization of a classical system on a coadjoint orbit of the Poincaré group in 1 + 1 dimensions. Journal of Mathematical Physics, 35(9):4405–4422, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4405_s1.

Aptekarev:1994:SEC

[2153] A. I. Aptekarev, J. S. Dehesa, and R. J. Yáñez. Spatial entropy of central potentials and strong asymptotics of orthogonal polynomials. Journal of Mathematical Physics, 35(9):4423–4428, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4423_s1.

Bollini:1994:STF

[2154] C. G. Bollini, L. E. Oxman, and M. Rocca. Space of test functions for higher-order field theories. Journal of Mathematical Physics, 35(9):4429– 4438, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4429_s1.

Borsari:1994:SOQ

[2155] I. Borsari and S. Graffi. Symbols of operators and quantum evolution. Journal of Mathematical Physics, 35(9):4439–4450, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4439_
s1.

Dattoli:1994:PSD

[2156] G. Dattoli, S. Lorenzutta, G. Maino, and A. Torre. Phase-space dynamics and Hermite polynomials of two variables and two indices. Journal of Mathematical Physics, 35(9):4451–4462, September 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4451_s1.

Duan:1994:GPD

[2157] Yi Shi Duan, Sheng Li Zhang, and Sze Shiang Feng. Gauge potential decomposition, space–time defects, and Planck’s constant. Journal of Mathematical Physics, 35(9):4463–4468, September 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4463_s1.

Ferretti:1994:UMP

[2158] G. Ferretti. The untruncated Marinari–Parisi superstring. Journal of Mathematical Physics, 35(9):4469–4476, September 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4469_s1.

Kachkachi:1994:RQL

[2159] H. Kachkachi and M. Kachkachi. Redundancy of the quantum level gauge condition. Journal of Mathematical Physics, 35(9):4477–4482, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4477_s1.

Kempf:1994:URQ

[2160] Achim Kempf. Uncertainty relation in quantum mechanics with quantum group symmetry. Journal of Mathematical Physics, 35(9):4483–4496, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4483_s1.

Lu:1994:ESE

[2161] Jing Fa Lü and S. I. Kruglov. The exact solutions of the equation for a composite particle in the external electromagnetic fields. Journal of Mathematical Physics, 35(9):4497–4504, September 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4497_s1.

Morato:1994:CBQ

[2162] L. M. Morato and S. Ugolini. A connection between quantum dynamics and approximation of Markov diffusions. Journal of Mathematical Physics, 35(9):4505–4516, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i9/p4505_s1.

Nedjadi:1994:SCF

[2163] Y. Nedjadi and R. C. Barrett. Solution of the central field problem for a Duffin–Kemmer–Petiau vector boson. Journal of Mathematical Physics, 35(9):4517–4533, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4517_s1.

Nistico:1994:KAN

[2164] Giuseppe Nisticò and Maria Concetta Romania. Knowledge about noncommuting quantum observables by means of Einstein–Podolsky–Rosen correlations. Journal of Mathematical Physics, 35(9):4534–4546, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p4534_s1.

Yasui:1994:UGT

[2165] Yukinori Yasui and Satoshi Takahashi. Unstable geodesics and topological field theory. Journal of Mathematical Physics, 35(9):4547–4567, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4547_s1.

Debecki:1994:ILT

[2166] Jacek Debecki, Jacek Gancarzewicz, Manuel de León, and Włodzimierz Mikulski. Invariants of Lagrangians and their classifications. Journal of Mathematical Physics, 35(9):4568–4593, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4568_s1.

Molino:1994:EWD

[2167] F. R. Molino and P. C. Sabatier. Elastic waves in discontinuous media: Three-dimensional scattering. Journal of Mathematical Physics, 35 (9):4594–4636, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4594_s1.

Karwowski:1994:HSA

[2168] Witold Karwowski and R. Vilela Mendes. Hierarchical structures and asymmetric stochastic processes on p-adics and adèles. Journal of Mathematical Physics, 35(9):4637–4650, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i9/p4637_s1.

Abdelwahid:1994:ETV

[2169] F. Abdelwahid and J. Burzlaff. Existence theorems for 90^{∘} vortex-vortex scattering. Journal of Mathematical Physics, 35(9):4651–4660, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p4651_s1.

Blaszak:1994:MAL

[2170] Maciej Błaszak and Krzysztof Marciniak. R-matrix approach to lattice integrable systems. Journal of Mathematical Physics, 35(9):4661–4682, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4661_s1.

Boiti:1994:PSK

[2171] M. Boiti, F. Pempinelli, and A. Pogrebkov. Properties of solutions of the Kadomtsev–Petviashvili I equation. Journal of Mathematical Physics, 35(9):4683–4718, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4683_s1.

Buyske:1994:BTL

[2172] Steven G. Buyske. Bäcklund transformations of linear Weingarten surfaces in Minkowski three-space. Journal of Mathematical Physics, 35 (9):4719–4724, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4719_s1.

Chen:1994:PCK

[2173] Deng Yuan Chen, Ning Sheng Zhu, and Min Zhu. The potential constraints of the KP system and the corresponding Hamiltonian equations. Journal of Mathematical Physics, 35(9):4725–4738, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4725_
s1.

Hu:1994:BTN

[2174] Xing Biao Hu. A Bäcklund transformation and nonlinear superposition formula of a modified Korteweg–de Vries-type bilinear equation. Journal of Mathematical Physics, 35(9):4739–4745, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4739_s1.

Radha:1994:SAL

[2175] R. Radha and M. Lakshmanan. Singularity analysis and localized coherent structures in (2 + 1)-dimensional generalized Korteweg–de Vries equations. Journal of Mathematical Physics, 35(9):4746–4756, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p4746_s1.

Romeiras:1994:DTW

[2176] Filipe J. Romeiras. Double three-wave interaction of four waves: Lax representations and exact solutions. Journal of Mathematical Physics, 35(9):4757–4778, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4757_s1.

Shivamoggi:1994:PFE

[2177] Bhimsen K. Shivamoggi and David K. Rollins. The Painlevé formulations and exact solutions of the nonlinear evolution equations for modulated gravity wave trains. Journal of Mathematical Physics, 35(9):4779–4798, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4779_s1.

Vaklev:1994:SSG

[2178] Y. Vaklev. Soliton solutions and gauge equivalence for the quadratic bundle. Journal of Mathematical Physics, 35(9):4799–4808, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p4799_s1.

Banach:1994:PTBa

[2179] Zbigniew Banach and Sławomir Piekarski. Perturbation theory based on the Einstein–Boltzmann system. I. Illustration of the theory for a Robertson–Walker geometry. Journal of Mathematical Physics, 35 (9):4809–4831, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4809_s1.

Carot:1994:MCI

[2180] J. Carot, J. da Costa, and E. G. L. R. Vaz. Matter collineations: The inverse “symmetry inheritance” problem. Journal of Mathematical Physics, 35(9):4832–4838, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p4832_s1.

Cheng:1994:DBH

[2181] Guang Jiun Cheng, Wei Fu Lin, and Rue Ron Hsu. Dyonic black holes in dilaton gravity. Journal of Mathematical Physics, 35(9):4839–4847, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4839_s1.

Hess:1994:EMS

[2182] Georg Heß. Exotic Majorana spinors in (3 + 1)-dimensional space–times. Journal of Mathematical Physics, 35(9):4848–4854, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4848_
s1.

Koike:1994:CHU

[2183] Tatsuhiko Koike, Masayuki Tanimoto, and Akio Hosoya. Compact homogeneous universes. Journal of Mathematical Physics, 35(9):4855–4888, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4855_s1.

McManus:1994:FDC

[2184] Des J. McManus. Five-dimensional cosmological models in induced matter theory. Journal of Mathematical Physics, 35(9):4889–4896, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p4889_s1.

Albeverio:1994:SRR

[2185] S. Albeverio and B. Torrésani. Some remarks on representations of jet groups and gauge groups. Journal of Mathematical Physics, 35(9):4897– 4908, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4897_s1.

Alekseevsky:1994:PSC

[2186] D. Alekseevsky, J. Grabowski, G. Marmo, and P. W. Michor. Poisson structures on the cotangent bundle of a Lie group or a principal bundle and their reductions. Journal of Mathematical Physics, 35(9):4909–4927, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p4909_s1.

Ballesteros:1994:FDQ

[2187] A. Ballesteros, F. J. Herranz, M. A. del Olmo, and M. Santander. Fourdimensional quantum affine algebras and space–time q-symmetries. Journal of Mathematical Physics, 35(9):4928–4940, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4928_s1.

Bender:1994:SSR

[2188] Carl M. Bender, Stefan Boettcher, and Moshe Moshe. Spherically symmetric random walks in noninteger dimension. Journal of Mathematical Physics, 35(9):4941–4963, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i9/p4941_s1.

Bordemann:1994:ACA

[2189] Martin Bordemann, Thomas Filk, and Claus Nowak. Algebraic classification of actions invariant under generalized flip moves of two-dimensional graphs. Journal of Mathematical Physics, 35(9):4964–4988, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p4964_s1.

Dowker:1994:FDS

[2190] J. S. Dowker. Functional determinants on spheres and sectors. Journal of Mathematical Physics, 35(9):4989–4999, September 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i9/p4989_s1. See er-

ratum [2308].

Hughes:1994:SWE

[2191] James Hughes. Spherical wave expansion for any spin. Journal of Mathematical Physics, 35(9):5000–5020, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i9/p5000_s1.

Juriev:1994:SHS

[2192] D. Juriev. Setting hidden symmetries free by the noncommutative Veronese mapping. Journal of Mathematical Physics, 35(9):5021–5024, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i9/p5021_s1.

Majid:1994:ESQ

[2193] S. Majid. q-Euclidean space and quantum Wick rotation by twisting. Journal of Mathematical Physics, 35(9):5025–5034, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p5025_
s1.

Maligranda:1994:IPD

[2194] Lech Maligranda, Lars Erik Persson, and John Wyller. Interpolation and partial differential equations. Journal of Mathematical Physics, 35 (9):5035–5046, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i9/p5035_s1.

Maslanka:1994:IRP

[2195] Paweł Maślanka. The induced representations of the κ-Poincaré group. Journal of Mathematical Physics, 35(9):5047–5056, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p5047_
s1.

Minnaert:1994:RSR

[2196] Pierre Minnaert and Stoyan Toshev. Racah sum rule and Biedenharn– Elliott identity for the super-rotation 6-j symbols. Journal of Mathematical Physics, 35(9):5057–5064, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i9/p5057_s1.

Pan:1994:GDA

[2197] Feng Pan. Generalized deformed algebras F (A_{1}) and their applications. Journal of Mathematical Physics, 35(9):5065–5073, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p5065_
s1.

ShafeiDehAbad:1994:AFQ

[2198] A. Shafei Deh Abad and V. Milani. Q-analytic functions on quantum spaces. Journal of Mathematical Physics, 35(9):5074–5086, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i9/p5074_s1.

Tagliani:1994:MEH

[2199] Aldo Tagliani. Maximum entropy in the Hamburger moments problem. Journal of Mathematical Physics, 35(9):5087–5096, September 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i9/p5087_
s1.

Witten:1994:SYM

[2200] Edward Witten. Supersymmetric Yang–Mills theory on a four-manifold. Topology and physics. Journal of Mathematical Physics, 35(10):5101– 5135, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i10/p5101_s1.

Crane:1994:FDT

[2201] Louis Crane and Igor B. Frenkel. Four-dimensional topological quantum field theory, Hopf categories, and the canonical bases. Topology and physics. Journal of Mathematical Physics, 35(10):5136–5154, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i10/p5136_s1.

Dai:1994:IDL

[2202] Xianzhe Dai and Daniel S. Freed. η-invariants and determinant lines. Topology and physics. Journal of Mathematical Physics, 35(10):5155– 5194, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i10/p5155_s1. See erratum [2322].

Chamseddine:1994:CSA

[2203] A. H. Chamseddine and J. Fröhlich. The Chern–Simons action in noncommutative geometry. Topology and physics. Journal of Mathematical Physics, 35(10):5195–5218, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i10/p5195_s1.

Rozansky:1994:RFJ

[2204] L. Rozansky. Reshetikhin’s formula for the Jones polynomial of a link: Feynman diagrams and Milnor’s linking numbers. Topology and physics. Journal of Mathematical Physics, 35(10):5219–5246, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i10/p5219_s1.

Bott:1994:SLK

[2205] Raoul Bott and Clifford Taubes. On the self-linking of knots. Topology and physics. Journal of Mathematical Physics, 35(10):5247–5287, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i10/p5247_s1.

Schwiebert:1994:ERE

[2206] Christian Schwiebert. Extended reflection equation algebras, the braid group on a handlebody, and associated link polynomials. Topology and physics. Journal of Mathematical Physics, 35(10):5288–5305, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i10/p5288_s1.

Durhuus:1994:CCU

[2207] Bergfinnur Durhuus and Thórdur Jónsson. Classification and construction of unitary topological field theories in two dimensions. Topology and physics. Journal of Mathematical Physics, 35(10):5306–5313, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i10/p5306_s1.

Broda:1994:PIA

[2208] Bogusław Broda. A path-integral approach to polynomial invariants of links. Topology and physics. Journal of Mathematical Physics, 35 (10):5314–5320, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i10/p5314_s1.

Aspinwall:1994:STT

[2209] Paul S. Aspinwall, Brian R. Greene, and David R. Morrison. Space–time topology change and stringy geometry. Topology and physics. Journal of Mathematical Physics, 35(10):5321–5337, October 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i10/p5321_s1.

King:1994:EDS

[2210] Christopher King and Ambar Sengupta. An explicit description of the symplectic structure of moduli spaces of flat connections. Topology and physics. Journal of Mathematical Physics, 35(10):5338–5353, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i10/p5338_s1.

King:1994:SLT

[2211] Christopher King and Ambar Sengupta. The semiclassical limit of the two-dimensional quantum Yang–Mills model. Topology and physics. Journal of Mathematical Physics, 35(10):5354–5361, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i10/
p5354_s1.

Thouless:1994:TIQ

[2212] D. J. Thouless. Topological interpretations of quantum Hall conductance. Topology and physics. Journal of Mathematical Physics, 35(10):5362– 5372, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i10/p5362_s1.

Bellissard:1994:NGQ

[2213] J. Bellissard, A. van Elst, and H. Schulz-Baldes. The noncommutative geometry of the quantum Hall effect. Topology and physics. Journal of Mathematical Physics, 35(10):5373–5451, October 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i10/p5373_s1.

Isham:1994:QTL

[2214] C. J. Isham and N. Linden. Quantum temporal logic and decoherence functionals in the histories approach to generalized quantum theory. Topology and physics. Journal of Mathematical Physics, 35(10):5452– 5476, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i10/p5452_s1.

Carlip:1994:TCD

[2215] S. Carlip and R. Cosgrove. Topology change in (2 + 1)-dimensional gravity. Topology and physics. Journal of Mathematical Physics, 35(10):5477– 5493, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i10/p5477_s1.

Brans:1994:ESP

[2216] Carl H. Brans. Exotic smoothness and physics. Topology and physics. Journal of Mathematical Physics, 35(10):5494–5506, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i10/
p5494_s1.

Gottlieb:1994:GLN

[2217] Daniel H. Gottlieb. A gravitational lens need not produce an odd number of images. Topology and physics. Journal of Mathematical Physics, 35(10):5507–5510, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i10/p5507_s1.

Damianou:1994:MHS

[2218] Pantelis A. Damianou. Multiple Hamiltonian structures for Toda-type systems. Topology and physics. Journal of Mathematical Physics, 35 (10):5511–5541, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i10/p5511_s1.

Adachi:1994:LRS

[2219] Tadayoshi Adachi. Long-range scattering for three-body Stark Hamiltonians. Journal of Mathematical Physics, 35(11):5547–5571, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5547_s1.

Burrows:1994:LAS

[2220] B. L. Burrows, M. Cohen, and Tova Feldmann. A Lie algebraic study of some Schrödinger equations. Journal of Mathematical Physics, 35 (11):5572–5581, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5572_s1.

Emch:1994:AAN

[2221] G. G. Emch, H. Narnhofer, W. Thirring, and G. L. Sewell. Anosov actions on noncommutative algebras. Journal of Mathematical Physics, 35(11):5582–5599, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5582_s1.

Friedberg:1994:NLF

[2222] R. Friedberg, T. D. Lee, and Y. Pang. A new lattice formulation of the continuum. Journal of Mathematical Physics, 35(11):5600–5629, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5600_s1.

Kholodnyi:1994:TPB

[2223] Valery A. Kholodnyi. On transformations preserving the basis conditions of a spin structure group in four-dimensional superstring theory in free fermionic formulation. Journal of Mathematical Physics, 35(11):5630– 5641, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p5630_s1.

Minzoni:1994:SYM

[2224] A. A. Minzoni, J. Muciño, and M. Rosenbaum. On the structure of Yang– Mills fields in compactified Minkowski space. Journal of Mathematical Physics, 35(11):5642–5659, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i11/p5642_s1.

Moses:1994:BEM

[2225] Harry E. Moses and Reese T. Prosser. Bounds for the exact matrix elements for radiation processes in hydrogenic atoms. II. Journal of Mathematical Physics, 35(11):5660–5668, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i11/p5660_s1.

Nergiz:1994:VOP

[2226] Serdar Nergiz. Vertex operators for physical states of bosonic string. Journal of Mathematical Physics, 35(11):5669–5673, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/
p5669_s1.

Slomczynski:1994:QCE

[2227] Wojciech Słomczyński and Karol Życzkowski. Quantum chaos: An entropy approach. Journal of Mathematical Physics, 35(11):5674–5700, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p5674_s1. See erratum [2314].

Wehrhahn:1994:PDG

[2228] Rodolfo Francisco Wehrhahn. Potentials derived geometrically for symmetry scattering. Journal of Mathematical Physics, 35(11):5701–5717, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p5701_s1.

Debbasch:1994:GPG

[2229] Fabrice Debbasch. Global potentials for general relativistic flows of ideal charged fluids. Journal of Mathematical Physics, 35(11):5718–5725, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p5718_s1.

Kichenassamy:1994:NME

[2230] S. Kichenassamy and R. A. Krikorian. Note on Maxwell’s equations in relativistically rotating frames. Journal of Mathematical Physics, 35 (11):5726–5733, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5726_s1.

Krause:1994:CSG

[2231] J. Krause. On the complete symmetry group of the classical Kepler system. Journal of Mathematical Physics, 35(11):5734–5748, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5734_s1.

Actor:1994:IPP

[2232] Alfred Arthur Actor. Infinite products, partition functions, and the Meinardus theorem. Journal of Mathematical Physics, 35(11):5749–5764, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p5749_s1.

Berge:1994:DSA

[2233] Luc Bergé. Dynamical stability analysis of strong/weak wave collapses. Journal of Mathematical Physics, 35(11):5765–5780, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/
p5765_s1.

Bonora:1994:TKV

[2234] L. Bonora and C. S. Xiong. The (N, M)th Korteweg–de Vries hierarchy and the associated W-algebra. Journal of Mathematical Physics, 35 (11):5781–5819, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5781_s1.

Pelinovsky:1994:RSK

[2235] Dmitry Pelinovsky. Rational solutions of the Kadomtsev–Petviashvili hierarchy and the dynamics of their poles. I. New form of a general rational solution. Journal of Mathematical Physics, 35(11):5820–5830, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p5820_s1.

Scott:1994:CFB

[2236] D. R. D. Scott. Classical functional Bethe ansatz for SL(N): Separation of variables for the magnetic chain. Journal of Mathematical Physics, 35 (11):5831–5843, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5831_s1.

Shiota:1994:CMH

[2237] Takahiro Shiota. Calogero–Moser hierarchy and KP hierarchy. Journal of Mathematical Physics, 35(11):5844–5849, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/p5844_s1.

Strampp:1994:KBS

[2238] Walter Strampp. The Kaup–Broer system and trilinear ODEs of Painlevé type. Journal of Mathematical Physics, 35(11):5850–5861, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5850_s1.

Tatar:1994:ASB

[2239] Liviu Tătar and Radu Tătar. Algebraic structure of the BRST symmetry. Journal of Mathematical Physics, 35(11):5862–5868, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5862_s1.

Zhang:1994:SVC

[2240] You Jin Zhang and Yi Cheng. Solutions for the vector k-constrained KP hierarchy. Journal of Mathematical Physics, 35(11):5869–5884, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5869_s1.

Banach:1994:PTBb

[2241] Zbigniew Banach and Sławomir Piekarski. Perturbation theory based on the Einstein–Boltzmann system. II. Illustration of the theory for an almost-Robertson–Walker geometry. Journal of Mathematical Physics, 35(11):5885–5907, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5885_s1.

Defever:1994:PST

[2242] Filip Defever, Ryszard Deszcz, Leopold Verstraelen, and Luc Vrancken. On pseudosymmetric space–times. Journal of Mathematical Physics, 35 (11):5908–5921, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5908_s1.

Dray:1994:PDT

[2243] Tevian Dray and Charles Hellaby. The patchwork divergence theorem. Journal of Mathematical Physics, 35(11):5922–5929, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/
p5922_s1.

Hall:1994:ACG

[2244] G. S. Hall, D. J. Low, and J. R. Pulham. Affine collineations in general relativity and their fixed point structure. Journal of Mathematical Physics, 35(11):5930–5944, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i11/p5930_s1.

Kelnhofer:1994:UBG

[2245] Gerald Kelnhofer. Universal bundle for gravity, local index theorem, and covariant gravitational anomalies. Journal of Mathematical Physics, 35 (11):5945–5968, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p5945_s1.

Naber:1994:TNF

[2246] M. G. Naber and E. N. Glass. Taub numbers at future null infinity. II. Journal of Mathematical Physics, 35(11):5969–5989, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/
p5969_s1.

Plebanski:1994:EHE

[2247] J. F. Plebański and M. Przanowski. Evolution hyperheavenly equations. Journal of Mathematical Physics, 35(11):5990–6000, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p5990_s1.

Zhytnikov:1994:WES

[2248] V. V. Zhytnikov. Wavelike exact solutions of R+R^{2}+Q^{2} gravity. Journal of Mathematical Physics, 35(11):6001–6017, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/p6001_s1.

Anderson:1994:SFQ

[2249] Arlen Anderson. Special functions from quantum canonical transformations. Journal of Mathematical Physics, 35(11):6018–6035, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p6018_s1.

Berman:1994:RME

[2250] C. Leonard Berman and Leslie Greengard. A renormalization method for the evaluation of lattice sums. Journal of Mathematical Physics, 35 (11):6036–6048, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p6036_s1.

Cortes:1994:LDE

[2251] J. L. Cortés and M. S. Plyushchay. Linear differential equations for a fractional spin field. Journal of Mathematical Physics, 35(11):6049– 6057, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p6049_s1.

Dobrev:1994:RTA

[2252] V. K. Dobrev, P. Truini, and L. C. Biedenharn. Representation theory approach to the polynomial solutions of q-difference equations: U_{q}(sl(3)) and beyond. Journal of Mathematical Physics, 35(11):6058– 6075, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p6058_s1.

Dowker:1994:ZFA

[2253] J. S. Dowker. A zeta function approach to the relation between the numbers of symmetry planes and axes of a polytope. Journal of Mathematical Physics, 35(11):6076–6095, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i11/p6076_s1.

ElHassouni:1994:DEA

[2254] A. El Hassouni, Y. Hassouni, and M. Zakkari. Deformation of the exterior algebra and the GL_{q}(r, C). Journal of Mathematical Physics, 35 (11):6096–6099, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p6096_s1.

Elizalde:1994:AIG

[2255] E. Elizalde. Analysis of an inhomogeneous generalized Epstein–Hurwitz zeta function with physical applications. Journal of Mathematical Physics, 35(11):6100–6122, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i11/p6100_s1.

Finkelstein:1994:OR

[2256] Robert J. Finkelstein. The orthogonality relations for SU_{q}(2). Journal of Mathematical Physics, 35(11):6123–6143, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/p6123_s1.

Fleury:1994:PRH

[2257] Norbert Fleury and Alexander Turbiner. Polynomial relations in the Heisenberg algebra. Journal of Mathematical Physics, 35(11):6144–6149, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p6144_s1.

Fu:1994:FDI

[2258] Hong Chen Fu and Xiang Sheng Li. Finite-dimensional irreducible representations of reflection algebra for su(1, 1)_{q} vertex model: classification, boson-fermion realizations, and Yang–Baxterization. Journal of Mathematical Physics, 35(11):6150–6157, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i11/p6150_s1.

Hasegawa:1994:OBM

[2259] Koji Hasegawa. L-operator for Belavin’s R-matrix acting on the space of theta functions. Journal of Mathematical Physics, 35(11):6158–6171, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p6158_s1.

Katriel:1994:NGT

[2260] Jacob Katriel and Allan I. Solomon. A no-go theorem for a Lieconsistent q-Campbell–Baker–Hausdorff expansion. Journal of Mathematical Physics, 35(11):6172–6178, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i11/p6172_s1.

Kowalenko:1994:AKF

[2261] Victor Kowalenko and N. E. Frankel. Asymptotics for the Kummer function of Bose plasmas. Journal of Mathematical Physics, 35(11):6179– 6198, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i11/p6179_s1.

Latrach:1994:SRE

[2262] Khalid Latrach. Some remarks on the essential spectrum of transport operators with abstract boundary conditions. Journal of Mathematical Physics, 35(11):6199–6212, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i11/p6199_s1.

Wambst:1994:QKC

[2263] Marc Wambst. Quantum Koszul complexes for Majid’s braided Lie algebras. Journal of Mathematical Physics, 35(11):6213–6223, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i11/p6213_s1.

Nelson:1994:EGL

[2264] Robert A. Nelson. Erratum: “Generalized Lorentz transformation for an accelerated, rotating frame of reference”, [J. Math. Phys. 28 (1987), no. 10, 2379–2383; MR 89c:83006]. Journal of Mathematical Physics, 35 (11):6224–6225, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p6224_s1. See [5].

Paveri-Fontana:1994:EHH

[2265] Stefano L. Paveri-Fontana and Paul F. Zweifel. Erratum: “The half-Hartley and half-Hilbert transforms”, [J. Math. Phys. 35, 2648 (1994)]. Journal of Mathematical Physics, 35(11):6226, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i11/

p6226_s1. See [2031].

Aktosun:1994:BSI

[2266] Tuncay Aktosun. Bound states and inverse scattering for the Schrödinger equation in one dimension. Journal of Mathematical Physics, 35(12): 6231–6236, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6231_s1.

Ali:1994:MAG

[2267] S. Twareque Ali, Amine M. El Gradechi, and Gérard G. Emch. Modular algebras in geometric quantization. Journal of Mathematical Physics, 35 (12):6237–6243, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6237_s1.

Bessis:1994:PFS

[2268] N. Bessis and G. Bessis. Perturbed factorization of the singular anharmonic-oscillator eigenequation. Journal of Mathematical Physics, 35(12):6244–6269, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6244_s1.

Cheng:1994:BSS

[2269] G. Cheng and T. K. Kuo. Bifurcation solutions of the Schwinger–Dyson equation. Journal of Mathematical Physics, 35(12):6270–6290, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i12/p6270_s1.

Gielerak:1994:KSM

[2270] R. Gielerak, L. Jakóbczyk, and R. Olkiewicz. W^{∗}–KMS structure from multi-time Euclidean Green functions. Journal of Mathematical Physics, 35(12):6291–6303, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6291_s1.

Grigore:1994:FFS

[2271] D. R. Grigore. Free fields for any spin in 1 + 2 dimensions. Journal of Mathematical Physics, 35(12):6304–6331, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6304_s1.

Gudder:1994:ESN

[2272] Stanley Gudder. Event structures in nonstandard quantum mechanics. Journal of Mathematical Physics, 35(12):6332–6343, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/
p6332_s1.

Haba:1994:SIF

[2273] Z. Haba. Stochastic interpretation of Feynman path integral. Journal of Mathematical Physics, 35(12):6344–6359, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6344_s1.

Isham:1994:CDF

[2274] C. J. Isham, N. Linden, and S. Schreckenberg. The classification of decoherence functionals: An analog of Gleason’s theorem. Journal of Mathematical Physics, 35(12):6360–6370, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i12/p6360_s1.

Monaco:1994:NAJ

[2275] R. L. Monaco and E. Capelas de Oliveira. A new approach for the Jeffreys–Wentzel–Kramers–Brillouin theory. Journal of Mathematical Physics, 35(12):6371–6378, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6371_s1.

Mourad:1994:TFR

[2276] J. Mourad and H. Sazdjian. The two-fermion relativistic wave equations of constraint theory in the Pauli–Schrödinger form. Journal of Mathematical Physics, 35(12):6379–6406, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i12/p6379_s1.

Navarro:1994:OMQ

[2277] Miguel Navarro, Julio Guerrero, and Víctor Aldaya. Optics, mechanics, and quantization of reparametrization-invariant systems. Journal of Mathematical Physics, 35(12):6407–6417, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6407_s1.

Niemi:1994:EBS

[2278] Antti J. Niemi and Olav Tirkkonen. Equivariance, BRST symmetry, and superspace. Journal of Mathematical Physics, 35(12):6418–6433, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6418_s1.

Ashtekar:1994:AED

[2279] Abhay Ashtekar and Ranjeet S. Tate. An algebraic extension of Dirac quantization: Examples. Journal of Mathematical Physics, 35(12):6434– 6470, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6434_s1.

Kamp:1994:FRP

[2280] Léon P. J. Kamp. Fully relativistic plasma dispersion function. Journal of Mathematical Physics, 35(12):6471–6488, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6471_s1.

Moorhouse:1994:NTP

[2281] T. Moorhouse and R. S. Ward. Numerical twistor procedure for solving a nonlinear field equation. Journal of Mathematical Physics, 35(12):6489– 6497, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6489_s1.

Daniel:1994:IIS

[2282] M. Daniel, K. Porsezian, and M. Lakshmanan. On the integrability of the inhomogeneous spherically symmetric Heisenberg ferromagnet in arbitrary dimensions. Journal of Mathematical Physics, 35(12):6498– 6510, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6498_s1. See erratum [2319].

Huillet:1994:CSS

[2283] Thierry Huillet and Bernard Jeannet. On a class of “skewed” self-similar and hyperbolic fractals. Journal of Mathematical Physics, 35(12):6511– 6524, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6511_s1.

Baikov:1994:COT

[2284] V. A. Baĭkov, R. K. Gazizov, N. H. Ibragimov, and F. M. Mahomed. Closed orbits and their stable symmetries. Journal of Mathematical Physics, 35(12):6525–6535, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6525_s1.

Chaichian:1994:NIC

[2285] Masud Chaichian and Domingo Louis Martinez. On the Noether identities for a class of systems with singular Lagrangians. Journal of Mathematical Physics, 35(12):6536–6545, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i12/p6536_s1.

deLeon:1994:SRH

[2286] M. de León, Paulo Pitanga, and P. R. Rodrigues. Symplectic reduction of higher order Lagrangian systems with symmetry. Journal of Mathematical Physics, 35(12):6546–6556, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v35/i12/p6546_s1.

Krupkova:1994:GSH

[2287] Olga Krupková. A geometric setting for higher-order Dirac–Bergmann theory of constraints. Journal of Mathematical Physics, 35(12):6557– 6576, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6557_s1.

Ranada:1994:NIT

[2288] Manuel F. Rañada. A new integrable Toda-related three-particle system. Journal of Mathematical Physics, 35(12):6577–6583, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i12/p6577_s1.

Sardanashvily:1994:CFS

[2289] G. Sardanashvily. Constraint field systems in multimomentum canonical variables. Journal of Mathematical Physics, 35(12):6584–6603, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i12/p6584_s1.

Steeb:1994:PLA

[2290] W.-H. Steeb and R. Stoop. Padé-like approximation and maximum entropy formalism applied to chaotic systems. Journal of Mathematical Physics, 35(12):6604–6608, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6604_s1.

Wu:1994:FSM

[2291] Ju Hao Wu and Ke Lin Wang. The function-series method and the coupled Boussinesq-like equations. Journal of Mathematical Physics, 35 (12):6609–6618, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6609_s1.

Bhui:1994:RSK

[2292] B. Bhui, Madhumita Barua Basu, and S. Chatterjee. Radiating sphere in a Kaluza–Klein space–time. Journal of Mathematical Physics, 35 (12):6619–6627, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6619_s1.

Cornish:1994:NGB

[2293] Neil J. Cornish and John W. Moffat. Nonsingular gravity without black holes. Journal of Mathematical Physics, 35(12):6628–6643, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i12/p6628_s1.

Manko:1994:MTA

[2294] V. S. Man′ko, J. Martín, and E. Ruiz. Metric of two arbitrary Kerr– Newman sources located on the symmetry axis. Journal of Mathematical Physics, 35(12):6644–6657, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6644_s1.

Arefeva:1994:BCP

[2295] I. Ya. Aref′eva, G. E. Arutyunov, and P. B. Medvedev. Bicovariant 3D calculus for SL_{q}(2) from Poisson–Lie structures. Journal of Mathematical Physics, 35(12):6658–6671, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6658_s1.

Biedenharn:1994:SPM

[2296] L. C. Biedenharn, M. A. Lohe, and H. T. Williams. Symmetry properties of matrix elements of canonical SU(3) tensor operators. Journal of Mathematical Physics, 35(12):6672–6684, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6672_s1.

Bremner:1994:UCE

[2297] Murray Bremner. Universal central extensions of elliptic affine Lie algebras. Journal of Mathematical Physics, 35(12):6685–6692, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i12/p6685_s1.

Cheng:1994:GPI

[2298] G. Cheng and T. K. Kuo. A generalized parameter imbedding method. Journal of Mathematical Physics, 35(12):6693–6702, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/
p6693_s1.

Dimakis:1994:DDC

[2299] Aristophanes Dimakis and Folkert Müller-Hoissen. Discrete differential calculus: Graphs, topologies, and gauge theory. Journal of Mathematical Physics, 35(12):6703–6735, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6703_s1.

Forrester:1994:CWM

[2300] P. J. Forrester and T. D. Hughes. Complex Wishart matrices and conductance in mesoscopic systems: Exact results. Journal of Mathematical Physics, 35(12):6736–6747, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6736_s1.

Frontini:1994:MEF

[2301] M. Frontini and A. Tagliani. Maximum entropy in the finite Stieltjes and Hamburger moment problem. Journal of Mathematical Physics, 35 (12):6748–6756, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6748_s1.

Gould:1994:QAL

[2302] Mark D. Gould and Yao Zhong Zhang. Quantum affine Lie algebras, Casimir invariants, and diagonalization of the braid generator. Journal of Mathematical Physics, 35(12):6757–6773, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6757_s1.

Grundland:1994:IKT

[2303] A. M. Grundland, M. Kovalyov, and M. Sussman. Interaction of kinktype solutions of the harmonic map equations. Journal of Mathematical Physics, 35(12):6774–6783, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6774_s1.

Isaev:1994:QGC

[2304] A. P. Isaev. Quantum group covariant noncommutative geometry. Journal of Mathematical Physics, 35(12):6784–6801, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6784_s1.

Kempf:1994:AIF

[2305] Achim Kempf and Shahn Majid. Algebraic q-integration and Fourier theory on quantum and braided spaces. Journal of Mathematical Physics, 35(12):6802–6837, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i12/p6802_s1.

Post:1994:CGL

[2306] Gerhard Post. A class of graded Lie algebras of vector fields and first order differential operators. Journal of Mathematical Physics, 35(12):6838– 6856, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v35/i12/p6838_s1.

Suzuki:1994:MRU

[2307] Takashi Suzuki. More on U_{q}(su(1, 1)) with q a root of unity. Journal of Mathematical Physics, 35(12):6857–6874, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v35/i12/p6857_s1.

Dowker:1995:EFD

[2308] J. S. Dowker. Erratum: “Functional determinants on spheres and sectors”. Journal of Mathematical Physics, 36(2):988, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2190].

Lammerzahl:1995:ADS

[2309] Claus Lämmerzahl. Addendum to: “On the dimensionality of space– time”, [J. Math. Phys. 34 (1993), no. 10, 4540–4553; MR 94e:81081] by Lämmerzahl and A. Macías. Journal of Mathematical Physics, 36(4): 2131, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1725].

Fatibene:1995:ENF

[2310] L. Fatibene, M. Ferraris, and M. Francaviglia. Erratum: “Nöther formalism for conserved quantities in classical gauge field theories”, [J. Math. Phys. 35 (1994), no. 4, 1644–1657; MR 95g:58064]. Journal of Mathematical Physics, 36(6):3183–3185, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1953].

LaChapelle:1995:EEG

[2311] J. LaChapelle. Erratum: “U(2) as the electroweak gauge group”, [J. Math. Phys. 35 (1994), no. 5, 2186–2198; MR 95e:81253]. Journal of Mathematical Physics, 36(6):3186, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1992].

LaChapelle:1995:EGM

[2312] J. LaChapelle. Erratum: “Generating mass without the Higgs particle”, [J. Math. Phys. 35 (1994), no. 5, 2199–2209; MR 95e:81257]. Journal of Mathematical Physics, 36(6):3187, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1993].

Chu:1995:EPF

[2313] Wen Chang Chu. Erratum: “Partial fractions and bilateral summations”, [J. Math. Phys. 35 (1994), no. 4, 2036–2042; MR 95f:33004]. Journal of Mathematical Physics, 36(9):5198–5199, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1982].

Slomczynski:1995:EQC

[2314] Wojciech Słomczyński and Karol Życzkowski. Erratum: “Quantum chaos: an entropy approach”, [J. Math. Phys. 35 (1994), no. 11, 5674– 5700; MR 95i:81062]. Journal of Mathematical Physics, 36(9):5201, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2227].

Navarro:1995:CGP

[2315] Miguel Navarro. Comments on Good’s proposal for new rules of quantization. Journal of Mathematical Physics, 36(12):6665–6672, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2081].

Bertolotti:1996:CRC

[2316] R. Bertolotti, G. Contreras, L. A. Núñez, U. Percoco, and J. Carot. Comment on: “Ricci collineations of static spherically symmetric spacetimes”, [J. Math. Phys. 35 (1994), no. 6, 3005–3012; MR 95d:83017] by M. J. Amir, A. H. Bokhari and A. Qadir. Journal of Mathematical Physics, 37(2):1086–1088, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2055].

Amir:1996:ERC

[2317] M. Jamil Amir, Ashfaque H. Bokhari, and Asghar Qadir. Erratum: “Ricci collineations of static spherically symmetric spacetimes”, [J. Math. Phys. 35 (1994), no. 6, 3005–3012; MR 95d:83017]. Journal of Mathematical Physics, 37(2):1089, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2055].

Enginer:1996:EFS

[2318] Y. Enginer, M. Hortaçsu, and N. Özdemir. Errata: “Fluctuations of a spherical gravitational impulsive wave”, [J. Math. Phys. 34 (1993), no. 2, 690–699; MR 93m:83024] by M. Hortaçsu and “Fluctuations of a rotating impulsive wave”, [ibid. 35 (1994), no. 6, 3043–3050; MR 95g:83036] by M. Hortaçsu and R. Kaya. Journal of Mathematical Physics, 37(8):4199, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1450, 2059].

Daniel:1996:EII

[2319] M. Daniel, K. Porsezian, and M. Lakshmanan. Erratum: “On the integrability of the inhomogeneous spherically symmetric Heisenberg ferromagnet in arbitrary dimensions”, [J. Math. Phys. 35 (1994), no. 12, 6498–6510; MR 95h:82022]. Journal of Mathematical Physics, 37(9):4768, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2282].

Halilsoy:2000:CCN

[2320] M. Halilsoy. Comment on “Colliding null shells of matter”, [J. Math. Phys. 33, 1065 (1992)]. Journal of Mathematical Physics, 41(12):8351– 8353, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1041, 2321].

Wang:2000:RCC

[2321] Anzhong Wang. Response to “Comment on ‘Colliding null shells of matter’ ”, [J. Math. Phys. 41, 8351 (2000)]. Journal of Mathematical Physics, 41(12):8354–8355, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1041, 2320].

Dai:2001:EID

[2322] Xianzhe Dai and Daniel S. Freed. Erratum: “η-invariants and determinant lines”, [J. Math. Phys. 35, 5155 (1994)]. Journal of Mathematical Physics, 42(5):2343–2344, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2202].